A Complete Bibliography of ACM Transactions on Software Engineering and Methodology

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

24 April 2024 Version 1.81

Title word cross-reference

1 [JFJ^{+}23]. k [GHX^{+}24]. N [RKBL19, JFJ^{+}23]. z [LCZL14].

-Equivalent [LCZL14]. -to- [JFJ^{+}23]. -way [RKBL19].

19 [FSZ^{+}22, MKR^{+}23, RHV24, RHvB24].

2-year [RHV24, RHvB24]. 2.0 [DKD21]. 2002 [Ano02]. 2013 [HP15].

Abandoned [KCAS23]. Abbreviated [SRTR17]. ABC [SF18]. Able [HNRA20]. Abnormal [ZCL^{+}24]. Abstract [PSZ21, XMA^{+}14, Jac95a, Pon02].

Abstracting [Gun00]. Abstraction [AEK^{+}16, ZNZ^{+}23, CF03, Egy02]. Abstractions [MCM24, PSZ21, BG98, DBGU13]. Abstractive [SFC^{+}24]. Academia [MS22]. Accelerating [LSX^{+}24, MUB^{+}23]. Acceleration [LHX21]. Access [LSX^{+}24, BDL06]. Accessibility [BXL^{+}22, SMY^{+}24]. accommodates [YHR92]. Accounting [SM12]. Accuracy [ASNB19, CWW^{+}20, CSFB24, ODE21, WYC^{+}24, WWAC23]. Accurate [HSB^{+}23, TLL^{+}23, ZLL^{+}23]. Achieving [BJMH02, HAB13, LBZ14]. Acknowledgement [ACM05]. ACM [NP08]. Acrobats [RFDS24]. across [CSV13, SXX^{+}24]. action [HN98]. Active [ASNB19, MS15]. ActivFORMS [WI23].

Activities [HGHM23, NKL^{+}24, SHLW21, WFW^{+}20]. Activity [KS22, Esh06]. Activity-Based [KS22]. activitycharts [BRG^{+}01]. Actor [CCRZ23]. Actor-Driven [CCRZ23]. Ada [Dil93, Dil97, DBDS94, YTL^{+}95]. ADAM [GL14]. adapt [DPT13]. Adaptation [SEM17, ZH24]. Adaptive [BS16, CLBY18, DR11, HDG^{+}24, RPT^{+}22, SMBO21, ZA22, HWH14, VTA04, WI23]. Addendum [HT98]. Address [Zav04]. Addresses [ZWGX22]. Adequacy [DDS23, GRS^{+}16, KFY23, KSD08]. Adequate [GGZ^{+}15]. Admitted [RXX^{+}19, GLL^{+}21]. Adonis [LYL^{+}24]. Adopting [WT24a, WT24b, SAB^{+}14]. Adoption [RSBO23]. Adversarial [KL21, ZCL^{+}24, ZZS^{+}22]. Affected [PLMC23, VCF21]. against [EWS14, GSC24, IC14, LZS^{+}24, ZSWS23]. Age [PHS^{+}24]. Agents [BT24, MPR06]. Aggregation [SPAS21]. aggregator [BPT10]. Agile [RFDS24, Rus21, CF10]. agnostic [DLSC23]. AI [MFBF^{+}22, OHDB92, PHS^{+}24]. AI-based [OHDB92, MFBF^{+}22]. Aide [PWB23]. Aide-mémoire [PWB23]. aided [SB06]. AIOps [LJL^{+}20, LLS^{+}21, LRL^{+}22]. algebra [ZB13]. Algebraic [SZH^{+}19, HRD08, PPP94]. algebras [BCD02]. Algorithm [CYM^{+}24, LD24, ZMC^{+}24, BS07, YHR92]. Algorithms [DBNG15, ARL^{+}15, HT17, HT98, HVT98, KK93]. aliasing [FYD^{+}08]. Alignment [CXH23, UFG14]. Alignments [TC20]. Alike [KFY24]. Allocation [SMBO21]. alloy [FPB^{+}05, MPF14, Jac02]. alone [UGKR22]. Alphabet [FDB^{+}12]. Alternative [TTL^{+}21]. Alternatives [KDM17, KDM23]. Always [HDKM23]. Always-on [HDKM23]. American [FMMB23b, FMMB23a]. Amoeba [DCS09]. among [GL11, KDM23]. Amplification [PLP^{+}23]. Amplifying [ZE14]. Analyses [CST16, CTA^{+}21, RKBL19, LS13]. Analysis [AMS^{+}18, AEK^{+}16, AGRR19, AFPB23, ASJA23, BK24, Bjø19, CCW^{+}23, CYM^{+}24, CTC^{+}23, CCXZ23, CFL^{+}16, DG17, EHF20, FCLL21, FTF22, FTF24, HNRA20, HLX23a, JFJ^{+}23, JZL^{+}19b, JA22, KJHY22, LKRF22, LCZL14, LASA24, LHX21, DDGR18, LM24, MGTR18, PHS^{+}24, PBCT10, RHV24, RHvB24, SEM17, SBS^{+}23, SGD15, SLB^{+}21, TDL^{+}24, VLJ^{+}18, WWY^{+}23, YXX^{+}23, YXK^{+}17, YBL15, ZHS^{+}23, ZZC^{+}23, ZHXZ24, ZWY^{+}21, BP98, BGO^{+}14, CS12, CK96, CK99, Cor00, CSX08, DRW96, Dev99, DBDS94, DCCN04, For94, GL14, GM01, GSH97, HKMB^{+}14, Hie06, HH95, HZZ13, HT98, HVT98, HN98, LH08, LH02, MRR05, NP08, OO92, PTY95, PGM12, RM03, Rob08, SGG^{+}14, Sne96, SRK06, TPT13, WP93, XCKX13, YTL^{+}95, YBL13, FPGA07, MVM07]. analyze [CFM00]. Analyzing [LTX19, OL22, DFG00]. Anatomy [SGhWG22]. Anchor [WWY^{+}23]. Android [ABC^{+}22, GHM18, GJS^{+}23, LVBBC^{+}18, LZF^{+}24, MCKG24, PLP^{+}23, PGZ^{+}20, RMCT22, SLB^{+}21, SCL^{+}23, WCG^{+}21, ZLW^{+}21, ZSL^{+}22, ZWY^{+}21]. annotation [KATS12]. annotation-based [KATS12]. Annotations [IC14, ODE21]. announcement [SPAK10]. Announcing [Pez23]. Anomalies [ZHZ^{+}21]. Anomaly [XAY23]. Ansible [RRPW21]. Answer [GXLG21, YHG^{+}24]. answering [KM10]. Anti [MS22]. Anti-patterns [MS22]. Apache [FAP22, MFH02]. API [Arc19, ARG17, HSX^{+}24, LLT^{+}24, PDHR23, SIB23, UKR21, WJZ^{+}23, XSA^{+}24, ZWY^{+}21]. APIRec [LLT^{+}24]. APIRO [SIB23]. APIs [AG22, BLS^{+}23, GZA24, HCJ^{+}24, LZF^{+}24, MZA22, SPK14, ZA22, ZBA23, ZA23, ZAL^{+}23]. Apollo [KSFC24]. App [Cai20, CCH^{+}21, GZX^{+}22, HDG^{+}24, JZL^{+}19a, NPB22, WCG^{+}21]. APPL [SHO95]. APPL/A [SHO95].

Applicability [RFR23]. Applicable [VHNF22]. Application [DBDS94, LWC^{+}23, MFLL12, MS15, SZZ^{+}23, ZE14, LH02]. Application-level [LWC^{+}23]. Applications [CGPP15, DG17, ETM22, GTD21, KAT12, LCL^{+}23, MBH^{+}17, NBB15, PBC24, YXM^{+}23, Yu23, ZOJH21, BM07, BCFM06, CPPRM03, CDP04, GGLT07, MZ09, MGMM11, PWX14, PBCT10, SS06, WGSD07, XM07]. Applying [FLM^{+}98, FTF22]. Approach [ASNB19, BCGB21, BZSW14, BMM^{+}17, BLBB23, BSA22, CLL^{+}22, CSW21, DL11, FDC^{+}21, GXSC21, HZZ^{+}14, HB22, JS23, MFLL12, LDUD13, SMBO21, SDG^{+}24, TG24, TCDZ19, WI23, YB20, YBL13, ZSL^{+}13, ZLL^{+}23, ZZC^{+}23, CTCC98, CRST12, CPPRM03, DFB99a, DFB99b, DvdHT05, DPT13, DF94, FK96, FPB^{+}05, Hen97, LASL13, MZ09, OHDB92, QRLV12, TD01, ZZL^{+}06]. Approaches [ABC^{+}22, PKMY24, XL20, ZLP^{+}24, YHC13]. approximation [JGB12]. Apps [CCH^{+}21, HHC^{+}24, LVBBC^{+}18, MCKG24, PLP^{+}23, RMCT22, SLB^{+}21, SCL^{+}23, ZSL^{+}22]. APTL [WME93]. Arachne [SKY23]. Architecting [BCD02, YBZK21]. Architectural [CSW21, MFLL12, AG97, AG98, LWF03, RVMRM04]. Architecture [AMGBK22, BNB14, HWH14, QWG^{+}23, SMBK22, YHG^{+}24, ZCT18, AAG95, BCC92, DvdHT05, GL14]. Architecture-centric [HWH14]. Architecture-Implementation [ZCT18]. Architecture-Level [BNB14]. Architecture-Oriented [QWG^{+}23]. Architectures [SBMK21, MRRR02]. ArchRepair [QWG^{+}23]. ARCTURUS [ZHXZ24]. Area [LWD^{+}24]. Aria [DRW96]. Arithmetic [GXSC21]. Array [NKWF14]. artifact [DFOT07]. ASM [BZSW14]. Aspect [Jac95a, SXX^{+}24, DR10, HE13, KF07, RS09, SGR^{+}10]. aspect- [RS09]. Aspect-level [SXX^{+}24]. aspect-oriented [DR10, HE13, SGR^{+}10]. AspectJ [HZS08, MHK11]. Aspects [GCX^{+}22, HGHM23, RSBO23, SZH^{+}19, BTI14]. Aspirations [CL23a]. Assembly [AM04, WMW^{+}23]. Assessing [ACF97, CM23, Cai20, CXH23, RST^{+}14, VBZ^{+}18, YHL^{+}22, ZWL^{+}24]. Assessment [CCRZ23, CMM^{+}15, CSFB24, HAY^{+}23, LWWM22, GJ08, PJRR10, SR05, Wey96]. assigned [BHL11]. Assignments [AFY^{+}22]. assistance [GN93]. assume [CAC08]. assume-guarantee [CAC08]. Assumptions [HGW^{+}16, IWY00]. Asteria [YDX^{+}24]. Asteria-Pro [YDX^{+}24]. ASTOOT [DF94]. Atomic [Liu22]. Atomicity [MWP^{+}21]. aToucan [YBL15]. Attack [WWZ^{+}22, YXM^{+}23, ZCL^{+}24, LLZ^{+}24]. Attacks [AMGBK22, EWS14, GSC24, PBC24, WCB^{+}20]. attribute [KK93]. Attributes [RFR23, GL14]. Augmenting [DPB17, GCX^{+}22]. Author [Ano96, TAW95]. Autoencoder [CCX^{+}20]. Autoencoders [KFY24]. Automated [Arc19, AG20, ASNB19, BLBB23, CSW21, DZV^{+}23, Egy02, EMM24, FA14, FSM^{+}15, FNT^{+}24, GM01, GN93, HJL96, HAB^{+}20, HCW^{+}22, HDG^{+}24, HGX^{+}24b, JO15, KSFC24, LKRF22, DGD^{+}19, LZL^{+}23, MZA22, MS14, NPB22, PGZ^{+}20, RR23, RKBL19, STDB23, STGR21, SIB23, TZZ^{+}23, TM14, WB13, WC23, XHL^{+}24, YLM^{+}23, YBL15, ZFM^{+}24, ZSL^{+}22, ZSW^{+}22, ZWT^{+}24, ZH24, BGO^{+}14, CS12, CAC08, FN03, FRB^{+}06, TBS92, XM07, XM08]. Automatic [BAJA23, CGPP15, CF03, DSV03, MGTR18, NBMK22, UKR21, BRRP05, DO93]. Automatically [CMM^{+}15, CDKP21, FFM24, LZF^{+}24, Mem08, YJW^{+}20, LS13]. Automating [GZX^{+}22]. Automation [NAF^{+}23]. Automaton [EM18]. Autonomous [AMV23, FDC^{+}21, GGPR24, LKA^{+}23, NA24, SER24]. Autotuning [ZHC24]. Avoiding [Hie06]. Aware [HAY^{+}23, HGC^{+}22, HQM^{+}23, KAT12, LWWM22, MS15, QHM^{+}23, VLJ^{+}18, WYW^{+}22, CYA22, DFB99a, DFB99b, HZB^{+}24, MGMM11, PKMY24, WYMW20, WXL^{+}21, WMW^{+}23, ZMC^{+}24]. Awareness [PHS^{+}24, SZZ^{+}23].

B [SB06]. back [Not13]. Backdoor [MRO^{+}23, SLWZ24]. Bad [NMA^{+}23, SWW^{+}24]. Barriers [GGHM24]. Based [AB12, ARL^{+}15, ASMP16, AFHC22, BMM^{+}17, BC23, CLL^{+}22, CYM^{+}24, DDE11, DTC^{+}24, EMM24, FFM24, GZX^{+}22, GGZ^{+}15, GCH^{+}24, JZL^{+}19b, KS22, KCMS24, LXL^{+}22, LHX21, MWP^{+}21, MFBF^{+}22, OKS^{+}16, SMBO21, SGD15, SNM23, SKY23, SVDHB23, TLP^{+}22, TSPRC18, WB13, WPX^{+}23, WJZ^{+}23, XHL^{+}24, XLD^{+}24, YXK^{+}17, ZTP^{+}23, ZA22, ZAL^{+}23, ZLW^{+}21, ZWY^{+}21, AAP^{+}20, AG22, ABC^{+}22, AVAS23, AFPB23, BGdV92, BCTW96, BLBB24, BT22, CCS24, CMP13, CDSM10, CY11, CCX^{+}20, CL23b, Cia93, CMCP^{+}99, CPPRM03, CW98, CZ19, DLP^{+}24, DBGU13, DBPU13, FPBS23, GGGU21, GHX^{+}24, GCC^{+}24, Ham09, HLX23a, HAB13, JWC^{+}24, KFY24, KATS12, KKLS02, KKP^{+}22, Kip92, KK04, Kuh99, LY05, LH08, LHW^{+}24, MMST14, Mem08, MB07, MS03, MG00, MPF14, NLR11, NAF^{+}23, OPK^{+}21, PBO07, PZS^{+}20, PBC24, QNR13, RXX^{+}19, RKBL19, SLWZ24, SS06, SGE00, SB02]. based [Sne96, TZZ09, TPT13, THHB06, TD01, TK02, UKM04, WAF00, WXL^{+}21, WSG^{+}23, WI23, XM07, XCKX13, XAY23, XL20, YDX^{+}24, YXZ^{+}23, YHC13, ZYL^{+}23, ZZC^{+}23, ZFM^{+}24, ZZX^{+}21, OHDB92]. Baseline [SP18, WOM15]. Bash [DLT^{+}23]. basis [AG97, AG98]. Battling [LZS^{+}24]. Bayesian [FTF22, PLM15, SMY19]. BDD [LH08, ZZC^{+}23]. BDD-based [LH08, ZZC^{+}23]. Be [CL23b]. Behavior [FDB^{+}12, DDGR18, MAKM22, SS02, DBGU13, LK14, MG00, PP93, UKM04]. Behavior-consistent [SS02]. Behavioral [CZ19, GLFW22]. Behaviors [TZW^{+}23, IWY00]. Behaviour [CCS24]. Being [RHV24, RHvB24]. Belief [BMC^{+}23]. benefits [LH08, MC08]. Best [TLL^{+}23]. Better [CCW^{+}23, LWWM22, YB20, ZLP^{+}24]. between [CXH23, Gun00]. Beyond [GWD^{+}21, WYC^{+}24]. Bias [CZSH23, HDKM23]. Biases [YLM^{+}23]. Big [LGC^{+}23]. Bilevel [SKBD14]. Binaries [JPS22, LWD^{+}24]. Binarized [ZZC^{+}23]. Binary [DGZ24, JFJ^{+}23, YXX^{+}23, YDX^{+}24, ZHXZ24]. Biofeedback [FHS^{+}24]. BIP [ABB^{+}18]. Bird [CM23, JPS22]. Black [AFPB23, CCW^{+}23, PGZ^{+}20, RMCT22, CTCC98]. Black-box [AFPB23, PGZ^{+}20, RMCT22, CCW^{+}23]. Blindspots [BLS^{+}23]. Bloat [NWB^{+}18, XMA^{+}14]. Block [QWG^{+}23]. Block-Level [QWG^{+}23]. Blockchain [AMGBK22, ASJA23, PORH23, YBZK21, ZOJH21]. Blockchain-Powered [ZOJH21]. Boa [DNRN15]. Bookkeeping [HSB^{+}23]. Boolean [CCX11, KB07]. Boosting [CS22, GXLG21]. Bootstrap [SMY19]. Bot [KCMS24, RCAH23]. Both [TLL^{+}23]. bound [CM08]. Boundaries [CSV13]. boundary [Hie06]. Bounded [PMS13]. Box [ZAL^{+}23, AFPB23, PGZ^{+}20, RMCT22, CCW^{+}23]. Brain [VBZ^{+}18]. Breaking [BKHT21, CAC08, VCP^{+}23]. Broke [VCP^{+}23]. Brutus [CJM00]. Buddy [ZWGX22]. Budget [SMBO21]. Bug [AFHC22, CUHH24, KKP^{+}23, LRCS14, LASA24, SNM23, TWB^{+}19, ZHV^{+}23, ZSL^{+}22, AM11, CSX08]. Bug-Fixing [TWB^{+}19]. Bug-report-driven [KKP^{+}23]. Bugginess [GLL^{+}23]. Bugs [AMS^{+}18, CLS^{+}23, CCVLW23, CCXZ23, DLT^{+}23, ZSL^{+}13, ZCW^{+}24, Jac95a]. Build [JS23, MAL24]. Building [ELN^{+}92, JCNS24, OFS24, WWZ^{+}22, KKLS02].

Business [LDUD13, DCS09, ODV^{+}09]. Bytecode [SVDHB23].

C [BCGB21, CWM^{+}20, DLRA15, SRTR17]. C/C [BCGB21, DLRA15]. Cache [WCB^{+}20]. Caching [TZD^{+}24]. CafeOBJ [RO18]. Calculus [EW11, DSV03]. Call [ARL^{+}15, ARG17, MNGL98]. calls [BHR95, MM13]. Can [CL22, CL23b]. capabilities [TS09]. capability [Kuh99, TK02]. Cardinality [CSFB24]. Cars [BKD^{+}23]. Case [Arc19, AVAS23, CCW^{+}23, CXH23, EM15, FTF22, GGHM24, HZZ^{+}14, KSFC24, OKS^{+}16, RFR23, RPT^{+}22, SMBO21, SMBK22, YBL15, ZOJH21, ZAL^{+}23, ZMM^{+}16, BJMH02, CD98, DO93, HAB13, MFH02, PSV01, SCK13, TBS92, YBL13]. Cases [BKD^{+}23, CMM^{+}15, IC14, TLP^{+}22]. Catalog [YBZK21]. Categorical [BLBB23]. Category [CXH^{+}21, YKY^{+}23]. Causal [FTF24]. Causality [CFP^{+}24, GGPR24]. Causality-driven [GGPR24]. Causing [NL11]. centered [ACF97, CDFG96]. Centralised [Hie14]. Centrality [ZCT18]. centred [LASL13]. centric [HWH14]. Certificate [NWZ^{+}23, TCDZ19]. Certification [HML23]. CFL [LHX21]. CFL-Reachability-Based [LHX21]. chaining [FK96]. Chains [GHXZ24]. Challenge [XCS^{+}22]. Challenges [MKR^{+}23, NKL^{+}24, PHS^{+}24, SLWZ24, XVN22, YXL^{+}22]. Challenging [SER24]. Change [AHGZ23, LWWM22, YQTR15]. Changes [BKHT21, DR11, DPB17, GLH^{+}24, VCP^{+}23, WHP^{+}23, PSV01]. Channel [EWS14, GSC24, KJHY22, GZSW19]. Characteristics [JCSA23, CF10, SM12]. Characterization [DZV^{+}23, SGHM23]. Characterizing [GHXZ24, LMY^{+}24, XLW^{+}22, YHXY24, ZCW^{+}24]. Charts [RGS12]. Chatbot [CLMPS^{+}24]. Chats [CDKP21]. Checker [MWP^{+}21, WYMW20, SGE00]. Checkers [LZL^{+}23]. Checking [ALMS22, BHB16, CK99, DPRO23, EBE^{+}14, DDGR18, ZZW^{+}23, dFLSV14, BS07, BRRP05, BGL00, CDEG03, Esh06, FM94, FGMP03, HJL96, IWY00, JGB12, KATS12, KF07, NEFE03, PBO07, PMS13, SGL12, SMAC08, WGSD07, XCCY10]. checkpoint [CY11]. Chef [RRPW21]. Chief [Pez19d, Ros19]. Choice [EW11]. Choices [HDKM23, XSA^{+}24]. Cities [FSZ^{+}22]. Class [AB23, DLY^{+}23, PKMY24, BM13, CTCC98, CTC01, CCX11, Egy02, LY05, PPP94, ZXLC14]. class-level [CTCC98]. Classes [AB12, VCF21, GRT09, HRD08, KB07, Kuh99, Pon02, Tiw08, TK02]. Classification [WCG^{+}21, DFB99a, DFB99b]. Classifications [TZW^{+}23]. Classified [WCG^{+}21]. Classifiers [CZSH23]. Classifying [ETM22, OSH04]. cleanroom [TBS92]. Clematis [ASMP16]. CLFuzz [ZMC^{+}24]. Client [MS14, VCP^{+}23, ZZW^{+}23]. Client-Specific [ZZW^{+}23]. Client-State [MS14]. Clone [DER10]. Clones [MZW^{+}23, YHL^{+}22]. Closer [ZWZ^{+}24]. Cloud [LJL^{+}20]. cluster [CTC01]. Clustering [AFPB23, CLMPS^{+}24]. Clusters [GHXZ24]. cmath [BCGB21]. Co [LSH23, MS22, SYL^{+}23, VD13]. Co-creation [MS22]. Co-evolution [LSH23, SYL^{+}23]. co-installability [VD13]. coarse [BRR01]. coarse- [BRR01]. Code [BUB23, BXX^{+}20, BFFG19, BC23, BLS^{+}23, CXH^{+}21, DLSC23, DLT^{+}23, DNRN15, FTF22, GXG^{+}20, GGH^{+}23, GXL^{+}23, GLL^{+}23, GHX^{+}24, GCC^{+}24, GCH^{+}24, HNRA20, HZBS14, HAB^{+}20, HCW^{+}22, HLX^{+}23b, HYX^{+}24, KS20, KKP^{+}22, KJHY22, LLL^{+}23, LLT^{+}24, LWWM22, LLZ^{+}22, LXL^{+}22, MB15, MZW^{+}23, OSD^{+}23, OKS^{+}16, PZS^{+}20, PGZ^{+}20, PLM15, RR23, SKBD14, STDB23, SBBL23, SGR^{+}15, SRTR17, SWW^{+}24, SURL11, SED14, SYL^{+}23, SFC^{+}24, SZZ^{+}23, WLS^{+}20, WPX^{+}23, XLD^{+}24, XVN22, YJW^{+}20, YKY^{+}23, YFD^{+}23, YXX^{+}23, YDX^{+}24, YZY^{+}24, YHL^{+}22, ZYL^{+}23, ZE14, ZZS^{+}22, ZLP^{+}24, ZWT^{+}24, LLZ^{+}24, Dev99, DER10, FMMH^{+}14, MPG^{+}13, PRM01, RM07, SGG^{+}14]. Code-Comment [YKY^{+}23]. Code-line-level [GLL^{+}23]. Code-Smell [SKBD14]. Coded [IC14]. CodeEditor [LLL^{+}23]. CodeMatcher [LXL^{+}22]. Coding [FTF24]. Cognitive [FDC^{+}21, SHLW21]. Cohesion [AB12, CBRO16, CSC06, KK04, MB07]. coincidental [Hie06, MA14]. Cold [LWC^{+}23]. Cold-start [LWC^{+}23]. Collaboration [MS22, PLMC23, Sin10, SB02]. collaboration-based [SB02]. collaborations [HWH14]. Collaborative [TLJZ24]. Collection [TDL^{+}24, TM14]. Collective [PWB23]. Collegiate [MAKM22]. Colossal [OBH24]. Combat [IRKY24]. Combating [NWB^{+}18]. Combatting [LCL^{+}23]. Combinatorial [NL11]. Combined [JS23]. Combining [DBNG15, Hie14, SMAC08, TLL^{+}23, XZZL18]. Come [GLL^{+}23]. COMET [LCT^{+}23]. Commands [YZP^{+}22]. Comment [CXH^{+}21, GCH^{+}24, HGD^{+}23, WHP^{+}23, YKY^{+}23, ZZS^{+}22]. Commenting [WHP^{+}23]. Comments [Bre95, Pet97, WHP^{+}23, ZWCH21]. Commercial [ZMM^{+}16]. Commit [WXL^{+}21]. Commitments [EBE^{+}14]. Commits [ZSW^{+}22]. Common [MGP^{+}22]. communicating [Bro93]. Communication [CCY^{+}21, ZQS^{+}23]. Communities [ZMM^{+}16]. Community [DR15]. Companies [ZLT^{+}22]. Comparative [HGD^{+}23, BRR01]. Comparing [EMM24, Hie02, PSW^{+}20, YHC13, XM07]. Comparison [WB13, ZA23, ZLP^{+}24]. Comparisons [GGZ^{+}15]. Compatibility [ZZW^{+}23]. Competition [MAKM22]. Competitions [FTF24]. Competitiveness [YXK^{+}17]. Compilation [MAL24]. Compiler [CS22, MHK11, ZHC24, DFG00]. Complementary [SLWZ24]. Complete [MWP^{+}21, XL20]. Completeness [BLBB24]. Completion [MUB^{+}23, PLM15, WYMW20, XSA^{+}24]. Completion-aware [WYMW20]. Complex [BS16]. Component [SEM17, BCC92, CMP13, Ham09, IWY00, VD13]. component-based [CMP13]. components [BO92, CFM00, DFB99a, DFB99b, ZW97]. Composing [BLW09]. Composite [BGL00, KDM17]. Composition [Ost99, QSZG24, Ham09, REM^{+}04, ZJ93]. Compositional [GSC24, HGW^{+}16, CK96, CK99]. Comprehensibility [SGR^{+}15, SGG^{+}14]. Comprehension [BC23, MTRK14, RST^{+}14, OSH04]. Comprehensive [CZSH23, GLL^{+}21, DvdHT05]. Compressed [TZW^{+}23]. Computational [LLG23]. Computing [BPSSA^{+}22, LJL^{+}20, LWC^{+}23, OHDB92, TC20, WCJL23, EF05, FGL^{+}12, MZ09, RMP97, XCCY10]. Concept [PGM12, Sne96]. Concepts [DG17, YXZ^{+}23, MG00, SGL12]. Conceptual [OFS24, QT12, TZZ09]. Concerns [DZV^{+}23, MVM07, RM07]. Concurrency [MQLR16, ZSL^{+}13, DL13, YTL^{+}95]. Concurrent [AGRR19, MNB^{+}22, SDG^{+}24, Cor00, DKM^{+}94, DCCN04, HZZ13, MRK^{+}97, PTY95]. Condition [LXW^{+}24]. Conditional [EBE^{+}14]. Conditions [IRKY24, TDL^{+}24, KB07, SRK06]. ConE [MNB^{+}22]. Conference [MP14]. Configurable [AMS^{+}18, CGZ23]. Configuration [BNB14, BSOM24, CL23a, ELvdH^{+}05, Gun00, Jéz99]. Configurations [OBH24, SYA21, ZKW^{+}23a, ZKW^{+}23b, Sne96]. Configuring [XZZL18]. Conflict [BFFG19, ZSHD20, FN03]. Conflicts [SGHM23]. Conformance [ZCT18, Ber94, LK14, PBO07, Pet97]. confounding [ZXLC14]. Conjunction [ZJ93]. ConMem [ZSL^{+}13]. connection [AG97, AG98]. connectors [LWF03]. Consent [RBPG23]. Considered [XCS^{+}22]. Consistency [SEM17, HJL96, NEFE03, PBO07, XCCY10]. Consistent [LRL^{+}22, SS02]. Consolidation [LDUD13]. Constitutes [BSOM24]. constrained [BM13, WRD^{+}22]. Constraint [DBNG15, EMM24, GWD^{+}21, SSB20, XCCY10]. Constraints [CZ19, DDMM22, MWK15, QT12, SGD15, CY11, CK96, OO92]. Constructed [LMY^{+}24]. constructing [Hen97]. Construction [ARL^{+}15]. Constructs [KS20]. Consumption [LVBBC^{+}18]. container [XR13]. Containers [WWY^{+}23]. containing [CFM00]. Context [CK96, GZW^{+}23, GCH^{+}24, HLX23a, KAT12, LWWM22, LHX21, SNM23, WXL^{+}21, WYW^{+}22, KGA^{+}12, LH08, XCCY10, vdBV96]. Context- [WYW^{+}22]. Context-Aware [KAT12, LWWM22, WXL^{+}21]. context-free [KGA^{+}12, vdBV96]. context-sensitive [LH08]. Contexts [ZLS^{+}23]. Contextual [GZX^{+}22]. Continuous [BZSW14, BK24, HMW23, JS23, SMBK22, ZTP^{+}23]. Contract [GCC^{+}24, HLX^{+}23b, YABLR20, ZOJH21]. Contracts [AGHC^{+}22, ASJA23, CXLG22, HSB^{+}23, MRO^{+}23, YQTR15]. Contradictions [Pez24a]. Contributions [KCAS23]. Contributor [KCAS23]. Contributor-Abandoned [KCAS23]. Control [BHB16, CH21, HB22, KAS20, LYL^{+}24, MSM^{+}24, MCM24, VHNF22, BDL06, DL13, MMST14, MU00, SHR01, TBS92]. Control-flow [HB22]. Controlled [FSM^{+}15, PSZ21, BFN^{+}14]. controllers [DBPU13]. controlling [HGS93]. Convolution [YFD^{+}23]. Convolutional [DLSC23, QSZG24]. cookie [TM14]. cooperative [HE13]. Coordinating [Cia93]. coordination [CFM00, MU00, MPR06, Tiw08]. CORBA [CPPRM03]. CORBA-based [CPPRM03]. Core [BAJA23]. Core-Developer [BAJA23]. corners [ZJ97]. Correct [TLL^{+}23]. Correctness [AB23, LWWM22, TLP^{+}22, Hie06, MA14]. Correlating [HCW^{+}22]. Correlations [GL11]. correspondence [CW99]. Corrigenda [DFB99a]. Cost [CST16, FCLL21, NPB22, ZOJH21, ATW94, Bre95, REM^{+}04, Wey96]. Cost-Effective [ZOJH21, FCLL21, NPB22, REM^{+}04]. Cost-Effectiveness [CST16]. Countermeasures [EWS14, GSC24]. Coupling [CBRO16, KK04, MB07, Off92]. Coverage [DDS23, GRS^{+}16, GGZ^{+}15, GSYT21, LKA^{+}23, LCT^{+}23, MWK15, MB20, NWZ^{+}23, PGZ^{+}20, SVDHB23, XLW^{+}22, ZHXZ24, YHC13]. Coverage-Based [GGZ^{+}15, SVDHB23, YHC13]. Coverage-directed [NWZ^{+}23]. Coverage-guided [LCT^{+}23]. COVID [FSZ^{+}22, RHV24, RHvB24, MKR^{+}23]. COVID-19 [FSZ^{+}22, RHV24, RHvB24, MKR^{+}23]. CPU [MPR^{+}13]. Craftsmanship [SGhWG22]. Crash [GWD^{+}21, ZSL^{+}13, ZSL^{+}22]. Crash-Triggering [ZSL^{+}13]. Crasher [CSX08]. Crayons [CCVLW23]. creation [MS22]. Criteria [MWK15, MB20, OKS^{+}16, PLMC23, TAA^{+}19, Hie02, KSD08, XL20]. Critical [SER24, FPBS23, GM01, MS94, NA24]. Cross [LLZ^{+}22, WSG^{+}23, ZYL^{+}18, DCS09]. Cross-Language [LLZ^{+}22]. Cross-modal [WSG^{+}23]. cross-organizational [DCS09]. Cross-Project [ZYL^{+}18]. crosscutting [SGL12, MVM07]. Crowdtesting [WYMW20]. Crowdworker [WYW^{+}22]. Crucial [BPSSA^{+}22]. Cryptographic [XSA^{+}24, ZMC^{+}24, DFG00]. CSCW [KAT12]. CSP [SLD^{+}13]. Curriculum [XAY23]. customizable [Dev99]. CVEs [XCS^{+}22]. Cyber [CYA22, MSM^{+}24, MCM24, SGR^{+}23, XAY23, XTW^{+}23, ZTP^{+}23]. Cyber-Physical [MCM24, ZTP^{+}23, CYA22, MSM^{+}24, XAY23, XTW^{+}23]. Cycle [VCF21]. Cycles [OL22, SS02].

Dahl [Ano02]. DAISY [ZHS^{+}23]. Dark [MCKG24, ZJ97]. DARWIN [QRLV12]. Data [BHB16, BLBB23, BLBB24, CYA22, CCY^{+}21, CTC^{+}23, CCXZ23, DPB17, DLY^{+}23, FTF22, FTF24, GSYT21, HPH23a, HPH23b, HGC^{+}22, KDM17, LM24, LLS^{+}21, NWB^{+}18, NBB15, SSB20, TG24, TTL^{+}21, ZHS^{+}23, BCC^{+}01, BG98, CW98, FK96, For94, OSH04, TZZ09, WGG13]. Data-Driven [TG24]. Data-flow [CCXZ23]. Data-Flow-Guided [HPH23a, HPH23b]. Data-Intensive [NWB^{+}18]. Database [MWK15, CF03, PWX14, WGSD07]. Databases [AG20]. dataflow [KSD08, HPH23a, HPH23b]. Dataset [YHL^{+}22]. Datasets [UGKR22]. Date [OL22]. David [Ros13b]. DC [GRS^{+}16]. Deadlines [DBNG15]. Deadlock [ABB^{+}18, DBDS94]. Dealing [BMC^{+}23]. Deallocation [CCXZ23]. Debloating [HLX23a, SVDHB23]. Debt [RXX^{+}19]. Debts [GLL^{+}21]. Debug [MWP^{+}21]. Debugger [PBC24]. Debugger-based [PBC24]. Debugging [CMM^{+}15, FPBS23, FSP^{+}13, JO15, MQLR16, ZNZ^{+}23, AM04, HRD08, OSH04, PBC24, QRLV12]. Deceiving [KFY24]. Decentralized [EHF20, ML00]. Deciding [SGD15]. Decision [HGW^{+}16, JWC^{+}24, KDM23, XLW^{+}22].

Decision-making [KDM23]. Decisions [AMGBK22, LLS^{+}21, SGR^{+}23, AM11]. Decompiled [DLY^{+}23]. Decomposition [BSA22, CCRZ23]. decoupling [BTI14]. deduction [FS93]. deductive [GM01]. Deep [BT24, CYA22, CWW^{+}20, CWM^{+}23, CLS^{+}23, CWH^{+}21, GHXZ24, GHX^{+}24, HGC^{+}22, HGX^{+}24b, HZB^{+}24, JYZ^{+}24, KFY23, LCT^{+}23, LLSM22, LGX^{+}22, LGC^{+}23, LHW^{+}24, MPT^{+}21, MZW^{+}23, NBMK22, QWG^{+}23, RMCT22, SKY23, WXL^{+}21, WYC^{+}24, WCP^{+}22, WWAC23, WT24a, WT24b, XLW^{+}22, XLD^{+}24, YDX^{+}24, YHL^{+}22, ZFL^{+}22, ZXS^{+}24, ZZS^{+}22, ZLP^{+}24, ZRGT23, ZZX^{+}21, LLZ^{+}24]. DeepPatch [WWAC23]. DeepWukong [CWH^{+}21]. Defect [FAP22, GLH^{+}24, KS20, NXL^{+}22, ZYL^{+}18, SM12]. Defective [VCF21]. Defects [AVY11, AHGZ23, FAP22, VCF21]. Defined [FFM24]. Defining [TG23]. deGraphCS [ZYL^{+}23]. Degree [FMMH^{+}14]. Degree-of-knowledge [FMMH^{+}14]. Delivery [GJS^{+}23, ZTP^{+}23]. Delta [HT98, HVT98]. DeltaDroid [GJS^{+}23]. Demystifying [SCL^{+}23]. Denoising [BXX^{+}20]. Depended [VCP^{+}23]. Dependence [FCLL21, PXJ17, Dil97, SHR01, SRK06]. dependences [Jac95a, OSH04]. dependencies [BGO^{+}14, Gun00, Rob08]. Dependency [JCSA23, RCAH23, SEM17, CY11, GL14]. dependency-based [CY11]. dependency-driven [GL14]. Deployed [AVY11]. Deployment [BSOM24]. Depth [XCS^{+}22, SBBL23, ZXLC14]. Derive [YBL15, XM08]. derived [IWY00]. Description [Bjø19, DvdHT05, DJ97]. Descriptions [GCX^{+}22, JZL^{+}19a, YXZ^{+}23, AAG95, BAD08, WJ10]. descriptors [DER10]. DESEN [KAS20]. Deserialization [SBBL23]. Desert [Rei99]. Design [AROK21, AMGBK22, BPT10, CCX^{+}20, EK11, HDKM23, KDM23, LSN^{+}23, LSBN24, Liu22, DDGR18, MFLL12, SGR^{+}15, TG24, BM07, BO92, BRRP05, BFN^{+}14, CSC06, CR94, FBC^{+}13, FP02, GGLT07, LL00, MRK^{+}97, RS09, SS06, SB06, SGR^{+}10, YTL^{+}95, ZB13]. Design-Pattern [SGR^{+}15]. Designers [CZ19]. Designing [BCC^{+}01, DL11, XM07, CPPRM03]. Designs [CLMPS^{+}24, SB02]. Destruct [CXLG22]. Detect [CSB^{+}24, ZCL^{+}24]. Detecting [AVY11, CWH^{+}21, CCXZ23, GCX^{+}22, LWD^{+}24, LZF^{+}24, DDGR18, MM13, WCB^{+}20, YHXY24, YHL^{+}22, ZSL^{+}13, ZCW^{+}24, Jac95a, LS13]. Detection [Cai20, CL22, CCY^{+}21, EKL^{+}19, GHM18, JPS22, LRCS14, MNB^{+}22, MS14, NBMK22, RD15, RXX^{+}19, SKBD14, SLWZ24, SXX^{+}24, SZZ^{+}23, UGKR22, WWZ^{+}22, XMA^{+}14, XAY23, XGF^{+}23, YXX^{+}23, YDX^{+}24, YXM^{+}23, ZAW92, ZFL^{+}22, ZHV^{+}23, ZLW^{+}21, ZCL^{+}24, ZMC^{+}24, ZZL^{+}24, ZWY^{+}21, LLZ^{+}24, FN03, Kuh99, SMT92, TK02, XR13]. Detector [ZZX^{+}21]. determination [OLR^{+}96]. deterministic [HT17]. Developer [BFFG19, BAJA23, CDKP21, CCY^{+}21, EHEZ21, GGHM24, KS22, SHLW21, WFW^{+}20, CF10, FMMH^{+}14, Sin10]. Developers [CXH23, FSZ^{+}22, NMA^{+}23, NKL^{+}24, PZS^{+}20, PLZ^{+}22, RHV24, RHvB24, YZP^{+}22, ZWCH21]. Developing [HRD08, TAA^{+}19, ZJW03, ZOJH21, MGMM11]. Development [BFFG19, BSA22, CFL^{+}16, GTD21, KRM24, KCMS24, LKRF22, LCS^{+}22, LSB^{+}23, MNB^{+}22, MKR^{+}23, MS15, PDHR23, RFDS24, SMY^{+}24, TG23, ZCT18, AM11, DvdHT05, EAS08, ELN^{+}92, GJ08, KK93, MFH02, PSV01, SCK13, Tiw08]. development-oriented [AM11]. Deviated [TZW^{+}23]. deviations [CDFG96]. device [HHC^{+}24, HDKM23]. Devices [LZS^{+}24]. diagnosis [NLR11]. diagram [BP05].

Diagrams

[AB23, BM13, Egy02, Esh06, LK14]. DiaPro [CST16]. Differ [JA22]. Differentiable [FYW24]. Differential [LHW^{+}24, MQLR16, NWZ^{+}23, TCDZ19]. DIG [NKWF14]. Digital [XAY23]. Digraphs [EM15]. Dijkstra [Ano02]. DIRE [DLY^{+}23]. Directed [SDLC23, YPRK14, NWZ^{+}23]. discipline [KLV05]. Discoverability [ACSA23]. Discovering [CW98, GL11]. Discovery [Böh18, SPK14, ZHS^{+}23, ZZW^{+}23, GL14]. Discrepancies [SXX^{+}24]. discrete [Ost99]. Discrimination [ZSWS23]. Discussions [EHEZ21]. Disengagement [GHXZ24]. Dissecting [FMMB23b, FMMB23a]. Distinguishing [HT17]. Distributed [BLX^{+}20, DDI22, FCLL21, GTD21, Hie14, LGC^{+}23, TG11, WME93, KK93, MU00]. Distribution [DDS23, HGC^{+}22, HZB^{+}24, TS09]. Distribution-Aware [HGC^{+}22, HZB^{+}24]. Diversifying [MJS^{+}21]. Diversity [XHL^{+}24, HAB13]. DNN [FPBS23, HGX^{+}24a, SLWZ24, TZW^{+}23]. DNN-based [FPBS23, SLWZ24]. DNNs [AFPB23, KFY24]. Do [CMM^{+}15, CXLG22, CL23a, JA22, LYYC14, PLZ^{+}22, SURL11, YZP^{+}22, ZWL^{+}24, CAC08]. Docker [RSBO23]. Dockerfile [ZZL^{+}24]. Documentation [CXH23, HCW^{+}22, UKR21, TBS92]. Documenting [SGR^{+}15]. documents [SMT92]. Does [CXH^{+}21, FSM^{+}15]. Domain [ASNB19, Bjø19, BMC^{+}23, HZS08, OFS24, YDX^{+}24, ZE14, ZH24, BJMH02, BAD08, Hie09, JW94, SS06, ZAW92]. Domain-Specific [OFS24, HZS08, BJMH02, SS06]. domain-testing [JW94]. Domains [GHXZ24, Hie02]. Done [PORH23]. Dormant [FAP22]. Double [For94]. DRIVE [ZZL^{+}24]. Driven [CCRZ23, CLBY18, DG17, FFM24, PVHW17, TG24, BDL06, GL14, GGPR24, GTD21, KKP^{+}23, NPB22]. Driving [GGPR24, KSFC24, LKA^{+}23, SER24, TZZ^{+}23, BKD^{+}23]. DSD [CSX08]. DSD-Crasher [CSX08]. Dual [CL23a, SNM23]. Duplicate [WYMW20, ZHV^{+}23]. Duplication [ZLW^{+}21, DER10]. During [MKR^{+}23, FHS^{+}24, FSZ^{+}22, NKL^{+}24, RHV24, RHvB24]. DynAlloy [FPGA07]. Dynamic [CST16, CYM^{+}24, FCLL21, GJS^{+}23, HB22, HSX^{+}24, HCJ^{+}24, KMYK19, LZS^{+}24, DDGR18, NKWF14, PXJ17, PSW^{+}20, XMA^{+}14, ZHS^{+}23, CY11, DR10, FC00, FPB^{+}05, GSH97, PJRR10]. Dynamic-Analysis-Induced [ZHS^{+}23]. Dynamic-typed [HCJ^{+}24]. dynamically [WGSD07]. Dynamics [BSA22, KCAS23]. Dynamite [MPF14].

E3 [JPL98]. Eagle [LHX21]. Early [CM23, CCS24, KDM17, KDM23]. EASE [PKMY24]. Easier [CMM^{+}15]. Easy [WWJ^{+}24, ZWCH21]. Ecosystem [ZWZ^{+}24]. Ecosystems [BKHT21]. Edit [LLL^{+}23, MNB^{+}22]. editing [BGdV92]. Editor [Pez19d, Ros19]. Editor-in-Chief [Pez19d, Ros19]. Editorial [DR15, GMRS03, Ghe05, Ghe07, Not07a, Not07b, Not07c, Not08a, Not08b, Not09, Not10, Not12, Not13, OGKW05, Pez19d, Pez19a, Pez19b, Pez19c, Pez20a, Pez20b, Pez21, Pez22, Pez24a, Pez24b, Ros13a, Ros14c, Ros14a, Ros14b, Ros16, Ros17, Ros18a, Ros18b, Ros19, Pez23]. Edsger [Ano02]. Education [SMY^{+}24]. Effect [GRS^{+}16, HZBS14, JFJ^{+}23, RST^{+}14, ZSL^{+}13, Off92, Sin10, ZXLC14]. Effect-Oriented [ZSL^{+}13]. Effective [DGZ24, FYD^{+}08, RD15, ZOJH21, ZBA23, ZRGT23, CF10, FCLL21, Hen97, NPB22, REM^{+}04]. Effectiveness [CST16, CMM^{+}15, GRS^{+}16, MWK15, ODE21, UGKR22, VR23, ZWL^{+}24, CM08].

Efficiency [CMM^{+}15, LH02]. Efficient [AGRR19, AVY11, CWW^{+}20, EMM24, FPGA07, HSB^{+}23, HQM^{+}23, LYL^{+}24, LSX^{+}24, QHM^{+}23, SRK06, WT24a, WT24b, ZRGT23, ZWGX22, ZNZ^{+}23, dFLSV14, RH97]. Efficiently [DL11]. Effort [KS22, PKMY24, SP18, SMY19, WOM15, AM11, BM07, GJ08, MY13]. Effort-aware [PKMY24]. Eight [Pez24b]. elaboration [UKM04]. Element [NAF^{+}23]. Elephant [YLM^{+}23]. Elevator [HAY^{+}23]. Eliciting [JPL98]. Elite [WFW^{+}20]. Embedded [BNB14, LZS^{+}24, MFLL12, GGLT07]. Embedding [LWWM22, WMW^{+}23, XSA^{+}24, YFD^{+}23, ZYL^{+}23]. Embeddings [DLSC23, ETM22, TLL^{+}23]. Emergent [LBZ14]. Emoji [CCY^{+}21]. Emoji-powered [CCY^{+}21]. Emotion [CCY^{+}21]. Empathy [GGHM24]. Empirical [AROK21, AAP^{+}20, BYTS23, BSOM24, CWM^{+}20, CZSH23, CL23a, EHEZ21, FSM^{+}15, FTF22, FTF24, GLL^{+}21, HGC^{+}22, KCMS24, LLSM22, LASA24, LLS^{+}21, MBH09, ODE21, PORH23, RSBP23, RWEB19, TWB^{+}19, UGKR22, VCP^{+}23, VLJ^{+}18, WYC^{+}24, YXK^{+}17, ZHO^{+}18, ZLP^{+}24, BM07, BGH07, GHK^{+}01, HT98, HVT98, MB07, MC08, MNGL98, SR05, Tiw08]. Emulation [DDI22, ZHXZ24]. emulators [MPR^{+}13]. Enable [ZBA23]. enabled [VTA04]. Enabledness [DBGU13, GGGU21]. Enabledness-based [DBGU13, GGGU21]. Enablers [GGHM24]. End [ZSL^{+}22, Dev99, LASL13]. End-to-End [ZSL^{+}22]. end-user [LASL13]. Energy [LCL^{+}23, LVBBC^{+}18]. Engagement [FHS^{+}24]. Engineer [WI23]. Engineered [TLL^{+}23]. Engineering [AAP^{+}20, CL22, CCS24, CL23b, DG17, DL11, DR15, FTF22, FTF24, GLFW22, HGHM23, LGX^{+}22, LGC^{+}23, MFBF^{+}22, MBH^{+}17, MP14, OKS^{+}16, Pez24a, STS^{+}18, SF18, UGKR22, WFF^{+}19, WCP^{+}22, YXL^{+}22, ACF97, CDP04, ELvdH^{+}05, KLV05, LASL13, RSB05, SR05, TBS92, UFG14, XL20, ZJ97]. Engineering-related [HGHM23]. Engineers [PLMC23, HBB^{+}09]. Engines [SURL11]. Enhanced [JWC^{+}24, YXM^{+}23, SS06]. Enhancement [HGC^{+}22, ZFL^{+}22]. Enhancing [AG22, YDX^{+}24, TS09]. Enough [CBRO16, MCB^{+}23]. Enriched [HLX^{+}23b]. Ensemble [UGKR22]. Ensuring [SEM17]. Entity [HSX^{+}24]. Entry [BLBB23, BLBB24]. environment [ATW94, Bre95, FGMP03, Kli93, MRK^{+}97, Rei99, RVMRM04, SN92, TY92]. Environments [BKD^{+}23, ACF97, DHW98, ELN^{+}92, KK93, Kli93, MGP^{+}13, PJRR10, PWD^{+}99]. equations [BRG^{+}01, KGA^{+}12]. Equivalence [LH02, MGTR18, DSV03]. Equivalent [LCZL14]. Era [LGC^{+}23]. ERC [MRO^{+}23]. Errata [AG98]. Error [TG24, Kuh99, TK02]. errors [TD01, ZAW92]. ESP [Cia93]. Essential [SLB^{+}21]. estimates [GJ08]. Estimating [GLH^{+}24, LLLM23, LSN^{+}23, MB15, PBU16, BM07]. Estimation [CWW^{+}20, KS22, LSBN24, PMM^{+}99, SP18, WOM15, ZFL^{+}22, MY13, TZZ09]. Estimator [KS20]. Ethereum [ASJA23, CXLG22, MRO^{+}23, PORH23, WWZ^{+}22, ZOJH21, ZCZ^{+}23]. Ethics [PHS^{+}24]. Evaluate [KDM23, MGP^{+}13]. Evaluating [KFY23, LH08, TAA^{+}19, WGG13]. Evaluation [AAP^{+}20, FMMB23b, FMMB23a, FA14, HCW^{+}22, KDM17, MS15, RWEB19, SBMK21, SMBK22, YLM^{+}23, YHL^{+}22, ZWL^{+}22, CAC08, DBDS94, KK93, MBH09, XCKX13]. evaluations [SM12]. Event [ASMP16, BCTW96, CW98, DBPU13, Mem08]. Event-Based [ASMP16, BCTW96, CW98, DBPU13]. Events [FPBS23]. every [LYYC14]. Evolution [Cai20, DR11, DKD21, KDM23, OSD^{+}23, LSH23, RM03, RVMRM04, SN92, SYL^{+}23, THHB06, WGG13]. Evolutionary [HLL^{+}16, YB20, Hen97, MBH09]. evolvability [CS12]. Evolving [OSD^{+}23, DCS09, QRLV12]. EvoMaster [Arc19, ZA22, ZAL^{+}23]. EvoSuite [FA14]. Exact [HKMB^{+}14]. Examination [ZYL^{+}18]. Examining [MAKM22]. Example [LKRF22, WWJ^{+}24]. Examples [BS16, ZCL^{+}24]. Exception [CMP13, ZE14, RM03]. Executables [AEK^{+}16]. Execution [CPCT21, KPC18, SBBL23, TDT^{+}22, WCB^{+}20, YPRK14, AM04, Dil93, DHW98, SMAC08]. Executions [BLX^{+}20, EM18, PSW^{+}20]. Exhaustive [HML23]. Existing [AG22]. Expansion [GCH^{+}24]. experience [CMCP^{+}99, YTL^{+}95]. Experiential [SMY^{+}24]. Experiment [OL22, PSZ21, BFN^{+}14]. Experimental [CMM^{+}15, DO93, DBDS94, OLR^{+}96, SMT92]. Experiments [SGR^{+}15, Ham09, YBL13]. expert [CF10, Kip92]. Explainability [RXX^{+}19]. Explainable [WPX^{+}23]. Explaining [JCNS24, NA24]. Explanation [FPBS23]. Explicit [BHB16]. Explicit-Data [BHB16]. Exploiting [CGPP15]. Exploits [SBBL23]. Exploration [CS22, ZRGT23, QNR13]. Explorations [PBU16]. Exploratory [GZW^{+}23, TTL^{+}21]. Exploring [CCW^{+}23, MZW^{+}23, RCAH23, SLWZ24]. Exposing [LBZ14]. expressions [KGA^{+}12]. Expressive [TG11, BLW09, WJ10]. Extended [EM15, EM18, ZE14, LY05]. extensibility [BJMH02]. Extensible [TG11]. Extension [PKMY24]. External [GL14]. Extracting [KM10]. Extraction [ASNB19, AFPB23, ASJA23, BXX^{+}20, GWD^{+}21, HSX^{+}24, YXZ^{+}23, MN96]. Extractive [SFC^{+}24]. Extractive-and-Abstractive [SFC^{+}24]. extractors [MNGL98].

FaaSLight [LWC^{+}23]. Faced [NKL^{+}24]. Facet [DGK21]. Facet-oriented [DGK21]. faceted [DFB99a, DFB99b]. Facilitating [RKBL19, YBL13]. Facing [SLWZ24]. factors [SAB^{+}14]. factory [BCC92, FLM^{+}98]. Failed [JA22]. Failing [TLP^{+}22]. Failure [JCNS24, NL11, Wey96]. Failure-Causing [NL11]. Failures [HMW23, JO15, JCNS24, LJL^{+}20]. Fair [MCB^{+}23]. Faire [LXW^{+}24]. Fairness [LXW^{+}24, MCB^{+}23, WYW^{+}22]. Fairness-Aware [WYW^{+}22]. families [BCD02]. Family [GHM18, RKBL19, SGR^{+}15, WFF^{+}19]. Family-based [RKBL19]. Fan [MVM07]. Fan-In [MVM07]. Far [CL22, GLL^{+}21, GLL^{+}23, HZX^{+}24, ZHV^{+}23, ZYL^{+}18]. Farewell [Ros19]. Fast [WWY^{+}23]. Fault [AROK21, EKL^{+}19, HMW23, JWC^{+}24, KKP^{+}23, KMYK19, Kuh99, NBMK22, TSPRC18, YHC13, YXK^{+}17, ZLS^{+}23, CCX11, Hie02, Hie09, KB07, LY05, MA14, SMT92, TK02, XCKX13, ZXLC14]. Fault-Prone [AROK21]. Faults [HZBS14, MZA22, SRTR17, LS13]. Faulty [SYA21]. Feasibility [EK11]. Feature [ABC^{+}22, AFPB23, CLBY18, DDS23, FN03, HLL^{+}16, RWEB19, WRD^{+}22, ZRGT23, ZS97, ZZL^{+}06]. Feature-Guided [CLBY18]. Features [JZL^{+}19a, KS22, KMYK19, NA24, TLL^{+}23, ZCZ^{+}23, DR10, Zav04]. Feedback [FHS^{+}24, SDLC23, GJ08]. Feedback-Directed [SDLC23]. Feedforward [BK23]. FEMOSAA [CLBY18]. Field [DPB17, JO15]. Fields [BLBB23]. File [DGZ24]. Filling [BLBB23]. Filtering [EMM24]. Finding [OBH24, TZW^{+}23, CSX08]. Fine [PGZ^{+}20, BRR01, DL13]. Fine-grained [PGZ^{+}20, BRR01, DL13]. Finite [BM13, DDMM22, EM18, Cor00]. Finite-State [EM18, Cor00]. Firmware [WRD^{+}22]. First [DR15, GSC24, HHC^{+}24, MCKG24]. First-Order [GSC24]. Fitness [FFM24]. Fix [LZL^{+}23]. Fixed [OL22]. Fixed-Date [OL22]. Fixes [SNM23]. Fixing [SRTR17, TWB^{+}19]. FlagRemover [BHL11]. flags [BHL11]. Flaky [PKHM22]. Flexible [NEFE03, BTI14]. Floating [BCGB21]. Floating-Point [BCGB21]. Flow [CH21, DCCN04, HPH23a, LYL^{+}24, TDL^{+}24, ZYL^{+}23, CCXZ23, For94, HPH23b, HB22, WWY^{+}23]. flow-sensitive [For94]. flows [MP09]. Focused [MJS^{+}21]. Fold2Vec [BC23]. Follow [STS^{+}18]. Foraging [FSP^{+}13]. Formal [BP05, CTA^{+}21, CR94, EWS14, GXSC21, RO18, YJW^{+}20, AG97, AG98, BRRP05, BKM07, CS12, CMCP^{+}99, CRST12, CPPRM03, FP02, MMST14, PGM12, SCK13, VTA04, SB06]. Formalizing [AAG95, CD98, BP98, CDFG96]. Formally [WI23, CMCP^{+}99]. FormatFuzzer [DGZ24]. Formats [DGZ24]. formatters [vdBV96]. Formed [TC20]. Forms [BLBB23, BLBB24]. formulas [XCKX13]. forward [Ros13a]. Found [MZA22]. Foundation [XTW^{+}23, SCK13]. Foundations [KF07]. Founded [WI23]. Four [ZJ97, CD98]. FQN [HYX^{+}24]. Fragile [HDKM23]. frames [KK04]. Framework [AEK^{+}16, BK24, CTA^{+}21, CLS^{+}23, DR11, JYZ^{+}24, KAT12, LWD^{+}24, MAL24, MS15, OFS24, SFC^{+}24, SIB23, TG24, WSG^{+}23, WWZ^{+}22, XW22, YBL15, BCTW96, CDP04, CDFG96, Dev99, For94, MS03, SGL12, WGG13]. Free [RCAH23, HGX^{+}24b, KGA^{+}12, vdBV96]. Freedom [ABB^{+}18]. front [Dev99]. front-end-retargetable [Dev99]. FSMs [HT17]. Full [SRTR17, ZHXZ24]. Full-Word [SRTR17]. Function [CXLG22, JFJ^{+}23, LWC^{+}23]. Function-as-a-Service [LWC^{+}23]. Functional [Bro93, GD08, LLLM23, RST^{+}14, MGP^{+}13].

Functions [BCGB21, FFM24, Hie09, MPG^{+}13, VKV03]. Future [Pez24b]. FuzzBench [FMMB23b, FMMB23a]. Fuzzer [HPH23a, HPH23b]. Fuzzers [LZS^{+}24]. Fuzzing [DGZ24, HQM^{+}23, LZS^{+}24, LHW^{+}24, LSX^{+}24, NKL^{+}24, QHM^{+}23, ZKW^{+}23a, ZKW^{+}23b, ZBA23, ZA23, ZAL^{+}23, ZMC^{+}24]. Fuzzy [FMMB23b, FMMB23a].

Gaia [DL11, ZJW03]. Gaia-PL [DL11]. Game [NMA^{+}23]. Gas [ZOJH21]. General [LWC^{+}23, CCX11]. Generalizable [DLSC23]. Generated [CMM^{+}15, FFM24, LS13, WGSD07]. Generating [ARG17, DRW96, GXG^{+}20, HT17, IC14, SSB20, YJW^{+}20, Kli93]. Generation [Arc19, AG20, BFFG19, DTC^{+}24, EMM24, FA14, FSM^{+}15, GZX^{+}22, GSYT21, GCH^{+}24, HLL^{+}20, HAB^{+}20, HCW^{+}22, JCNS24, KFY24, LCT^{+}23, LHW^{+}24, MZA22, MGTR18, SMBO21, SWW^{+}24, WXL^{+}21, XHL^{+}24, XVN22, YZY^{+}24, ZA22, ZZS^{+}22, ZWT^{+}24, vdBV96, EF05, FK96, FRB^{+}06, HZS08, PWX14]. Generation-based [LHW^{+}24]. Generative [KAT12]. Generator [NKWF14, DO93]. Generic [CWM^{+}20, LL00]. Genetic [DBNG15, YXK^{+}17]. GENOA [Dev99]. Git [YZP^{+}22]. GitHub [BYTS23, BAJA23, WBRH23]. Gitter [EHEZ21]. Global [ABB^{+}18, CFL^{+}16, WFF^{+}19, ZLS^{+}23]. Go [CL22, GLL^{+}23]. Goals [BBS16, DBPU13]. Good [SWW^{+}24, YZY^{+}24]. Google [CCH^{+}21]. governed [MU00]. GQM [FLM^{+}98]. Graded [Ban23]. grained [BRR01, DL13, PGZ^{+}20]. grammars [KGA^{+}12]. grammarware [KLV05]. Graph [ARL^{+}15, CWH^{+}21, DLP^{+}24, DLSC23, HLX^{+}23b, LD24, NBMK22, PTY95, TG23, WPX^{+}23, YFD^{+}23, ZYL^{+}23, MNGL98]. graphical [DKM^{+}94, MRK^{+}97].

GraphPrior [DLP^{+}24]. graphs [SRK06]. GreASE [dFLSV14]. Greenkeeper [RCAH23]. Grid [EMM24]. Grid-Based [EMM24]. Group [BSA22, ZSWS23]. growth [JMS08]. guarantee [CAC08]. Guaranteeing [ALMS22]. Guarantees [OBH24]. GUI [ABC^{+}22, Mem08, XM07, XM08]. GUI-based [XM07]. Guide [HAB^{+}20]. Guided [CLBY18, GGH^{+}23, HPH23a, PWX14, SDG^{+}24, WRD^{+}22, ZLS^{+}23, ZZW^{+}23, HPH23b, LCT^{+}23, TCDZ19, ZHXZ24]. Guidelines [FTF22, GGZ^{+}15, GLFW22]. Guiding [CL23a]. GUIs [HDG^{+}24, LVBBC^{+}18].

HAMPI [KGA^{+}12]. handle [LYYC14]. handlers [CMP13]. Handling [AG20, ZE14]. Hard [LSBN24, CAC08]. Harmful [CL23b]. Hazard [FPBS23]. Hazard-triggering [FPBS23]. HCSP [YJW^{+}20]. healing [CMP13]. Health [LM24]. Heap [LKR22]. Help [FSM^{+}15]. Helpfulness [KCMS24]. Heterogeneous [CLMPS^{+}24, SXX^{+}24, WSG^{+}23, MU00]. Heuristic [CM23, ZHO^{+}18, ZZX^{+}21]. Heuristics [ZBA23]. Hidden [SCL^{+}23]. Hierarchical [CGZ23, HZB^{+}24, JEB^{+}23, YWC16, BO92, SLD^{+}13, WJ10]. hierarchies [CCX11]. hierarchy [BM13, DFB99a, DFB99b, LY05]. hierarchy-aware [DFB99a, DFB99b]. high [CF03]. high-quality [CF03]. Higher [GXSC21, LWF03, XTW^{+}23]. Higher-Order [GXSC21, LWF03, XTW^{+}23]. Highly [AMS^{+}18]. HINNPerf [CGZ23]. Hippodrome [CTC^{+}23]. History [ARG17, OPK^{+}21, FM94]. History-based [OPK^{+}21]. history-checking [FM94]. Hoc [CDKP21]. Holistic [FDC^{+}21]. Home [FSZ^{+}22, MKR^{+}23]. Horus [LSX^{+}24]. Host [LSX^{+}24]. hosts [MPR06]. HOTTest [SS06]. Human [HGHM23, HCW^{+}22, NKL^{+}24, SPAS21, YXK^{+}17, CDFG96]. human-centered [CDFG96]. Humans [KFY24]. Hybrid [GXSC21, GSH97, JEB^{+}23, ZMM^{+}16, CRST12, CSX08]. HybridCISave [JS23]. Hybridized [BBS16]. Hyper [ZHO^{+}18]. Hyper-Heuristic [ZHO^{+}18]. Hypermutation [ZA22]. Hyperparameter [LLSM22]. hypotheses [Hie02, Hie09]. Hypothesis [WSG^{+}23].

iBiR [KKP^{+}23]. ICSE [MP14, Pez24a]. IDE [XVN22]. Identification [BAJA23, DZV^{+}23, GHM18, GLL^{+}23, PKMY24, STDB23, SYL^{+}23, WRD^{+}22, WC23, ZCZ^{+}23, ZSW^{+}22]. Identifier [SRTR17, ZLL^{+}23]. Identify [HNRA20]. Identifying [CDKP21, GLL^{+}21, MVM07, NA24]. Identity [ZQS^{+}23]. IFDS [HLX23a]. IFDS-based [HLX23a]. III [MKS^{+}15]. Image [CCX^{+}20, TZW^{+}23]. Images [RSBO23]. Impact [AHGZ23, CST16, ELvdH^{+}05, FAP22, FTF24, LLSM22, LLS^{+}21, MZW^{+}23, PVHW17, Tiw08, ZLW^{+}21, EAS08, MA14, RSB05, SGG^{+}14]. Impact-Driven [PVHW17]. Impacts [TDT^{+}22]. Implementation [KDM17, Liu22, ZCT18, ZMC^{+}24, BO92, BPT10, LH08, SB02, WWL^{+}23]. Implementations [GZSW19, NWZ^{+}23, TTL^{+}21, TCDZ19]. implementing [CDP04]. implicit [SPAK10]. implied [UKM04]. Important [LXL^{+}22, YZY^{+}24]. Improve [ABC^{+}22, CCH^{+}21, VHNF22]. Improved [CST16]. Improvement [CXH^{+}21, WWAC23, SR05]. Improving [ASNB19, BGO^{+}14, Cai20, DTC^{+}24, JYZ^{+}24, PWB23, WYMW20, YHL^{+}22, ZH24, DPT13, GJ08, LH02]. In-Depth [XCS^{+}22, SBBL23, ZXLC14]. In-IDE [XVN22]. In-Process [WYW^{+}22].

Inclusive [SMY^{+}24]. Incoming [Pez19d]. Incompatibility [XGF^{+}23]. Incompatible [LZF^{+}24]. Inconsistencies [WBRH23, CDFG96, GZ05]. inconsistent [HN98]. Incorporating [SZZ^{+}23, YDX^{+}24]. Incredible [Pez24a]. Incremental [DPB17, UKM04, YPRK14, KK93, KF07]. incrementally [KKLS02]. Independent [WHP^{+}23]. Index [Ano96, TPT13]. index-sensitive [TPT13]. Indicators [AAP^{+}20]. Induced [ZHS^{+}23]. Inductive [ASJDW21, BG96]. Industrial [HAY^{+}23, OKS^{+}16, ZAL^{+}23, CMCP^{+}99, FLM^{+}98, SR05]. Industrially [VHNF22]. Industry [MS22, SER24]. Industry-Academia [MS22]. Inference [CSFB24, HYX^{+}24, NBB15, SMY19]. Inferred [LZL^{+}23]. Inferring [EM18, MG00, RO18]. Inflow [ZMM^{+}16]. Influence [HGHM23, RSBO23, Sin10]. Influential [ZLS^{+}23]. Information [CXH23, FSP^{+}13, RFR23, SXX^{+}24, WCB^{+}20, ZLP^{+}24, DFOT07, FBC^{+}13, GSH97, MP09, PGM12, TZZ09, THHB06, Wey96, YHC13]. information-theoretic [YHC13]. informed [YHG^{+}24]. infrastructures [BDL06]. inheritance [KKLS02]. inheritance-based [KKLS02]. Injection [KKP^{+}23]. Inline [HGD^{+}23]. Inlining [JFJ^{+}23]. inner [SAB^{+}14]. Input [DLP^{+}24, DDS23, EMM24, KFY24]. Insight [SLWZ24]. inspections [BFN^{+}14, PSMV98]. Inspired [FNT^{+}24]. installability [VD13]. Instances [SGR^{+}15]. instruction [AM04]. IntDroid [ZWY^{+}21]. Integer [DLRA15, XL20]. Integrated [YB20, CTCC98, ELN^{+}92, PWD^{+}99, SCK13]. Integrating [BFFG19, FRB^{+}06, SBS^{+}23, GSH97]. Integration [HMW23, JS23, ZTP^{+}23, BCTW96, BHR95, SN92, YHR92]. Integrity [MWK15]. Intellectual [CSV13]. Intelligent [PLM15]. Intended [GZW^{+}23]. Intensive [NWB^{+}18]. Intent [YXM^{+}23].

Inter [SEM17]. Inter- [SEM17]. Interacting [GRT09, SYA21]. Interaction [AB12, CGZ23, DDS23, MU00]. Interaction-Based [AB12]. Interactions [ASMP16, GGHM24, PWX14]. interchangeability [DPT13]. Interface [PSZ21, CL94]. Interfaces [IC14, BRRP05, BTI14]. Interleaving [SDG^{+}24]. international [NP08, MP14]. Internet [YBZK21]. Interpretable [YFD^{+}23, ZSWS23]. Interpretation [LRL^{+}22, WCG^{+}21, CSC06]. interpreters [BP05]. Interpreting [ZZX^{+}21]. Interprocedural [OO92, SHR01, For94]. Interval [SMY19, CDSM10, DKM^{+}94]. interval-based [CDSM10]. Interventions [RPT^{+}22]. Interview [ZTP^{+}23]. Interview-Based [ZTP^{+}23]. Interviews [FHS^{+}24]. Intimacy [ZWY^{+}21]. Intra [SEM17]. Intra-Component [SEM17]. Intrinsic [CGPP15]. Introduced [RCAH23]. Introduction [GLFW22, HP15, MP14, NP08]. Invariant [NKWF14]. Invariants [NKWF14]. investigating [HBB^{+}09, CXLG22, JFJ^{+}23]. Investigation [CL23a, LRCS14, WWZ^{+}22, MC08]. Investigations [Off92]. invocation [SPAK10]. Involvement [ZMM^{+}16]. iOS [HHC^{+}24]. IoT [SMBK22]. IP [MR99]. iSENSE2.0 [WYMW20]. Isolation [JZL^{+}19b]. ISSTA [NP08, HP15]. Issue [AHGZ23, HP15, MP14, PWB23]. Issues [LCL^{+}23, Pez23, Pez24b]. items [Gun00]. iterative [For94].

J [TS09]. J-Orchestra [TS09]. Java [BS07, BLS^{+}23, Cor00, HRD08, KM10, LTX19, MRR05, MAL24, PBC24, RD15, SBBL23, SRTR17, SGHM23, SVDHB23, TS09, WHP^{+}23, XR13]. JavaScript [ASMP16, NXL^{+}22, ZBA23]. Johan [Ano02]. Join [BTI14]. Joint [HSX^{+}24]. Journal [DR15]. Journal-First [DR15].

Journey [ZYL^{+}18]. JUnit [WC23].

Jupyter [LASA24]. Just

[GLH^{+}24, NXL^{+}22]. Just-In-Time

[GLH^{+}24, NXL^{+}22].

KAPE [GHX^{+}24]. Katana [SNM23]. KBSE [DJ97]. Kernel [LSX^{+}24]. Key [GCX^{+}22, NA24, PKMY24, SAB^{+}14]. Killing [CPCT21]. KLEESpectre [WCB^{+}20]. Knee [CLBY18]. Knee-Driven [CLBY18]. Know [GXL^{+}23, YZP^{+}22, ZWL^{+}24]. Knowledge [CH21, GZX^{+}22, HLX^{+}23b, MS22, TG23, WPX^{+}23, WJZ^{+}23, YDX^{+}24, ZZW^{+}23, FMMH^{+}14, KK04, MG00]. knowledge-based [KK04, MG00]. Knowledge-Guided [ZZW^{+}23]. Kristen [Ano02]. Kubernetes [RSBP23].

L2S [XW22]. Label [VCF21, ML00].

Labeled [GLH^{+}24]. Labeling

[HGX^{+}24b, YXZ^{+}23]. Labeling-free [HGX^{+}24b]. LaF [HGX^{+}24b]. Landscape [LM24]. Landscapes [TZZ^{+}23]. Language [AROK21, CSFB24, DLT^{+}23, HSX^{+}24, HYX^{+}24, KJHY22, LMY^{+}24, LLZ^{+}22, WB13, WJZ^{+}23, XVN22, BGdV92, CL94, CFM00, GZ05, JPL98, SHO95, TY92, WAF00, MRRR02]. language-based [BGdV92, WAF00]. Languages [Bjø19, CXH23, FTF22, FTF24, BJMH02, BHR95, CDSM10, DvdHT05, HZS08, KSD08, RSB05, vdBV96]. Larch [CL94]. Larch/Smalltalk [CL94]. Large [BNB14, DNRN15, FA14, KCAS23, KCMS24, LJL^{+}20, LLZ^{+}22, MNB^{+}22, MAL24, RFDS24, Rus21, WT24a, WT24b, YZP^{+}22, MC08, PSV01]. Large-Scale [BNB14, FA14, LLZ^{+}22, MAL24, RFDS24, WT24a, WT24b, MNB^{+}22, Rus21, YZP^{+}22, PSV01]. Latency [LWC^{+}23]. latent [BGO^{+}14]. Lattice [DDE11]. Lattice-Based [DDE11]. Law [MU00]. Law-governed [MU00]. laws [LSV08]. layers [SB02]. Layout [LKR22]. lazy [FC00]. Leads [DPRO23]. Leads-to [DPRO23]. leak [XR13]. Leakage [WCB^{+}20]. Learn [KMYK19]. Learn-to-Rank [KMYK19]. Learned [RCAH23, TLL^{+}23]. Learning [ASNB19, BS16, BLBB23, BLBB24, BT22, BT24, CSB^{+}24, CYA22, CZSH23, CWM^{+}23, CLS^{+}23, DKD21, DLSC23, FCLL21, GHXZ24, GCC^{+}24, HGW^{+}16, HZX^{+}24, HGC^{+}22, HZB^{+}24, HCJ^{+}24, HDKM23, JYZ^{+}24, KKP^{+}22, KFY23, LLLM23, LCT^{+}23, LLL^{+}23, LGX^{+}22, LGC^{+}23, LHW^{+}24, LM24, MPT^{+}21, MZW^{+}23, MS15, NBMK22, OFS24, RFR23, RMCT22, SWW^{+}24, SNM23, TWB^{+}19, WLS^{+}20, WYC^{+}24, WCP^{+}22, WWAC23, XLD^{+}24, XAY23, YDX^{+}24, YHL^{+}22, ZFM^{+}24, ZLW^{+}21, ZHC24, ZLP^{+}24, ZRGT23, ZZX^{+}21, BG96, MY13]. Learning-based [BLBB24, BT22, YDX^{+}24, ZFM^{+}24, ZZX^{+}21]. legacy [THHB06]. Less [PBU16]. Lessons [RCAH23]. Level [BNB14, DG18, QWG^{+}23, AM04, CCRZ23, CTCC98, CUHH24, GLL^{+}23, KSD08, LWC^{+}23, LYL^{+}24, MMST14, Sin10, SXX^{+}24]. levels [CTC01]. Leveraging [AHGZ23, CH21, HCJ^{+}24, LLT^{+}24, VCF21]. lexical [MN96]. LibAM [LWD^{+}24]. Libraries [LWD^{+}24, LHW^{+}24, ZW95]. Library [DKD21, LCT^{+}23, OHDB92]. License [WBRH23, XGF^{+}23]. LiDetector [XGF^{+}23]. Life [VCF21, SS02]. Lightweight [GHM18, MN96, Jac02]. LIME [MPR06]. Line [CTA^{+}21, DL11, ZCT18, GLL^{+}23]. Linear [SP18, ZAW92]. Lines [HLL^{+}20, HAB^{+}20, XZZL18, XHL^{+}24, BJMH02, KATS12, MPG^{+}13]. Linking [KS20, SZH^{+}19, FC00]. Links [PWB23, DFOT07]. Literature [BUB23, CCS24, LCS^{+}22, SBMK21, TWS^{+}22, WCP^{+}22]. Little [LM24]. Live [TLJZ24]. liveness [DBPU13, SGE00].

Local [ABB^{+}18, TC20, ZLS^{+}23, ZWGX22]. Localisation [YXK^{+}17]. Localization [AFHC22, HMW23, JWC^{+}24, KMYK19, NAF^{+}23, TSPRC18, ZLS^{+}23, MA14, XCKX13, YHC13]. locating [TD01]. Location [RWEB19, PGM12, ZZL^{+}06]. Lockdowns [MKR^{+}23]. Locking [CCVLW23]. Log [CCW^{+}23]. LoGenText [DTC^{+}24]. LoGenText-Plus [DTC^{+}24]. Logging [DTC^{+}24]. Logic [BMM^{+}17, XLW^{+}22, DKM^{+}94, PMS13, TPT13, ZS97]. Logic-Based [BMM^{+}17]. logical [FGL^{+}12, MS94]. Logics [DDMM22, DJ97]. Longitudinal [JEB^{+}23, RHV24, RHvB24]. Look [HHC^{+}24, MCKG24, ZWZ^{+}24]. looking [Not13, Ros13a]. loop [BHL11]. loop-assigned [BHL11]. Loops [FYW24, MSM^{+}24, RD15]. Lop [FMMB23b, FMMB23a]. Lost [YHXY24]. LSCs [MHK11]. LTL [BLS11]. Lunch [RCAH23].

Machine

[BLBB23, CLL^{+}22, CZSH23, DKD21, DTC^{+}24, HDKM23, LLLM23, OFS24, RFR23, SWW^{+}24, TWB^{+}19, ZLW^{+}21]. Machine-Learning-Based [ZLW^{+}21]. Machines [KFY24, PSZ21, WJ10]. macro [Sin10]. macro-level [Sin10]. Made [WWJ^{+}24]. Mae [RVMRM04]. Maintainability [KDM17, KDM23]. Maintaining [WWAC23, ZCT18]. Maintenance [PLMC23]. majority [MM13]. Make [BKHT21, CMM^{+}15, SGR^{+}23]. Making [SGR^{+}23, KDM23]. Malware [Cai20, GHM18, WCG^{+}21, ZLW^{+}21, ZWY^{+}21]. Managed [NWB^{+}18]. Management [MGP^{+}22, WYMW20, ZQS^{+}23, DFOT07, ELvdH^{+}05, Jéz99]. Managers [SGR^{+}23]. Managing [HN98, RVMRM04]. Manifesting [VCP^{+}23]. Manifests [RSBP23]. Manipulation [MS14, TTL^{+}21, BG98]. Manual [BFFG19].

Manually [FFM24]. Many [HLL^{+}16, HLL^{+}20, MKS^{+}15, XZZL18]. Many-Objective [HLL^{+}16, HLL^{+}20, MKS^{+}15, XZZL18]. Mapping [HDG^{+}24, HCJ^{+}24, KRM24]. Markov [HGW^{+}16, WP93]. Marple [LS13]. marts [BCC^{+}01]. Masked [GZSW19, GXSC21]. Masking [GSC24]. Mastering [SPAS21]. Matching [ABC^{+}22, LKRF22, LWD^{+}24, WAH23, ZW95, ZW97]. Matching-based [ABC^{+}22]. Math [SPK14]. math.h [BCGB21]. math.h [?]Bagnara:2021:PAV. mathematical [Sne96]. Matter [CL23a]. Maturity [VHNF22]. MC [GRS^{+}16]. MC/DC [GRS^{+}16]. MDD [MGP^{+}13]. MDE [AB23]. MDESL [SZH^{+}19]. Me [VCP^{+}23]. Means [KKP^{+}22, BG96]. Measure [SHLW21]. Measurement [PGZ^{+}20, XSA^{+}24, GD08, MGP^{+}13]. Measures [MCB^{+}23, PSW^{+}20]. Measuring [BSA22, CLMPS^{+}24, DDS23, MP09, ZSHD20, CW99]. mechanism [MU00, WAF00]. mémoire [PWB23]. memoriam [Ros13b]. Memory [AEK^{+}16, CSB^{+}24, CYM^{+}24, CCXZ23, JPS22, LSX^{+}24, NWB^{+}18, PWB23, WWY^{+}23, XCS^{+}22, XR13]. Memory-related [CSB^{+}24]. Memory-Safety [XCS^{+}22]. Merge [SGHM23]. Merging [FDB^{+}12, RKBL19, LDUD13]. Message [RGS12, TG24, WXL^{+}21, CR94]. Meta [ZHO^{+}18, Kli93, HZS08]. Meta- [ZHO^{+}18]. meta-AspectJ [HZS08]. meta-environment [Kli93]. Metaconstraints [DDMM22]. metalocking [BS07]. Metamodel [LKRF22, PBO07]. Metamodel-based [PBO07]. Metamodels [AB23]. Metamorphic [SDLC23, SDG^{+}24]. Method [AB12, CUHH24, CF10, HGD^{+}23, OL22, WWL^{+}23, YZY^{+}24, BRRP05, MG00, MM13]. Method-level [CUHH24].

Method-Method [AB12].

Methodological [GLFW22]. Methodology [LBZ14, CTC01, DCS09, FGL^{+}12, FRB^{+}06, HGS93, KSD08, MPR^{+}13, MGMM11, RBL^{+}01, SCK13, ZJW03]. Methods [CZSH23, GGHM24, KRM24, KCAS23, SXX^{+}24, VBZ^{+}18, ZCL^{+}24, CMCP^{+}99, DFOT07, DBDS94, Rus21]. Metric [AB12, VHNF22, GCC^{+}24, PMS13]. Metrics [KS20, KS22, ZQS^{+}23, CSC06, KK04, MSW12, MB07]. MICOSE4aPS [VHNF22]. Microservice [ZQS^{+}23]. Microservices [CCRZ23]. middleware [EAS08, MPR06, VTA04]. Migration [CSW21, SPK14]. millions [MPG^{+}13]. Mined [GXG^{+}20]. Minimal [NL11]. Minimization [XL20]. Mining [BBS16, DNRN15, KL21, LCS^{+}22, PZS^{+}20, SPK14, SBS^{+}23, ZZL^{+}24]. Mining-based [PZS^{+}20]. Misconfigurations [RSBP23]. Missing [GCX^{+}22, SWW^{+}24, ZWL^{+}22, MM13]. Mitigating [PBC24]. Mitigation [CZSH23, MA14]. Mixed [GGHM24, KCAS23, Rus21]. Mixed-Methods [KCAS23, Rus21]. Mixin [SB02]. Mobile [ETM22, JZL^{+}19a, LCL^{+}23, RMP97, ZE14, CFM00, FGMP03, FC00, MZ09, MR99, PRM01]. Mobility [JZL^{+}19b, MPR06, PRM01]. Mock [ZWT^{+}24]. Mockups [RST^{+}14]. modal [WSG^{+}23]. Mode [MCKG24]. Model [ALMS22, AFHC22, ASNB19, BS16, BDL06, BS07, BHB16, CCS24, CXH^{+}21, CSFB24, DG17, DPRO23, EBE^{+}14, GRS^{+}16, GZW^{+}23, GTD21, HGX^{+}24b, HSX^{+}24, HYX^{+}24, JYZ^{+}24, LKRF22, DGD^{+}19, LCT^{+}23, LLSM22, LGC^{+}23, DDGR18, MMST14, NBB15, OPK^{+}21, PVHW17, LDUD13, Rus21, TSPRC18, WWAC23, WI23, WOM15, YZY^{+}24, BKM07, BGL00, CS12, Cal95, CDEG03, CW99, Dil93, Esh06, FGMP03, HAB13, JGB12, KF07, LL00, MS03, MN96, MPR06, ML00, NLR11, PBO07, RVMRM04, SMAC08, SS06, SGE00, TZZ09, VTA04, XM08]. Model-Based [AFHC22, CCS24, MMST14, WI23, HAB13, MS03, SS06, TZZ09]. model-checking [BGL00, CDEG03, FGMP03, KF07]. Model-Driven [DG17, GTD21]. Modeling [BRG^{+}01, BSA22, FMMH^{+}14, JZL^{+}19b, MFLL12, MR99, MRRR02, SLD^{+}13, SZZ^{+}23, BCFM06, BAD08, CDP04, DCS09, DHW98, PWD^{+}99, SB06]. Modelling [BZSW14, Bjø19, CFP^{+}24, DGC14, DGK21, Jac02]. Models [BMM^{+}17, BT24, BMC^{+}23, CYA22, DG18, EM18, FDB^{+}12, GZW^{+}23, HLL^{+}16, HHC^{+}24, JPS22, JYZ^{+}24, JCNS24, DGD^{+}19, LLL^{+}23, LLT^{+}24, LRL^{+}22, QSZG24, RGCS14, TC20, TZW^{+}23, WB13, YJW^{+}20, YXL^{+}22, YBL15, ZFL^{+}22, LLZ^{+}24, BDL06, CMCP^{+}99, CW98, Cor00, JPL98, MGP^{+}13, MG00, MPF14, ODV^{+}09, PTY95, SGG^{+}14, SCK13, UKM04, YBL13]. Models-Detection [ZFL^{+}22]. Modern [BUB23, RSB05]. modifiability [SGG^{+}14]. Modular [SGR^{+}10, WLS^{+}20, DvdHT05, FC00]. modularity [CS12, HE13, SPAK10]. Modularization [QSZG24, BGO^{+}14]. Molecular [EKL^{+}19]. monadic [TPT13]. Monitoring [CYM^{+}24, DDMM22, DDE11, EHF20, RBPG23]. Monotone [PSW^{+}20]. Most [XW22]. motivations [HBB^{+}09]. Mozilla [MFH02]. MPI [GSYT21]. Multi [AROK21, BS16, BKD^{+}23, CCRZ23, CDEG03, CLBY18, CL23b, DG18, GCC^{+}24, LVBBC^{+}18, OKS^{+}16, SYA21, XL20, ZHO^{+}18]. Multi-Criteria [OKS^{+}16, XL20]. Multi-Language [AROK21]. Multi-Level [DG18, CCRZ23]. Multi-metric [GCC^{+}24]. Multi-Objective [CLBY18, LVBBC^{+}18, ZHO^{+}18, BKD^{+}23, CL23b, SYA21, XL20]. Multi-Step [BS16]. Multi-valued [CDEG03]. Multiagent [DL11, EMM24, ZJW03]. Multidimensional [GL11]. Multilevel [DGC14]. Multilingual [LMY^{+}24].

Multimodal [MHK11]. multinational [Tiw08]. multiobjective [MY13]. Multiple [ZHC24]. Multiple-phase [ZHC24]. Multitolerant [EK11]. multiuser [KK93]. multiview [PBO07]. Mutant [ZWL^{+}22, OLR^{+}96]. Mutants [CPCT21, OSD^{+}23]. Mutation [DLP^{+}24, KJHY22, OSD^{+}23]. Mutation-based [DLP^{+}24]. My [CCH^{+}21, CXH^{+}21, PORH23].

Name [WWL^{+}23]. Names [SRTR17, YZY^{+}24]. Natural [KJHY22, WJZ^{+}23, XVN22, GZ05]. Near [OBH24]. Near-optimal [OBH24]. need [LYYC14]. Needs [CXH23]. nesting [MBH09]. Nets [RFDS24]. Network [CWW^{+}20, CWH^{+}21, CGZ23, DDS23, HQM^{+}23, HGX^{+}24b, QSZG24, QHM^{+}23, RXX^{+}19, WLS^{+}20, YFD^{+}23]. Network-based [RXX^{+}19]. Networked [LZS^{+}24]. Networks [BK23, DLP^{+}24, DLSC23, LD24, LXW^{+}24, LLSM22, PLM15, QWG^{+}23, SKY23, WT24a, WT24b, XLW^{+}22, ZZC^{+}23, ZSWS23, ZXS^{+}24, Sin10]. Neural [BK23, CWW^{+}20, CWH^{+}21, CGZ23, DLP^{+}24, DTC^{+}24, DDS23, DLY^{+}23, HGX^{+}24b, LD24, LXW^{+}24, LLSM22, QWG^{+}23, QSZG24, RXX^{+}19, SWW^{+}24, SKY23, TWB^{+}19, WT24a, WT24b, XLW^{+}22, YFD^{+}23, YZY^{+}24, ZYL^{+}23, ZZC^{+}23, ZSWS23, ZXS^{+}24]. Neuron [LXW^{+}24, XLW^{+}22]. next [HKMB^{+}14]. NN [GHX^{+}24]. NN-based [GHX^{+}24]. no [RCAH23]. Node [LJL^{+}20]. Node.js [ACSA23]. NodeJS [ZBA23]. Non [GGZ^{+}15, HT17, MWP^{+}21]. Non-Adequate [GGZ^{+}15]. Non-deterministic [HT17]. Non-Serializable [MWP^{+}21]. nonanomalous [DBPU13]. Nonequivalence [dFLSV14]. noninteractive [ZZL^{+}06]. noninterference [DFG00]. notation [FP02, Jac02]. notations [BP05, CDP04]. Notebook [LASA24]. Notebooks [LLG23]. Notkin [Ros13b]. Novice [CZ19]. NPC [XLW^{+}22]. npm [ACSA23, RCAH23]. NSFuzz [HQM^{+}23, QHM^{+}23]. NSGA [MKS^{+}15]. NSGA-III [MKS^{+}15]. Nudge [MUB^{+}23]. numerical [SMAC08]. Nygaard [Ano02].

Obfuscation [GHM18]. Obfuscation-Resilient [GHM18]. Obituary [Ano02]. Object [AB12, GGGU21, HLX23a, LHX21, MS94, TG11, WWL^{+}23, Cal95, CTCC98, CTC01, CSC06, DFB99a, DF94, Jac02, MRR05, RS09, RM03, SS02, SB02]. Object-Oriented [AB12, MS94, WWL^{+}23, CTCC98, CTC01, CSC06, DFB99a, DF94, RS09, RM03, SB02]. Object-Sensitive [HLX23a, LHX21]. objected [DFB99b]. objected-oriented [DFB99b]. Objective [CLBY18, HLL^{+}16, HLL^{+}20, LVBBC^{+}18, MKS^{+}15, SHLW21, TAA^{+}19, XZZL18, ZHO^{+}18, BKD^{+}23, CL23b, SYA21, XL20]. Objectives [CL23a]. Objects [ZWT^{+}24]. obliviousness [HE13]. OBSERV [TY92]. Observability [ZQS^{+}23]. Observable [HT17]. observational [PSV01]. OCL [QT12]. Octopus [TDL^{+}24]. Ole [Ano02]. Ole-Johan [Ano02]. On-device [HHC^{+}24, HDKM23]. One [Liu22]. Opacity [ALMS22]. Open [AHGZ23, BSA22, BKHT21, CWM^{+}20, JEB^{+}23, KCAS23, KCMS24, LMY^{+}24, PLZ^{+}22, RSBP23, RGCS14, TWS^{+}22, WFW^{+}20, WYC^{+}24, WBRH23, XGF^{+}23, ZA23, MFH02, Sin10]. Open-Source [BSA22, KCAS23, KCMS24, LMY^{+}24, PLZ^{+}22, RGCS14, JEB^{+}23, WYC^{+}24, Sin10]. OpenStack [BSOM24, ZLT^{+}22]. Operational [SZH^{+}19]. Operations [SCL^{+}23]. operators [OLR^{+}96]. Opinion [LCS^{+}22, CF10].

Opportunities [YXL^{+}22]. Optimal [HLL^{+}16, Liu22, OBH24]. Optimization [CLBY18, CS22, DPRO23, GCC^{+}24, HLL^{+}16, HGX^{+}24a, LLSM22, LVBBC^{+}18, LWC^{+}23, XZZL18, XHL^{+}24, AGHC^{+}22]. Options [ZKW^{+}23a, ZKW^{+}23b]. Oracles [WPB19, XM07]. Orchestra [TS09]. Order [GXSC21, GSC24, SGD15, LWF03, TPT13, XTW^{+}23]. Organizational [JEB^{+}23, DCS09]. Organized [ZWCH21]. Orientation [WWY^{+}23]. Oriented [AB12, QWG^{+}23, ZSL^{+}13, AM11, CTCC98, CTC01, CWM^{+}23, CSC06, DFB99a, DFB99b, DF94, DR10, FGL^{+}12, HE13, DGK21, MS94, ODV^{+}09, RS09, RM03, SB02, SGR^{+}10, WWL^{+}23]. OS-level [LYL^{+}24]. OSS [KS22, ZMM^{+}16]. outcome [GJ08]. Outgoing [Ros19]. output [KM10, QNR13]. Overdue [MUB^{+}23]. Overfitting [IRKY24]. Overflow [DLRA15, ARG17, GXG^{+}20, GXL^{+}23, HZX^{+}24, WJZ^{+}23, YZP^{+}22, ZWCH21]. Overhead [RCAH23]. overlapping [HaK92].

PACE [BK24]. Pacemaker [BZSW14]. Package [GHXZ24, JCSA23]. Packages [VCP^{+}23]. Pan [BGdV92]. Pandemic [FSZ^{+}22, RHV24, RHvB24]. Pandemic-RCR [RHvB24]. Parallel [HT17, KK93, PSV01, RD15, TDL^{+}24, ZCH22, SMAC08]. Parameter [LKA^{+}23, TG11]. parameterization [BAD08]. Parameterized [MRR05]. Parameters [RGCS14]. Parametric [ALMS22, WAH23]. Pareto [CL23b]. Part [ELN^{+}92]. Partial [FDB^{+}12, HYX^{+}24, LHX21, PBU16, SGD15, XCCY10]. Partial-Order [SGD15]. Participation [TWS^{+}22]. Particular [XL20]. Party [LWD^{+}24]. Passing [TG11]. Patch [HML23, LWWM22, MGTR18, STGR21, TLP^{+}22, YXX^{+}23]. PatchCensor [HML23]. Patches [IRKY24, KPC18, TLL^{+}23, TWB^{+}19, ZSW^{+}22]. Patching [CCVLW23]. Path [DDE11, GSYT21, QNR13, SGD15, TDL^{+}24, TPT13, XLW^{+}22, LS13, SRK06]. Path- [TPT13]. Path-Sensitive [SGD15]. Paths [YWC16]. Pattern [CZ19, SGR^{+}15, WAH23]. Pattern-based [CZ19]. Patterns [KAS20, LZL^{+}23, DDGR18, MS22, ZB13]. Peer [RGCS14]. Penetration [MAKM22]. Perceive [NMA^{+}23]. Perception [KSFC24]. Performance [BK24, CL23a, CGZ23, CFL^{+}16, GHX^{+}24, JYZ^{+}24, LLSM22, LLS^{+}21, RXX^{+}19, ZHZ^{+}21, Tiw08]. Person [ZSHD20]. Personality [CL22]. Personalized [ZL13]. Perspective [BXL^{+}22, FSP^{+}13, HGHM23, LGC^{+}23, YZY^{+}24]. pervasive [MZ09, XCCY10]. phase [JGB12, ZHC24]. philosophies [MSW12]. Phones [HDG^{+}24]. Phrase [YXZ^{+}23]. Phrase-based [YXZ^{+}23]. Physical [MCM24, ZTP^{+}23, CYA22, MSM^{+}24, XAY23, XTW^{+}23]. Pied [MRO^{+}23]. Pied-Piper [MRO^{+}23]. pilot [XM08]. Piper [MRO^{+}23]. PL [DL11]. Place [MS15]. Place-Aware [MS15]. Planet [WCJL23]. Planning [CSW21, OL22, ZHO^{+}18]. Plasticity [BT22]. Platform [EHEZ21, LJL^{+}20, LSH23, PORH23, ZOJH21]. Platforms [DZV^{+}23]. Platys [MS15]. Play [CCH^{+}21]. Plugins [LSH23]. Plus [DTC^{+}24]. Point [BCGB21, BTI14]. Pointer [HLX23a, LHX21]. pointers [OSH04]. points [LH08, MRR05]. points-to [LH08, MRR05]. Poison [LLZ^{+}24]. Policies [BBS16, BKHT21, BLW09]. Polychronous [GGLT07]. Polynomial [NKWF14]. Ponzi [ZCZ^{+}23]. Popular [CCH^{+}21]. Poracle [IRKY24]. Portfolio [MPG^{+}13]. Post [CDKP21, JMS08, JYZ^{+}24]. Post-release [JMS08]. Post-training [JYZ^{+}24]. Posteriori [DG17]. Posts [ARG17, GXL^{+}23, HZX^{+}24]. potentially [ZXLC14]. Power [GSC24, LSV08].

Powered [ZOJH21, CCY^{+}21]. Practical [BCGB21, CWW^{+}20, HB22, LYL^{+}24, SZH^{+}19, SSB20, YXX^{+}23]. Practice [BUB23, BXL^{+}22, OFS24, ELvdH^{+}05]. Practices [BKHT21, SER24, WHP^{+}23, ZTP^{+}23]. Practitioner [BXL^{+}22, HGHM23, PHS^{+}24]. pragmatic [HW12]. Pre [GZW^{+}23, LLL^{+}23, LLT^{+}24, LLZ^{+}22, WWL^{+}23]. Pre-implementation [WWL^{+}23]. Pre-Trained [GZW^{+}23, LLL^{+}23, LLT^{+}24]. Pre-Training [LLZ^{+}22]. Precise [AB12, KMYK19, LRCS14, WWY^{+}23, XR13, ZZC^{+}23]. Precision [LHX21, PSW^{+}20]. Precision-Preserving [LHX21]. Predict [FHS^{+}24]. Predicting [AHGZ23, CM23, LJL^{+}20, LM24, MBH^{+}17, TLP^{+}22, ZHZ^{+}21]. Prediction [BK24, CYA22, CXH^{+}21, CGZ23, CUHH24, FAP22, GLH^{+}24, NXL^{+}22, SMY19, TLL^{+}23, WWL^{+}23, YKY^{+}23, ZLL^{+}23, ZL13, ZYL^{+}18, ZXLC14]. Predictions [ZZX^{+}21]. Predictive [KJHY22, LRCS14, YXL^{+}22, HZZ13]. Preference [PLP^{+}23]. Preference-wise [PLP^{+}23]. preliminary [YTL^{+}95]. presence [FYD^{+}08, FC00, Hie02, OSH04]. Preservation [IRKY24]. Preserving [LHX21, YHR92]. Prevalence [MA14, ZLT^{+}22]. PRIME [PWD^{+}99, MGMM11]. Principles [Bjø19]. Prioritisation [RFR23]. Prioritization [BKD^{+}23, CCW^{+}23, DLP^{+}24, EM15, HZZ^{+}14, ZZY^{+}21, ZCH22, YHC13]. Privacy [BBS16, ML00]. Pro [YDX^{+}24]. Proactive [SMBK22]. Probabilistic [EM18, LSN^{+}23, LSBN24, PBU16, JGB12]. Probable [XW22]. Problem [IRKY24, SKBD14, XL20, HKMB^{+}14, MY13]. Problematizing [RFDS24]. Problems [CUHH24, ZA23]. procedure [BHR95, MGP^{+}13]. Procedures [LSX^{+}24]. Process [BCFM06, PVHW17, LDUD13, TG23, TC20, WYW^{+}22, ACF97, BCD02, BAD08, CW99, DHW98, GRT09, JPL98, ODV^{+}09, PWD^{+}99, SR05, SHO95]. process-centered [ACF97]. process-integrated [PWD^{+}99]. process-oriented [ODV^{+}09]. Processes [HGW^{+}16, Cia93, CW98, DCS09, FGMP03]. Processing [PORH23, ZFL^{+}22, LLZ^{+}24, ATW94, Bre95]. Product [CTA^{+}21, DL11, FHS^{+}24, HLL^{+}16, HLL^{+}20, KRM24, XZZL18, XHL^{+}24, ZCT18, BJMH02, KATS12, MC08]. product-lines [BJMH02]. Production [SYL^{+}23, TBS92]. Productivity [RHV24, RHvB24]. products [JMS08]. profiling [XR13]. Program [ASJDW21, BHR95, BK24, CH21, GWD^{+}21, GRS^{+}16, IRKY24, LZL^{+}23, MTRK14, RKBL19, SBS^{+}23, TZD^{+}24, XW22, YWC16, YLM^{+}23, YB20, ZFL^{+}22, ZKW^{+}23a, ZKW^{+}23b, ZFM^{+}24, ZH24, BG96, BGH07, BG98, DBGU13, EF05, GN93, HZS08, KM10, LH02, OSH04, SLB^{+}21, YHR92]. Programmed [EKL^{+}19]. Programmer [DAL^{+}23]. Programming [AFY^{+}22, DBNG15, BXX^{+}20, CWM^{+}20, CXH23, FYW24, FTF22, FTF24, KS20, KAT12, MZ09, SP18, TLJZ24, WWL^{+}23, WWJ^{+}24, XL20, YXK^{+}17, HE13, Kli93, RSB05, SHO95]. Programs [BCGB21, BK23, CCXZ23, EK11, GXSC21, GSYT21, HCJ^{+}24, IC14, KMYK19, LKR22, NBMK22, SDG^{+}24, WWAC23, Yu23, CTCC98, CF03, Cor00, DF94, FC00, HZZ13, MP09, PTY95, QRLV12, SMAC08, TS09, YTL^{+}95]. Progressed [GLL^{+}21, ZYL^{+}18]. Project [CM23, LM24, PWB23, ZYL^{+}18, Sin10]. Projections [MQLR16]. Projects [ACSA23, BSOM24, FAP22, JA22, JEB^{+}23, KCAS23, KCMS24, LASA24, NXL^{+}22, PLZ^{+}22, RGCS14, SGHM23, WFW^{+}20, WHP^{+}23, WYC^{+}24, ZMM^{+}16]. Promise [XVN22]. Promises [CUHH24]. Prompt [HSX^{+}24, HYX^{+}24]. Prompt-tuned [HSX^{+}24, HYX^{+}24]. Prone [AROK21].

Proof [FC00, RO18]. proofs [KKLS02]. Propagation [HDKM23]. Properties [DPRO23, EHF20, LBZ14, LLSM22, CK99, DCCN04, FPB^{+}05, JGB12, Sin10, SGE00]. Property [CSV13, DDE11, NBB15]. Prospective [Pez22]. Protecting [LZS^{+}24, ML00, ZWGX22]. Protocol [LZS^{+}24, Ber94, Pet97]. Protocols [GGGU21, CJM00, DFG00, PBCT10]. prototyping [TY92]. Prove [RO18]. provenance [MGMM11]. provenance-aware [MGMM11]. Proving [AB23]. psc2code [BXX^{+}20]. Psychometrics [GLFW22]. PTM [LLT^{+}24]. PTM-APIRec [LLT^{+}24]. Publication [DR15]. Pull [BYTS23, KCAS23, KCMS24, MUB^{+}23, PWB23]. Pull-Based [KCMS24]. purely [FPB^{+}05]. PVS [MPF14]. PyPI [GHXZ24]. Python [BLS^{+}23, HAB^{+}20].

Q&A [UKR21, YHG^{+}24]. Q8A [GXLG21]. QoS [VTA04]. QoS-enabled [VTA04]. Qualitative [AMS^{+}18, DZV^{+}23]. Quality [AAP^{+}20, CM23, CDKP21, CWM^{+}23, FTF22, HCW^{+}22, MBH^{+}17, PLMC23, RSBO23, XHL^{+}24, CF03, GL14, MGP^{+}13]. Quality-Diversity [XHL^{+}24]. Quality-oriented [CWM^{+}23]. Quantifying [GZSW19]. Quantitative [ZZC^{+}23]. quantitatively [CW99]. Quantum [FYW24, Yu23]. Queries [WJZ^{+}23, WGSD07]. Query [LXL^{+}22, MBH^{+}17, RR23, WPX^{+}23, PJRR10, WGG13]. Question [GXG^{+}20, GXLG21]. questions [KM10]. Quo [WFF^{+}19]. QuoTe [CWM^{+}23]. QVM [AVY11].

Race [CTC^{+}23, JPS22, RD15]. Random [EMM24, TAA^{+}19]. Randomization [ZWGX22]. Ranges [LSN^{+}23]. Rank [KMYK19]. Rationale [ZLT^{+}22, FBC^{+}13, YTL^{+}95]. RCR [FMMB23a, HQM^{+}23, RHvB24, WT24b, ZKW^{+}23b]. Re [ZWGX22, KRM24]. Re-Randomization [ZWGX22]. Reachability [LHX21, ZHXZ24, CK96, CK99, PTY95]. Reachability-guided [ZHXZ24]. Reactions [BYTS23]. reactive [FS93]. Read [Liu22]. Real [LSN^{+}23, LSBN24, MCKG24, TLJZ24, Cal95, FM94, FP02, GGLT07, MRK^{+}97, Ost99, Pon02, SLD^{+}13, WME93]. Real-Time [LSN^{+}23, TLJZ24, LSBN24, Cal95, FM94, FP02, GGLT07, MRK^{+}97, Ost99, Pon02, SLD^{+}13, WME93]. Real-world [MCKG24]. Reality [KRM24]. Realizable [TDL^{+}24]. Really [FSM^{+}15, YZP^{+}22]. Reasoning [EBE^{+}14, FPB^{+}05, GZ05, PRM01, CAC08, HN98, RMP97, VTA04]. Recall [PSW^{+}20]. ReCDroid [ZSL^{+}22]. Recognition [HSB^{+}23]. Recommendation [GXLG21, GXL^{+}23, GCC^{+}24, LLT^{+}24, SIB23, WYW^{+}22, YHG^{+}24]. Recommendations [OPK^{+}21, PDHR23]. Recommenders [AM11]. Recommending [DR11, JZL^{+}19a, SYA21, STS^{+}18]. recompilation [ATW94, Bre95]. Recomposition [BBS16]. Reconciling [SN92]. Recovering [DFOT07]. Recovery [LYL^{+}24]. reduce [Cor00]. reduced [Ber94, Pet97]. Reducing [AM11]. Reduction [TZD^{+}24, ZWL^{+}22, DBDS94]. Redundancy [CGPP15, SLWZ24, HZZ13]. Reengineering [Sne96, CF03]. Refactor [PZS^{+}20]. Refactoring [AB23, DG18, FSP^{+}13, LLG23, OKS^{+}16, TDT^{+}22, VBZ^{+}18, SGL12]. Refactorings [PLZ^{+}22]. referees [ACM05]. reference [BCC92]. Refinement [ASJDW21, Ban23, ZNZ^{+}23, Ost99]. refinements [SB02]. Reflection [LTX19, Pez22, SLB^{+}21]. reflective [LL00]. Reformulations [RR23]. region [DER10]. RegionTrack [MWP^{+}21]. regression [BRR01, GHK^{+}01, Mem08, RH97, REM^{+}04]. regular [KGA^{+}12]. Regulation [KAS20]. Regulatory [GL11]. Reifying [Jéz99]. reimplementation [CF03]. Reinforcement [BT22, BT24, FCLL21, GCC^{+}24, RMCT22]. Reject [PLZ^{+}22]. related [CSB^{+}24, HGHM23]. Relatedness [MB15]. Relation [HSX^{+}24]. Relational [MWK15, FPB^{+}05]. Relationships [HB22]. Relative [HAB^{+}20, MB20]. Relaxation [BLBB24]. Relaxed [JPS22]. Release [OL22, ZHO^{+}18, HKMB^{+}14, JMS08]. Relevance [OSD^{+}23]. relevant [MPG^{+}13]. Reliability [KDM23, ZL13, JMS08, PMM^{+}99, Wey96]. Reliable [LZL^{+}23]. Remodularization [CBRO16, MKS^{+}15]. Remote [SBBL23, WT24a, WT24b]. Remote-Code [SBBL23]. Removal [WRD^{+}22]. Removing [LCZL14, HZZ13]. Renaming [ZLL^{+}23]. Renamings [DLY^{+}23]. Repair [AFY^{+}22, CTC^{+}23, FNT^{+}24, GWD^{+}21, GCC^{+}24, IRKY24, LZL^{+}23, NBB15, OPK^{+}21, PVHW17, SKY23, YLM^{+}23, YABLR20, YB20, ZFM^{+}24, ZWT^{+}24, ZH24]. Repairing [LXW^{+}24, QWG^{+}23, Mem08]. Replicability [LGX^{+}22]. Replication [NXL^{+}22, RRPW21]. Report [FMMB23a, HQM^{+}23, RHvB24, WT24b, ZHV^{+}23, ZKW^{+}23b, AM11, KKP^{+}23]. Reported [DZV^{+}23]. Reports [AHGZ23, KS20, SXX^{+}24, ZSL^{+}22]. Repositories [GZW^{+}23, Hen97]. Repository [DNRN15, SBS^{+}23]. Representation [BC23, EW11, HZX^{+}24, KKP^{+}22, LLZ^{+}22, WLS^{+}20, ZLS^{+}23]. representations [BGL00]. Representing [RM07, DER10]. Reproducibility [LGX^{+}22]. Reproducing [JO15]. Reproduction [ZSL^{+}22]. Request [BYTS23, PWB23]. Request-Issue [PWB23]. Requests [KCAS23, MUB^{+}23]. Required [LK14]. Requirements [BLBB24, DPB17, GL11, HGHM23, LKRF22, DGD^{+}19, RST^{+}14, WFF^{+}19, ZZY^{+}21, CRST12, CD98, GM01, GZ05, HJL96, SMT92, SR05, UFG14, ZJ97]. Resampling [SMY19]. Research [CL22, MS22, PDHR23, SF18, WCP^{+}22, EAS08, ELvdH^{+}05, RSB05]. Residual [LRCS14]. Resilient [GHM18]. Resistance [GZSW19]. Resolution [BFFG19]. Resolving [AHGZ23]. Resource [WRD^{+}22]. Resource-constrained [WRD^{+}22]. Respect [DLY^{+}23]. Response [GZX^{+}22, TAW95]. Responses [LSB^{+}23]. REST [MZA22, ZA22]. RESTful [Arc19, GZA24, ZA23]. restructuring [BG98, GN93]. Result [BLS^{+}23]. results [DO93, PJRR10]. Retain [WWAC23]. retargetable [Dev99]. Retention [ZMM^{+}16]. Retraining [AFPB23]. Retrenchment [Ban23]. Retrieval [MBH^{+}17, SURL11, WXL^{+}21, WSG^{+}23, ZWCH21, ZLP^{+}24, DFOT07, PGM12]. Retrieval-based [WXL^{+}21]. Retrieving [PP93, WJZ^{+}23]. Retrospective [Pez22]. Return [ZWGX22]. reusable [BO92, PP93]. Reuse [FSP^{+}13, HLX^{+}23b, DGD^{+}19, VHNF22, EF05, Hen97, HW12, MC08, OHDB92]. Reusing [HGX^{+}24b, QSZG24]. Reveal [HLX^{+}23b]. Revealing [MRO^{+}23]. reverse [AM04]. Review [AMV23, CCS24, GZX^{+}22, LCS^{+}22, RR23, RGCS14, SBMK21, TG23, WCP^{+}22, WCJL23, HBB^{+}09]. Reviewers [ACM03]. Reviews [BUB23, CCH^{+}21, STDB23]. revisit [CCX11]. revisited [GD08]. Revisiting [MB20, SYL^{+}23]. Rewards [PBU16]. rewriting [VKV03]. RFC [TCDZ19]. RFC-guided [TCDZ19]. rich [RKBL19]. Rigorous [CSFB24]. Rise [LGC^{+}23, WCJL23]. Risk [GL11, LBZ14, MGP^{+}22, XCKX13]. Risks [ZWZ^{+}24]. Roads [LSM23]. Robotic [AMV23]. Robust [NAF^{+}23]. Robustness [ABC^{+}22, HAY^{+}23, HML23, WWAC23, YZY^{+}24, ZFL^{+}22, ZZS^{+}22]. Role [SPAS21, GJ08]. Route [LSM23]. router [CR94]. RPC [ZAL^{+}23]. RPC-Based [ZAL^{+}23]. Rule [ZZL^{+}24, Cia93, Kip92, MM13]. rule-based [Cia93, Kip92]. Rules [ARG17, MFLL12, KK04]. Run [ZHZ^{+}21]. Run-time [ZHZ^{+}21]. Runtime [AVY11, BLS11, BSOM24, EKL^{+}19, XMA^{+}14, ZCW^{+}24, ZWGX22, BLW09]. Rust [CXH23, CCXZ23, XCS^{+}22, ZWZ^{+}24].

SAEO [GSYT21]. Safe

[LSN^{+}23, LSBN24, BRR01, BTI14, RH97]. SafeDrop [CCXZ23]. Safety [AFPB23, CYM^{+}24, DZV^{+}23, FPBS23, KSFC24, NA24, RFDS24, XCS^{+}22, BFN^{+}14, CK99, SGE00, SRK06]. Safety-critical [FPBS23, NA24]. SAFKASI [WAF00]. Saltzer [PDHR23]. Sample [ZLW^{+}21]. Sampling [DDE11, PP93, PMM^{+}99]. Sanity [WYMW20]. SAT [AGRR19, XZZL18]. SAT/SMT [AGRR19]. satisfiability [BM13, PMS13]. Sator [BPT10]. SBST [ZBA23]. Scala [ARL^{+}15]. Scalability [CCRZ23]. Scalable [FCLL21, WWZ^{+}22, XMA^{+}14, BRRP05, HKMB^{+}14, HAB13]. Scale [BNB14, DNRN15, FA14, LJL^{+}20, LLZ^{+}22, MAL24, RFDS24, VLJ^{+}18, WT24a, WT24b, MNB^{+}22, PSV01, Rus21, YZP^{+}22]. Scaling [HZZ13, LCZL14, TDL^{+}24]. Scanner [YXM^{+}23]. Scenario [UKR21, UKM04, WJ10]. scenario-based [UKM04]. Scenarios [MHK11, NA24, SER24, UKM04]. Schedule [MQLR16]. Schema [MWK15, NL11]. Schemas [QT12]. Schemes [TZD^{+}24, ZCZ^{+}23]. Schroeder [PDHR23]. Scientific [CFP^{+}24, CY11, EF05, LYYC14]. Scope [MB20]. Scoping [WPX^{+}23]. Scores [RO18]. Screen [RST^{+}14]. Screencasts [BXX^{+}20]. Scripts [RRPW21, RO18].

Scrum [VR23]. SEADS [FCLL21]. SEAL [SBS^{+}23]. Search [AAP^{+}20, AG22, AVAS23, BS16, CCX^{+}20, CL23b, FFM24, GHX^{+}24, KFY24, OKS^{+}16, RR23, SYA21, SMBO21, SURL11, SKY23, SED14, TC20, WPX^{+}23, XLD^{+}24, XL20, ZYL^{+}23, ZHO^{+}18, ZZY^{+}21, ZA22]. Search-Based [FFM24, OKS^{+}16, SMBO21, SKY23, ZA22, AAP^{+}20, AG22, AVAS23, CL23b, KFY24, XL20]. Searching [GXL^{+}23, LXL^{+}22, MCB^{+}23, MPG^{+}13, ZZX^{+}21]. second [TPT13]. second-order [TPT13]. section [NP08]. Secure [ZQS^{+}23]. Securing [ZCZ^{+}23]. Security [AMGBK22, LSB^{+}23, PDHR23, RRPW21, RSBP23, RPT^{+}22, SGR^{+}23, SIB23, ZQS^{+}23, ZWZ^{+}24, ZSW^{+}22, BDL06, BLW09, CJM00, WAF00]. See [KKP^{+}22]. Seed [ZXS^{+}24]. Seeding [AVAS23]. Seeds [AVAS23]. Seeing [YLM^{+}23]. segments [LS13]. Selecting [PLMC23]. Selection [AVAS23, HLL^{+}16, HGC^{+}22, HGX^{+}24b, JS23, LD24, LMY^{+}24, MPT^{+}21, ZXS^{+}24, BRR01, CY11, GHK^{+}01, RH97]. selective [ATW94, Bre95]. Self [BKD^{+}23, CLBY18, CXLG22, GLL^{+}21, PBC24, RXX^{+}19, WI23, PJRR10]. Self-Adaptive [CLBY18, WI23]. Self-Admitted [RXX^{+}19, GLL^{+}21]. self-assessment [PJRR10]. Self-debugging [PBC24]. Self-Destruct [CXLG22]. Self-driving [BKD^{+}23]. Selfdestruct [CXLG22]. sem2vec [WMW^{+}23]. Semantic [BAD08, CLL^{+}22, HLX^{+}23b, MB15, PJRR10, SLWZ24, WSG^{+}23, YXX^{+}23, YHL^{+}22, ZMC^{+}24, MG00, KKP^{+}22]. Semantic-aware [ZMC^{+}24]. Semantic-Based [CLL^{+}22]. Semantic-Enriched [HLX^{+}23b]. Semantics [EHF20, LXL^{+}22, LK14, SZH^{+}19, WMW^{+}23, XTW^{+}23, YFD^{+}23, HN96, YHR92]. Semantics-aware [WMW^{+}23]. semantics-preserving [YHR92]. SemMT [CLL^{+}22]. Sense [SGR^{+}23]. Sensing [BZSW14]. Sensitive [HLX23a, LKR22, LHX21, SGD15, SCL^{+}23, ZHS^{+}23, Bro93, Cal95, For94, LH08, TPT13]. Sensitivity [LHX21, HKMB^{+}14, MRR05]. Sentiment [CCY^{+}21, UGKR22]. Sequence [RGS12, LK14, Mem08]. sequence-based [Mem08]. Sequences [HT17]. sequencing [OO92]. Sequential [DDI22, LXL^{+}22]. Serializable [MWP^{+}21]. Serverless [BPSSA^{+}22, LWC^{+}23, WCJL23]. Service [HQM^{+}23, LWC^{+}23, QHM^{+}23, BPT10, DPT13, FGL^{+}12, HWH14, PBCT10]. service-oriented [FGL^{+}12]. Services [ZL13, BKM07]. sets [Hie02]. Severity [SXX^{+}24]. Shadow [KPC18]. shape [Cor00]. Share [TLJZ24]. Should [CCH^{+}21, ZWL^{+}24]. shuttle [CD98]. Side [EWS14, GZSW19, GSC24, NKL^{+}24]. Side-Channel [EWS14, GSC24, GZSW19]. Signal [SZZ^{+}23, BRG^{+}01]. Signature [ZW95]. Significance [YKY^{+}23]. Significant [HZBS14]. Similarity [JFJ^{+}23, KS22, NAF^{+}23, TLP^{+}22, YXX^{+}23, YDX^{+}24, ZHXZ24, OHDB92]. Similarity-based [NAF^{+}23]. simplified [JW94]. Simulating [FS93, KDM23]. Simulation [Ban23, EHF20, KKLS02]. Simulator [FPBS23]. Simulator-based [FPBS23]. Single [BKD^{+}23, ZCT18]. SIP [HLL^{+}16]. Site [GXLG21, YHG^{+}24]. Sites [UKR21]. Six [Pez23]. Size [LLLM23, BGH07, GD08, HGS93, MGP^{+}13, TZZ09, ZXLC14]. slice [BGH07, MB07]. slice-based [MB07]. slices [BFN^{+}14]. Slicing [SNM23, XMA^{+}14, GSH97, LH02, TD01]. slicing-based [TD01]. SLR [PDHR23]. Small [HZBS14, Sin10]. small-world [Sin10]. Smalltalk [CL94]. Smart [AGHC^{+}22, ASJA23, CXLG22, CYM^{+}24, GCC^{+}24, HSB^{+}23, HLX^{+}23b, YABLR20, ZOJH21, ZCZ^{+}23]. SMC [SGE00]. Smell [SKBD14]. Smells [AROK21, DLT^{+}23, HZBS14, NMA^{+}23, RRPW21, YHXY24]. SMT [AGRR19, CH21]. SNIAFL [ZZL^{+}06]. Snippet [GXL^{+}23, GCH^{+}24]. Snippets [GXG^{+}20]. SNOW [Liu22]. SNOW-optimal [Liu22]. Socio [JA22]. Socio-Technical [JA22]. Sociotechnical [KAS20]. Software [AAP^{+}20, AHGZ23, BNB14, BXL^{+}22, BSA22, BKHT21, Böh18, CL22, CBRO16, CTA^{+}21, CLBY18, CCY^{+}21, CL23a, CL23b, CWH^{+}21, CSW21, CFP^{+}24, CW99, CFL^{+}16, CZ19, DZV^{+}23, DHW98, DLY^{+}23, DR15, DNRN15, EWS14, EM18, EF05, EW11, FSZ^{+}22, FFM24, FSM^{+}15, FTF22, FTF24, GZSW19, GLFW22, GGHM24, HGHM23, HLL^{+}20, HH95, JPS22, JEB^{+}23, KS22, KDM23, KRM24, KPC18, LLLM23, LD24, LMY^{+}24, LCS^{+}22, LGX^{+}22, LGC^{+}23, LSB^{+}23, LM24, MNB^{+}22, MFBF^{+}22, MGP^{+}22, MFLL12, MBH^{+}17, MY13, MB20, MKS^{+}15, MZW^{+}23, MKR^{+}23, MP14, OKS^{+}16, Pez24a, RKBL19, RGCS14, RFDS24, SP18, SLWZ24, STS^{+}18, SMY^{+}24, SBMK21, SMBK22, SMY19, SF18, SGhWG22, TDT^{+}22, TWS^{+}22, UGKR22, VHNF22, WB13, WPB19, WCP^{+}22, WSG^{+}23, WOM15, XZZL18, XHL^{+}24, XGF^{+}23, XL20, YXL^{+}22, YQTR15, ZSHD20, ZHZ^{+}21, AAG95, ACF97, BCTW96, BO92, BGO^{+}14, BCD02]. software [CS12, CTC01, CM08, Cia93, CW98, CDP04, CD98, DvdHT05, DFOT07, DCCN04, ELN^{+}92, ELvdH^{+}05, FK96, FLM^{+}98, GJ08, Gun00, HBB^{+}09, Hen97, HW12, JPL98, JMS08, KK93, LASL13, LSV08, MMST14, MRRR02, MSW12, MFH02, MC08, NLR11, NP08, Off92, ODV^{+}09, PSV01, PP93, PMM^{+}99, PSMV98, Rob08, RSB05, SRK06, SN92, SHO95, TZZ09, Tiw08, TBS92, UFG14, VD13, WP93, WGG13, XM07, XR13, ZW95, ZW97, DKD21]. Software-2.0 [DKD21]. Software-engineering [XL20].

Solution [LJL^{+}20]. Solutions

[LLS^{+}21, XL20]. Solved [XCS^{+}22]. solver [KGA^{+}12]. Solvers [XZZL18]. Solving [AGRR19, CH21, EMM24, SSB20, SED14, XL20]. Some [AVAS23, HZBS14]. Sound [MWP^{+}21, XL20]. Source [AHGZ23, BSA22, BKHT21, CWM^{+}20, DNRN15, GGH^{+}23, HNRA20, KS20, KCAS23, KCMS24, KJHY22, LLL^{+}23, LMY^{+}24, LLT^{+}24, MB15, PZS^{+}20, PLZ^{+}22, RSBP23, RR23, RGCS14, SGR^{+}15, SRTR17, SED14, SFC^{+}24, SZZ^{+}23, TWS^{+}22, WLS^{+}20, WFW^{+}20, WBRH23, XGF^{+}23, ZHS^{+}23, LLZ^{+}24, Dev99, DER10, JEB^{+}23, MFH02, MN96, RM07, SGG^{+}14, Sin10, SAB^{+}14, WYC^{+}24]. Source-Code [DNRN15, SGR^{+}15, SGG^{+}14]. SourcererJBF [MAL24]. sources [PSMV98]. Space [STS^{+}18, ZRGT23, CD98, DBDS94]. Spaces [OBH24, WSG^{+}23]. Special [HP15, MP14, NP08]. specialization [SS02]. Species [Böh18]. Specific [OFS24, ZZW^{+}23, BJMH02, BGL00, HZS08, SS06]. Specification [KAS20, KL21, XW22, ZW97, Bro93, CDSM10, CL94, CR94, ELN^{+}92, FN03, Kuh99, LY05, MS94, PPP94, RMP97, TK02, WME93]. specification-based [Kuh99, LY05, TK02]. Specifications [EHF20, PSW^{+}20, CCX11, DSV03, FM94, HJL96, HRD08, HN98, Jac95b, KB07, MMST14, MS03, Pon02, PMS13, UKM04, WP93, FPGA07]. specify [CFM00]. Specifying [PSZ21, DKM^{+}94]. spectra [NLR11]. spectra-based [NLR11]. Spectrum [TSPRC18, YXK^{+}17, MSW12, XCKX13]. Spectrum-Based [TSPRC18, YXK^{+}17, XCKX13]. Speculative [WCB^{+}20]. Speed [ODE21, TZD^{+}24]. Speeding [TTL^{+}21]. spi [DSV03, ZSW^{+}22]. Splitting [LLS^{+}21]. spreadsheet [FRB^{+}06]. spreadsheets [RBL^{+}01]. Spurious [JCNS24]. SQL [AG20, TG24]. SSL [NWZ^{+}23, TCDZ19]. SSL/TLS [NWZ^{+}23, TCDZ19]. Stack [AEK^{+}16, ARG17, GXG^{+}20, GXL^{+}23, HZX^{+}24, WJZ^{+}23, YZP^{+}22, ZWCH21]. Stacks [ZWGX22]. STADS [Böh18]. Stages [LSN^{+}23, LSBN24]. Stale [KCMS24]. Stand [UGKR22]. Stand-alone [UGKR22]. Standard [WWAC23]. start [LWC^{+}23]. State [ASJA23, EM18, HQM^{+}23, MS14, OFS24, PSZ21, QHM^{+}23, RWEB19, WB13, Cor00, DBDS94, WJ10]. State-Aware [HQM^{+}23, QHM^{+}23]. State-Based [WB13]. statecharts [BRG^{+}01, HaK92, HN96]. Stateful [BPSSA^{+}22, SLD^{+}13]. STATEMATE [HN96]. Statement [BC23]. Statement-Based [BC23]. States [LCZL14]. Static [CTC^{+}23, CCXZ23, HNRA20, IWY00, KMYK19, LZL^{+}23, RD15, RWEB19, RM03, VLJ^{+}18, WGSD07, ZCZ^{+}23, BGH07, FPB^{+}05, GSH97, MNGL98, OO92, ZZL^{+}06]. Statically [CWH^{+}21, CCVLW23]. Statistical [OBH24, RGCS14, ZNZ^{+}23]. Status [CYM^{+}24, WFF^{+}19]. Step [BS16, SLB^{+}21]. Stepwise [EK11]. Stochastic [CFL^{+}16]. Storage [ASJA23, ZWGX22]. Strategies [AVAS23, JCSA23, JCNS24]. strategy [JW94]. Stratified [DPRO23, PMM^{+}99]. StreamGen [GTD21]. Streaming [GTD21]. strength [MP09]. Stress [DBNG15, MSM^{+}24]. string [TPT13]. strings [KGA^{+}12]. Strong [AVAS23]. Structural [Kip92]. Structure [GGH^{+}23, GRS^{+}16, WB13, RM03]. Structure-Guided [GGH^{+}23]. Structured [Yu23, BP98]. Structures [JEB^{+}23, KDM17]. Structuring [Jac95b]. Stub [ZWT^{+}24]. Stubborn [CPCT21]. StubCoder [ZWT^{+}24]. Studies [Cai20, DAL^{+}23, SPAS21, BM07, CD98, HBB^{+}09, MFH02]. Studio [TLJZ24]. Study [AROK21, AMGBK22, BYTS23, BSOM24, BAJA23, CWM^{+}20, CCH^{+}21, CZSH23, CXH23, DKD21, EHEZ21, FAP22, FSM^{+}15, GZW^{+}23, GGHM24, GLL^{+}21, HGC^{+}22, HGD^{+}23, JEB^{+}23, KRM24, KCAS23, KCMS24, KSFC24, LLSM22, LASA24, LLS^{+}21, NXL^{+}22, ODE21, OKS^{+}16, PORH23, PZS^{+}20, RRPW21, RSBP23, Rus21, SBBL23, SGHM23, TLJZ24, TTL^{+}21, TWB^{+}19, UGKR22, VCP^{+}23, VLJ^{+}18, WYC^{+}24, XCS^{+}22, YZP^{+}22, ZTP^{+}23, ZOJH21, ZE14, ZHO^{+}18, ZA22, ZAL^{+}23, ZMM^{+}16, BJMH02, BRR01, BGH07, CF10, GHK^{+}01, MB07, MNGL98, PSV01, SMT92, SR05, Tiw08, TBS92, XM08, ZXLC14]. Studying [OSD^{+}23]. style [AAG95]. Subject [DPB17]. Suboptimal [WHP^{+}23]. Substantial [WWAC23]. Subtle [ZSWS23]. Success [Rus21, Sin10]. Successful [JA22]. Such [RCAH23]. Sufficient [MCB^{+}23, OLR^{+}96]. Suggestion [ODE21, ZLL^{+}23]. Suite [HLL^{+}20, XHL^{+}24, Ber94, HGS93, Pet97, REM^{+}04, XL20, YTL^{+}95]. Suites [GGZ^{+}15, ZWL^{+}24, Mem08]. Summaries [CTC^{+}23, PXJ17]. Summarization [CXH^{+}21, GGH^{+}23, SFC^{+}24, ZLP^{+}24]. Super [AGHC^{+}22]. Super-optimization [AGHC^{+}22]. Supervisors [WT24a, WT24b]. Supply [GHXZ24]. Support [DBNG15, JO15, SURL11, ZCT18, BFN^{+}14, HWH14, MS03, RM03]. Supporting [BG98, DR10, SMY^{+}24, Ham09, MPR06]. Surprise [KFY23]. Surrogate [BT24]. Survey [BUB23, GZA24, HGX^{+}24a, MFBF^{+}22, PKHM22, TZZ^{+}23, TWS^{+}22, XLD^{+}24, ZFM^{+}24]. Surveys [WFF^{+}19]. Sustainability [Cai20]. Symbolic [AEK^{+}16, BHB16, CPCT21, Esh06, KPC18, LCZL14, RGS12, WCB^{+}20, YPRK14, BGL00, CDEG03, QNR13, SMAC08]. symmetry [SGE00]. symmetry-based [SGE00]. symposium [NP08]. Synchronisation [AGRR19]. Synchronization [YKY^{+}23, YXM^{+}23].

Syntactic [DTC^{+}24]. Synthesis [LXW^{+}24, MMST14]. synthesized [PWX14]. Synthesizing [DBPU13, WJ10, XW22, DL13]. Synthetic [SMY19]. SysML [BFN^{+}14]. System [AG20, BLX^{+}20, BSOM24, CYA22, CCS24, KFY23, LBZ14, PBU16, SSB20, TZZ^{+}23, ZA22, ZCT18, BGdV92, CDSM10, IWY00, MMST14, MG00, OHDB92, RVMRM04, TBS92, WME93]. system-level [MMST14]. Systematic [AMV23, CCS24, LCS^{+}22, RR23, SBMK21, TG23, VHNF22, WCP^{+}22, WCJL23, HBB^{+}09, MS03]. Systematically [YLM^{+}23]. Systematizing [HW12]. SystemC [YJW^{+}20]. Systems [AMS^{+}18, AGRR19, AMV23, AFHC22, AVY11, AHGZ23, BNB14, BT22, CLL^{+}22, CCS24, CWM^{+}20, CWM^{+}23, CGZ23, DL11, DPB17, DDI22, EKL^{+}19, EMM24, FPBS23, FDC^{+}21, FCLL21, GGPR24, HAY^{+}23, KAS20, KSFC24, LKA^{+}23, LSN^{+}23, LSBN24, LMY^{+}24, MPT^{+}21, MSM^{+}24, MCM24, MFBF^{+}22, NWB^{+}18, OSD^{+}23, PSW^{+}20, SYA21, SER24, TG11, WI23, XAY23, XTW^{+}23, YBZK21, ZTP^{+}23, ZHZ^{+}21, ZRGT23, BO92, BCD02, Bro93, Cal95, CMP13, CY11, CFM00, CRST12, CDFG96, DFOT07, DJ97, DKM^{+}94, DCCN04, FM94, FP02, FS93, GM01, JGB12, Kip92, KK04, LYYC14, MU00, MS94, MRK^{+}97, Ost99, ODV^{+}09, Pon02, RM03, SLD^{+}13, TZZ09, THHB06, WAF00, ZJW03]. SZZ [GLH^{+}24].

TACCLE [CTC01]. Tactics [YBZK21]. Tagger [WYMW20]. tailoring [CF10]. Taint [HB22]. Taken [LSM23]. Tale [FSZ^{+}22]. Taming [SLB^{+}21]. Target [KMYK19]. Task [DBNG15, BBS16, Dil97, DLSC23, ZSHD20, SCK13]. Task-agnostic [DLSC23]. tasking [Dil93]. Tasks [FSP^{+}13, MBH^{+}17, PLMC23, SWW^{+}24, TTL^{+}21, XSA^{+}24]. Taxonomy [RFR23, UFG14]. Team [VR23]. Teams [CFL^{+}16]. Technical [GXLG21, GLL^{+}21, JA22, RXX^{+}19, UKR21, YHG^{+}24]. technique [KKLS02, RH97, SS06, SB02]. Techniques [Bjø19, DPRO23, RD15, BRR01, GHK^{+}01, SM12]. technology [EAS08]. telecom [MC08]. telecommunication [Zav04]. Templates [CWM^{+}20, DTC^{+}24]. Temporal [CY11, DDMM22, Pon02, LYYC14, PMS13]. Tensor [BSA22, YFD^{+}23]. Term [VKV03]. Termination [TAA^{+}19, Dil97]. Terminology [MGP^{+}22]. Terms [WB13]. Test [Arc19, AG20, AVAS23, BKD^{+}23, CMM^{+}15, CCW^{+}23, DLP^{+}24, DPT13, EM15, EMM24, FA14, FSM^{+}15, GRS^{+}16, GGZ^{+}15, GSYT21, HZZ^{+}14, HLL^{+}20, HAB^{+}20, HGC^{+}22, HGX^{+}24a, IC14, JS23, JCNS24, KFY24, KB07, MPT^{+}21, MZA22, MWK15, MGTR18, NAF^{+}23, PLP^{+}23, RFR23, SYA21, SMBO21, SSB20, SYL^{+}23, TLP^{+}22, XHL^{+}24, XL20, YHXY24, ZA22, ZWL^{+}24, ZCH22, Ber94, BRR01, DO93, FK96, FRB^{+}06, GHK^{+}01, HGS93, HAB13, Hie02, Hie09, KSD08, Mem08, PWX14, Pet97, RH97, REM^{+}04, SS06, UFG14, XM07]. Test-and-adapt [DPT13]. Test-Equivalence [MGTR18]. Test-Generation [EMM24]. test-selection [BRR01]. Test-suite [XL20]. Testability [AG22, BHL11, MBH09]. Tester [MAKM22]. Testers [FSM^{+}15]. Testing [DBNG15, AMV23, AG22, ABC^{+}22, BG96, BT22, BT24, Böh18, BK23, CLL^{+}22, CWW^{+}20, CS22, CWM^{+}23, CFP^{+}24, DPB17, DDS23, FFM24, GJS^{+}23, GGPR24, GGGU21, GZA24, GHX^{+}24, Hie14, HGX^{+}24a, HML23, HZB^{+}24, IRKY24, KFY23, KPC18, LKA^{+}23, LCT^{+}23, LSM23, MSM^{+}24, MCM24, MJS^{+}21, MAKM22, MB20, NPB22, NL11, NWZ^{+}23, OSD^{+}23, PLP^{+}23, PGZ^{+}20, RMCT22, SDLC23, SDG^{+}24, TZZ^{+}23, TCDZ19, TAA^{+}19, WPB19, WYC^{+}24, WSG^{+}23, ZSHD20, ZSWS23, ZXS^{+}24, ZRGT23, Ber94, CTCC98, CTC01, CM08, DRW96, DF94, DSV03, FRB^{+}06, Ham09, HAB13, Hie09, JW94, KSD08, Kip92, Kuh99, LY05, MPR^{+}13, MBH09, Mem08, MS03, NP08, Off92, OSH04, Pet97, RBL^{+}01, REM^{+}04, SS06, SM12, TM14, TK02, Wey96, XM08, ZAW92]. Testing-based [WSG^{+}23, Ham09]. Tests [GWD^{+}21, PKHM22, SPK14, WC23, ZE14]. TestSGD [ZSWS23]. Text [HSX^{+}24, MBH^{+}17]. Texts [DTC^{+}24]. Their [AMGBK22, WB13, MPG^{+}13]. Them [NMA^{+}23]. Theorem [Yu23]. theoretic [YHC13]. Theoretical [SZH^{+}19, YXK^{+}17, XCKX13]. Theory [FSP^{+}13, RGCS14, VR23, WFF^{+}19, HBB^{+}09, Ham09, PPP94]. There [RCAH23, ZWL^{+}22]. Thing [RCAH23]. Things [YBZK21]. Third [LWD^{+}24]. Third-Party [LWD^{+}24]. Thread [ZWGX22]. Thread-Local [ZWGX22]. Threats [AMGBK22, MRO^{+}23]. Three [BM07, ZMM^{+}16, CSC06]. Time [GLH^{+}24, LSN^{+}23, NXL^{+}22, TLJZ24, TDT^{+}22, WWZ^{+}22, Bro93, Cal95, FM94, FP02, GGLT07, LSBN24, MS94, MRK^{+}97, Ost99, Pon02, SLD^{+}13, WME93, ZHZ^{+}21]. time-critical [MS94]. time-sensitive [Bro93, Cal95]. Time-travel [WWZ^{+}22]. Timed [ALMS22, BMM^{+}17, WAH23, PBCT10, SLD^{+}13]. Timeliness [WPB19]. Times [PORH23]. Tiny [HDKM23]. Titles [GXG^{+}20]. TLS [NWZ^{+}23, TCDZ19]. TLTL [BLS11]. Token [HSB^{+}23, MRO^{+}23]. TokenAware [HSB^{+}23]. Tool [MNB^{+}22, dFLSV14, CSX08, ELN^{+}92, MS03, MPF14, YTL^{+}95, ZW95]. Tools [CL22, FSP^{+}13, GLH^{+}24, Ham09, SIB23, UGKR22, ZA23, DRW96, Tiw08]. Topic [YHG^{+}24]. Topic-informed [YHG^{+}24]. TopicAns [YHG^{+}24]. Topics [AHGZ23, BGO^{+}14]. Topology [Rob08]. TOSEM [Pez23]. TOTA [MZ09]. Toxic [STDB23]. ToxiCR [STDB23]. Trace [MWP^{+}21, ZLS^{+}23, HZZ13]. Trace-Based [MWP^{+}21]. Traceability [BFN^{+}14, DFOT07]. Tracelet [WMW^{+}23]. Traces [DDMM22, LYL^{+}24, MWP^{+}21]. Tracking [HB22, DER10]. Trading [HE13]. Trained [GZW^{+}23, LLL^{+}23, LLT^{+}24]. Training [BK23, LLZ^{+}22, LGC^{+}23, JYZ^{+}24]. Trait [PSZ21]. Transaction [PORH23]. Transactional [MWP^{+}21]. Transactions [Liu22, ZOJH21]. Transfer [HCJ^{+}24, KKP^{+}22]. Transformation [JYZ^{+}24, LKRF22, Rus21, BHL11, CF03, MBH09]. Transformations [ASJDW21, AG22, BS16, DGD^{+}19, NBMK22, TSPRC18, YHR92]. Transformer [FNT^{+}24, GGH^{+}23]. Transformers [HML23]. Transforming [MHK11, BHL11]. transition [YBL13]. Translation [CLL^{+}22, DTC^{+}24, TWB^{+}19, Zav04]. Transplantation [STGR21]. travel [WWZ^{+}22]. traversal [VKV03]. Tree [JWC^{+}24, WLS^{+}20]. Trends [TZZ^{+}23]. triage [AM11]. Triggering [ZSL^{+}13, FPBS23]. Triggers [SLWZ24]. TRIO [FM94]. trustability [HH95]. tuned [HSX^{+}24, HYX^{+}24]. Tuning [CL23a, LLSM22]. Turnover [ZLT^{+}22]. Tutorials [WJZ^{+}23]. TVs [HDG^{+}24]. Twin [XAY23]. Twin-based [XAY23]. Twitter [STS^{+}18]. Two [FSZ^{+}22, MFH02, WT24a, WT24b, JGB12]. two-phase [JGB12]. Type [ARL^{+}15, KATS12, ODE21, SGD15, BGL00, TD01]. Type-Based [ARL^{+}15, SGD15]. type-specific [BGL00]. typed [HCJ^{+}24]. Types [ASJDW21, SPAK10, SXX^{+}24]. typestate [FYD^{+}08]. Typing [DG17, DGD^{+}19].

UAV [DZV^{+}23]. UI

[CCX^{+}20, CCH^{+}21, LSM23]. Ultra [DNRN15, LJL^{+}20]. Ultra-Large-Scale [DNRN15, LJL^{+}20]. UML [BM13, BMM^{+}17, BDL06, Esh06, MFLL12, QT12, SGG^{+}14, SB06, YBL15]. UML-B [SB06]. Unbounded [FYW24, JGB12]. Uncertainty [BMC^{+}23, CYA22, GLH^{+}24, HAY^{+}23, LKA^{+}23, OL22, SBMK21, WPB19, ZZY^{+}21, GJ08]. Uncertainty-Aware [HAY^{+}23, CYA22]. Uncertainty-wise [ZZY^{+}21]. Uncovering [YLM^{+}23]. Understand [GL11, AAG95]. Understandable [CZ19]. Understanding [ASMP16, CLS^{+}23, DLRA15, DKD21, KCAS23, KCMS24, LTX19, NWB^{+}18, PSMV98, RHV24, RHvB24, TLJZ24, YLM^{+}23]. Undiscovered [YHXY24]. Unified [HZZ^{+}14, HMW23, ZS97, MRRR02]. Unifying [CST16, RS09]. UniLoc [HMW23]. Uniqueness [WC23]. Unit [FA14, FSM^{+}15, KSD08, MJS^{+}21, SPK14, WYC^{+}24]. Unit-level [KSD08]. UNITY [MR99, PRM01, RMP97]. Unknown [SGR^{+}23]. Unknowns [HLX^{+}23b]. Unnecessary [HNRA20]. Unsupervised [PKMY24, YXZ^{+}23]. until [JGB12]. Unveiling [WFW^{+}20]. Update [JCSA23, NPB22]. Update-driven [NPB22]. Upgrade [ZZW^{+}23]. upper [CM08]. Usage [DKD21, DLT^{+}23, GZW^{+}23, KS20, MB20, UKR21, ZOJH21]. usages [MPG^{+}13]. Use [CDKP21, DGC14, LMY^{+}24, SCK13, WCP^{+}22, WT24a, WT24b, YZP^{+}22, YBL15, DJ97, HBB^{+}09, YBL13]. Used [CWM^{+}20]. User [CCH^{+}21, DAL^{+}23, FHS^{+}24, GGHM24, BRRP05, LASL13, SMT92]. Using [AGRR19, BBS16, CL22, CBRO16, CWH^{+}21, CFM00, Cor00, CTC^{+}23, DLSC23, ETM22, EM15, FHS^{+}24, FA14, GSYT21, HLL^{+}16, HAB^{+}20, KFY24, KS22, KMYK19, LKRF22, LD24, MVM07, MGP^{+}13, MKS^{+}15, NBMK22, OKS^{+}16, SYA21, STDB23, Wey96, XM08, XMA^{+}14, XTW^{+}23, YZP^{+}22, ZCL^{+}24, ZCZ^{+}23, ALMS22, CK99, CSFB24, DFOT07, DFG00, IWY00, KK04, LS13, LH08, ML00, PGM12, SLWZ24, SLD^{+}13, TC20, UKM04, XR13, ZW95]. utility [CSC06]. UTP [XTW^{+}23].

Validate [ZE14]. Validating [FM94, MSW12]. Validation [AMV23, BAJA23, CCS24, CRST12, FDC^{+}21, NWZ^{+}23, PDHR23, QT12, TCDZ19, CW99, DBGU13]. Validator [CYA22]. Value [LM24, TDL^{+}24, WWY^{+}23, FBC^{+}13, Hie06]. Value-Flow [TDL^{+}24, WWY^{+}23]. valued [CDEG03]. Variability [AMS^{+}18, VLJ^{+}18]. Variability-Aware [VLJ^{+}18]. Variable [CST16, DLY^{+}23, JWC^{+}24, ZYL^{+}23]. Variable-based [JWC^{+}24, ZYL^{+}23]. Variant [RKBL19]. Variant-rich [RKBL19]. variants [Jéz99]. Variation [EW11, SPAS21, PSMV98]. Variational [KFY24]. Various [GZW^{+}23]. Vehicles [NA24]. Verdict [Hie09]. Verification [ASJDW21, AGRR19, AMV23, BCGB21, BMM^{+}17, BLS11, BDD^{+}22, CCS24, CSV13, CH21, DDI22, EWS14, FDC^{+}21, GXSC21, GSC24, HGW^{+}16, LKR22, LD24, Liu22, NBB15, QT12, RBPG23, YXX^{+}23, BGL00, CDSM10, CY11, DSV03, FGL^{+}12, FGMP03, FYD^{+}08, FC00, MPF14, SGE00, WME93]. Verified [AFY^{+}22]. Verifix [AFY^{+}22]. verify [SMAC08]. Verifying [CJM00, GZSW19, DCCN04, SLD^{+}13]. Version [ARG17]. versioning [ZS97]. Versions [VCF21]. versus [CL23b, EMM24]. Very [LM24]. via [BGO^{+}14, CS22, CCW^{+}23, CCXZ23, DDI22, FCLL21, GWD^{+}21, GXLG21, HML23, HSX^{+}24, JYZ^{+}24, JWC^{+}24, KAS20, KJHY22, LLLM23, LXW^{+}24, NA24, PWX14, PLP^{+}23, PWB23, SMY19, TDL^{+}24, TWB^{+}19, WCB^{+}20, WWY^{+}23, XLW^{+}22, YXX^{+}23, YHL^{+}22, ZAW92, ZSW^{+}22, ZMC^{+}24, ZZW^{+}23]. Video [NMA^{+}23]. Views [DL13, Jac95b].

Violation [ZZL^{+}24, LYYC14]. Violations [MWP^{+}21, MM13]. Virtual [BFFG19, BKD^{+}23, KRM24, DHW98, Pon02]. Vision [FNT^{+}24]. Visual [ABC^{+}22, CDP04, Dil93, KSD08, MG00, TLJZ24]. Visualization [KKP^{+}22, BG98]. Visualizing [BLX^{+}20]. VM [LSX^{+}24]. Voice [FHS^{+}24]. vs [SRTR17]. Vulnerabilities [ACSA23, CSB^{+}24, CWH^{+}21, MS14, SBBL23]. Vulnerability [FNT^{+}24, GWD^{+}21, GCX^{+}22, SLWZ24, SXX^{+}24, SZZ^{+}23, YXX^{+}23, YXM^{+}23, YXZ^{+}23, ZMC^{+}24, ZZX^{+}21]. Vulnerable [BLS^{+}23].

warehouses [BCC^{+}01]. Wasted [KCAS23]. Way [DPRO23, RKBL19]. WCET [LSN^{+}23, LSBN24]. Weak [FDB^{+}12]. Weakly [LSBN24]. Web [ZBA23, LASL13, BM07, BCFM06, BPT10, CGPP15, NAF^{+}23, NBB15, SURL11, YXM^{+}23, ZL13]. web-centred [LASL13]. WebAssembly [ZCW^{+}24]. Weighted [CL23b, HGW^{+}16]. Weights [CL23b]. Well [RHV24, RHvB24, SURL11, TC20, ZWCH21]. Well-Being [RHV24, RHvB24]. Well-Formed [TC20]. White [ZAL^{+}23, CTCC98]. White-Box [ZAL^{+}23]. Who [STS^{+}18]. Whole [SLB^{+}21, YLM^{+}23]. Whole-program [SLB^{+}21]. Wild [AFHC22, DLT^{+}23, TWB^{+}19]. Wireframe [CCX^{+}20]. Wireframe-based [CCX^{+}20]. wise [PLP^{+}23, ZZY^{+}21]. within [DHW98]. Witnesses [BDD^{+}22]. Women [TWS^{+}22]. Word [ETM22, SRTR17, KGA^{+}12]. WordPress [LSH23]. Words [LXL^{+}22]. Work [CXH^{+}21]. Workarounds [CGPP15]. workflow [CY11, LYYC14]. Workflows [HDKM23, OFS24]. Working [DAL^{+}23, FSZ^{+}22, MKR^{+}23]. World [YHXY24, MCKG24, Sin10]. Worlds [TLL^{+}23]. Wrapper [THHB06]. Wrapper-based [THHB06]. Wybe [Ano02]. WYSIWYT [FRB^{+}06].

X.509 [NWZ^{+}23]. XCode [LLZ^{+}22]. XCoS [WPX^{+}23]. XP [CF10]. XPIs [SGR^{+}10].

Year [Pez23, Pez24b, RHV24, RHvB24].

Years [PDHR23].

Z [Jac95b].

References

Abowd:1995:FSU

[AAG95] Gregory D. Abowd, Robert Allen, and David Garlan. Formalizing style to understand descriptions of software architecture. ACM Transactions on Software Engineering and Methodology, 4(4):319–364, October 1995. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

                                                    www.acm.org/pubs/articles/
                                                    journals/tosem/1995-4-4/p319-
                                                    abowd/p319-abowd.pdf;   http:
                                                    //www.acm.org/pubs/citations/
                                                    journals/tosem/1995-4-4/p319-
                                                    abowd/.

Ali:2020:QIS

[AAP^{+}20] Shaukat Ali, Paolo Arcaini, Dipesh Pradhan, Safdar Aqeel Safdar, and Tao Yue. Quality indicators in search-based software engineering: an empirical evaluation. ACM Transactions on Software Engineering and Methodology, 29(2):10:1– 10:29, April 2020. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

                                                    URL https://dl.acm.org/
                                                    doi/abs/10.1145/3375636.

AlDallal:2012:PMM

[AB12] Jehad Al Dallal and Lionel C. Briand. A precise method-method interactionbased cohesion metric for object-oriented classes. ACM Transactions on Software Engineering and Methodology, 21 (2):8:1–8:??, March 2012. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Altoyan:2023:PCR

[AB23] Najd Altoyan and Don Batory. On proving the correctness of refactoring class diagrams of MDE metamodels. ACM Transactions on Software Engineering and Methodology, 32 (2):44:1–44:??, April 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3549541.

Attie:2018:GLD

[ABB^{+}18] Paul C. Attie, Saddek Bensalem, Marius Bozga, Mohamad Jaber, Joseph Sifakis, and Fadi A. Zaraket. Global and local deadlock freedom in BIP. ACM Transactions on Software Engineering and Methodology, 26(3):9:1–9:??, January 2018. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Ardito:2022:FMB

[ABC^{+}22] Luca Ardito, Andrea Bottino, Riccardo Coppola, Fabrizio Lamberti, Francesco Manigrasso, Lia Morra, and Marco Torchiano. Feature matchingbased approaches to improve the robustness of Android visual GUI testing. ACM Transactions on Software Engineering and Methodology, 31(2): 21:1–21:32, April 2022. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3477427.

Ambriola:1997:APC

[ACF97] Vincenzo Ambriola, Reidar Conradi, and Alfonso Fuggetta. Assessing process-centered software engineering environments. ACM Transactions on Software Engineering and Methodology, 6(3):283–328, July 1997. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1997-6-3/p283-
ambriola/p283-ambriola.pdf;
 http://www.acm.org/pubs/
citations/journals/tosem/
1997-6-3/p283-ambriola/.

Staff:2003:R

[ACM03] ACM Transactions on Software Engineering and Methodology staff. Reviewers 2002. ACM Transactions on Software Engineering and Methodology, 12(1): 105, 2003. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Staff:2005:AR

[ACM05] ACM Transactions on Software Engineering and Methodology staff. Acknowledgement of referees 2004. ACM Transactions on Software Engineering and Methodology, 14(2):246, April 2005. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Alfadel:2023:DNV

[ACSA23] Mahmoud Alfadel, Diego Elias Costa, Emad Shihab, and Bram Adams. On the discoverability of npm vulnerabilities in node.js projects. ACM Transactions on Software Engineering and Methodology, 32(4): 91:1–91:??, July 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3571848.

Anand:2016:SMA

[AEK^{+}16] Kapil Anand, Khaled Elwazeer, Aparna Kotha, Matthew Smithson, Rajeev Barua, and Angelos Keromytis. A stack memory abstraction and symbolic analysis framework for executables. ACM Transactions on Software Engineering and Methodology, 25(2):19:1–19:??, May 2016. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Arcega:2022:BLM

[AFHC22] Lorena Arcega, Jaime Font, Øystein Haugen, and Carlos Cetina. Bug localization in model-based systems in the wild. ACM Transactions on Software Engineering and Methodology, 31(1):10:1– 10:32, January 2022. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3472616.

Attaoui:2023:BBS

[AFPB23] Mohammed Attaoui, Hazem Fahmy, Fabrizio Pastore, and Lionel Briand. Black-box safety analysis and retraining of DNNs based on feature extraction and clustering. ACM Transactions on Software Engineering and Methodology, 32(3):79:1–79:??, July 2023. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3550271.

Ahmed:2022:VVR

[AFY^{+}22] Umair Z. Ahmed, Zhiyu Fan, Jooyong Yi, Omar I. Al-Bataineh, and Abhik Roychoudhury. Verifix: Verified repair of programming assignments. ACM Transactions on Software Engineering and Methodology, 31(4):74:1–74:??, October 2022. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3510418.

Allen:1997:FBA

[AG97] Robert Allen and David Garlan. A formal basis for architectural connection. ACM Transactions on Software Engineering and Methodology, 6(3):213–249, July 1997. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1997-6-3/p213-
allen/p213-allen.pdf;   http:
//www.acm.org/pubs/citations/
journals/tosem/1997-6-3/p213-

allen/. See errata [AG98].

Allen:1998:EFB

[AG98] Robert Allen and David Garlan. Errata: “A formal basis for architectural connection”. ACM Transactions on Software Engineering and Methodology, 7(3):333–334, July 1998. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1998-7-3/p333-
allen/p333-allen.pdf;   http:
//www.acm.org/pubs/citations/
journals/tosem/1998-7-3/p333-

allen/. See [AG97].

Arcuri:2020:HSD

[AG20] Andrea Arcuri and Juan P.

Galeotti. Handling SQL databases in automated system test generation. ACM Transactions on Software Engineering and Methodology, 29(4): 22:1–22:31, October 2020. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3391533.

Arcuri:2022:ESB

[AG22] Andrea Arcuri and Juan P. Galeotti. Enhancing search-based testing with testability transformations for existing APIs. ACM Transactions on Software Engineering and Methodology, 31(1): 1:1–1:34, January 2022. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3477271.

Albert:2022:SOS

[AGHC^{+}22] Elvira Albert, Pablo Gordillo, Alejandro Hernández-Cerezo, Albert Rubio, and Maria A. Schett. Super-optimization of smart contracts. ACM Transactions on Software Engineering and Methodology, 31(4): 70:1–70:??, October 2022. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3506800.

Antonino:2019:EVC

[AGRR19] Pedro Antonino, Thomas Gibson-Robinson, and A. W. Roscoe. Efficient verification of concurrent systems using synchronisation analysis and SAT/ SMT solving. ACM Transactions on Software Engineering and Methodology, 28(3): 18:1–18:??, August 2019. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.
                                        acm.org/ft_gateway.cfm?id=
                                        3335149.

Assi:2023:PCI

[AHGZ23] Maram Assi, Safwat Hassan, Stefanos Georgiou, and Ying Zou. Predicting the change impact of resolving defects by leveraging the topics of issue reports in open source software systems. ACM Transactions on Software Engineering and Methodology, 32(6):141:1– 141:??, November 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3593802.

Andre:2022:GTO

[ALMS22] Étienne André, Didier Lime, Dylan Marinho, and Jun Sun. Guaranteeing timed opacity using parametric timed model checking. ACM Transactions on Software Engineering and Methodology, 31(4):64:1– 64:??, October 2022. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3502851.

Akgul:2004:AIL

[AM04] Tankut Akgul and Vincent J.

Mooney III. Assembly instruction level reverse execution for debugging. ACM Transactions on Software Engineering and Methodology, 13(2):149– 198, April 2004. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Anvik:2011:REB

[AM11] John Anvik and Gail C. Murphy. Reducing the effort of bug report triage: Recommenders for developmentoriented decisions. ACM Transactions on Software Engineering and Methodology, 20(3):10:1– 10:??, August 2011. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Ahmadjee:2022:SBA

[AMGBK22] Sabreen Ahmadjee, Carlos Mera-Gómez, Rami Bahsoon, and Rick Kazman. A study on blockchain architecture design decisions and their security attacks and threats. ACM Transactions on Software Engineering and Methodology, 31(2): 36e:1–36e:45, April 2022. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3502740.

Abal:2018:VBH

[AMS^{+}18] Iago Abal, Jean Melo, Stefan Stanciulescu, Claus Brabrand, Márcio Ribeiro, and Andrzej Wasowski. Variability bugs in highly configurable systems: a qualitative analysis. ACM Transactions on Software Engineering and Methodology, 26(3): 10:1–10:??, January 2018. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Araujo:2023:TVV

[AMV23] Hugo Araujo, Mohammad Reza Mousavi, and Mahsa Varshosaz. Testing, validation, and verification of robotic and autonomous systems: a systematic review. ACM Transactions on Software Engineering and Methodology, 32(2):51:1–51:??, April 2023. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3542945.

Anonymous:1996:AI

[Ano96] Anonymous. Author index.

ACM Transactions on Software Engineering and Methodology, 5(4):422–423, October 1996. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1996-5-4/p422-
authorindex/p422-authorindex.
pdf;   http://www.acm.org/
pubs/citations/journals/tosem/
1996-5-4/p422-author\_index/
; http://www.acm.org/pubs/
citations/journals/tosem/
1996-5-4/p422-author_index/

.

Anonymous:2002:OOJ

[Ano02] Anonymous. Obituary: Ole-Johan Dahl, 1931–2002; Edsger Wybe Dijkstra, 1930–2002; Kristen Nygaard, 1926–2002. ACM Transactions on Software Engineering and Methodology, 11(4):385, 2002. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Arcuri:2019:RAA

[Arc19] Andrea Arcuri. RESTful API automated test case generation with EvoMaster. ACM Transactions on Software Engineering and Methodology, 28(1):3:1– 3:??, February 2019. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Azad:2017:GAC

[ARG17] Shams Azad, Peter C. Rigby, and Latifa Guerrouj. Generating API call rules from version history and Stack Overflow posts. ACM Transactions on Software Engineering and Methodology, 25(4):29:1–29:??, May 2017. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Ali:2015:TBC

[ARL^{+}15] Karim Ali, Marianna Rapoport, Ondrej Lhoták, Julian Dolby, and Frank Tip. Type-based call graph construction algorithms for Scala. ACM Transactions on Software Engineering and Methodology, 25(1):9:1– 9:??, December 2015. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Abidi:2021:MLD

[AROK21] Mouna Abidi, Md Saidur Rahman, Moses Openja, and Foutse Khomh. Are multi-language design smells fault-prone? An empirical study. ACM Transactions on Software Engineering and Methodology, 30(3): 29:1–29:56, May 2021. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3432690.

Ayub:2023:SSA

[ASJA23] Maha Ayub, Tania Saleem, Muhammad Janjua, and Talha Ahmad. Storage state analysis and extraction of Ethereum blockchain smart contracts. ACM Transactions on Software Engineering and Methodology, 32(3):57:1–57:??, July 2023. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3548683.

Al-Sibahi:2021:VPT

[ASJDW21] Ahmad Salim Al-Sibahi, Thomas P.

Jensen, Aleksandar S. Dimovski, and Andrzej Wasowski. Verification of program transformations with inductive refinement types. ACM Transactions on Software Engineering and Methodology, 30(1): 5:1–5:33, January 2021. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3409805.

Alimadadi:2016:UJE

[ASMP16] Saba Alimadadi, Sheldon Sequeira, Ali Mesbah, and Karthik Pattabiraman. Understanding JavaScript eventbased interactions with Clematis. ACM Transactions on Software Engineering and Methodology, 25(2):12:1–12:??, May 2016. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Arora:2019:ALA

[ASNB19] Chetan Arora, Mehrdad Sabetzadeh, Shiva Nejati, and Lionel Briand. An active learning approach for improving the accuracy of automated domain model extraction. ACM Transactions on Software Engineering and Methodology, 28(1):4:1– 4:??, February 2019. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Adams:1994:CSR

[ATW94] Rolf Adams, Walter Tichy, and Annette Weinert. The cost of selective recompilation and environment processing. ACM Transactions on Software Engineering and Methodology, 3(1):3–28, January 1994. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

                                        www.acm.org/pubs/articles/
                                        journals/tosem/1994-3-1/p3-
                                        adams/p3-adams.pdf;   http:
                                        //www.acm.org/pubs/citations/
                                        journals/tosem/1994-3-1/p3-

adams/. See comment [Bre95, TAW95].

Arrieta:2023:SSS

[AVAS23] Aitor Arrieta, Pablo Valle, Joseba A. Agirre, and Goiuria Sagardui. Some seeds are strong: Seeding strategies for search-based test case selection. ACM Transactions on Software Engineering and Methodology, 32(1):17:1–17:??, January 2023. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3532182.

Arnold:2011:QER

[AVY11] Matthew Arnold, Martin Vechev, and Eran Yahav. QVM: An efficient runtime for detecting defects in deployed systems. ACM Transactions on Software Engineering and Methodology, 21(1): 2:1–2:??, December 2011. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Breaux:2008:SPP

[BAD08] Travis D. Breaux, Annie I.

Antón, and Jon Doyle. Semantic parameterization: a process for modeling domain descriptions. ACM Transactions on Software Engineering and Methodology, 18(2):5:1–5:??, November 2008. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Bock:2023:ACD

[BAJA23] Thomas Bock, Nils Alznauer, Mitchell Joblin, and Sven Apel. Automatic core-developer identification on GitHub: a validation study. ACM Transactions on Software Engineering and Methodology, 32(6):138:1– 138:??, November 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3593803.

Banach:2023:GRR

[Ban23] Richard Banach. Graded refinement, retrenchment, and simulation. ACM Transactions on Software Engineering and Methodology, 32(2): 29:1–29:??, April 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3534116.

Bhatia:2016:MPG

[BBS16] Jaspreet Bhatia, Travis D. Breaux, and Florian Schaub. Mining privacy goals from privacy policies using hybridized task recomposition. ACM Transactions on Software Engineering and Methodology, 25(3): 22:1–22:??, August 2016. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Bertolotti:2023:FTS

[BC23] Francesco Bertolotti and Walter Cazzola. Fold2Vec: Towards a statement-based representation of code for code comprehension. ACM Transactions on Software Engineering and Methodology, 32(1):6:1–6:??, January 2023. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3514232.

Basili:1992:RAC

[BCC92] Victor R. Basili, Gianluigi Caldiera, and Giovanni Cantone. A reference architecture for the component factory. ACM Transactions on Software Engineering and Methodology, 1(1):53–80, January 1992. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1992-1-1/p53-
basili/p53-basili.pdf;   http:
//www.acm.org/pubs/citations/
journals/tosem/1992-1-1/p53-
basili/.

Bonifati:2001:DDM

[BCC^{+}01] Angela Bonifati, Fabiano Cattaneo, Stefano Ceri, Alfonso Fuggetta, and Stefano Paraboschi. Designing data marts for data warehouses. ACM Transactions on Software Engineering and Methodology, 10(4): 452–483, October 2001. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Bernardo:2002:AFS

[BCD02] Marco Bernardo, Paolo Ciancarini, and Lorenzo Donatiello. Architecting families of software systems with process algebras. ACM Transactions on Software Engineering and Methodology, 11(4):386–426, 2002. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Brambilla:2006:PMW

[BCFM06] Marco Brambilla, Stefano Ceri, Piero Fraternali, and Ioana Manolescu. Process modeling in Web applications. ACM Transactions on Software Engineering and Methodology, 15(4):360– 409, October 2006. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Bagnara:2021:PAV

[BCGB21] Roberto Bagnara, Michele Chiari, Roberta Gori, and Abramo Bagnara. A practical approach to verification of floating-point C/C++ programs with math.h/cmath functions. ACM Transactions on Software Engineering and Methodology, 30(1): 9:1–9:53, January 2021. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3410875.

Barrett:1996:FEB

[BCTW96] Daniel J. Barrett, Lori A. Clarke, Peri L. Tarr, and Alexander E. Wise. A framework for event-based software integration. ACM Transactions on Software Engineering and Methodology, 5(4):378–421, October 1996. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

                                        www.acm.org/pubs/articles/
                                        journals/tosem/1996-5-4/p378-
                                        barrett/p378-barrett.pdf;
                                         http://www.acm.org/pubs/

citations/journals/tosem/
1996-5-4/p378-barrett/.

Beyer:2022:VW

[BDD^{+}22] Dirk Beyer, Matthias Dangl, Daniel Dietsch, Matthias Heizmann, Thomas Lemberger, and Michael Tautschnig. Verification witnesses. ACM Transactions on Software Engineering and Methodology, 31(4): 57:1–57:??, October 2022. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3477579.

Basin:2006:MDS

[BDL06] David Basin, Jürgen Doser, and Torsten Lodderstedt. Model driven security: From UML models to access control infrastructures. ACM Transactions on Software Engineering and Methodology, 15(1):39–91, January 2006. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Bernhard:1994:RTS

[Ber94] Philip J. Bernhard. A reduced test suite for protocol conformance testing. ACM Transactions on Software Engineering and Methodology, 3(3):201–220, July 1994. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1994-3-3/p201-
bernhard/p201-bernhard.pdf;
http://www.acm.org/pubs/
citations/journals/tosem/

1994-3-3/p201-bernhard/. See

comments [Pet97].

Bernaschina:2019:VDI

[BFFG19] Carlo Bernaschina, Emanuele Falzone, Piero Fraternali, and Sergio Luis Herrera Gonzalez. The virtual developer: Integrating code generation and manual development with conflict resolution. ACM Transactions on Software Engineering and Methodology, 28(4): 20:1–20:??, October 2019. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?id=
3340545.

Briand:2014:TSD

[BFN^{+}14] Lionel Briand, Davide Falessi, Shiva Nejati, Mehrdad Sabetzadeh, and Tao Yue. Traceability and SysML design slices to support safety inspections: a controlled experiment. ACM Transactions on Software Engineering and Methodology, 23(1): 9:1–9:??, February 2014. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Bergadano:1996:TMI

[BG96] Francesco Bergadano and Daniele Gunetti. Testing by means of inductive program learning. ACM Transactions on Software Engineering and Methodology, 5(2):119–145, April 1996. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1996-5-2/p119-
bergadano/p119-bergadano.
pdf;   http://www.acm.org/
pubs/citations/journals/tosem/
1996-5-2/p119-bergadano/.

Bowdidge:1998:SRD

[BG98] Robert W. Bowdidge and William G. Griswold. Supporting the restructuring of data abstractions through manipulation of a program visualization. ACM Transactions on Software Engineering and Methodology, 7(2):109–157, April 1998. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1998-7-2/p109-
bowdidge/p109-bowdidge.pdf;
 http://www.acm.org/pubs/
citations/journals/tosem/
1998-7-2/p109-bowdidge/.

Ballance:1992:PLB

[BGdV92] Robert A. Ballance, Susan L.

Graham, and Michael L. Van de Vanter. The Pan languagebased editing system. ACM Transactions on Software Engineering and Methodology, 1(1):95–127, January 1992. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1992-1-1/p95-
ballance/p95-ballance.pdf;
 http://www.acm.org/pubs/

                                        citations/journals/tosem/
                                        1992-1-1/p95-ballance/.

Binkley:2007:ESS

[BGH07] David Binkley, Nicolas Gold, and Mark Harman. An empirical study of static program slice size. ACM Transactions on Software Engineering and Methodology, 16(2):8:1– 8:??, April 2007. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Bultan:2000:CMC

[BGL00] Tevfik Bultan, Richard Gerber, and Christopher League. Composite model-checking: verification with type-specific symbolic representations. ACM Transactions on Software Engineering and Methodology, 9(1):3– 50, January 2000. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

                                        URL http://www.acm.org/
                                        pubs/articles/journals/tosem/
                                        2000-9-1/p3-bultan/p3-bultan.
                                        pdf;   http://www.acm.org/
                                        pubs/citations/journals/tosem/
                                        2000-9-1/p3-bultan/.

Bavota:2014:ISM

[BGO^{+}14] Gabriele Bavota, Malcom Gethers, Rocco Oliveto, Denys Poshyvanyk, and Andrea de Lucia. Improving software modularization via automated analysis of latent topics and dependencies. ACM Transactions on Software Engineering and Methodology, 23(1):4:1– 4:??, February 2014. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Bauch:2016:CED

[BHB16] Petr Bauch, Vojtech Havel, and Jirí Barnat. Control explicitdata symbolic model checking. ACM Transactions on Software Engineering and Methodology, 25(2):15:1–15:??, May 2016. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Binkley:2011:FTT

[BHL11] David W. Binkley, Mark Harman, and Kiran Lakhotia. FlagRemover: a testability transformation for transforming loop-assigned flags. ACM Transactions on Software Engineering and Methodology, 20(3): 12:1–12:??, August 2011. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Binkley:1995:PIL

[BHR95] David Binkley, Susan Horwitz, and Thomas Reps. Program integration for languages with procedure calls. ACM Transactions on Software Engineering and Methodology, 4(1):3–35, January 1995. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1995-4-1/p3-
binkley/p3-binkley.pdf;
http://www.acm.org/pubs/citations/

                                        journals/tosem/1995-4-1/p3-
                                        binkley/.

Batory:2002:AET

[BJMH02] Don Batory, Clay Johnson, Bob Macdonald, and Dale Von Heeder. Achieving extensibility through product-lines and domain-specific languages: a case study. ACM Transactions on Software Engineering and Methodology, 11(2):191– 214, 2002. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Bjorner:2019:DAD

[Bjø19] Dines Bjørner. Domain analysis and description principles, techniques, and modelling languages. ACM Transactions on Software Engineering and Methodology, 28(2): 8:1–8:??, April 2019. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.
                                        acm.org/ft_gateway.cfm?id=
                                        3295738.

Braiek:2023:TFN

[BK23] Houssem Ben Braiek and Foutse Khomh. Testing feedforward neural networks training programs. ACM Transactions on Software Engineering and Methodology, 32(4): 105:1–105:??, July 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3529318.

Biringa:2024:PPA

[BK24] Chidera Biringa and Gökhan Kul. PACE: a program analysis framework for continuous performance prediction. ACM Transactions on Software Engineering and Methodology, 33 (4):85:1–85:??, May 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3637230.

Birchler:2023:SMO

[BKD^{+}23] Christian Birchler, Sajad Khatiri, Pouria Derakhshanfar, Sebastiano Panichella, and Annibale Panichella. Single and multiobjective test cases prioritization for self-driving cars in virtual environments. ACM Transactions on Software Engineering and Methodology, 32(2): 28:1–28:??, April 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3533818.

Bogart:2021:WHM

[BKHT21] Chris Bogart, Christian Kästner, James Herbsleb, and Ferdian Thung. When and how to make breaking changes: Policies and practices in 18 open source software ecosystems. ACM Transactions on Software Engineering and Methodology, 30(4): 42:1–42:56, July 2021. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3447245.

Broy:2007:FMS

[BKM07] Manfred Broy, Ingolf H. Krüger, and Michael Meisinger. A formal model of services. ACM Transactions on Software Engineering and Methodology, 16(1): 5:1–5:??, February 2007. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Belgacem:2023:MLA

[BLBB23] Hichem Belgacem, Xiaochen Li, Domenico Bianculli, and Lionel Briand. A machine learning approach for automated filling of categorical fields in data entry forms. ACM Transactions on Software Engineering and Methodology, 32(2): 47:1–47:??, April 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3533021.

Belgacem:2024:LBR

[BLBB24] Hichem Belgacem, Xiaochen Li, Domenico Bianculli, and Lionel Briand. Learning-based relaxation of completeness requirements for data entry forms. ACM Transactions on Software Engineering and Methodology, 33(3):77:1–77:??, March 2024. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3635708.

Bauer:2011:RVL

[BLS11] Andreas Bauer, Martin Leucker, and Christian Schallhart. Runtime verification for LTL and TLTL. ACM Transactions on Software Engineering and Methodology, 20(4):14:1–14:??, September 2011. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Brun:2023:BPJ

[BLS^{+}23] Yuriy Brun, Tian Lin, Jessie Elise Somerville, Elisha M. Myers, and Natalie Ebner. Blindspots in Python and Java APIs result in vulnerable code. ACM Transactions on Software Engineering and Methodology, 32 (3):76:1–76:??, July 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3571850.

Bauer:2009:CER

[BLW09] Lujo Bauer, Jay Ligatti, and David Walker. Composing expressive runtime security policies. ACM Transactions on Software Engineering and Methodology, 18(3): 9:1–9:??, May 2009. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Beschastnikh:2020:VDS

[BLX^{+}20] Ivan Beschastnikh, Perry Liu, Albert Xing, Patty Wang, Yuriy Brun, and Michael D. Ernst. Visualizing distributed system executions. ACM Transactions on Software Engineering and Methodology, 29(2): 9:1–9:38, April 2020. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3375633.

Baresi:2007:TES

[BM07] Luciano Baresi and Sandro Morasca. Three empirical studies on estimating the design effort of Web applications. ACM Transactions on Software Engineering and Methodology, 16 (4):15:1–15:??, September 2007. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Balaban:2013:FSU

[BM13] Mira Balaban and Azzam Maraee. Finite satisfiability of UML class diagrams with constrained class hierarchy. ACM Transactions on Software Engineering and Methodology, 22 (3):24:1–24:??, July 2013. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Burgueno:2023:DBU

[BMC^{+}23] Lola Burgueño, Paula Muñoz, Robert Clarisó, Jordi Cabot, Sébastien Gérard, and Antonio Vallecillo. Dealing with belief uncertainty in domain models. ACM Transactions on Software Engineering and Methodology, 32(2):31:1–31:??, April 2023. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3542947.

Baresi:2017:LBA

[BMM^{+}17] Luciano Baresi, Angelo Morzenti, Alfredo Motta, Mohammad Mehdi Pourhashem K., and Matteo Rossi. A logic-based approach for the verification of UML timed models. ACM Transactions on Software Engineering and Methodology, 26(2): 7:1–7:??, October 2017. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Behjati:2014:ALC

[BNB14] Razieh Behjati, Shiva Nejati, and Lionel C. Briand. Architecture-level configuration of large-scale embedded software systems. ACM Transactions on Software Engineering and Methodology, 23(3):25:1– 25:??, May 2014. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Batory:1992:DIH

[BO92] Don Batory and Sean O’Malley.

The design and implementation of hierarchical software systems with reusable components. ACM Transactions on Software Engineering and Methodology, 1(4):355–398, October 1992. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1992-1-4/p355-
batory/p355-batory.pdf;
http://www.acm.org/pubs/citations/
journals/tosem/1992-1-4/p355-
batory/.

Bohme:2018:SST

[Böh18] Marcel Böhme. STADS: Software testing as species discovery. ACM Transactions on Software Engineering and Methodology, 27(2):7:1–7:??, July 2018. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Baresi:1998:TFS

[BP98] Luciano Baresi and Mauro Pezzè. Toward formalizing structured analysis. ACM Transactions on Software Engineering and Methodology, 7(1):80–107, January 1998. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

                                        www.acm.org/pubs/articles/
                                        journals/tosem/1998-7-1/p80-
                                        baresi/p80-baresi.pdf;   http:
                                        //www.acm.org/pubs/citations/
                                        journals/tosem/1998-7-1/p80-
                                        baresi/.

Baresi:2005:FID

[BP05] Luciano Baresi and Mauro Pezzè. Formal interpreters for diagram notations. ACM Transactions on Software Engineering and Methodology, 14(1):42– 84, January 2005. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Barcelona-Pons:2022:SSC

[BPSSA^{+}22] Daniel Barcelona-Pons, Pierre Sutra, Marc Sánchez-Artigas, Gerard París, and Pedro García-López. Stateful serverless computing with crucial. ACM Transactions on Software Engineering and Methodology, 31 (3):39:1–39:38, July 2022. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3490386.

Brogi:2010:DIS

[BPT10] Antonio Brogi, Razvan Popescu, and Matteo Tanca. Design and implementation of Sator: a Web service aggregator. ACM Transactions on Software Engineering and Methodology, 19(3): 10:1–10:??, January 2010. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Brett:1995:CCS

[Bre95] Bevin R. Brett. Comments on “The cost of selective recompilation and environment processing”. ACM Transactions on Software Engineering and Methodology, 4(2):214–216, April 1995. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1995-4-2/p214-
brett/p214-brett.pdf;   http:
//www.acm.org/pubs/citations/
journals/tosem/1995-4-2/p214-

brett/. See [ATW94, TAW95].

Beauvais:2001:MSA

[BRG^{+}01] J.-R. Beauvais, E. Rutten, T. Gautier, R. Houdebine, P. Le Guernic, and Y.-M. Tang. Modeling statecharts and activitycharts as signal equations. ACM Transactions on Software Engineering and Methodology, 10(4): 397–451, October 2001. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Broy:1993:FST

[Bro93] Manfred Broy. Functional specification of time-sensitive communicating systems. ACM Transactions on Software Engineering and Methodology, 2(1): 1–46, January 1993. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

                                        URL http://www.acm.org/
                                        pubs/articles/journals/tosem/
                                        1993-2-1/p1-broy/p1-broy.
                                        pdf;   http://www.acm.org/
                                        pubs/citations/journals/tosem/
                                        1993-2-1/p1-broy/.

Bible:2001:CSC

[BRR01] John Bible, Gregg Rothermel, and David S. Rosenblum. A comparative study of coarseand fine-grained safe regression test-selection techniques. ACM Transactions on Software Engineering and Methodology, 10(2):149–183, April 2001. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

                                        www.acm.org/pubs/articles/
                                        journals/tosem/2001-10-2/
                                        p149-bible/p149-bible.pdf;
                                         http://www.acm.org/pubs/
                                        citations/journals/tosem/
                                        2001-10-2/p149-bible/.

Berstel:2005:SFM

[BRRP05] Jean Berstel, Stefano Crespi Reghizzi, Gilles Roussel, and Pierluigi San Pietro. A scalable formal method for design and automatic checking of user interfaces. ACM Transactions on Software Engineering and Methodology, 14(2):124– 167, April 2005. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Basu:2007:MCJ

[BS07] Samik Basu and Scott A.

Smolka. Model checking the Java metalocking algorithm. ACM Transactions on Software Engineering and Methodology, 16(3):12:1–12:??, July 2007. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Baki:2016:MSL

[BS16] Islem Baki and Houari Sahraoui.

Multi-step learning and adaptive search for learning complex model transformations from examples. ACM Transactions on Software Engineering and Methodology, 25(3):20:1– 20:??, August 2016. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Bock:2022:MMG

[BSA22] Thomas Bock, Angelika Schmid, and Sven Apel. Measuring and modeling group dynamics in open-source software development: a tensor decomposition approach. ACM Transactions on Software Engineering and Methodology, 31(2): 19:1–19:50, April 2022. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3473139.

Bessghaier:2024:WCD

[BSOM24] Narjes Bessghaier, Mohammed Sayagh, Ali Ouni, and Mohamed Wiem Mkaouer. What constitutes the deployment and runtime configuration system? An empirical study on Open-Stack projects. ACM Transactions on Software Engineering and Methodology, 33(1): 5:1–5:??, January 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3607186.

Biagiola:2022:TPR

[BT22] Matteo Biagiola and Paolo Tonella. Testing the plasticity of reinforcement learningbased systems. ACM Transactions on Software Engineering and Methodology, 31(4): 80:1–80:??, October 2022. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3511701.

Biagiola:2024:TDR

[BT24] Matteo Biagiola and Paolo Tonella. Testing of deep reinforcement learning agents with surrogate models. ACM Transactions on Software Engineering and Methodology, 33(3): 73:1–73:??, March 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3631970.

Bodden:2014:JPI

[BTI14] Eric Bodden, Éric Tanter, and Milton Inostroza. Join point interfaces for safe and flexible decoupling of aspects. ACM Transactions on Software Engineering and Methodology, 23(1): 7:1–7:??, February 2014. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Badampudi:2023:MCR

[BUB23] Deepika Badampudi, Michael Unterkalmsteiner, and Ricardo Britto. Modern code reviews — survey of literature and practice. ACM Transactions on Software Engineering and Methodology, 32(4): 107:1–107:??, July 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3585004.

Bi:2022:ASP

[BXL^{+}22] Tingting Bi, Xin Xia, David Lo, John Grundy, Thomas Zimmermann, and Denae Ford. Accessibility in software practice: a practitioner’s perspective. ACM Transactions on Software Engineering and Methodology, 31(4): 66:1–66:??, October 2022. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3503508.

Bao:2020:PDC

[BXX^{+}20] Lingfeng Bao, Zhenchang Xing, Xin Xia, David Lo, Minghui Wu, and Xiaohu Yang. psc2code: Denoising code extraction from programming screencasts. ACM Transactions on Software Engineering and Methodology, 29(3): 21:1–21:38, July 2020. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3392093.

Batoun:2023:ESG

[BYTS23] Mohamed Amine Batoun, Ka Lai Yung, Yuan Tian, and Mohammed Sayagh. An empirical study on GitHub pull requests’ reactions. ACM Transactions on Software Engineering and Methodology, 32(6):146:1– 146:??, November 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3597208.

Banach:2014:CAM

[BZSW14] Richard Banach, Huibiao Zhu, Wen Su, and Xiaofeng Wu. A continuous ASM modelling approach to pacemaker sensing. ACM Transactions on Software Engineering and Methodology, 24(1):2:1–2:??, September 2014. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Cobleigh:2008:BHD

[CAC08] Jamieson M. Cobleigh, George S.

Avrunin, and Lori A. Clarke. Breaking up is hard to do: an evaluation of automated assume-guarantee reasoning. ACM Transactions on Software Engineering and Methodology, 17(2):7:1–7:??, April 2008. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Cai:2020:AIM

[Cai20] Haipeng Cai. Assessing and improving malware detection sustainability through app evolution studies. ACM Transactions on Software Engineering and Methodology, 29(2):8:1– 8:28, April 2020. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3371924.

Callison:1995:TSO

[Cal95] H. Rebecca Callison. A timesensitive object model for realtime systems. ACM Transactions on Software Engineering and Methodology, 4(3):287–317, July 1995. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1995-4-3/p287-
callison/p287-callison.pdf;
http://www.acm.org/pubs/

citations/journals/tosem/
1995-4-3/p287-callison/.

Candela:2016:UCC

[CBRO16] Ivan Candela, Gabriele Bavota, Barbara Russo, and Rocco Oliveto. Using cohesion and coupling for software remodularization: Is it enough? ACM Transactions on Software Engineering and Methodology, 25(3): 24:1–24:??, August 2016. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Chen:2021:HSI

[CCH^{+}21] Qiuyuan Chen, Chunyang Chen, Safwat Hassan, Zhengchang Xing, Xin Xia, and Ahmed E. Hassan. How should I improve the UI of my app?: a study of user reviews of popular apps in the Google Play. ACM Transactions on Software Engineering and Methodology, 30 (3):37:1–37:38, May 2021. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3447808.

Camilli:2023:ADD

[CCRZ23] Matteo Camilli, Carmine Colarusso, Barbara Russo, and Eugenio Zimeo. Actor-driven decomposition of microservices through multi-level scalability assessment. ACM Transactions on Software Engineering and Methodology, 32(5):117:1– 117:??, September 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3583563.

Cederbladh:2024:EVV

[CCS24] Johan Cederbladh, Antonio Cicchetti, and Jagadish Suryadevara. Early validation and verification of system behaviour in model-based systems engineering: a systematic literature review. ACM Transactions on Software Engineering and Methodology, 33(3): 81:1–81:??, March 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3631976.

Cruz-Carlon:2023:PLB

[CCVLW23] Juan Cruz-Carlon, Mahsa Varshosaz, Claire Le Goues, and Andrzej Wasowski. Patching locking bugs statically with crayons. ACM Transactions on Software Engineering and Methodology, 32(3):56:1–56:??, July 2023. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3548684.

Chen:2023:EBB

[CCW^{+}23] Zhichao Chen, Junjie Chen, Weijing Wang, Jianyi Zhou, Meng Wang, Xiang Chen, Shan Zhou, and Jianmin Wang. Exploring better black-Box test case prioritization via log analysis. ACM Transactions on Software Engineering and Methodology, 32(3):72:1–72:??, July 2023. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3569932.

Chen:2011:RFC

[CCX11] Zhenyu Chen, Tsong Yueh Chen, and Baowen Xu. A revisit of fault class hierarchies in general Boolean specifications. ACM Transactions on Software Engineering and Methodology, 20(3):13:1–13:??, August 2011. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Chen:2020:WBU

[CCX^{+}20] Jieshan Chen, Chunyang Chen, Zhenchang Xing, Xin Xia, Liming Zhu, John Grundy, and Jinshui Wang. Wireframe-based UI design search through image autoencoder. ACM Transactions on Software Engineering and Methodology, 29(3):19:1– 19:31, July 2020. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3391613.

Cui:2023:SDM

[CCXZ23] Mohan Cui, Chengjun Chen, Hui Xu, and Yangfan Zhou. SafeDrop: Detecting memory deallocation bugs of Rust programs via static data-flow analysis. ACM Transactions on Software Engineering and Methodology, 32(4):82:1–82:??, July 2023. CODEN ATSMER.

ISSN 1049-331X (print), 1557-7392 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3542948.

Chen:2021:EPS

[CCY^{+}21] Zhenpeng Chen, Yanbin Cao, Huihan Yao, Xuan Lu, Xin Peng, Hong Mei, and Xuanzhe Liu. Emoji-powered sentiment and emotion detection from software developers’ communication data. ACM Transactions on Software Engineering and Methodology, 30(2): 18:1–18:48, March 2021. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3424308.

Crow:1998:FSS

[CD98] Judith Crow and Ben Di Vito.

Formalizing space shuttle software requirements: four case studies. ACM Transactions on Software Engineering and Methodology, 7(3):296–332, July 1998. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1998-7-3/p296-
crow/p296-crow.pdf;   http:
//www.acm.org/pubs/citations/
journals/tosem/1998-7-3/p296-
crow/.

Chechik:2003:MVS

[CDEG03] Marsha Chechik, Benet Devereux, Steve Easterbrook, and Arie Gurfinkel. Multi-valued symbolic model-checking. ACM Transactions on Software Engineering and Methodology, 12(4): 371–408, October 2003. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Cugola:1996:FFI

[CDFG96] Gianpaolo Cugola, Elisabetta Di Nitto, Alfonso Fuggetta, and Carlo Ghezzi. A framework for formalizing inconsistencies and deviations in human-centered systems. ACM Transactions on Software Engineering and Methodology, 5(3):191–230, July 1996. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

                                        www.acm.org/pubs/articles/
                                        journals/tosem/1996-5-3/p191-
                                        cugola/p191-cugola.pdf;
                                        http://www.acm.org/pubs/citations/
                                        journals/tosem/1996-5-3/p191-
                                        cugola/.

Chatterjee:2021:AIQ

[CDKP21] Preetha Chatterjee, Kostadin Damevski, Nicholas A. Kraft, and Lori Pollock. Automatically identifying the quality of developer chats for post hoc use. ACM Transactions on Software Engineering and Methodology, 30(4):48:1–48:28, July 2021. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3450503.

Costagliola:2004:FMI

[CDP04] Gennaro Costagliola, Vincenzo Deufemia, and Giuseppe Polese.

A framework for modeling and implementing visual notations with applications to software engineering. ACM Transactions on Software Engineering and Methodology, 13(4):431– 487, October 2004. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Chen:2010:VSI

[CDSM10] Chunqing Chen, Jin Song Dong, Jun Sun, and Andrew Martin. A verification system for interval-based specification languages. ACM Transactions on Software Engineering and Methodology, 19(4):13:1–13:??, April 2010. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Cohen:2003:AHQ

[CF03] Yossi Cohen and Yishai A. Feldman. Automatic high-quality reengineering of database programs by abstraction, transformation and reimplementation. ACM Transactions on Software Engineering and Methodology, 12(3):285–316, July 2003. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Conboy:2010:MDC

[CF10] Kieran Conboy and Brian Fitzgerald. Method and developer characteristics for effective agile method tailoring: a study of XP expert opinion. ACM Transactions on Software Engineering and Methodology, 20 (1):2:1–2:??, June 2010. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Czekster:2016:SPA

[CFL^{+}16] Ricardo M. Czekster, Paulo Fernandes, Lucelene Lopes, Afonso Sales, Alan R. Santos, and Thais Webber. Stochastic performance analysis of global software development teams. ACM Transactions on Software Engineering and Methodology, 25(3): 26:1–26:??, August 2016. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Ciancarini:2000:UCL

[CFM00] P. Ciancarini, F. Franzé, and C. Mascolo. Using a coordination language to specify and analyze systems containing mobile components. ACM Transactions on Software Engineering and Methodology, 9(2):167–198, April 2000. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

                                        www.acm.org/pubs/articles/
                                        journals/tosem/2000-9-2/p167-
                                        ciancarini/p167-ciancarini.
                                        pdf;   http://www.acm.org/
                                        pubs/citations/journals/tosem/
                                        2000-9-2/p167-ciancarini/.

Clark:2024:TCS

[CFP^{+}24] Andrew G. Clark, Michael Foster, Benedikt Prifling, Neil Walkinshaw, Robert M. Hierons, Volker Schmidt, and Robert D. Turner. Testing causality in scientific modelling software. ACM Transactions on Software Engineering and Methodology, 33(1): 10:1–10:??, January 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3607184.

Carzaniga:2015:AWE

[CGPP15] Antonio Carzaniga, Alessandra Gorla, Nicolò Perino, and Mauro Pezzè. Automatic workarounds: Exploiting the intrinsic redundancy of Web applications. ACM Transactions on Software Engineering and Methodology, 24(3):16:1–16:??, May 2015. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Cheng:2023:HHI

[CGZ23] Jiezhu Cheng, Cuiyun Gao, and Zibin Zheng. HINNPerf: Hierarchical interaction neural network for performance prediction of configurable systems. ACM Transactions on Software Engineering and Methodology, 32 (2):46:1–46:??, April 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3528100.

Chen:2021:LCF

[CH21] Jianhui Chen and Fei He. Leveraging control flow knowledge in SMT solving of program verification. ACM Transactions on Software Engineering and Methodology, 30(4): 41:1–41:26, July 2021. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3446211.

Ciancarini:1993:CRB

[Cia93] Paolo Ciancarini. Coordinating rule-based software processes with ESP. ACM Transactions on Software Engineering and Methodology, 2(3):203–227, July 1993. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

                                        www.acm.org/pubs/articles/
                                        journals/tosem/1993-2-3/p203-
                                        ciancarini/p203-ciancarini.
                                        pdf;   http://www.acm.org/
                                        pubs/citations/journals/tosem/
                                        1993-2-3/p203-ciancarini/.

Clarke:2000:VSP

[CJM00] E. M. Clarke, S. Jha, and W. Marrero. Verifying security protocols with Brutus. ACM Transactions on Software Engineering and Methodology, 9(4):443–487, October 2000. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

                                        www.acm.org/pubs/articles/
                                        journals/tosem/2000-9-4/p443-
                                        clarke/p443-clarke.pdf;
                                        http://www.acm.org/pubs/citations/
                                        journals/tosem/2000-9-4/p443-
                                        clarke/.

Cheung:1996:CCC

[CK96] Shing Chi Cheung and Jeff Kramer. Context constraints for compositional reachability analysis. ACM Transactions on Software Engineering and Methodology, 5(4):334–377, October 1996. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1996-5-4/p334-
cheung/p334-cheung.pdf;
http://www.acm.org/pubs/citations/
journals/tosem/1996-5-4/p334-
cheung/.

Cheung:1999:CSP

[CK99] Shing Chi Cheung and Jeff Kramer. Checking safety properties using compositional reachability analysis. ACM Transactions on Software Engineering and Methodology, 8(1):49–78, January 1999. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1999-8-1/p49-
cheung/p49-cheung.pdf;   http:
//www.acm.org/pubs/citations/
journals/tosem/1999-8-1/p49-
cheung/.

Cheon:1994:LSI

[CL94] Yoonsik Cheon and Gary T.

Leavens. The Larch/Smalltalk interface specification language. ACM Transactions on Software Engineering and Methodology, 3(3):221–253, July 1994. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/

                                        journals/tosem/1994-3-3/p221-
                                        cheon/p221-cheon.pdf;   http:
                                        //www.acm.org/pubs/citations/
                                        journals/tosem/1994-3-3/p221-
                                        cheon/.

Calefato:2022:UPD

[CL22] Fabio Calefato and Filippo Lanubile. Using personality detection tools for software engineering research: How far can we go? ACM Transactions on Software Engineering and Methodology, 31(3): 42:1–42:48, July 2022. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3491039.

Chen:2023:DPA

[CL23a] Tao Chen and Miqing Li. Do performance aspirations matter for guiding software configuration tuning? An empirical investigation under dual performance objectives. ACM Transactions on Software Engineering and Methodology, 32 (3):68:1–68:??, July 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3571853.

Chen:2023:WCH

[CL23b] Tao Chen and Miqing Li. The weights can be harmful: Pareto search versus weighted search in multi-objective search-based software engineering. ACM Transactions on Software Engineering and Methodology, 32(1):

5:1–5:??, January 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3514233.

Chen:2018:FFG

[CLBY18] Tao Chen, Ke Li, Rami Bahsoon, and Xin Yao. FEMOSAA: Feature-guided and knee-driven multi-objective optimization for self-adaptive software. ACM Transactions on Software Engineering and Methodology, 27 (2):5:1–5:??, July 2018. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Cao:2022:SSB

[CLL^{+}22] Jialun Cao, Meiziniu Li, Yeting Li, Ming Wen, Shing-Chi Cheung, and Haiming Chen. SemMT: a semantic-based testing approach for machine translation systems. ACM Transactions on Software Engineering and Methodology, 31(2): 34e:1–34e:36, April 2022. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3490488.

Canizares:2024:MCH

[CLMPS^{+}24] Pablo C. Cañizares, Jose María López-Morales, Sara Pérez-Soler, Esther Guerra, and Juan de Lara. Measuring and clustering heterogeneous chatbot designs. ACM Transactions on Software Engineering and Methodology, 33(4):90:1–90:??, May 2024. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3637228.

Chen:2023:TUD

[CLS^{+}23] Junjie Chen, Yihua Liang, Qingchao Shen, Jiajun Jiang, and Shuochuan Li. Toward understanding deep learning framework bugs. ACM Transactions on Software Engineering and Methodology, 32(6):135:1– 135:??, November 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3587155.

Chen:2008:UBS

[CM08] Tsong Yueh Chen and Robert Merkel. An upper bound on software testing effectiveness. ACM Transactions on Software Engineering and Methodology, 17(3):16:1–16:??, June 2008. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

C:2023:AEB

[CM23] Shrikanth N. C. and Tim Menzies. Assessing the early bird heuristic (for predicting project quality). ACM Transactions on Software Engineering and Methodology, 32(5):116:1– 116:??, September 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3583565.

Ciapessoni:1999:FMF

[CMCP^{+}99] Emanuele Ciapessoni, Piergiorgio Mirandola, Alberto Coen-Porisini, Dino Mandrioli, and Angelo Morzenti. From formal models to formally based methods: an industrial experience. ACM Transactions on Software Engineering and Methodology, 8(1):79–113, January 1999. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1999-8-1/p79-
ciapessoni/p79-ciapessoni.
pdf;   http://www.acm.org/
pubs/citations/journals/tosem/
1999-8-1/p79-ciapessoni/.

Ceccato:2015:DAG

[CMM^{+}15] Mariano Ceccato, Alessandro Marchetto, Leonardo Mariani, Cu D. Nguyen, and Paolo Tonella. Do automatically generated test cases make debugging easier? An experimental assessment of debugging effectiveness and efficiency. ACM Transactions on Software Engineering and Methodology, 25(1): 5:1–5:??, December 2015. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Chang:2013:EHH

[CMP13] Hervé Chang, Leonardo Mariani, and Mauro Pezzè. Exception handlers for healing component-based systems. ACM Transactions on Software Engineering and Methodology, 22(4):30:1–30:??, October 2013. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Corbett:2000:USA

[Cor00] James C. Corbett. Using shape analysis to reduce finite-state models of concurrent Java programs. ACM Transactions on Software Engineering and Methodology, 9(1):51–93, January 2000. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

                                        www.acm.org/pubs/articles/
                                        journals/tosem/2000-9-1/p51-
                                        corbett/p51-corbett.pdf;
                                        http://www.acm.org/pubs/citations/
                                        journals/tosem/2000-9-1/p51-
                                        corbett/.

Chekam:2021:KSM

[CPCT21] Thierry Titcheu Chekam, Mike Papadakis, Maxime Cordy, and Yves Le Traon. Killing stubborn mutants with symbolic execution. ACM Transactions on Software Engineering and Methodology, 30(2): 19:1–19:23, March 2021. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3425497.

Coen-Porisini:2003:FAD

[CPPRM03] Alberto Coen-Porisini, Matteo Pradella, Matteo Rossi, and Dino Mandrioli. A formal approach for designing CORBAbased applications. ACM Transactions on Software Engineering and Methodology, 12(2):107– 151, April 2003. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Creveuil:1994:FSD

[CR94] Christian Creveuil and Gruia-Catalin Roman. Formal specification and design of a message router. ACM Transactions on Software Engineering and Methodology, 3(4):271–307, October 1994. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1994-3-4/p271-
creveuil/p271-creveuil.pdf;
http://www.acm.org/pubs/
citations/journals/tosem/
1994-3-4/p271-creveuil/.

Cimatti:2012:VRH

[CRST12] Alessandro Cimatti, Marco Roveri, Angelo Susi, and Stefano Tonetta. Validation of requirements for hybrid systems: a formal approach. ACM Transactions on Software Engineering and Methodology, 21(4):22:1– 22:??, November 2012. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Cai:2012:FMA

[CS12] Yuanfang Cai and Kevin Sullivan. A formal model for automated software modularity and evolvability analysis. ACM Transactions on Software Engineering and Methodology, 21 (4):21:1–21:??, November 2012.

CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Chen:2022:BCT

[CS22] Junjie Chen and Chenyao Suo. Boosting compiler testing via compiler optimization exploration. ACM Transactions on Software Engineering and Methodology, 31(4):72:1– 72:??, October 2022. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3508362.

Cao:2024:LDM

[CSB^{+}24] Sicong Cao, Xiaobing Sun, Lili Bo, Rongxin Wu, Bin Li, Xiaoxue Wu, Chuanqi Tao, Tao Zhang, and Wei Liu. Learning to detect memory-related vulnerabilities. ACM Transactions on Software Engineering and Methodology, 33(2):43:1– 43:??, February 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3624744.

Counsell:2006:IUT

[CSC06] Steve Counsell, Stephen Swift, and Jason Crampton. The interpretation and utility of three cohesion metrics for objectoriented design. ACM Transactions on Software Engineering and Methodology, 15(2):123– 149, April 2006. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Clun:2024:RAM

[CSFB24] Donato Clun, Donghwan Shin, Antonio Filieri, and Domenico Bianculli. Rigorous assessment of model inference accuracy using language cardinality. ACM Transactions on Software Engineering and Methodology, 33 (4):95:1–95:??, May 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3640332.

Cai:2016:DUD

[CST16] Haipeng Cai, Raul Santelices, and Douglas Thain. DiaPro: Unifying dynamic impact analyses for improved and variable cost-effectiveness. ACM Transactions on Software Engineering and Methodology, 25 (2):18:1–18:??, May 2016. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Chaki:2013:VAI

[CSV13] Sagar Chaki, Christian Schallhart, and Helmut Veith. Verification across intellectual property boundaries. ACM Transactions on Software Engineering and Methodology, 22(2):15:1– 15:??, March 2013. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Chondamrongkul:2021:SAM

[CSW21] Nacha Chondamrongkul, Jing Sun, and Ian Warren. Software architectural migration: an automated planning approach.

ACM Transactions on Software Engineering and Methodology, 30(4):50:1–50:35, July 2021. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3461011.

Csallner:2008:DCH

[CSX08] Christoph Csallner, Yannis Smaragdakis, and Tao Xie. DSD-Crasher: a hybrid analysis tool for bug finding. ACM Transactions on Software Engineering and Methodology, 17 (2):8:1–8:??, April 2008. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Castro:2021:FFS

[CTA^{+}21] Thiago Castro, Leopoldo Teixeira, Vander Alves, Sven Apel, Maxime Cordy, and Rohit Gheyi. A formal framework of software product line analyses. ACM Transactions on Software Engineering and Methodology, 30(3):34:1–34:37, May 2021. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3442389.

Chen:2001:TMO

[CTC01] Huo Yan Chen, T. H. Tse, and T. Y. Chen. TAC-CLE: a methodology for objectoriented software testing at the class and cluster levels. ACM Transactions on Software Engineering and Methodology, 10(1):56–109, January 2001. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/2001-10-1/
p56-chen/p56-chen.pdf;   http:
//www.acm.org/pubs/citations/
journals/tosem/2001-10-1/
p56-chen/.

Costea:2023:HDR

[CTC^{+}23] Andreea Costea, Abhishek Tiwari, Sigmund Chianasta, Kishore R., Abhik Roychoudhury, and Ilya Sergey. Hippodrome: Data race repair using static analysis summaries. ACM Transactions on Software Engineering and Methodology, 32 (2):41:1–41:??, April 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3546942.

Chen:1998:BWI

[CTCC98] Huo Yan Chen, T. H. Tse, F. T. Chan, and T. Y. Chen. In black and white: an integrated approach to class-level testing of object-oriented programs. ACM Transactions on Software Engineering and Methodology, 7(3):250–295, July 1998. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1998-7-3/p250-
chen/p250-chen.pdf;   http:
//www.acm.org/pubs/citations/
journals/tosem/1998-7-3/p250-
chen/.

Chowdhury:2024:MLB

[CUHH24] Shaiful Chowdhury, Gias Uddin, Hadi Hemmati, and Reid Holmes. Method-level bug prediction: Problems and promises. ACM Transactions on Software Engineering and Methodology, 33(4):98:1–98:??, May 2024. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3640331.

Cook:1998:DMS

[CW98] Jonathan E. Cook and Alexander L. Wolf. Discovering models of software processes from event-based data. ACM Transactions on Software Engineering and Methodology, 7(3):215–249, July 1998. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

                                        www.acm.org/pubs/articles/
                                        journals/tosem/1998-7-3/p215-
                                        cook/p215-cook.pdf;   http:
                                        //www.acm.org/pubs/citations/
                                        journals/tosem/1998-7-3/p215-
                                        cook/.

Cook:1999:SPV

[CW99] Jonathan E. Cook and Alexander L. Wolf. Software process validation: quantitatively measuring the correspondence of a process to a model. ACM Transactions on Software Engineering and Methodology, 8(2):147–176, April 1999. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

                                        www.acm.org/pubs/articles/
                                        journals/tosem/1999-8-2/p147-

cook/p147-cook.pdf;   http:
//www.acm.org/pubs/citations/
journals/tosem/1999-8-2/p147-
cook/.

Cheng:2021:DSD

[CWH^{+}21] Xiao Cheng, Haoyu Wang, Jiayi Hua, Guoai Xu, and Yulei Sui. DeepWukong: Statically detecting software vulnerabilities using deep graph neural network. ACM Transactions on Software Engineering and Methodology, 30(3): 38:1–38:33, May 2021. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3436877.

Chen:2020:HCT

[CWM^{+}20] Lin Chen, Di Wu, Wanwangying Ma, Yuming Zhou, Baowen Xu, and Hareton Leung. How C++ templates are used for generic programming: an empirical study on 50 open source systems. ACM Transactions on Software Engineering and Methodology, 29(1):3:1– 3:49, February 2020. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3356579.

Chen:2023:QQO

[CWM^{+}23] Jialuo Chen, Jingyi Wang, Xingjun Ma, Youcheng Sun, Jun Sun, Peixin Zhang, and Peng Cheng. QuoTe: Qualityoriented testing for deep learning systems. ACM Transactions on Software Engineering and Methodology, 32(5):125:1– 125:??, September 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3582573.

Chen:2020:PAE

[CWW^{+}20] Junjie Chen, Zhuo Wu, Zan Wang, Hanmo You, Lingming Zhang, and Ming Yan. Practical accuracy estimation for efficient deep neural network testing. ACM Transactions on Software Engineering and Methodology, 29(4):30:1–30:35, October 2020. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3394112.

Chen:2021:WMC

[CXH^{+}21] Qiuyuan Chen, Xin Xia, Han Hu, David Lo, and Shanping Li. Why my code summarization model does not work: Code comment improvement with category prediction. ACM Transactions on Software Engineering and Methodology, 30(2): 25:1–25:29, March 2021. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3434280.

Cogo:2023:AAB

[CXH23] Filipe Roseiro Cogo, Xin Xia, and Ahmed E. Hassan. Assessing the alignment between the information needs of developers and the documentation of programming languages: a case study on Rust. ACM Transactions on Software Engineering and Methodology, 32 (2):43:1–43:??, April 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3546945.

Chen:2022:WDS

[CXLG22] Jiachi Chen, Xin Xia, David Lo, and John Grundy. Why do smart contracts self-destruct? Investigating the selfdestruct function on Ethereum. ACM Transactions on Software Engineering and Methodology, 31(2): 30:1–30:37, April 2022. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3488245.

Chen:2011:TDB

[CY11] Jinjun Chen and Yun Yang.

Temporal dependency-based checkpoint selection for dynamic verification of temporal constraints in scientific workflow systems. ACM Transactions on Software Engineering and Methodology, 20(3):9:1– 9:??, August 2011. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Catak:2022:UAP

[CYA22] Ferhat Ozgur Catak, Tao Yue, and Shaukat Ali. Uncertaintyaware prediction validator in deep learning models for cyberphysical system data. ACM Transactions on Software Engineering and Methodology, 31(4): 79:1–79:??, October 2022. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3527451.

Chen:2024:SSB

[CYM^{+}24] Zhe Chen, Rui Yan, Yingzi Ma, Yulei Sui, and Jingling Xue. A smart status based monitoring algorithm for the dynamic analysis of memory safety. ACM Transactions on Software Engineering and Methodology, 33 (4):89:1–89:??, May 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3637227.

Czepa:2019:HUP

[CZ19] Christoph Czepa and Uwe Zdun. How understandable are pattern-based behavioral constraints for novice software designers? ACM Transactions on Software Engineering and Methodology, 28(2): 11:1–11:??, April 2019. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.
                                        acm.org/ft_gateway.cfm?id=
                                        3306608.

Chen:2023:CES

[CZSH23] Zhenpeng Chen, Jie M. Zhang, Federica Sarro, and Mark Harman. A comprehensive empirical study of bias mitigation methods for machine learning classifiers. ACM Transactions on Software Engineering and Methodology, 32(4): 106:1–106:??, July 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3583561.

Davis:2023:WWP

[DAL^{+}23] Matthew C. Davis, Emad Aghayi, Thomas D. Latoza, Xiaoyin Wang, Brad A. Myers, and Joshua Sunshine. What’s (not) working in programmer user studies? ACM Transactions on Software Engineering and Methodology, 32(5):120:1– 120:??, September 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3587157.

Duri:1994:AEE

[DBDS94] S. Duri, U. Buy, R. Devarapalli, and S. M. Shatz. Application and experimental evaluation of state space reduction methods for deadlock analysis in Ada. ACM Transactions on Software Engineering and Methodology, 3(4):340–380, October 1994. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1994-3-4/p340-
duri/p340-duri.pdf;   http:
//www.acm.org/pubs/citations/
journals/tosem/1994-3-4/p340-
duri/.

DeCaso:2013:EBP

[DBGU13] Guido De Caso, Victor Braberman, Diego Garbervetsky, and Sebastian Uchitel. Enablednessbased program abstractions for behavior validation. ACM Transactions on Software Engineering and Methodology, 22 (3):25:1–25:??, July 2013. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Alesio:2015:CGA

[DBNG15] Stefano Di Alesio, Lionel C. Briand, Shiva Nejati, and Arnaud Gotlieb. Combining genetic algorithms and constraint programming to support stress testing of task deadlines. ACM Transactions on Software Engineering and Methodology, 25(1): 4:1–4:??, December 2015. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Dippolito:2013:SNE

[DBPU13] Nicolás D’ippolito, Victor Braberman, Nir Piterman, and Sebastián Uchitel. Synthesizing nonanomalous event-based controllers for liveness goals. ACM Transactions on Software Engineering and Methodology, 22(1): 9:1–9:??, February 2013. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Dwyer:2004:FAV

[DCCN04] Matthew B. Dwyer, Lori A. Clarke, Jamieson M. Cobleigh, and Gleb Naumovich. Flow analysis for verifying properties of concurrent software systems. ACM Transactions on Software Engineering and Methodology, 13(4):359–430, October 2004. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Desai:2009:AMM

[DCS09] Nirmit Desai, Amit K. Chopra, and Munindar P. Singh. Amoeba: a methodology for modeling and evolving cross-organizational business processes. ACM Transactions on Software Engineering and Methodology, 19(2):6:1– 6:??, October 2009. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Diep:2011:LBS

[DDE11] Madeline M. Diep, Matthew B.

Dwyer, and Sebastian Elbaum. Lattice-based sampling for path property monitoring. ACM Transactions on Software Engineering and Methodology, 21(1): 5:1–5:??, December 2011. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Lucia:2018:DBD

[DDGR18] Andrea De Lucia, Vincenzo Deufemia, Carmine Gravino, and Michele Risi. Detecting the behavior of design patterns through model checking and dynamic analysis. ACM Transactions on Software Engineering and Methodology, 26(4):

13:1–13:??, February 2018. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

DiStefano:2022:VDS

[DDI22] Luca Di Stefano, Rocco De Nicola, and Omar Inverso. Verification of distributed systems via sequential emulation. ACM Transactions on Software Engineering and Methodology, 31 (3):37:1–37:41, July 2022. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3490387.

DeGiacomo:2022:MCM

[DDMM22] Giuseppe De Giacomo, Riccardo De Masellis, Fabrizio Maria Maggi, and Marco Montali. Monitoring constraints and metaconstraints with temporal logics on finite traces. ACM Transactions on Software Engineering and Methodology, 31(4): 68:1–68:??, October 2022. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3506799.

Dola:2023:IDC

[DDS23] Swaroopa Dola, Matthew B. Dwyer, and Mary Lou Soffa. Input distribution coverage: Measuring feature interaction adequacy in neural network testing. ACM Transactions on Software Engineering and Methodology, 32(3):81:1–81:??, July 2023. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3576040.

Duala-Ekoko:2010:CRD

[DER10] Ekwa Duala-Ekoko and Martin P. Robillard. Clone region descriptors: Representing and tracking duplication in source code. ACM Transactions on Software Engineering and Methodology, 20(1):3:1– 3:??, June 2010. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Devanbu:1999:GCF

[Dev99] Premkumar T. Devanbu. GENOA — a customizable, front-end-retargetable source code analysis framework. ACM Transactions on Software Engineering and Methodology, 8(2):177–212, April 1999. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1999-8-2/p177-
devanbu/p177-devanbu.pdf;
http://www.acm.org/pubs/
citations/journals/tosem/
1999-8-2/p177-devanbu/.

Doong:1994:AAT

[DF94] Roong-Ko Doong and Phyllis G. Frankl. The ASTOOT approach to testing objectoriented programs. ACM Transactions on Software Engineering and Methodology, 3(2):101–130, April 1994. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1994-3-2/p101-
doong/p101-doong.pdf;   http:
//www.acm.org/pubs/citations/
journals/tosem/1994-3-2/p101-
doong/.

Damiani:1999:CHA

[DFB99a] E. Damiani, M. G. Fugini, and C. Bellettini. Corrigenda: a hierarchy-aware approach to faceted classification of object-oriented components. ACM Transactions on Software Engineering and Methodology, 8(4):425–472, October 1999. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1999-8-4/p425-
damiani/p425-damiani.pdf;
 http://www.acm.org/pubs/
citations/journals/tosem/
1999-8-4/p425-damiani/.

Damiani:1999:HAA

[DFB99b] E. Damiani, M. G. Fugini, and C. Bellettini. A hierarchyaware approach to faceted classification of objected-oriented components. ACM Transactions on Software Engineering and Methodology, 8(3):215–262, July 1999. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1999-8-3/p215-
damiani/p215-damiani.pdf;
 http://www.acm.org/pubs/
citations/journals/tosem/
1999-8-3/p215-damiani/.

Durante:2000:CAC

[DFG00] Antonio Durante, Riccardo Focardi, and Roberto Gorrieri. A compiler for analyzing cryptographic protocols using noninterference. ACM Transactions on Software Engineering and Methodology, 9(4):488–528, October 2000. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/2000-9-4/p488-
durante/p488-durante.pdf;
 http://www.acm.org/pubs/
citations/journals/tosem/
2000-9-4/p488-durante/.

deFrancesco:2014:GTE

[dFLSV14] Nicoletta de Francesco, Giuseppe Lettieri, Antonella Santone, and Gigliola Vaglini. GreASE: a tool for efficient “Nonequivalence” checking. ACM Transactions on Software Engineering and Methodology, 23(3):24:1–24:??, May 2014. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

DeLucia:2007:RTL

[DFOT07] Andrea De Lucia, Fausto Fasano, Rocco Oliveto, and Genoveffa Tortora. Recovering traceability links in software artifact management systems using information retrieval methods. ACM Transactions on Software Engineering and Methodology, 16(4):13:1–13:??, September 2007. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

DeLara:2017:PTM

[DG17] Juan De Lara and Esther Guerra. A posteriori typing for model-driven engineering: Concepts, analysis, and applications. ACM Transactions on Software Engineering and Methodology, 25(4):31:1–31:??, May 2017. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

DeLara:2018:RML

[DG18] Juan De Lara and Esther Guerra. Refactoring multilevel models. ACM Transactions on Software Engineering and Methodology, 27(4):17:1– 17:??, November 2018. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

DeLara:2014:WHU

[DGC14] Juan De Lara, Esther Guerra, and Jesús Sánchez Cuadrado. When and how to use multilevel modelling. ACM Transactions on Software Engineering and Methodology, 24(2):12:1– 12:??, December 2014. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Lara:2019:ARM

[DGD^{+}19] Juan De Lara, Esther Guerra, Davide Di Ruscio, Juri Di Rocco, Jesús S’nchez Cuadrado, Ludovico Iovino, and Alfonso Pierantonio. Automated reuse of model transformations through typing requirements models. ACM Transactions on Software Engineering and Methodology, 28(4): 21:1–21:??, October 2019. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?id=
3340108.

Lara:2021:FOM

[DGK21] Juan De Lara, Esther Guerra, and Jörg Kienzle. Facetoriented modelling. ACM Transactions on Software Engineering and Methodology, 30 (3):27:1–27:59, May 2021. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3428076.

Dutra:2024:FEF

[DGZ24] Rafael Dutra, Rahul Gopinath, and Andreas Zeller. Format-Fuzzer: Effective fuzzing of binary file formats. ACM Transactions on Software Engineering and Methodology, 33(2):53:1– 53:??, February 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3628157.

Doppke:1998:SPM

[DHW98] John C. Doppke, Dennis Heimbigner, and Alexander L. Wolf. Software process modeling and execution within virtual environments. ACM Transactions on Software Engineering and Methodology, 7(1):1–40, January 1998. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1998-7-1/p1-
doppke/p1-doppke.pdf;   http:
//www.acm.org/pubs/citations/
journals/tosem/1998-7-1/p1-
doppke/.

Dillon:1993:VEM

[Dil93] Laura K. Dillon. A visual execution model for Ada tasking. ACM Transactions on Software Engineering and Methodology, 2(4):311–345, October 1993. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1993-2-4/p311-
dillon/p311-dillon.pdf;
http://www.acm.org/pubs/citations/
journals/tosem/1993-2-4/p311-
dillon/.

Dillon:1997:TDT

[Dil97] Laura K. Dillon. Task dependence and termination in Ada. ACM Transactions on Software Engineering and Methodology, 6(1):80–110, January 1997. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1997-6-1/p80-
dillon/p80-dillon.pdf;   http:
//www.acm.org/pubs/citations/
journals/tosem/1997-6-1/p80-
dillon/.

Devanbu:1997:UDL

[DJ97] Premkumar Devanbu and Mark A.

Jones. The use of description logics in KBSE systems. ACM Transactions on Software Engineering and Methodology, 6(2):141–172, April 1997. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1997-6-2/p141-
devalbu/p141-devalbu.pdf;
 http://www.acm.org/pubs/
citations/journals/tosem/
1997-6-2/p141-devanbu/.

Dilhara:2021:USS

[DKD21] Malinda Dilhara, Ameya Ketkar, and Danny Dig. Understanding Software-2.0: a study of machine learning library usage and evolution. ACM Transactions on Software Engineering and Methodology, 30(4): 55:1–55:42, July 2021. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3453478.

Dillon:1994:GIL

[DKM^{+}94] L. K. Dillon, G. Kutty, L. E.

Moser, P. M. Melliar-Smith, and Y. S. Ramakrishna. A graphical interval logic for specifying concurrent systems. ACM Transactions on Software Engineering and Methodology, 3(2):131–165, April 1994. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/

                                        journals/tosem/1994-3-2/p131-
                                        dillon/p131-dillon.pdf;
                                        http://www.acm.org/pubs/citations/
                                        journals/tosem/1994-3-2/p131-
                                        dillon/.

Dehlinger:2011:GPP

[DL11] Josh Dehlinger and Robyn R. Lutz. Gaia-PL: a product line engineering approach for efficiently designing multiagent systems. ACM Transactions on Software Engineering and Methodology, 20(4):17:1–17:??, September 2011. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Demsky:2013:VSF

[DL13] Brian Demsky and Patrick Lam. Views: Synthesizing finegrained concurrency control. ACM Transactions on Software Engineering and Methodology, 22(1):4:1–4:??, February 2013. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Dang:2024:GMB

[DLP^{+}24] Xueqi Dang, Yinghua Li, Mike Papadakis, Jacques Klein, Tegawendé F. Bissyandé, and Yves Le Traon. GraphPrior: Mutation-based test input prioritization for graph neural networks. ACM Transactions on Software Engineering and Methodology, 33(1):22:1– 22:??, January 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3607191.

Dietz:2015:UIO

[DLRA15] Will Dietz, Peng Li, John Regehr, and Vikram Adve. Understanding integer overflow in C/C++. ACM Transactions on Software Engineering and Methodology, 25(1):2:1– 2:29, December 2015. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Ding:2023:TLG

[DLSC23] Zishuo Ding, Heng Li, Weiyi Shang, and Tse-Hsun (Peter) Chen. Towards learning generalizable code embeddings using task-agnostic graph convolutional networks. ACM Transactions on Software Engineering and Methodology, 32(2): 48:1–48:??, April 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3542944.

Dong:2023:BWL

[DLT^{+}23] Yiwen Dong, Zheyang Li, Yongqiang Tian, Chengnian Sun, Michael W. Godfrey, and Meiyappan Nagappan. Bash in the wild: Language usage, code smells, and bugs. ACM Transactions on Software Engineering and Methodology, 32(1): 8:1–8:22, January 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3517193.

Dramko:2023:DDN

[DLY^{+}23] Luke Dramko, Jeremy Lacomis, Pengcheng Yin, Ed Schwartz, Miltiadis Allamanis, Graham Neubig, Bogdan Vasilescu, and Claire Le Goues. DIRE and its data: Neural decompiled variable renamings with respect to software class. ACM Transactions on Software Engineering and Methodology, 32(2): 39:1–39:??, April 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3546946.

Dyer:2015:BUL

[DNRN15] Robert Dyer, Hoan Anh Nguyen, Hridesh Rajan, and Tien N. Nguyen. Boa: Ultralarge-scale software repository and source-code mining. ACM Transactions on Software Engineering and Methodology, 25(1): 7:1–7:??, December 2015. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

DeMillo:1993:ERA

[DO93] Richard A. DeMillo and A. Jefferson Offutt. Experimental results from an automatic test case generator. ACM Transactions on Software Engineering and Methodology, 2(2):109–127, April 1993. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1993-2-2/p109-
demillo/p109-demillo.pdf;

 http://www.acm.org/pubs/
citations/journals/tosem/
1993-2-2/p109-demillo/.

DiNardo:2017:AFD

[DPB17] Daniel Di Nardo, Fabrizio Pastore, and Lionel Briand. Augmenting field data for testing systems subject to incremental requirements changes. ACM Transactions on Software Engineering and Methodology, 26 (1):1:1–1:??, July 2017. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Do:2023:OTM

[DPRO23] Canh Minh Do, Yati Phyo, Adrián Riesco, and Kazuhiro Ogata. Optimization techniques for model checking leadsto properties in a stratified way. ACM Transactions on Software Engineering and Methodology, 32(6):151:1–151:??, November 2023. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3604610.

Denaro:2013:TAA

[DPT13] Giovanni Denaro, Mauro Pezzè, and Davide Tosi. Test-and-adapt: an approach for improving service interchangeability. ACM Transactions on Software Engineering and Methodology, 22(4):28:1–28:??, October 2013. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Dyer:2010:SDA

[DR10] Robert Dyer and Hridesh Rajan. Supporting dynamic aspect-oriented features. ACM Transactions on Software Engineering and Methodology, 20 (2):7:1–7:??, August 2010. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Dagenais:2011:RAC

[DR11] Barthélémy Dagenais and Martin P. Robillard. Recommending adaptive changes for framework evolution. ACM Transactions on Software Engineering and Methodology, 20(4):19:1– 19:??, September 2011. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Dwyer:2015:EJF

[DR15] Matthew B. Dwyer and David S. Rosenblum. Editorial: Journalfirst publication for the software engineering community. ACM Transactions on Software Engineering and Methodology, 25(1): 1:1–1:??, December 2015. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Devanbu:1996:GTA

[DRW96] Premkumar T. Devanbu, David S. Rosenblum, and Alexander L. Wolf. Generating testing and analysis tools with Aria. ACM Transactions on Software Engineering and Methodology, 5(1):42–62, January 1996.

CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1996-5-1/p42-
devanbu/p42-devanbu.pdf;
http://www.acm.org/pubs/citations/
journals/tosem/1996-5-1/p42-
devanbu/.

Durante:2003:ATE

[DSV03] Luca Durante, Riccardo Sisto, and Adriano Valenzano. Automatic testing equivalence verification of spi calculus specifications. ACM Transactions on Software Engineering and Methodology, 12(2): 222–284, April 2003. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Ding:2024:LPI

[DTC^{+}24] Zishuo Ding, Yiming Tang, Xiaoyu Cheng, Heng Li, and Weiyi Shang. LoGenText-Plus: Improving neural machine translation based logging texts generation with syntactic templates. ACM Transactions on Software Engineering and Methodology, 33(2):38:1– 38:??, February 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3624740.

Dashofy:2005:CAD

[DvdHT05] Eric M. Dashofy, André van der Hoek, and Richard N. Taylor. A comprehensive approach for the development of modular software architecture description languages. ACM Transactions on Software Engineering and Methodology, 14 (2):199–245, April 2005. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

DiSorbo:2023:AIQ

[DZV^{+}23] Andrea Di Sorbo, Fiorella Zampetti, Aaron Visaggio, Massimiliano Di Penta, and Sebastiano Panichella. Automated identification and qualitative characterization of safety concerns reported in UAV software platforms. ACM Transactions on Software Engineering and Methodology, 32(3): 67:1–67:??, July 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3564821.

Emmerich:2008:IRD

[EAS08] Wolfgang Emmerich, Mikio Aoyama, and Joe Sventek. The impact of research on the development of middleware technology. ACM Transactions on Software Engineering and Methodology, 17(4):19:1–19:??, August 2008. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

ElKholy:2014:CCR

[EBE^{+}14] Warda El Kholy, Jamal Bentahar, Mohamed El Menshawy, Hongyang Qu, and Rachida Dssouli. Conditional commitments: Reasoning and model checking. ACM Transactions on Software Engineering and Methodology, 24(2):9:1– 9:??, December 2014. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Erwig:2005:SRS

[EF05] Martin Erwig and Zhe Fu.

Software reuse for scientific computing through program generation. ACM Transactions on Software Engineering and Methodology, 14(2): 168–198, April 2005. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Egyed:2002:AAC

[Egy02] Alexander Egyed. Automated abstraction of class diagrams. ACM Transactions on Software Engineering and Methodology, 11(4):449–491, 2002. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Ehsan:2021:ESD

[EHEZ21] Osama Ehsan, Safwat Hassan, Mariam El Mezouar, and Ying Zou. An empirical study of developer discussions in the Gitter platform. ACM Transactions on Software Engineering and Methodology, 30(1): 8:1–8:39, January 2021. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3412378.

El-Hokayem:2020:MDS

[EHF20] Antoine El-Hokayem and Yliès Falcone. On the monitoring of decentralized specifications: Semantics, properties, analysis, and simulation. ACM Transactions on Software Engineering and Methodology, 29(1):1:1– 1:57, February 2020. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3355181.

Ebnenasir:2011:FSD

[EK11] Ali Ebnenasir and Sandeep S. Kulkarni. Feasibility of stepwise design of multitolerant programs. ACM Transactions on Software Engineering and Methodology, 21(1):1:1– 1:??, December 2011. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Ellis:2019:RFD

[EKL^{+}19] Samuel J. Ellis, Titus H. Klinge, James I. Lathrop, Jack H. Lutz, Robyn R. Lutz, Andrew S. Miner, and Hugh D. Potter. Runtime fault detection in programmed molecular systems. ACM Transactions on Software Engineering and Methodology, 28(2):6:1–6:??, April 2019. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.
                                        acm.org/ft_gateway.cfm?id=
                                        3295740.

Engels:1992:BIS

[ELN^{+}92] G. Engels, C. Lewerentz, M. Nagl, W. Schäfer, and A. Schürr. Building integrated software development environments. Part I tool specification. ACM Transactions on Software Engineering and Methodology, 1(2):135–167, April 1992. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1992-1-2/p135-
engels/p135-engels.pdf;
http://www.acm.org/pubs/citations/
journals/tosem/1992-1-2/p135-
engels/.

Estublier:2005:ISE

[ELvdH^{+}05] Jacky Estublier, David Leblang, André van der Hoek, Reidar Conradi, Geoffrey Clemm, Walter Tichy, and Darcy Wiborg-Weber. Impact of software engineering research on the practice of software configuration management. ACM Transactions on Software Engineering and Methodology, 14(4):383– 430, October 2005. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Emam:2015:TCP

[EM15] Seyedeh Sepideh Emam and James Miller. Test case prioritization using extended digraphs. ACM Transactions on Software Engineering and Methodology, 25(1):6:1–6:??, December 2015. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Emam:2018:IEP

[EM18] S. S. Emam and J. Miller. Inferring extended probabilistic finite-state automaton models from software executions. ACM Transactions on Software Engineering and Methodology, 27 (1):4:1–4:??, June 2018. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Entekhabi:2024:AET

[EMM24] Sina Entekhabi, Wojciech Mostowski, and Mohammad Reza Mousavi. Automated and efficient test-generation for grid-based multiagent systems: Comparing random input filtering versus constraint solving. ACM Transactions on Software Engineering and Methodology, 33(1):12:1–12:??, January 2024. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3624736.

Eshuis:2006:SMC

[Esh06] Rik Eshuis. Symbolic model checking of UML activity diagrams. ACM Transactions on Software Engineering and Methodology, 15(1):1–38, January 2006. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Ebrahimi:2022:CMA

[ETM22] Fahimeh Ebrahimi, Miroslav Tushev, and Anas Mahmoud. Classifying mobile applications using word embeddings. ACM Transactions on Software Engineering and Methodology, 31(2): 20:1–20:30, April 2022. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3474827.

Erwig:2011:CCR

[EW11] Martin Erwig and Eric Walkingshaw. The choice calculus: a representation for software variation. ACM Transactions on Software Engineering and Methodology, 21(1):6:1– 6:??, December 2011. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Eldib:2014:FVS

[EWS14] Hassan Eldib, Chao Wang, and Patrick Schaumont. Formal verification of software countermeasures against sidechannel attacks. ACM Transactions on Software Engineering and Methodology, 24(2):11:1– 11:??, December 2014. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Fraser:2014:LSE

[FA14] Gordon Fraser and Andrea Arcuri. A large-scale evaluation of automated unit test generation using EvoSuite. ACM Transactions on Software Engineering and Methodology, 24(2): 8:1–8:??, December 2014. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Falessi:2022:IDD

[FAP22] Davide Falessi, Aalok Ahluwalia, and Massimiliano DI Penta. The impact of dormant defects on defect prediction: a study of 19 Apache projects. ACM Transactions on Software Engineering and Methodology, 31(1): 4:1–4:26, January 2022. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3467895.

Falessi:2013:VDR

[FBC^{+}13] Davide Falessi, Lionel C. Briand, Giovanni Cantone, Rafael Capilla, and Philippe Kruchten. The value of design rationale information. ACM Transactions on Software Engineering and Methodology, 22 (3):21:1–21:??, July 2013. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Fong:2000:PLM

[FC00] Philip W. L. Fong and Robert D. Cameron. Proof linking: modular verification of mobile programs in the presence of lazy, dynamic linking. ACM Transactions on Software Engineering and Methodology, 9(4):379–409, October 2000. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/2000-9-4/p379-
fong/p379-fong.pdf;   http:
//www.acm.org/pubs/citations/
journals/tosem/2000-9-4/p379-
fong/.

Fu:2021:SSC

[FCLL21] Xiaoqin Fu, Haipeng Cai, Wen Li, and Li Li. SEADS: Scalable and cost-effective dynamic dependence analysis of distributed systems via reinforcement learning. ACM Transactions on Software Engineering and Methodology, 30(1):10:1–10:45, January 2021. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3379345.

Fischbein:2012:WAM

[FDB^{+}12] Dario Fischbein, Nicolas D’Ippolito, Greg Brunet, Marsha Chechik, and Sebastian Uchitel. Weak alphabet merging of partial behavior models. ACM Transactions on Software Engineering and Methodology, 21(2):9:1– 9:??, March 2012. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Ferrando:2021:THA

[FDC^{+}21] Angelo Ferrando, Louise A.

Dennis, Rafael C. Cardoso, Michael Fisher, Davide Ancona, and Viviana Mascardi. Toward a holistic approach to verification and validation of autonomous cognitive systems. ACM Transactions on Software Engineering and Methodology, 30(4):43:1–43:43, July 2021. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3447246.

Formica:2024:SBS

[FFM24] Federico Formica, Tony Fan, and Claudio Menghi. Searchbased software testing driven by automatically generated and manually defined fitness functions. ACM Transactions on Software Engineering and Methodology, 33(2):40:1– 40:??, February 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3624745.

Fantechi:2012:LVM

[FGL^{+}12] Alessandro Fantechi, Stefania Gnesi, Alessandro Lapadula, Franco Mazzanti, Rosario Pugliese, and Francesco Tiezzi. A logical verification methodology for service-oriented computing. ACM Transactions on Software Engineering and Methodology, 21(3):16:1–16:??, June 2012. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Ferrari:2003:MCV

[FGMP03] Gian-Luigi Ferrari, Stefania Gnesi, Ugo Montanari, and Marco Pistore. A modelchecking verification environment for mobile processes. ACM Transactions on Software Engineering and Methodology, 12(4): 440–473, October 2003. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Ferrari:2024:UVB

[FHS^{+}24] Alessio Ferrari, Thaide Huichapa, Paola Spoletini, Nicole Novielli, Davide Fucci, and Daniela Girardi. Using voice and biofeedback to predict user engagement during product feedback interviews. ACM Transactions on Software Engineering and Methodology, 33(4):87:1–87:??, May 2024. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3635712.

Ferguson:1996:CAS

[FK96] Roger Ferguson and Bogdan Korel. The chaining approach for software test data generation. ACM Transactions on Software Engineering and Methodology, 5(1):63–86, January 1996. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1996-5-1/p63-
ferguson/p63-ferguson.pdf;
http://www.acm.org/pubs/
citations/journals/tosem/
1996-5-1/p63-ferguson/.

Fuggetta:1998:AGI

[FLM^{+}98] Alfonso Fuggetta, Luigi Lavazza, Sandro Morasca, Stefano Cinti, Giandomenico Oldano, and Elena Orazi. Applying GQM in an industrial software factory. ACM Transactions on Software Engineering and Methodology, 7(4):411–448, October 1998. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1998-7-4/p411-
fuggetta/p411-fuggetta.pdf;
 http://www.acm.org/pubs/
citations/journals/tosem/
1998-7-4/p411-fuggetta/.

Felder:1994:VRT

[FM94] Miguel Felder and Angelo Morzenti. Validating real-time systems by history-checking TRIO specifications. ACM Transactions on Software Engineering and Methodology, 3 (4):308–339, October 1994. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1994-3-4/p308-
felder/p308-felder.pdf;
http://www.acm.org/pubs/citations/
journals/tosem/1994-3-4/p308-
felder/.

Fioraldi:2023:DAFb

[FMMB23a] Andrea Fioraldi, Alessandro Mantovani, Dominik Maier, and Davide Balzarotti. Dissecting American Fuzzy Lop — a FuzzBench evaluation — RCR report. ACM Transactions on Software Engineering and Methodology, 32(2):54:1–54:??, April 2023. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3580600.

Fioraldi:2023:DAFa

[FMMB23b] Andrea Fioraldi, Alessandro Mantovani, Dominik Maier, and Davide Balzarotti. Dissecting American fuzzy lop: a FuzzBench evaluation. ACM Transactions on Software Engineering and Methodology, 32 (2):52:1–52:??, April 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3580596.

Fritz:2014:DKM

[FMMH^{+}14] Thomas Fritz, Gail C. Murphy, Emerson Murphy-Hill, Jingwen Ou, and Emily Hill. Degree-of-knowledge: Modeling a developer’s knowledge of code. ACM Transactions on Software Engineering and Methodology, 23(2): 14:1–14:??, March 2014. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Felty:2003:FSA

[FN03] Amy P. Felty and Kedar S.

Namjoshi. Feature specification and automated conflict detection. ACM Transactions on Software Engineering and Methodology, 12(1):3–27, 2003. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Fu:2024:VTI

[FNT^{+}24] Michael Fu, Van Nguyen, Chakkrit Tantithamthavorn, Dinh Phung, and Trung Le. Vision transformer inspired automated vulnerability repair. ACM Transactions on Software Engineering and Methodology, 33(3):78:1–78:??, March 2024. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3632746.

Forgacs:1994:DIF

[For94] István Forgács. Double iterative framework for flow-sensitive interprocedural data flow analysis. ACM Transactions on Software Engineering and Methodology, 3(1):29–55, January 1994. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

                                        www.acm.org/pubs/articles/
                                        journals/tosem/1994-3-1/p29-
                                        forgacs/p29-forgacs.pdf;
                                        http://www.acm.org/pubs/citations/
                                        journals/tosem/1994-3-1/p29-
                                        forgacs/.

Felder:2002:FDN

[FP02] Miguel Felder and Mauro Pezzè. A formal design notation for real-time systems. ACM Transactions on Software Engineering and Methodology, 11(2):149– 190, 2002. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Frias:2005:RAS

[FPB^{+}05] Marcelo F. Frias, Carlos G. López Pombo, Gabriel A. Baum, Nazareno M. Aguirre, and Thomas S. E. Maibaum. Reasoning about static and dynamic properties in alloy: a purely relational approach. ACM Transactions on Software Engineering and Methodology, 14(4):478–526, October 2005. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Fahmy:2023:SBE

[FPBS23] Hazem Fahmy, Fabrizio Pastore, Lionel Briand, and Thomas Stifter. Simulatorbased explanation and debugging of hazard-triggering events in DNN-based safetycritical systems. ACM Transactions on Software Engineering and Methodology, 32(4): 104:1–104:??, July 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3569935.

Frias:2007:EAD

[FPGA07] Marcelo F. Frias, Carlos G. Lopez Pombo, Juan P. Galeotti, and Nazareno M. Aguirre. Efficient Analysis of DynAlloy Specifications. ACM Transactions on Software Engineering and Methodology, 17(1):4:1– 4:34, December 2007. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Fisher:2006:IAT

[FRB^{+}06] Marc Fisher II, Gregg Rothermel, Darren Brown, Mingming Cao, Curtis Cook, and Margaret Burnett. Integrating automated test generation into the WYSIWYT spreadsheet testing methodology. ACM Transactions on Software Engineering and Methodology, 15(2):150– 194, April 2006. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Feldman:1993:SRS

[FS93] Yishai A. Feldman and Haim Schneider. Simulating reactive systems by deduction. ACM Transactions on Software Engineering and Methodology, 2(2):128–175, April 1993. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

                                        www.acm.org/pubs/articles/
                                        journals/tosem/1993-2-2/p128-
                                        feldman/p128-feldman.pdf;
                                         http://www.acm.org/pubs/
                                        citations/journals/tosem/
                                        1993-2-2/p128-feldman/.

Fraser:2015:DAU

[FSM^{+}15] Gordon Fraser, Matt Staats, Phil McMinn, Andrea Arcuri, and Frank Padberg. Does automated unit test generation really help software testers? A controlled empirical study. ACM Transactions on Software Engineering and Methodology, 24(4):23:1–23:??, August 2015. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Fleming:2013:IFT

[FSP^{+}13] Scott D. Fleming, Chris Scaffidi, David Piorkowski, Margaret Burnett, Rachel Bellamy, Joseph Lawrance, and Irwin Kwan. An information foraging theory perspective on tools for debugging, refactoring, and reuse tasks. ACM Transactions on Software Engineering and Methodology, 22(2):14:1– 14:??, March 2013. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Ford:2022:TTC

[FSZ^{+}22] Denae Ford, Margaret-Anne Storey, Thomas Zimmermann, Christian Bird, Sonia Jaffe, Chandra Maddila, Jenna L. Butler, Brian Houck, and Nachiappan Nagappan. A tale of two cities: Software developers working from home during the COVID-19 pandemic. ACM Transactions on Software Engineering and Methodology, 31 (2):27:1–27:37, April 2022. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3487567.

Furia:2022:ABA

[FTF22] Carlo A. Furia, Richard Torkar, and Robert Feldt. Applying Bayesian analysis guidelines to empirical software engineering data: The case of programming languages and code quality.

ACM Transactions on Software Engineering and Methodology, 31(3):40:1–40:38, July 2022. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3490953.

Furia:2024:TCA

[FTF24] Carlo A. Furia, Richard Torkar, and Robert Feldt. Towards causal analysis of empirical software engineering data: The impact of programming languages on coding competitions. ACM Transactions on Software Engineering and Methodology, 33(1): 13:1–13:??, January 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3611667.

Fink:2008:ETV

[FYD^{+}08] Stephen J. Fink, Eran Yahav, Nurit Dor, G. Ramalingam, and Emmanuel Geay. Effective typestate verification in the presence of aliasing. ACM Transactions on Software Engineering and Methodology, 17(2):9:1– 9:??, April 2008. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Fang:2024:DQP

[FYW24] Wang Fang, Mingsheng Ying, and Xiaodi Wu. Differentiable quantum programming with unbounded loops. ACM Transactions on Software Engineering and Methodology, 33(1): 19:1–19:??, January 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3617178.

Guo:2024:SCC

[GCC^{+}24] Hanyang Guo, Yingye Chen, Xiangping Chen, Yuan Huang, and Zibin Zheng. Smart contract code repair recommendation based on reinforcement learning and multi-metric optimization. ACM Transactions on Software Engineering and Methodology, 33(4): 106:1–106:??, May 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3637229.

Guo:2024:SCG

[GCH^{+}24] Hanyang Guo, Xiangping Chen, Yuan Huang, Yanlin Wang, Xi Ding, Zibin Zheng, Xiaocong Zhou, and Hong-Ning Dai. Snippet comment generation based on code context expansion. ACM Transactions on Software Engineering and Methodology, 33(1): 24:1–24:??, January 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3611664.

Guo:2022:DAM

[GCX^{+}22] Hao Guo, Sen Chen, Zhenchang Xing, Xiaohong Li, Yude Bai, and Jiamou Sun. Detecting and augmenting missing key aspects in vulnerability descriptions. ACM Transactions on Software Engineering and Methodology, 31(3): 49:1–49:27, July 2022. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3498537.

Gencel:2008:FSM

[GD08] Cigdem Gencel and Onur Demirors. Functional size measurement revisited. ACM Transactions on Software Engineering and Methodology, 17 (3):15:1–15:??, June 2008. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Godoy:2021:EBT

[GGGU21] Javier Godoy, Juan Pablo Galeotti, Diego Garbervetsky, and Sebastián Uchitel. Enabledness-based testing of object protocols. ACM Transactions on Software Engineering and Methodology, 30(2): 12:1–12:36, March 2021. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3415153.

Gao:2023:CSG

[GGH^{+}23] Shuzheng Gao, Cuiyun Gao, Yulan He, Jichuan Zeng, Lunyiu Nie, Xin Xia, and Michael Lyu. Code structure-guided transformer for source code summarization. ACM Transactions on Software Engineering and Methodology, 32(1):

23:1–23:??, January 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3522674.

Gunatilake:2024:EBE

[GGHM24] Hashini Gunatilake, John Grundy, Rashina Hoda, and Ingo Mueller. Enablers and barriers of empathy in software developer and user interactions: a mixed methods case study. ACM Transactions on Software Engineering and Methodology, 33(4):109:1–109:??, May 2024. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3641849.

Gamatie:2007:PDE

[GGLT07] Abdoulaye Gamatié, Thierry Gautier, Paul Le Guernic, and Jean-Pierre Talpin. Polychronous design of embedded real-time applications. ACM Transactions on Software Engineering and Methodology, 16 (2):9:1–9:??, April 2007. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Giamattei:2024:CDT

[GGPR24] Luca Giamattei, Antonio Guerriero, Roberto Pietrantuono, and Stefano Russo. Causalitydriven testing of autonomous driving systems. ACM Transactions on Software Engineering and Methodology, 33(3): 74:1–74:??, March 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3635709.

Gligoric:2015:GCB

[GGZ^{+}15] Milos Gligoric, Alex Groce, Chaoqiang Zhang, Rohan Sharma, Mohammad Amin Alipour, and Darko Marinov. Guidelines for coverage-based comparisons of non-adequate test suites. ACM Transactions on Software Engineering and Methodology, 24(4): 22:1–22:??, August 2015. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Ghezzi:2005:E

[Ghe05] Carlo Ghezzi. Editorial. ACM Transactions on Software Engineering and Methodology, 14 (2):119–123, April 2005. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Ghezzi:2007:E

[Ghe07] Carlo Ghezzi. Editorial. ACM Transactions on Software Engineering and Methodology, 16(1): 2:1–2:??, February 2007. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Graves:2001:ESR

[GHK^{+}01] Todd L. Graves, Mary Jean Harrold, Jung-Min Kim, Adam Porter, and Gregg Rothermel. An empirical study of regression test selection techniques.

ACM Transactions on Software Engineering and Methodology, 10(2):184–208, April 2001. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/2001-10-2/
p184-graves/p184-graves.pdf;
http://www.acm.org/pubs/
citations/journals/tosem/
2001-10-2/p184-graves/.

Garcia:2018:LOR

[GHM18] Joshua Garcia, Mahmoud Hammad, and Sam Malek. Lightweight, obfuscation-resilient detection and family identification of Android malware. ACM Transactions on Software Engineering and Methodology, 26(3): 11:1–11:??, January 2018. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Guo:2024:KKN

[GHX^{+}24] Yuejun Guo, Qiang Hu, Xiaofei Xie, Maxime Cordy, Mike Papadakis, and Yves Le Traon. KAPE: k NN-based performance testing for deep code search. ACM Transactions on Software Engineering and Methodology, 33(2):48:1– 48:??, February 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3624735.

Gao:2024:CDL

[GHXZ24] Kai Gao, Runzhi He, Bing Xie, and Minghui Zhou. Characterizing deep learning package supply chains in PyPI: Domains, clusters, and disengagement. ACM Transactions on Software Engineering and Methodology, 33(4): 97:1–97:??, May 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3640336.

Gruschke:2008:ROF

[GJ08] Tanja M. Gruschke and Magne Jørgensen. The role of outcome feedback in improving the uncertainty assessment of software development effort estimates. ACM Transactions on Software Engineering and Methodology, 17(4):20:1–20:??, August 2008. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Ghorbani:2023:DDD

[GJS^{+}23] Negar Ghorbani, Reyhaneh Jabbarvand, Navid Salehnamadi, Joshua Garcia, and Sam Malek. DeltaDroid: Dynamic delivery testing in Android. ACM Transactions on Software Engineering and Methodology, 32(4):84:1–84:??, July 2023. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3563213.

Gandhi:2011:DMC

[GL11] R. A. Gandhi and S. W. Lee. Discovering multidimensional correlations among regulatory requirements to understand risk. ACM Transactions on Software Engineering and Methodology, 20(4):16:1– 16:??, September 2011. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Ganesan:2014:AED

[GL14] Dharmalingam Ganesan and Mikael Lindvall. ADAM: External dependency-driven architecture discovery and analysis of quality attributes. ACM Transactions on Software Engineering and Methodology, 23(2): 17:1–17:??, March 2014. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Graziotin:2022:PBS

[GLFW22] Daniel Graziotin, Per Lenberg, Robert Feldt, and Stefan Wagner. Psychometrics in behavioral software engineering: a methodological introduction with guidelines. ACM Transactions on Software Engineering and Methodology, 31(1): 7:1–7:36, January 2022. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3469888.

Guo:2024:EUL

[GLH^{+}24] Shikai Guo, Dongmin Li, Lin Huang, Sijia Lv, Rong Chen, Hui Li, Xiaochen Li, and He Jiang. Estimating uncertainty in labeled changes by SZZ tools on just-in-time defect prediction. ACM Transactions on Software Engineering and Methodology, 33(4): 105:1–105:??, May 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3637226.

Guo:2021:HFW

[GLL^{+}21] Zhaoqiang Guo, Shiran Liu, Jinping Liu, Yanhui Li, Lin Chen, Hongmin Lu, and Yuming Zhou. How far have we progressed in identifying self-admitted technical debts? A comprehensive empirical study. ACM Transactions on Software Engineering and Methodology, 30 (4):45:1–45:56, July 2021. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3447247.

Guo:2023:CLL

[GLL^{+}23] Zhaoqiang Guo, Shiran Liu, Xutong Liu, Wei Lai, Mingliang Ma, Xu Zhang, Chao Ni, Yibiao Yang, Yanhui Li, Lin Chen, Guoqiang Zhou, and Yuming Zhou. Code-line-level bugginess identification: How far have we come, and how far have we yet to go? ACM Transactions on Software Engineering and Methodology, 32(4): 102:1–102:??, July 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3582572.

Gargantini:2001:ADR

[GM01] Angelo Gargantini and Angelo Morzenti. Automated deductive requirements analysis of critical systems. ACM Transactions on Software Engineering and Methodology, 10(3): 255–307, July 2001. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Ghezzi:2003:E

[GMRS03] Carlo Ghezzi, Jeffrey N. Magee, Dieter Rombach, and Mary Lou Soffa. Editorial. ACM Transactions on Software Engineering and Methodology, 12(1):1– 2, 2003. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Griswold:1993:AAP

[GN93] William G. Griswold and David Notkin. Automated assistance for program restructuring. ACM Transactions on Software Engineering and Methodology, 2(3):228–269, July 1993. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1993-2-3/p228-
griswold/p228-griswold.pdf;
http://www.acm.org/pubs/
citations/journals/tosem/
1993-2-3/p228-griswold/.

Gay:2016:EPM

[GRS^{+}16] Gregory Gay, Ajitha Rajan, Matt Staats, Michael Whalen, and Mats P. E. Heimdahl. The effect of program and model structure on the effectiveness of MC/DC test adequacy coverage. ACM Transactions on Software Engineering and Methodology, 25(3):25:1–25:??, August 2016. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Goel:2009:IPC

[GRT09] Ankit Goel, Abhik Roychoudhury, and P. S. Thiagarajan. Interacting process classes. ACM Transactions on Software Engineering and Methodology, 18 (4):13:1–13:??, July 2009. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Gao:2024:CVF

[GSC24] Pengfei Gao, Fu Song, and Taolue Chen. Compositional verification of first-order masking countermeasures against power side-channel attacks. ACM Transactions on Software Engineering and Methodology, 33(3):79:1–79:??, March 2024. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3635707.

Gupta:1997:HSI

[GSH97] Rajiv Gupta, Mary Lou Soffa, and John Howard. Hybrid slicing: integrating dynamic information with static analysis. ACM Transactions on Software Engineering and Methodology, 6(4):370–397, October 1997. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1997-6-4/p370-
gupta/p370-gupta.pdf;   http:
//www.acm.org/pubs/citations/
journals/tosem/1997-6-4/p370-
gupta/.

Gong:2021:TDG

[GSYT21] Dunwei Gong, Baicai Sun, Xiangjuan Yao, and Tian Tian. Test data generation for path coverage of MPI programs using SAEO. ACM Transactions on Software Engineering and Methodology, 30(2): 17:1–17:37, March 2021. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3423132.

Guerriero:2021:SMD

[GTD21] Michele Guerriero, Damian Andrew Tamburri, and Elisabetta Di Nitto. StreamGen: Modeldriven development of distributed streaming applications. ACM Transactions on Software Engineering and Methodology, 30(1):1:1–1:30, January 2021. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3408895.

Gunter:2000:ADB

[Gun00] Carl A. Gunter. Abstracting dependencies between software configuration items. ACM Transactions on Software Engineering and Methodology, 9(1):94–131, January 2000. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

                                        www.acm.org/pubs/articles/
                                        journals/tosem/2000-9-1/p94-
                                        gunter/p94-gunter.pdf;   http:
                                        //www.acm.org/pubs/citations/
                                        journals/tosem/2000-9-1/p94-
                                        gunter/.

Gao:2021:BTP

[GWD^{+}21] Xiang Gao, Bo Wang, Gregory J. Duck, Ruyi Ji, Yingfei Xiong, and Abhik Roychoudhury. Beyond tests: Program vulnerability repair via crash constraint extraction. ACM Transactions on Software Engineering and Methodology, 30(2): 14:1–14:27, March 2021. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3418461.

Gao:2020:GQT

[GXG^{+}20] Zhipeng Gao, Xin Xia, John Grundy, David Lo, and Yuan-Fang Li. Generating question titles for Stack Overflow from mined code snippets. ACM Transactions on Software Engineering and Methodology, 29(4): 26:1–26:37, October 2020. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3401026.

Gao:2023:KWY

[GXL^{+}23] Zhipeng Gao, Xin Xia, David Lo, John Grundy, Xindong Zhang, and Zhenchang Xing. I know what you are searching for: Code snippet recommendation from Stack Overflow posts. ACM Transactions on Software Engineering and Methodology, 32(3): 80:1–80:??, July 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3550150.

Gao:2021:TQS

[GXLG21] Zhipeng Gao, Xin Xia, David Lo, and John Grundy. Technical Q8A site answer recommendation via question boosting. ACM Transactions on Software Engineering and Methodology, 30(1):11:1–11:34, January 2021. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3412845.

Gao:2021:HAF

[GXSC21] Pengfei Gao, Hongyi Xie, Fu Song, and Taolue Chen. A hybrid approach to formal verification of higher-order masked arithmetic programs. ACM Transactions on Software Engineering and Methodology, 30 (3):26:1–26:42, May 2021. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3428015.

Gervasi:2005:RAI

[GZ05] Vincenzo Gervasi and Didar Zowghi. Reasoning about inconsistencies in natural language requirements. ACM Transactions on Software Engineering and Methodology, 14(3): 277–330, July 2005. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Golmohammadi:2024:TRA

[GZA24] Amid Golmohammadi, Man Zhang, and Andrea Arcuri. Testing RESTful APIs: a survey. ACM Transactions on Software Engineering and Methodology, 33(1):27:1–27:??, January 2024. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3617175.

Gao:2019:VQS

[GZSW19] Pengfei Gao, Jun Zhang, Fu Song, and Chao Wang. Verifying and quantifying sidechannel resistance of masked software implementations. ACM Transactions on Software Engineering and Methodology, 28(3): 16:1–16:??, August 2019. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.
                                        acm.org/ft_gateway.cfm?id=
                                        3330392.

Gong:2023:WIU

[GZW^{+}23] Lina Gong, Jingxuan Zhang, Mingqiang Wei, Haoxiang Zhang, and Zhiqiu Huang. What is the intended usage context of this model? An exploratory study of pre-trained models on various model repositories. ACM Transactions on Software Engineering and Methodology, 32(3):69:1–69:??, July 2023. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3569934.

Gao:2022:AAR

[GZX^{+}22] Cuiyun Gao, Wenjie Zhou, Xin Xia, David Lo, Qi Xie, and Michael R. Lyu. Automating app review response generation based on contextual knowledge. ACM Transactions on Software Engineering and Methodology, 31(1):11:1–11:36, January 2022. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3464969.

Hemmati:2013:ASM

[HAB13] Hadi Hemmati, Andrea Arcuri, and Lionel Briand. Achieving scalable model-based testing through test case diversity. ACM Transactions on Software Engineering and Methodology, 22(1):6:1–6:??, February 2013. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Holmes:2020:URL

[HAB^{+}20] Josie Holmes, Iftekhar Ahmed, Caius Brindescu, Rahul Gopinath, He Zhang, and Alex Groce. Using relative lines of code to guide automated test generation for Python. ACM Transactions on Software Engineering and Methodology, 29(4): 28:1–28:38, October 2020. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3408896.

Harel:1992:SO

[HaK92] David Harel and Chaim arie Kahana. On statecharts with overlapping. ACM Transactions on Software Engineering and Methodology, 1(4):399–421, October 1992. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

                                        www.acm.org/pubs/articles/
                                        journals/tosem/1992-1-4/p399-
                                        harel/p399-harel.pdf;   http:
                                        //www.acm.org/pubs/citations/
                                        journals/tosem/1992-1-4/p399-
                                        harel/.

Hamlet:2009:TES

[Ham09] Dick Hamlet. Tools and experiments supporting a testingbased theory of component composition. ACM Transactions on Software Engineering and Methodology, 18(3):12:1–12:??, May 2009. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Han:2023:UAR

[HAY^{+}23] Liping Han, Shaukat Ali, Tao Yue, Aitor Arrieta, and Maite Arratibel. Uncertainty-aware robustness assessment of industrial elevator systems. ACM Transactions on Software Engineering and Methodology, 32 (4):95:1–95:??, July 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3576041.

Hough:2022:PAD

[HB22] Katherine Hough and Jonathan Bell. A practical approach for dynamic taint tracking with control-flow relationships. ACM Transactions on Software Engineering and Methodology, 31(2): 26:1–26:43, April 2022. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3485464.

Hall:2009:SRT

[HBB^{+}09] Tracy Hall, Nathan Baddoo, Sarah Beecham, Hugh Robinson, and Helen Sharp. A systematic review of theory use in studies investigating the motivations of software engineers. ACM Transactions on Software Engineering and Methodology, 18(3):10:1–10:??, May 2009. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Huang:2024:MAD

[HCJ^{+}24] Zhenfei Huang, Junjie Chen, Jiajun Jiang, Yihua Liang, Hanmo You, and Fengjie Li. Mapping APIs in dynamictyped programs by leveraging transfer learning. ACM Transactions on Software Engineering and Methodology, 33(4): 102:1–102:??, May 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3641848.

Hu:2022:CAH

[HCW^{+}22] Xing Hu, Qiuyuan Chen, Haoye Wang, Xin Xia, David Lo, and Thomas Zimmermann. Correlating automated and human evaluation of code documentation generation quality. ACM Transactions on Software Engineering and Methodology, 31(4): 63:1–63:??, October 2022. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3502853.

Hu:2024:AMA

[HDG^{+}24] Han Hu, Ruiqi Dong, John Grundy, Thai Minh Nguyen, Huaxiao Liu, and Chunyang Chen. Automated mapping of adaptive app GUIs from phones to TVs. ACM Transactions on Software Engineering and Methodology, 33(2):47:1– 47:??, February 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3631968.

Hutiri:2023:TAF

[HDKM23] Wiebke (Toussaint) Hutiri, Aaron Yi Ding, Fahim Kawsar, and Akhil Mathur. Tiny, always-on, and fragile: Bias propagation through design choices in on-device machine learning workflows. ACM Transactions on Software Engineering and Methodology, 32(6): 155:1–155:??, November 2023. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3591867.

Hoffman:2013:TOM

[HE13] Kevin Hoffman and Patrick Eugster. Trading obliviousness for modularity with cooperative aspect-oriented programming. ACM Transactions on Software Engineering and Methodology, 22(3):22:1–22:??, July 2013. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Henninger:1997:EAC

[Hen97] Scott Henninger. An evolutionary approach to constructing effective software reuse repositories. ACM Transactions on Software Engineering and Methodology, 6(2):111–140, April 1997. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1997-6-2/p111-

henninger/p111-henninger.
pdf;   http://www.acm.org/
pubs/citations/journals/tosem/
1997-6-2/p111-henninger/.

Hu:2022:ESD

[HGC^{+}22] Qiang Hu, Yuejun Guo, Maxime Cordy, Xiaofei Xie, Lei Ma, Mike Papadakis, and Yves Le Traon. An empirical study on data distribution-aware test selection for deep learning enhancement. ACM Transactions on Software Engineering and Methodology, 31(4):78:1– 78:??, October 2022. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3511598.

Huang:2023:CSM

[HGD^{+}23] Yuan Huang, Hanyang Guo, Xi Ding, Junhuai Shu, Xiangping Chen, Xiapu Luo, Zibin Zheng, and Xiaocong Zhou. A comparative study on method comment and inline comment. ACM Transactions on Software Engineering and Methodology, 32(5):126:1–126:??, September 2023. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3582570.

Hidellaarachchi:2023:IHA

[HGHM23] Dulaji Hidellaarachchi, John Grundy, Rashina Hoda, and Ingo Mueller. The influence of human aspects on requirements engineering-related activities: Software practitioners’ perspective. ACM Transactions on Software Engineering and Methodology, 32(5):108:1– 108:??, September 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3546943.

Harrold:1993:MCS

[HGS93] M. Jean Harrold, Rajiv Gupta, and Mary Lou Soffa. A methodology for controlling the size of a test suite. ACM Transactions on Software Engineering and Methodology, 2(3):270–285, July 1993. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1993-2-3/p270-
harrold/p270-harrold.pdf;
http://www.acm.org/pubs/
citations/journals/tosem/
1993-2-3/p270-harrold/.

He:2016:LWA

[HGW^{+}16] Fei He, Xiaowei Gao, Miaofei Wang, Bow-Yaw Wang, and Lijun Zhang. Learning weighted assumptions for compositional verification of Markov decision processes. ACM Transactions on Software Engineering and Methodology, 25(3):21:1– 21:??, August 2016. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Hu:2024:TOD

[HGX^{+}24a] Qiang Hu, Yuejun Guo, Xiaofei Xie, Maxime Cordy, Lei Ma, Mike Papadakis, and Yves Le Traon. Test optimization in DNN testing: a survey. ACM Transactions on Software Engineering and Methodology, 33(4): 111:1–111:??, May 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3643678.

Hu:2024:LLF

[HGX^{+}24b] Qiang Hu, Yuejun Guo, Xiaofei Xie, Maxime Cordy, Mike Papadakis, and Yves Le Traon. LaF: Labeling-free model selection for automated deep neural network reusing. ACM Transactions on Software Engineering and Methodology, 33(1): 25:1–25:??, January 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3611666.

Howden:1995:STA

[HH95] W. E. Howden and Yudong Huang. Software trustability analysis. ACM Transactions on Software Engineering and Methodology, 4(1):36–64, January 1995. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1995-4-1/p36-
howden/p36-howden.pdf;   http:
//www.acm.org/pubs/citations/
journals/tosem/1995-4-1/p36-
howden/.

Hu:2024:FLD

[HHC^{+}24] Han Hu, Yujin Huang, Qiuyuan Chen, Terry Yue Zhuo, and Chunyang Chen. A first look at on-device models in iOS apps. ACM Transactions on Software Engineering and Methodology, 33(1):26:1–26:??, January 2024. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3617177.

Hierons:2002:CTS

[Hie02] R. M. Hierons. Comparing test sets and criteria in the presence of test hypotheses and fault domains. ACM Transactions on Software Engineering and Methodology, 11(4):427– 448, 2002. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Hierons:2006:ACC

[Hie06] R. M. Hierons. Avoiding coincidental correctness in boundary value analysis. ACM Transactions on Software Engineering and Methodology, 15(3): 227–241, July 2006. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Hierons:2009:VFT

[Hie09]  Robert M. Hierons.  Ver-

dict functions in testing with a fault domain or test hypotheses. ACM Transactions on Software Engineering and Methodology, 18(4):14:1–14:??, July 2009. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Hierons:2014:CCD

[Hie14] Robert M. Hierons. Combining centralised and distributed testing. ACM Transactions on Software Engineering and Methodology, 24(1):5:1–5:??, September 2014. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Heitmeyer:1996:ACC

[HJL96] Constance L. Heitmeyer, Ralph D. Jeffords, and Bruce G. Labaw. Automated consistency checking of requirements specifications. ACM Transactions on Software Engineering and Methodology, 5(3):231–261, July 1996. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

                                        www.acm.org/pubs/articles/
                                        journals/tosem/1996-5-3/p231-
                                        heitmeyer/p231-heitmeyer.
                                        pdf;   http://www.acm.org/
                                        pubs/citations/journals/tosem/
                                        1996-5-3/p231-heitmeyer/.

Harman:2014:ESS

[HKMB^{+}14] Mark Harman, Jens Krinke, Inmaculada Medina-Bulo, Francisco Palomo-Lozano, Jian Ren, and Shin Yoo. Exact scalable sensitivity analysis for the next release problem. ACM Transactions on Software Engineering and Methodology, 23(2): 19:1–19:??, March 2014. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Hierons:2016:SOP

[HLL^{+}16] Robert M. Hierons, Miqing Li, Xiaohui Liu, Sergio Segura, and Wei Zheng. SIP: Optimal product selection from feature models using many-objective evolutionary optimization. ACM Transactions on Software Engineering and Methodology, 25 (2):17:1–17:??, May 2016. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Hierons:2020:MOT

[HLL^{+}20] Robert M. Hierons, Miqing Li, Xiaohui Liu, Jose Antonio Parejo, Sergio Segura, and Xin Yao. Many-objective test suite generation for software product lines. ACM Transactions on Software Engineering and Methodology, 29(1):2:1– 2:46, February 2020. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3361146.

He:2023:IBC

[HLX23a] Dongjie He, Jingbo Lu, and Jingling Xue. IFDS-based context debloating for object-sensitive pointer analysis. ACM Transactions on Software Engineering and Methodology, 32(4): 101:1–101:??, July 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3579641.

Huang:2023:SEC

[HLX^{+}23b] Qing Huang, Dianshu Liao, Zhenchang Xing, Zhengkang Zuo, Changjing Wang, and Xin Xia. Semantic-enriched code knowledge graph to reveal unknowns in smart contract code reuse. ACM Transactions on Software Engineering and Methodology, 32(6):147:1– 147:??, November 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3597206.

Huang:2023:PPR

[HML23] Yuheng Huang, Lei Ma, and Yuanchun Li. PatchCensor: Patch robustness certification for transformers via exhaustive testing. ACM Transactions on Software Engineering and Methodology, 32(6):154:1– 154:??, November 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3591870.

Hassan:2023:UUF

[HMW23] Foyzul Hassan, Na Meng, and Xiaoyin Wang. UniLoc: Unified fault localization of continuous integration failures. ACM Transactions on Software Engineering and Methodology, 32(6): 136:1–136:??, November 2023. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3593799.

Harel:1996:SSS

[HN96] David Harel and Amnon Naamad. The STATEMATE semantics of statecharts. ACM Transactions on Software Engineering and Methodology, 5 (4):293–333, October 1996. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1996-5-4/p293-
harel/p293-harel.pdf;   http:
//www.acm.org/pubs/citations/
journals/tosem/1996-5-4/p293-
harel/.

Hunter:1998:MIS

[HN98] Anthony Hunter and Bashar Nuseibeh. Managing inconsistent specifications: reasoning, analysis, and action. ACM Transactions on Software Engineering and Methodology, 7 (4):335–367, October 1998. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1998-7-4/p335-
hunter/p335-hunter.pdf;
http://www.acm.org/pubs/citations/
journals/tosem/1998-7-4/p335-
hunter/.

Haas:2020:SAA

[HNRA20] Roman Haas, Rainer Niedermayr, Tobias Roehm, and Sven Apel. Is static analysis able to identify unnecessary source code? ACM Transactions on Software Engineering and Methodology, 29(1):6:1– 6:23, February 2020. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3368267.

Harman:2015:ISI

[HP15] Mark Harman and Mauro Pezzé. Introduction to the special issue on ISSTA 2013. ACM Transactions on Software Engineering and Methodology, 24(4): 21:1–21:??, August 2015. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Herrera:2023:DTDa

[HPH23a] Adrian Herrera, Mathias Payer, and Antony L. Hosking. DatAFLow: Toward a data-flow-guided fuzzer. ACM Transactions on Software Engineering and Methodology, 32(5):132:1–132:??, September 2023. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3587156.

Herrera:2023:DTDb

[HPH23b] Adrian Herrera, Mathias Payer, and Antony L. Hosking. DatAFLow: Toward a data-flow-guided fuzzer. ACM Transactions on Software Engineering and Methodology, 32(5):133:1–133:??, September 2023. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3587159.

Hu:2023:NTE

[HQM^{+}23] Fan Hu, Shisong Qin, Zheyu Ma, Bodong Zhao, Tingting Yin, and Chao Zhang. NSFuzz: Towards efficient and stateaware network service fuzzing — RCR report. ACM Transactions on Software Engineering and Methodology, 32(6):161:1– 161:??, November 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3580599.

Henkel:2008:DDA

[HRD08] Johannes Henkel, Christoph Reichenbach, and Amer Diwan. Developing and debugging algebraic specifications for Java classes. ACM Transactions on Software Engineering and Methodology, 17(3):14:1–14:??, June 2008. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

He:2023:TAE

[HSB^{+}23] Zheyuan He, Shuwei Song, Yang Bai, Xiapu Luo, Ting Chen, Wensheng Zhang, Peng He, Hongwei Li, Xiaodong Lin, and Xiaosong Zhang. TokenAware: Accurate and efficient bookkeeping recognition for token smart contracts. ACM Transactions on Software Engineering and Methodology, 32(1): 26:1–26:??, January 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3560263.

Huang:2024:AER

[HSX^{+}24] Qing Huang, Yanbang Sun, Zhenchang Xing, Min Yu, Xiwei Xu, and Qinghua Lu. API entity and relation joint extraction from text via dynamic prompttuned language model. ACM Transactions on Software Engineering and Methodology, 33(1): 6:1–6:??, January 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3607188.

Hunt:1998:ADA

[HT98] James J. Hunt and Walter F. Tichy. Addendum to “Delta algorithms: an empirical analysis”. ACM Transactions on Software Engineering and Methodology, 7(4):449, October 1998. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

                                        www.acm.org/pubs/articles/
                                        journals/tosem/1998-7-4/p449-
                                        hunt/p449-hunt.pdf;   http:
                                        //www.acm.org/pubs/citations/
                                        journals/tosem/1998-7-4/p449-

hunt/. See [HVT98].

Hierons:2017:PAG

[HT17] Robert M. Hierons and Uraz Cengiz Türker. Parallel algorithms for generating distinguishing sequences for observable nondeterministic FSMs. ACM Transactions on Software Engineering and Methodology, 26 (1):5:1–5:??, July 2017. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Hunt:1998:DAE

[HVT98] James J. Hunt, Kiem-Phong Vo, and Walter F. Tichy. Delta algorithms an empirical analysis. ACM Transactions on Software Engineering and Methodology, 7(2):192–214, April 1998. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1998-7-2/p192-
hunt/p192-hunt.pdf;   http:
//www.acm.org/pubs/citations/
journals/tosem/1998-7-2/p192-

hunt/. See addendum [HT98].

Holmes:2012:SPS

[HW12] Reid Holmes and Robert J.

Walker. Systematizing pragmatic software reuse. ACM Transactions on Software Engineering and Methodology, 21 (4):20:1–20:??, November 2012. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Haesevoets:2014:ACS

[HWH14] Robrecht Haesevoets, Danny Weyns, and Tom Holvoet. Architecture-centric support for adaptive service collaborations. ACM Transactions on Software Engineering and Methodology, 23(1):2:1–2:??, February 2014. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Huang:2024:FIP

[HYX^{+}24] Qing Huang, Zhiqiang Yuan, Zhenchang Xing, Xin Peng, Xiwei Xu, and Qinghua Lu. FQN inference in partial code by prompt-tuned language model of code. ACM Transactions on Software Engineering and Methodology, 33(2):31:1– 31:??, February 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3617174.

Huang:2024:HDA

[HZB^{+}24] Wei Huang, Xingyu Zhao, Alec Banks, Victoria Cox, and Xiaowei Huang. Hierarchical distribution-aware testing of deep learning. ACM Transactions on Software Engineering and Methodology, 33(2):42:1– 42:??, February 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3625290.

Hall:2014:SCS

[HZBS14] Tracy Hall, Min Zhang, David Bowes, and Yi Sun. Some code smells have a significant but small effect on faults. ACM Transactions on Software Engineering and Methodology, 23(4): 33:1–33:??, August 2014. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Huang:2008:DSL

[HZS08] Shan Shan Huang, David Zook, and Yannis Smaragdakis. Domain-specific languages and program generation with meta-AspectJ. ACM Transactions on Software Engineering and Methodology, 18(2): 6:1–6:??, November 2008. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

He:2024:RLS

[HZX^{+}24] Junda He, Xin Zhou, Bowen Xu, Ting Zhang, Kisub Kim, Zhou Yang, Ferdian Thung, Ivana Clairine Irsan, and David Lo. Representation learning for Stack Overflow posts: How far are we? ACM Transactions on Software Engineering and Methodology, 33(3): 69:1–69:??, March 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3635711.

Huang:2013:SPA

[HZZ13] Jeff Huang, Jinguo Zhou, and Charles Zhang. Scaling predictive analysis of concurrent programs by removing trace redundancy. ACM Transactions on Software Engineering and Methodology, 22(1):8:1– 8:??, February 2013. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Hao:2014:UTC

[HZZ^{+}14] Dan Hao, Lingming Zhang, Lu Zhang, Gregg Rothermel, and Hong Mei. A unified test case prioritization approach. ACM Transactions on Software Engineering and Methodology, 24(2):10:1–10:??, December 2014. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Islam:2014:GTC

[IC14] Mainul Islam and Christoph Csallner. Generating test cases for programs that are coded against interfaces and annotations. ACM Transactions on Software Engineering and Methodology, 23(3):21:1–21:??, May 2014. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Ismayilzada:2024:PTP

[IRKY24] Elkhan Ismayilzada, Md Mazba Ur Rahman, Dongsun Kim, and Jooyong Yi. Poracle: Testing patches under preservation conditions to combat the overfitting problem of program repair. ACM Transactions on Software Engineering and Methodology, 33(2):44:1–44:??, February 2024. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3625293.

Inverardi:2000:SCS

[IWY00] Paola Inverardi, Alexander L. Wolf, and Daniel Yankelevich.

Static checking of system behaviors using derived component assumptions. ACM Transactions on Software Engineering and Methodology, 9(3):239–272, July 2000. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/2000-9-3/p239-
inverardi/p239-inverardi.
pdf;   http://www.acm.org/
pubs/citations/journals/tosem/
2000-9-3/p239-inverardi/.

Joblin:2022:HDS

[JA22] Mitchell Joblin and Sven Apel.

How do successful and failed projects differ? A sociotechnical analysis. ACM Transactions on Software Engineering and Methodology, 31(4): 67:1–67:??, October 2022. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3504003.

Jackson:1995:ADB

[Jac95a] Daniel Jackson. Aspect: detecting bugs with abstract dependences. ACM Transactions on Software Engineering and Methodology, 4(2):109–145, April 1995. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1995-4-2/p109-
jackson/p109-jackson.pdf;
 http://www.acm.org/pubs/
citations/journals/tosem/
1995-4-2/p109-jackson/.

Jackson:1995:SZS

[Jac95b] Daniel Jackson. Structuring Z specifications with views. ACM Transactions on Software Engineering and Methodology, 4(4):365–389, October 1995. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

                                        www.acm.org/pubs/articles/
                                        journals/tosem/1995-4-4/p365-
                                        jackson/p365-jackson.pdf;
                                         http://www.acm.org/pubs/
                                        citations/journals/tosem/
                                        1995-4-4/p365-jackson/.

Jackson:2002:ALO

[Jac02] Daniel Jackson. Alloy: a lightweight object modelling notation. ACM Transactions on Software Engineering and Methodology, 11(2):256–290, 2002. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Jodat:2024:TGS

[JCNS24] Baharin A. Jodat, Abhishek Chandar, Shiva Nejati, and Mehrdad Sabetzadeh. Test generation strategies for building failure models and explaining spurious failures. ACM Transactions on Software Engineering and Methodology, 33 (4):93:1–93:??, May 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3638246.

Jafari:2023:DUS

[JCSA23] Abbas Javan Jafari, Diego Elias Costa, Emad Shihab, and Rabe Abdalkareem. Dependency update strategies and package characteristics. ACM Transactions on Software Engineering and Methodology, 32(6):149:1– 149:??, November 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3603110.

Joblin:2023:HHO

[JEB^{+}23] Mitchell Joblin, Barbara Eckl, Thomas Bock, Angelika Schmid, Janet Siegmund, and Sven Apel. Hierarchical and hybrid organizational structures in opensource software projects: a longitudinal study. ACM Transactions on Software Engineering and Methodology, 32(4): 86:1–86:??, July 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3569949.

Jezequel:1999:RVC

[Jéz99] Jean-Marc Jézéquel. Reifying variants in configuration management. ACM Transactions on Software Engineering and Methodology, 8(3):284–295, July 1999. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1999-8-3/p284-
jezequel/p284-jezequel.pdf;
 http://www.acm.org/pubs/

                                        citations/journals/tosem/
                                        1999-8-3/p284-jezequel/.

Jia:2023:IEF

[JFJ^{+}23] Ang Jia, Ming Fan, Wuxia Jin, Xi Xu, Zhaohui Zhou, Qiyi Tang, Sen Nie, Shi Wu, and Ting Liu. 1-to-1 or 1to-n ? Investigating the effect of function inlining on binary similarity analysis. ACM Transactions on Software Engineering and Methodology, 32 (4):87:1–87:??, July 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3561385.

Jennings:2012:TPA

[JGB12] Paul Jennings, Arka P. Ghosh, and Samik Basu. A two-phase approximation for model checking probabilistic unbounded until properties of probabilistic systems. ACM Transactions on Software Engineering and Methodology, 21(3):18:1–18:??, June 2012. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Jalote:2008:PRR

[JMS08] Pankaj Jalote, Brendan Murphy, and Vibhu Saujanya Sharma. Post-release reliability growth in software products. ACM Transactions on Software Engineering and Methodology, 17(4):17:1–17:??, August 2008. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Jin:2015:ASR

[JO15] Wei Jin and Alessandro Orso.

Automated support for reproducing and debugging field failures. ACM Transactions on Software Engineering and Methodology, 24(4):26:1– 26:??, August 2015. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Jaccheri:1998:ESP

[JPL98] Maria Letizia Jaccheri, Gian Pietro Picco, and Patricia Lago. Eliciting software process models with the E3 language. ACM Transactions on Software Engineering and Methodology, 7(4):368–410, October 1998. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1998-7-4/p368-
jaccheri/p368-jaccheri.pdf;
http://www.acm.org/pubs/
citations/journals/tosem/
1998-7-4/p368-jaccheri/.

Jain:2022:BRD

[JPS22] Ridhi Jain, Rahul Purandare, and Subodh Sharma. BiRD: Race detection in software binaries under relaxed memory models. ACM Transactions on Software Engineering and Methodology, 31(4):58:1–58:??, October 2022. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3498538.

Jin:2023:HCB

[JS23] Xianhao Jin and Francisco Servant. HybridCISave: a combined build and test selection approach in continuous integration. ACM Transactions on Software Engineering and Methodology, 32(4): 93:1–93:??, July 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3576038.

Jeng:1994:SDT

[JW94] Bingchiang Jeng and Elaine J. Weyuker. A simplified domaintesting strategy. ACM Transactions on Software Engineering and Methodology, 3(3):254–270, July 1994. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1994-3-3/p254-
jeng/p254-jeng.pdf;   http:
//www.acm.org/pubs/citations/
journals/tosem/1994-3-3/p254-
jeng/.

Jiang:2024:VBF

[JWC^{+}24] Jiajun Jiang, Yumeng Wang, Junjie Chen, Delin Lv, and Mengjiao Liu. Variable-based fault localization via enhanced decision tree. ACM Transactions on Software Engineering and Methodology, 33(2):41:1– 41:??, February 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3624741.

Jiang:2024:PTF

[JYZ^{+}24] Jiajun Jiang, Junjie Yang, Yingyi Zhang, Zan Wang, Hanmo You, and Junjie Chen. A post-training framework for improving the performance of deep learning models via model transformation. ACM Transactions on Software Engineering and Methodology, 33(3): 61:1–61:??, March 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3630011.

Jiang:2019:RNF

[JZL^{+}19a] He Jiang, Jingxuan Zhang, Xiaochen Li, Zhilei Ren, David Lo, Xindong Wu, and Zhongxuan Luo. Recommending new features from mobile app descriptions. ACM Transactions on Software Engineering and Methodology, 28(4):22:1– 22:??, October 2019. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?id=
3344158.

Jiang:2019:IMA

[JZL^{+}19b] Jian-Min Jiang, Huibiao Zhu, Qin Li, Yongxin Zhao, Zhong Hong, Shi Zhang, and Ping Gong. Isolation modeling and analysis based on mobility. ACM Transactions on Software Engineering and Methodology, 28(2):10:1–10:??, April 2019. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.
                                        acm.org/ft_gateway.cfm?id=
                                        3306606.

Kafali:2020:DSS

[KAS20] Özgür Kafali, Nirav Ajmeri, and Munindar P. Singh. DESEN: Specification of sociotechnical systems via patterns of regulation and control. ACM Transactions on Software Engineering and Methodology, 29(1):7:1– 7:50, February 2020. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3365664.

Kulkarni:2012:GPF

[KAT12] Devdatta Kulkarni, Tanvir Ahmed, and Anand Tripathi. A generative programming framework for context-aware CSCW applications. ACM Transactions on Software Engineering and Methodology, 21(2):11:1– 11:??, March 2012. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Kastner:2012:TCA

[KATS12] Christian Kästner, Sven Apel, Thomas Thüm, and Gunter Saake. Type checking annotationbased product lines. ACM Transactions on Software Engineering and Methodology, 21 (3):14:1–14:??, June 2012. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Kapoor:2007:TCF

[KB07] Kalpesh Kapoor and Jonathan P.

Bowen. Test conditions for fault classes in Boolean specifications. ACM Transactions on Software Engineering and Methodology, 16(3):10:1–10:??, July 2007. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Khatoonabadi:2023:WCU

[KCAS23] Sayedhassan Khatoonabadi, Diego Elias Costa, Rabe Abdalkareem, and Emad Shihab. On wasted contributions: Understanding the dynamics of contributor-abandoned pull requests — a mixed-methods study of 10 large open-source projects. ACM Transactions on Software Engineering and Methodology, 32(1):15:1– 15:??, January 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3530785.

Khatoonabadi:2024:UHS

[KCMS24] Sayedhassan Khatoonabadi, Diego Elias Costa, Suhaib Mujahid, and Emad Shihab. Understanding the helpfulness of stale bot for pull-based development: an empirical study of 20 large open-source projects. ACM Transactions on Software Engineering and Methodology, 33(2):36:1–36:??, February 2024. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3624739.

Karanikolas:2017:EEI

[KDM17] Chris Karanikolas, Grigoris Dimitroulakos, and Konstantinos Masselos. Early evaluation of implementation alternatives of composite data structures toward maintainability. ACM Transactions on Software Engineering and Methodology, 26(2): 8:1–8:??, October 2017. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Karanikolas:2023:SSE

[KDM23] Chris Karanikolas, Grigoris Dimitroulakos, and Konstantinos Masselos. Simulating software evolution to evaluate the reliability of early decisionmaking among design alternatives toward maintainability. ACM Transactions on Software Engineering and Methodology, 32(3):70:1–70:??, July 2023. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3569931.

Krishnamurthi:2007:FIA

[KF07] Shriram Krishnamurthi and Kathi Fisler. Foundations of incremental aspect modelchecking. ACM Transactions on Software Engineering and Methodology, 16(2):7:1–7:??, April 2007. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Kim:2023:ESA

[KFY23] Jinhan Kim, Robert Feldt, and Shin Yoo. Evaluating surprise adequacy for deep learning system testing. ACM Transactions on Software Engineering and Methodology, 32(2): 42:1–42:??, April 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3546947.

Kang:2024:DHM

[KFY24] Sungmin Kang, Robert Feldt, and Shin Yoo. Deceiving humans and machines alike: Search-based test input generation for DNNs using variational autoencoders. ACM Transactions on Software Engineering and Methodology, 33(4): 103:1–103:??, May 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3635706.

Kiezun:2012:HSW

[KGA^{+}12] Adam Kiezun, Vijay Ganesh, Shay Artzi, Philip J. Guo, Pieter Hooimeijer, and Michael D. Ernst. HAMPI: a solver for word equations over strings, regular expressions, and contextfree grammars. ACM Transactions on Software Engineering and Methodology, 21(4):25:1– 25:??, November 2012. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Kiper:1992:STR

[Kip92]  James D. Kiper.  Struc-

tural testing of rule-based expert systems. ACM Transactions on Software Engineering and Methodology, 1(2):168–187, April 1992. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1992-1-2/p168-
kiper/p168-kiper.pdf;   http:
//www.acm.org/pubs/citations/
journals/tosem/1992-1-2/p168-
kiper/.

Kim:2022:PMA

[KJHY22] Jinhan Kim, Juyoung Jeon, Shin Hong, and Shin Yoo. Predictive mutation analysis via the natural language channel in source code. ACM Transactions on Software Engineering and Methodology, 31(4): 73:1–73:??, October 2022. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3510417.

Kaiser:1993:PDI

[KK93] Gail E. Kaiser and Simon M. Kaplan. Parallel and distributed incremental attribute evaluation algorithms for multiuser software development environments. ACM Transactions on Software Engineering and Methodology, 2(1):47–92, January 1993. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/

journals/tosem/1993-2-1/p47-
kaiser/p47-kaiser.pdf;   http:
//www.acm.org/pubs/citations/
journals/tosem/1993-2-1/p47-
kaiser/.

Kramer:2004:CCM

[KK04] Stefan Kramer and Hermann Kaindl. Coupling and cohesion metrics for knowledgebased systems using frames and rules. ACM Transactions on Software Engineering and Methodology, 13(3): 332–358, July 2004. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Keidar:2002:IBT

[KKLS02] Idit Keidar, Roger Khazan, Nancy Lynch, and Alex Shvartsman. An inheritance-based technique for building simulation proofs incrementally. ACM Transactions on Software Engineering and Methodology, 11 (1):63–91, January 2002. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Keller:2022:WYS

[KKP^{+}22] Patrick Keller, Abdoul Kader Kaboré, Laura Plein, Jacques Klein, Yves Le Traon, and Tegawendé F. Bissyandé. What you see is what it means! Semantic representation learning of code based on visualization and transfer learning. ACM Transactions on Software Engineering and Methodology, 31(2): 31:1–31:34, April 2022. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3485135.

Khanfir:2023:IBR

[KKP^{+}23] Ahmed Khanfir, Anil Koyuncu, Mike Papadakis, Maxime Cordy, Tegawende F. Bissyandé, Jacques Klein, and Yves Le Traon. iBiR: Bug-report-driven fault injection. ACM Transactions on Software Engineering and Methodology, 32(2):33:1–33:??, April 2023. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3542946.

Kang:2021:ASM

[KL21] Hong Jin Kang and David Lo. Adversarial specification mining. ACM Transactions on Software Engineering and Methodology, 30(2):16:1–16:40, March 2021. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3424307.

Klint:1993:MEG

[Kli93] P. Klint. A meta-environment for generating programming environments. ACM Transactions on Software Engineering and Methodology, 2(2):176–201, April 1993. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1993-2-2/p176-
klint/p176-klint.pdf;   http:

//www.acm.org/pubs/citations/
journals/tosem/1993-2-2/p176-
klint/.

Klint:2005:TED

[KLV05] Paul Klint, Ralf Lämmel, and Chris Verhoef. Toward an engineering discipline for grammarware. ACM Transactions on Software Engineering and Methodology, 14(3):331–380, July 2005. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Ko:2010:EAW

[KM10] Andrew J. Ko and Brad A. Myers. Extracting and answering why and why not questions about Java program output. ACM Transactions on Software Engineering and Methodology, 20(2):4:1–4:??, August 2010. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Kim:2019:PLR

[KMYK19] Yunho Kim, Seokhyeon Mun, Shin Yoo, and Moonzoo Kim. Precise learn-to-rank fault localization using dynamic and static features of target programs. ACM Transactions on Software Engineering and Methodology, 28(4):23:1–23:??, October 2019. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?id=
3345628.

Kuchta:2018:SSE

[KPC18] Tomasz Kuchta, Hristina Palikareva, and Cristian Cadar. Shadow symbolic execution for testing software patches. ACM Transactions on Software Engineering and Methodology, 27(3): 10:1–10:??, October 2018. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Karre:2024:RMV

[KRM24] Sai Anirudh Karre, Y. Raghu Reddy, and Raghav Mittal. RE methods for virtual reality software product development: a mapping study. ACM Transactions on Software Engineering and Methodology, 33(4): 88:1–88:??, May 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3649595.

Kapur:2020:DES

[KS20] Ritu Kapur and Balwinder Sodhi. A defect estimator for source code: Linking defect reports with programming constructs usage metrics. ACM Transactions on Software Engineering and Methodology, 29(2): 12:1–12:35, April 2020. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3384517.

Kapur:2022:OEE

[KS22] Ritu Kapur and Balwinder Sodhi. OSS effort estimation using software features similarity and developer activitybased metrics. ACM Transactions on Software Engineering and Methodology, 31(2): 33:1–33:35, April 2022. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3485819.

Karam:2008:ULT

[KSD08] Marcel R. Karam, Trevor J.

Smedley, and Sergiu M. Dascalu. Unit-level test adequacy criteria for visual dataflow languages and a testing methodology. ACM Transactions on Software Engineering and Methodology, 18(1):1:1–1:??, September 2008. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Kochanthara:2024:SPS

[KSFC24] Sangeeth Kochanthara, Tajinder Singh, Alexandru Forrai, and Loek Cleophas. Safety of perception systems for automated driving: a case study on Apollo. ACM Transactions on Software Engineering and Methodology, 33(3): 64:1–64:??, March 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3631969.

Kuhn:1999:FCE

[Kuh99] D. Richard Kuhn. Fault classes and error detection capability of specification-based testing. ACM Transactions on Software Engineering and Methodology, 8(4):411–424, October 1999. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

                                        www.acm.org/pubs/articles/
                                        journals/tosem/1999-8-4/p411-
                                        kuhn/p411-kuhn.pdf;   http:
                                        //www.acm.org/pubs/citations/
                                        journals/tosem/1999-8-4/p411-
                                        kuhn/.

LoiolaDeSantana:2024:BAJ

[LASA24] Taijara Loiola De Santana, Paulo Anselmo Da Mota Silveira Neto, Eduardo Santana De Almeida, and Iftekhar Ahmed. Bug analysis in Jupyter Notebook projects: an empirical study. ACM Transactions on Software Engineering and Methodology, 33(4): 101:1–101:??, May 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3641539.

Lizcano:2013:WCA

[LASL13] David Lizcano, Fernando Alonso, Javier Soriano, and Genoveva Lopez. A web-centred approach to end-user software engineering. ACM Transactions on Software Engineering and Methodology, 22(4):36:1–36:??, October 2013. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Layman:2014:MER

[LBZ14] Lucas Layman, Victor R. Basili, and Marvin V. Zelkowitz. A methodology for exposing risk in achieving emergent system properties. ACM Transactions on Software Engineering and Methodology, 23(3):22:1–22:??, May 2014. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Li:2023:CEI

[LCL^{+}23] Xueliang Li, Junyang Chen, Yepang Liu, Kaishun Wu, and John P. Gallagher. Combatting energy issues for mobile applications. ACM Transactions on Software Engineering and Methodology, 32(1):13:1– 13:??, January 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3527851.

Lin:2022:OMS

[LCS^{+}22] Bin Lin, Nathan Cassee, Alexander Serebrenik, Gabriele Bavota, Nicole Novielli, and Michele Lanza. Opinion mining for software development: a systematic literature review. ACM Transactions on Software Engineering and Methodology, 31 (3):38:1–38:41, July 2022. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3490388.

Li:2023:CCG

[LCT^{+}23] Meiziniu Li, Jialun Cao, Yongqiang Tian, Tsz On Li, Ming Wen, and Shing-Chi Cheung. COMET: Coverage-guided model generation for deep learning library testing. ACM Transactions on Software Engineering and Methodology, 32(5):127:1– 127:??, September 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3583566.

Li:2014:SSA

[LCZL14] Yueqi Li, S. C. Cheung, Xiangyu Zhang, and Yepang Liu. Scaling up symbolic analysis by removing z-equivalent states. ACM Transactions on Software Engineering and Methodology, 23(4):34:1–34:??, August 2014. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Leeson:2024:ASS

[LD24] Will Leeson and Matthew B. Dwyer. Algorithm selection for software verification using graph neural networks. ACM Transactions on Software Engineering and Methodology, 33(3): 76:1–76:??, March 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3637225.

Rosa:2013:BPM

[LDUD13] Marcello La Rosa, Marlon Dumas, Reina Uba, and Remco Dijkman. Business process model merging: An approach to business process consolidation. ACM Transactions on Software Engineering and Methodology, 22(2):11:1–11:??, March 2013. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Liu:2023:RDD

[LGC^{+}23] Xuanzhe Liu, Diandian Gu, Zhenpeng Chen, Jinfeng Wen, Zili Zhang, Yun Ma, Haoyu Wang, and Xin Jin. Rise of distributed deep learning training in the big model era: From a software engineering perspective. ACM Transactions on Software Engineering and Methodology, 32(6):156:1– 156:??, November 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3597204.

Liu:2022:RRD

[LGX^{+}22] Chao Liu, Cuiyun Gao, Xin Xia, David Lo, John Grundy, and Xiaohu Yang. On the reproducibility and replicability of deep learning in software engineering. ACM Transactions on Software Engineering and Methodology, 31(1):15:1–15:46, January 2022. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3477535.

Liang:2002:EAA

[LH02] Donglin Liang and Mary Jean Harrold. Equivalence analysis and its application in improving the efficiency of program slicing. ACM Transactions on Software Engineering and Methodology, 11(3):347–383, 2002. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Lhotak:2008:EBC

[LH08] Ondřej Lhoták and Laurie Hendren. Evaluating the benefits of context-sensitive pointsto analysis using a BDD-based implementation. ACM Transactions on Software Engineering and Methodology, 18(1):3:1– 3:??, September 2008. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Liu:2024:GBD

[LHW^{+}24] Jiawei Liu, Yuheng Huang, Zhijie Wang, Lei Ma, Chunrong Fang, Mingzheng Gu, Xufan Zhang, and Zhenyu Chen. Generation-based differential fuzzing for deep learning libraries. ACM Transactions on Software Engineering and Methodology, 33(2):50:1– 50:??, February 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3628159.

Lu:2021:ECR

[LHX21] Jingbo Lu, Dongjie He, and Jingling Xue. Eagle: CFLreachability-based precisionpreserving acceleration of objectsensitive pointer analysis with partial context sensitivity. ACM Transactions on Software Engineering and Methodology, 30 (4):46:1–46:46, July 2021. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3450492.

Liu:2022:AOD

[Liu22] Si Liu. All in one: Design, verification, and implementation of SNOW-optimal read atomic transactions. ACM Transactions on Software Engineering and Methodology, 31 (3):43:1–43:44, July 2022. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3494517.

Li:2020:PNF

[LJL^{+}20] Yangguang Li, Zhen Ming (Jack)

Jiang, Heng Li, Ahmed E. Hassan, Cheng He, Ruirui Huang, Zhengda Zeng, Mian Wang, and Pinan Chen. Predicting node failures in an ultra-large-scale cloud computing platform: an AIOps solution. ACM Transactions on Software Engineering and Methodology, 29(2):13:1– 13:24, April 2020. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3385187.

Lu:2014:RBS

[LK14] Lunjin Lu and Dae-Kyoo Kim.

Required behavior of sequence diagrams: Semantics and conformance. ACM Transactions on Software Engineering and Methodology, 23(2):15:1– 15:??, March 2014. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Laurent:2023:PCT

[LKA^{+}23] Thomas Laurent, Stefan Klikovits, Paolo Arcaini, Fuyuki Ishikawa, and Anthony Ventresque. Parameter coverage for testing of autonomous driving systems under uncertainty. ACM Transactions on Software Engineering and Methodology, 32(3): 58:1–58:??, July 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3550270.

Lauko:2022:VPS

[LKR22] Henrich Lauko, Lukás Korencik, and Petr Rockai. Verification of programs sensitive to heap layout. ACM Transactions on Software Engineering and Methodology, 31(4): 71:1–71:??, October 2022. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3508363.

Lano:2022:MTD

[LKRF22] K. Lano, S. Kolahdouz-Rahimi, and S. Fang. Model transformation development using automated requirements analysis, metamodel matching, and transformation by example. ACM Transactions on Software Engineering and Methodology, 31(2):18:1–18:71, April 2022.

CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3471907.

Louridas:2000:GMR

[LL00] Panagiotis Louridas and Pericles Loucopoulos. A generic model for reflective design. ACM Transactions on Software Engineering and Methodology, 9(2):199–237, April 2000. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/2000-9-2/p199-
louridas/p199-louridas.pdf;
http://www.acm.org/pubs/
citations/journals/tosem/
2000-9-2/p199-louridas/.

Liu:2023:RCN

[LLG23] Eric S. Liu, Dylan A. Lukes, and William G. Griswold. Refactoring in computational notebooks. ACM Transactions on Software Engineering and Methodology, 32(3):77:1–77:??, July 2023. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3576036.

Li:2023:CLE

[LLL^{+}23] Jia Li, Ge Li, Zhuo Li, Zhi Jin, Xing Hu, Kechi Zhang, and Zhiyi Fu. CodeEditor: Learning to edit source code with pretrained models. ACM Transactions on Software Engineering and Methodology, 32(6):143:1– 143:??, November 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3597207.

Lavazza:2023:ESF

[LLLM23] Luigi Lavazza, Angela Locoro, Geng Liu, and Roberto Meli. Estimating software functional size via machine learning. ACM Transactions on Software Engineering and Methodology, 32(5): 114:1–114:??, September 2023. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3582575.

Lyu:2021:ESI

[LLS^{+}21] Yingzhe Lyu, Heng Li, Mohammed Sayagh, Zhen Ming (Jack) Jiang, and Ahmed E. Hassan. An empirical study of the impact of data splitting decisions on the performance of AIOps solutions. ACM Transactions on Software Engineering and Methodology, 30(4): 54:1–54:38, July 2021. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3447876.

Liao:2022:ESI

[LLSM22] Lizhi Liao, Heng Li, Weiyi Shang, and Lei Ma. An empirical study of the impact of hyperparameter tuning and model optimization on the performance properties of deep neural networks. ACM Transactions on Software Engineering and Methodology, 31(3): 53:1–53:40, July 2022. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3506695.

Li:2024:PAL

[LLT^{+}24] Zhihao Li, Chuanyi Li, Ze Tang, Wanhong Huang, Jidong Ge, Bin Luo, Vincent Ng, Ting Wang, Yucheng Hu, and Xiaopeng Zhang. PTM-APIRec: Leveraging pre-trained models of source code in API recommendation. ACM Transactions on Software Engineering and Methodology, 33(3): 72:1–72:??, March 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3632745.

Lin:2022:XTC

[LLZ^{+}22] Zehao Lin, Guodun Li, Jingfeng Zhang, Yue Deng, Xiangji Zeng, Yin Zhang, and Yao Wan. XCode: Towards cross-language code representation with largescale pre-training. ACM Transactions on Software Engineering and Methodology, 31(3): 52:1–52:44, July 2022. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3506696.

xexx:2024:PAP

[LLZ^{+}24] Jia Li, Zhuo Li, Huangzhao Zhang, Ge Li, Zhi Jin, Xing Hu, and Xin Xia. Poison attack and poison detection on deep source code processing models. ACM Transactions on Software Engineering and Methodology, 33(3): 62:1–62:??, March 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3630008.

Lustosa:2024:LVL

[LM24] Andre Lustosa and Tim Menzies. Learning from very little data: On the value of landscape analysis for predicting software project health. ACM Transactions on Software Engineering and Methodology, 33(3): 58:1–58:??, March 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3630252.

Li:2024:HMS

[LMY^{+}24] Wen Li, Austin Marino, Haoran Yang, Na Meng, Li Li, and Haipeng Cai. How are multilingual systems constructed: Characterizing language use and selection in open-source multilingual software. ACM Transactions on Software Engineering and Methodology, 33(3): 63:1–63:??, March 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3631967.

Li:2014:RIP

[LRCS14] Kaituo Li, Christoph Reichenbach, Christoph Csallner, and Yannis Smaragdakis. Residual investigation: Predictive and precise bug detection. ACM Transactions on Software Engineering and Methodology, 24(2): 7:1–7:??, December 2014. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Lyu:2022:TCI

[LRL^{+}22] Yingzhe Lyu, Gopi Krishnan Rajbahadur, Dayi Lin, Boyuan Chen, and Zhen Ming (Jack) Jiang. Towards a consistent interpretation of AIOps models. ACM Transactions on Software Engineering and Methodology, 31(1):16:1–16:38, January 2022. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3488269.

Le:2013:MDF

[LS13] Wei Le and Mary Lou Soffa.

Marple: Detecting faults in path segments using automatically generated analyses. ACM Transactions on Software Engineering and Methodology, 22 (3):18:1–18:??, July 2013. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Lopez:2023:SRS

[LSB^{+}23] Tamara Lopez, Helen Sharp, Arosha Bandara, Thein Tun, Mark Levine, and Bashar Nuseibeh. Security responses in software development. ACM Transactions on Software Engineering and Methodology, 32 (3):64:1–64:??, July 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3563211.

Lee:2024:PSW

[LSBN24] Jaekwon Lee, Seung Yeob Shin, Lionel C. Briand, and Shiva Nejati. Probabilistic safe WCET estimation for weakly hard realtime systems at design stages. ACM Transactions on Software Engineering and Methodology, 33(2):32:1–32:??, February 2024. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3617176.

Lin:2023:CEW

[LSH23] Jiahuei Lin, Mohammed Sayagh, and Ahmed E. Hassan. The co-evolution of the WordPress platform and its plugins. ACM Transactions on Software Engineering and Methodology, 32(1): 19:1–19:??, January 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3533700.

Lin:2023:RRT

[LSM23] Jun-Wei Lin, Navid Salehnamadi, and Sam Malek. Route: Roads not taken in UI testing. ACM Transactions on Software Engineering and Methodology, 32(3):71:1–71:??, July 2023. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3571851.

Lee:2023:EPS

[LSN^{+}23] Jaekwon Lee, Seung Yeob Shin, Shiva Nejati, Lionel Briand, and Yago Isasi Parache. Estimating probabilistic safe WCET ranges of real-time systems at design stages. ACM Transactions on Software Engineering and Methodology, 32(2): 37:1–37:??, April 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3546941.

Louridas:2008:PLS

[LSV08] Panagiotis Louridas, Diomidis Spinellis, and Vasileios Vlachos. Power laws in software. ACM Transactions on Software Engineering and Methodology, 18(1): 2:1–2:??, September 2008. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Liu:2024:HAK

[LSX^{+}24] Jianzhong Liu, Yuheng Shen, Yiru Xu, Hao Sun, and Yu Jiang. Horus: Accelerating kernel fuzzing through efficient Host–VM memory access procedures. ACM Transactions on Software Engineering and Methodology, 33(1): 11:1–11:??, January 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3611665.

Li:2019:UAJ

[LTX19] Yue Li, Tian Tan, and Jingling Xue. Understanding and analyzing Java reflection. ACM Transactions on Software Engineering and Methodology, 28 (2):7:1–7:??, April 2019. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?id=
3295739.

Linares-Vasquez:2018:MOO

[LVBBC^{+}18] Mario Linares-Vásquez, Gabriele Bavota, Carlos Bernal-Cárdenas, Massimiliano Di Penta, Rocco Oliveto, and Denys Poshyvanyk. Multi-objective optimization of energy consumption of GUIs in Android apps. ACM Transactions on Software Engineering and Methodology, 27(3):14:1– 14:??, October 2018. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Liu:2023:FGA

[LWC^{+}23] Xuanzhe Liu, Jinfeng Wen, Zhenpeng Chen, Ding Li, Junkai Chen, Yi Liu, Haoyu Wang, and Xin Jin. FaaSLight: General application-level coldstart latency optimization for function-as-a-service in serverless computing. ACM Transactions on Software Engineering and Methodology, 32(5):119:1– 119:??, September 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3585007.

Li:2024:LAM

[LWD^{+}24] Siyuan Li, Yongpan Wang, Chaopeng Dong, Shouguo Yang, Hong Li, Hao Sun, Zhe Lang, Zuxin Chen, Weijie Wang, Hongsong Zhu, and Limin Sun. LibAM: an area matching framework for detecting thirdparty libraries in binaries. ACM Transactions on Software Engineering and Methodology, 33 (2):52:1–52:??, February 2024. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3625294.

Lopes:2003:HOA

[LWF03] Antónia Lopes, Michel Wermelinger, and José Luiz Fiadeiro. Higher-order architectural connectors. ACM Transactions on Software Engineering and Methodology, 12(1):64– 104, 2003. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Lin:2022:CAC

[LWWM22] Bo Lin, Shangwen Wang, Ming Wen, and Xiaoguang Mao. Context-aware code change embedding for better patch correctness assessment. ACM Transactions on Software Engineering and Methodology, 31 (3):51:1–51:29, July 2022. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3505247.

Liu:2022:CSC

[LXL^{+}22] Chao Liu, Xin Xia, David Lo, Zhiwe Liu, Ahmed E. Hassan, and Shanping Li. Code-Matcher: Searching code based on sequential semantics of important query words. ACM Transactions on Software Engineering and Methodology, 31(1): 12:1–12:37, January 2022. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3465403.

Li:2024:FRF

[LXW^{+}24] Tianlin Li, Xiaofei Xie, Jian Wang, Qing Guo, Aishan Liu, Lei Ma, and Yang Liu. Faire: Repairing fairness of neural networks via neuron condition synthesis. ACM Transactions on Software Engineering and Methodology, 33(1): 21:1–21:??, January 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3617168.

Lau:2005:EFC

[LY05] Man F. Lau and Yuen T. Yu. An extended fault class hierarchy for specification-based testing. ACM Transactions on Software Engineering and Methodology, 14(3):247–276, July 2005. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Liu:2024:APE

[LYL^{+}24] Xuanzhe Liu, Chengxu Yang, Ding Li, Yuhan Zhou, Shaofei Li, Jiali Chen, and Zhenpeng Chen. Adonis: Practical and efficient control flow recovery through OS-level traces. ACM Transactions on Software Engineering and Methodology, 33(1): 2:1–2:??, January 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3607187.

Liu:2014:DWN

[LYYC14] Xiao Liu, Yun Yang, Dong Yuan, and Jinjun Chen. Do we need to handle every temporal violation in scientific workflow systems? ACM Transactions on Software Engineering and Methodology, 23(1):5:1– 5:??, February 2014. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Liu:2024:ADI

[LZF^{+}24] Pei Liu, Yanjie Zhao, Mattia Fazzini, Haipeng Cai, John Grundy, and Li Li. Automatically detecting incompatible Android APIs. ACM Transactions on Software Engineering and Methodology, 33(1): 15:1–15:??, January 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3624737.

Liu:2023:RFP

[LZL^{+}23] Kui Liu, Jingtang Zhang, Li Li, Anil Koyuncu, Dongsun Kim, Chunpeng Ge, Zhe Liu, Jacques Klein, and Tegawendé F. Bissyandé. Reliable fix patterns inferred from static checkers for automated program repair. ACM Transactions on Software Engineering and Methodology, 32(4):96:1–96:??, July 2023. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3579637.

Liu:2024:BAP

[LZS^{+}24] Puzhuo Liu, Yaowen Zheng, Chengnian Sun, Hong Li, Zhi Li, and Limin Sun. Battling against protocol fuzzing: Protecting networked embedded devices from dynamic fuzzers. ACM Transactions on Software Engineering and Methodology, 33(4):108:1–108:??, May 2024. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3641847.

Masri:2014:PCC

[MA14] Wes Masri and Rawad Abou Assi. Prevalence of coincidental correctness and mitigation of its impact on fault localization. ACM Transactions on Software Engineering and Methodology, 23(1):8:1–8:??, February 2014. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Meyers:2022:EPT

[MAKM22] Benjamin S. Meyers, Sultan Fahad Almassari, Brandon N. Keller, and Andrew Meneely. Examining penetration tester behavior in the collegiate penetration testing competition. ACM Transactions on Software Engineering and Methodology, 31(3):55:1–55:25, July 2022. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3514040.

Misu:2024:SJB

[MAL24] Md Rakib Hossain Misu, Rohan Achar, and Cristina V. Lopes. SourcererJBF: a Java build framework for large-scale compilation. ACM Transactions on Software Engineering and Methodology, 33(3): 71:1–71:??, March 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3635710.

Meyers:2007:ESS

[MB07] Timothy M. Meyers and David Binkley. An empirical study of slice-based cohesion and coupling metrics. ACM Transactions on Software Engineering and Methodology, 17(1):2:1– 2:27, December 2007. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Mahmoud:2015:ESR

[MB15] Anas Mahmoud and Gary Bradshaw. Estimating semantic relatedness in source code. ACM Transactions on Software Engineering and Methodology, 25 (1):10:1–10:??, December 2015. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Miranda:2020:TRU

[MB20] Breno Miranda and Antonia Bertolino. Testing relative to usage scope: Revisiting software coverage criteria. ACM Transactions on Software Engineering and Methodology, 29(3): 18:1–18:24, July 2020. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3389126.

McMinn:2009:EEN

[MBH09] Phil McMinn, David Binkley, and Mark Harman. Empirical evaluation of a nesting testability transformation for evolutionary testing. ACM Transactions on Software Engineering and Methodology, 18(3):11:1–11:??, May 2009. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Mills:2017:PQQ

[MBH^{+}17] Chris Mills, Gabriele Bavota, Sonia Haiduc, Rocco Oliveto, Andrian Marcus, and Andrea De Lucia. Predicting query quality for applications of text retrieval to software engineering tasks. ACM Transactions on Software Engineering and Methodology, 26(1):3:1– 3:??, July 2017. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Mohagheghi:2008:EIS

[MC08] Parastoo Mohagheghi and Reidar Conradi. An empirical investigation of software reuse benefits in a large telecom product. ACM Transactions on Software Engineering and Methodology, 17(3):13:1–13:??, June 2008. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Majumder:2023:FES

[MCB^{+}23] Suvodeep Majumder, Joymallya Chakraborty, Gina R. Bai, Kathryn T. Stolee, and Tim Menzies. Fair enough: Searching for sufficient measures of fairness. ACM Transactions on Software Engineering and Methodology, 32(6):134:1– 134:??, November 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3585006.

Ma:2024:FLD

[MCKG24] Suyu Ma, Chunyang Chen, Hourieh Khalajzadeh, and John Grundy. A first look at dark mode in real-world Android apps. ACM Transactions on Software Engineering and Methodology, 33(1): 3:1–3:??, January 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3604607.

Mandrioli:2024:TAC

[MCM24] Claudio Mandrioli, Max Nyberg Carlsson, and Martina Maggio. Testing abstractions for cyberphysical control systems. ACM Transactions on Software Engineering and Methodology, 33(1): 18:1–18:??, January 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3617170.

Memon:2008:ARE

[Mem08] Atif M. Memon. Automatically repairing event sequencebased GUI test suites for regression testing. ACM Transactions on Software Engineering and Methodology, 18(2):4:1– 4:??, November 2008. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Martinez-Fernandez:2022:SEA

[MFBF^{+}22] Silverio Martínez-Fernández, Justus Bogner, Xavier Franch, Marc Oriol, Julien Siebert, Adam Trendowicz, Anna Maria Vollmer, and Stefan Wagner. Software engineering for AIbased systems: a survey. ACM Transactions on Software Engineering and Methodology, 31(2): 37e:1–37e:59, April 2022. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3487043.

Mockus:2002:TCS

[MFH02] Audris Mockus, Roy T. Fielding, and James D. Herbsleb. Two case studies of open source software development: Apache and Mozilla. ACM Transactions on Software Engineering and Methodology, 11(3):309– 346, 2002. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Mattsson:2012:AMA

[MFLL12] Anders Mattsson, Brian Fitzgerald, Björn Lundell, and Brian Lings. An approach for modeling architectural design rules in UML and its application to embedded software. ACM Transactions on Software Engineering and Methodology, 21(2):10:1– 10:??, March 2012. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Mills:2000:KBM

[MG00] Kevin L. Mills and Hassan Gomaa. A knowledge-based method for inferring semantic concepts from visual models of system behavior. ACM Transactions on Software Engineering and Methodology, 9(3):306–337, July 2000. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/2000-9-3/p306-
mills/p306-mills.pdf;   http:
//www.acm.org/pubs/citations/
journals/tosem/2000-9-3/p306-
mills/.

Miles:2011:PMD

[MGMM11] Simon Miles, Paul Groth, Steve Munroe, and Luc Moreau. PrIMe: a methodology for developing provenance-aware applications. ACM Transactions on Software Engineering and Methodology, 20(3):8:1– 8:??, August 2011. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Marin:2013:UFS

[MGP^{+}13] Beatriz Marín, Giovanni Giachetti, Oscar Pastor, Tanja E. J. Vos, and Alain Abran. Using a functional size measurement procedure to evaluate the quality of models in MDD environments. ACM Transactions on Software Engineering and Methodology, 22(3):26:1–26:??, July 2013. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Masso:2022:CTS

[MGP^{+}22] Jhon Masso, Félix García, César Pardo, Francisco J. Pino, and Mario Piattini. A common terminology for software risk management. ACM Transactions on Software Engineering and Methodology, 31(4):59:1– 59:??, October 2022. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3498539.

Mechtaev:2018:TEA

[MGTR18] Sergey Mechtaev, Xiang Gao, Shin Hwei Tan, and Abhik Roychoudhury. Test-equivalence analysis for automatic patch generation. ACM Transactions on Software Engineering and Methodology, 27(4):15:1– 15:??, November 2018. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Maoz:2011:CMS

[MHK11] Shahar Maoz, David Harel, and Asaf Kleinbort. A compiler for multimodal scenarios: Transforming LSCs into AspectJ. ACM Transactions on Software Engineering and Methodology, 20(4):18:1–18:??, September 2011. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Menendez:2021:DFT

[MJS^{+}21] Héctor D. Menéndez, Gunel Jahangirova, Federica Sarro, Paolo Tonella, and David Clark. Diversifying focused testing for unit testing. ACM Transactions on Software Engineering and Methodology, 30(4): 44:1–44:24, July 2021. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3447265.

Muller:2023:CWH

[MKR^{+}23] Katharina Müller, Christian Koch, Dirk Riehle, Michael Stops, and Nikolay Harutyunyan. Challenges of working from home in software development during Covid-19 lockdowns. ACM Transactions on Software Engineering and Methodology, 32(5):111:1– 111:??, September 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3579636.

Mkaouer:2015:MOS

[MKS^{+}15] Wiem Mkaouer, Marouane Kessentini, Adnan Shaout, Patrice Koligheu, Slim Bechikh, Kalyanmoy Deb, and Ali Ouni. Many-objective software remodularization using NSGA-III. ACM Transactions on Software Engineering and Methodology, 24(3):17:1–17:??, May 2015. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Myers:2000:PPU

[ML00] Andrew C. Myers and Barbara Liskov. Protecting privacy using the decentralized label model. ACM Transactions on Software Engineering and Methodology, 9(4):410–442, October 2000. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

                                        www.acm.org/pubs/articles/
                                        journals/tosem/2000-9-4/p410-
                                        myers/p410-myers.pdf;   http:
                                        //www.acm.org/pubs/citations/
                                        journals/tosem/2000-9-4/p410-
                                        myers/.

Monperrus:2013:DMM

[MM13] Martin Monperrus and Mira Mezini. Detecting missing method calls as violations of the majority rule. ACM Transactions on Software Engineering and Methodology, 22(1):7:1– 7:??, February 2013. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Mari:2014:MBS

[MMST14] Federico Mari, Igor Melatti, Ivano Salvo, and Enrico Tronci. Model-based synthesis of control software from system-level formal specifications. ACM Transactions on Software Engineering and Methodology, 23(1): 6:1–6:??, February 2014. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Murphy:1996:LLS

[MN96] Gail C. Murphy and David Notkin. Lightweight lexical source model extraction. ACM Transactions on Software Engineering and Methodology, 5(3):262–292, July 1996. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1996-5-3/p262-
murphy/p262-murphy.pdf;
http://www.acm.org/pubs/citations/
journals/tosem/1996-5-3/p262-
murphy/.

Maddila:2022:CCE

[MNB^{+}22] Chandra Maddila, Nachiappan Nagappan, Christian Bird, Georgios Gousios, and Arie van Deursen. ConE: a concurrent edit detection tool for largescale software development. ACM Transactions on Software Engineering and Methodology, 31(2):22:1–22:26, April 2022. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3478019.

Murphy:1998:ESS

[MNGL98] Gail C. Murphy, David Notkin, William G. Griswold, and Erica S. Lan. An empirical study of static call graph extractors. ACM Transactions on Software Engineering and Methodology, 7(2):158–191, April 1998. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

                                        www.acm.org/pubs/articles/
                                        journals/tosem/1998-7-2/p158-
                                        murphy/p158-murphy.pdf;
                                        http://www.acm.org/pubs/citations/
                                        journals/tosem/1998-7-2/p158-
                                        murphy/.

Masri:2009:MSI

[MP09] Wes Masri and Andy Podgurski. Measuring the strength of information flows in programs. ACM Transactions on Software Engineering and Methodology, 19(2): 5:1–5:??, October 2009. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Murphy:2014:ISI

[MP14] Gail C. Murphy and Mauro Pezzé. Introduction to the special issue International Conference on Software Engineering (ICSE 2012). ACM Transactions on Software Engineering and Methodology, 23(4):28:1– 28:??, August 2014. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Moscato:2014:DTV

[MPF14] Mariano M. Moscato, Carlos G. Lopez Pombo, and Marcelo F. Frias. Dynamite: a tool for the verification of alloy models based on PVS. ACM Transactions on Software Engineering and Methodology, 23(2): 20:1–20:??, March 2014. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Mcmillan:2013:PSR

[MPG^{+}13] Collin Mcmillan, Denys Poshyvanyk, Mark Grechanik, Qing Xie, and Chen Fu. Portfolio: Searching for relevant functions and their usages in millions of lines of code. ACM Transactions on Software Engineering and Methodology, 22(4): 37:1–37:??, October 2013. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Murphy:2006:LCM

[MPR06] Amy L. Murphy, Gian Pietro Picco, and Gruia-Catalin Roman. LIME: a coordination model and middleware supporting mobility of hosts and agents. ACM Transactions on Software Engineering and Methodology, 15(3):279–328, July 2006. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Martignoni:2013:MTC

[MPR^{+}13] Lorenzo Martignoni, Roberto Paleari, Alessandro Reina, Giampaolo Fresi Roglia, and Danilo Bruschi. A methodology for testing CPU emulators. ACM Transactions on Software Engineering and Methodology, 22(4):29:1–29:??, October 2013. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Ma:2021:TSD

[MPT^{+}21] Wei Ma, Mike Papadakis, Anestis Tsakmalis, Maxime Cordy, and Yves Le Traon. Test selection for deep learning systems. ACM Transactions on Software Engineering and Methodology, 30(2): 13:1–13:22, March 2021. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3417330.

Machado:2016:CDD

[MQLR16] Nuno Machado, Daniel Quinta, Brandon Lucia, and Luís Rodrigues. Concurrency debugging with differential schedule projections. ACM Transactions on Software Engineering and Methodology, 25(2):14:1–14:??, May 2016. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

McCann:1999:MMI

[MR99] Peter J. McCann and Gruia-Catalin Roman. Modeling mobile IP in mobile UNITY. ACM Transactions on Software Engineering and Methodology, 8(2):115–146, April 1999. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1999-8-2/p115-
mccann/p115-mccann.pdf;
http://www.acm.org/pubs/citations/
journals/tosem/1999-8-2/p115-
mccann/.

Moser:1997:GED

[MRK^{+}97] L. E. Moser, Y. S. Ramakrishna, G. Kutty, P. M. Melliar-Smith, and L. K. Dillon. A graphical environment for the design of concurrent real-time systems. ACM Transactions on Software Engineering and Methodology, 6(1):31–79, January 1997. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1997-6-1/p31-
moser/p31-moser.pdf;   http:
//www.acm.org/pubs/citations/
journals/tosem/1997-6-1/p31-
moser/.

Ma:2023:PPR

[MRO^{+}23] Fuchen Ma, Meng Ren, Lerong Ouyang, Yuanliang Chen, Juan Zhu, Ting Chen, Yingli Zheng, Xiao Dai, Yu Jiang, and Jiaguang Sun. Pied-piper: Revealing the backdoor threats in Ethereum ERC token contracts. ACM Transactions on Software Engineering and Methodology, 32(3):61:1–61:??, July 2023. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3560264.

Milanova:2005:POS

[MRR05] Ana Milanova, Atanas Rountev, and Barbara G. Ryder. Parameterized object sensitivity for points-to analysis for Java. ACM Transactions on Software Engineering and Methodology, 14(1):1–41, January 2005. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Medvidovic:2002:MSA

[MRRR02] Nenad Medvidovic, David S. Rosenblum, David F. Redmiles, and Jason E. Robbins. Modeling software architectures in the Unified Modeling Language. ACM Transactions on Software Engineering and Methodology, 11(1):2–57, January 2002. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Morzenti:1994:OOL

[MS94] Angelo Morzenti and Pierluigi San Pietro. Object-oriented logical specification of time-critical systems. ACM Transactions on Software Engineering and Methodology, 3(1):56–98, January 1994. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1994-3-1/p56-
morzenti/p56-morzenti.pdf;
 http://www.acm.org/pubs/
citations/journals/tosem/
1994-3-1/p56-morzenti/.

Miller:2003:FTS

[MS03] Tim Miller and Paul Strooper.

A framework and tool support for the systematic testing of model-based specifications. ACM Transactions on Software Engineering and Methodology, 12(4):409–439, October 2003. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Moller:2014:ADC

[MS14] Anders Møller and Mathias Schwarz. Automated detection of client-state manipulation vulnerabilities. ACM Transactions on Software Engineering and Methodology, 23(4):29:1– 29:??, August 2014. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Murukannaiah:2015:PAL

[MS15] Pradeep K. Murukannaiah and Munindar P. Singh. Platys: an active learning framework for place-aware application development and its evaluation. ACM Transactions on Software Engineering and Methodology, 24(3):19:1–19:??, May 2015. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Marijan:2022:IAR

[MS22] Dusica Marijan and Sagar Sen. Industry-academia research collaboration and knowledge cocreation: Patterns and antipatterns. ACM Transactions on Software Engineering and Methodology, 31(3):45:1–45:52, July 2022. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3494519.

Mandrioli:2024:STC

[MSM^{+}24] Claudio Mandrioli, Seung Yeob Shin, Martina Maggio, Domenico Bianculli, and Lionel Briand. Stress testing control loops in cyber-physical systems. ACM Transactions on Software Engineering and Methodology, 33 (2):35:1–35:??, February 2024. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3624742.

Meneely:2012:VSM

[MSW12] Andrew Meneely, Ben Smith, and Laurie Williams. Validating software metrics: a spectrum of philosophies. ACM Transactions on Software Engineering and Methodology, 21 (4):24:1–24:??, November 2012. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Maalej:2014:CPC

[MTRK14] Walid Maalej, Rebecca Tiarks, Tobias Roehm, and Rainer Koschke. On the comprehension of program comprehension. ACM Transactions on Software Engineering and Methodology, 23(4):31:1–31:??, August 2014. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Minsky:2000:LGI

[MU00] Naftaly H. Minsky and Victoria Ungureanu. Law-governed interaction: a coordination and control mechanism for heterogeneous distributed systems. ACM Transactions on Software Engineering and Methodology, 9(3):273–305, July 2000. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/2000-9-3/p273-
minsky/p273-minsky.pdf;
http://www.acm.org/pubs/citations/
journals/tosem/2000-9-3/p273-
minsky/.

Maddila:2023:NAO

[MUB^{+}23] Chandra Maddila, Sai Surya Upadrasta, Chetan Bansal, Nachiappan Nagappan, Georgios Gousios, and Arie van Deursen. Nudge: Accelerating overdue pull requests toward completion. ACM Transactions on Software Engineering and Methodology, 32(2): 35:1–35:??, April 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3544791.

Marin:2007:ICC

[MVM07] Marius Marin, Arie Van Deursen, and Leon Moonen. Identifying Crosscutting Concerns Using Fan-In Analysis. ACM Transactions on Software Engineering and Methodology, 17(1): 3:1–3:37, December 2007. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Mcminn:2015:ETC

[MWK15] Phil Mcminn, Chris J. Wright, and Gregory M. Kapfhammer. The effectiveness of test coverage criteria for relational database schema integrity constraints. ACM Transactions on Software Engineering and Methodology, 25(1):8:1– 8:??, December 2015. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Ma:2021:RTB

[MWP^{+}21] Xiaoxue Ma, Shangru Wu, Ernest Pobee, Xiupei Mei, Hao Zhang, Bo Jiang, and Wing-Kwong Chan. RegionTrack: a trace-based sound and complete checker to debug transactional atomicity violations and non-serializable traces. ACM Transactions on Software Engineering and Methodology, 30(1): 7:1–7:49, January 2021. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3412377.

Minku:2013:SEE

[MY13] Leandro L. Minku and Xin Yao. Software effort estimation as a multiobjective learning problem. ACM Transactions on Software Engineering and Methodology, 22(4):35:1– 35:??, October 2013. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Mamei:2009:PPM

[MZ09] Marco Mamei and Franco Zambonelli. Programming pervasive and mobile computing applications: The TOTA approach. ACM Transactions on Software Engineering and Methodology, 18(4):15:1–15:??, July 2009. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Marculescu:2022:FFR

[MZA22] Bogdan Marculescu, Man Zhang, and Andrea Arcuri. On the faults found in REST APIs by automated test generation. ACM Transactions on Software Engineering and Methodology, 31(3):41:1–41:43, July 2022. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3491038.

Mo:2023:EIC

[MZW^{+}23] Ran Mo, Yao Zhang, Yushuo Wang, Siyuan Zhang, Pu Xiong, Zengyang Li, and Yang Zhao. Exploring the impact of code clones on deep learning software. ACM Transactions on Software Engineering and Methodology, 32(6):153:1–153:??, November 2023. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

URL https://dl.acm.org/
doi/10.1145/3607181.

Neelofar:2024:IES

[NA24] Neelofar Neelofar and Aldeida Aleti. Identifying and explaining safety-critical scenarios for autonomous vehicles via key features. ACM Transactions on Software Engineering and Methodology, 33(4): 94:1–94:??, May 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3640335.

Nass:2023:SBW

[NAF^{+}23] Michel Nass, Emil Alégroth, Robert Feldt, Maurizio Leotta, and Filippo Ricca. Similaritybased Web element localization for robust test automation. ACM Transactions on Software Engineering and Methodology, 32(3):75:1–75:??, July 2023. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3571855.

Nijjar:2015:DMP

[NBB15] Jaideep Nijjar, Ivan Bocić, and Tevfik Bultan. Data model property inference, verification, and repair for Web applications. ACM Transactions on Software Engineering and Methodology, 24(4):25:1–25:??, August 2015. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Nikanjam:2022:AFD

[NBMK22] Amin Nikanjam, Houssem Ben Braiek, Mohammad Mehdi Morovati, and Foutse Khomh. Automatic fault detection for deep learning programs using graph transformations. ACM Transactions on Software Engineering and Methodology, 31(1): 14:1–14:27, January 2022. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3470006.

Nentwich:2003:FCC

[NEFE03] Christian Nentwich, Wolfgang Emmerich, Anthony Finkelstein, and Ernst Ellmer. Flexible consistency checking. ACM Transactions on Software Engineering and Methodology, 12 (1):28–63, 2003. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Nourry:2024:HSF

[NKL^{+}24] Olivier Nourry, Yutaro Kashiwa, Bin Lin, Gabriele Bavota, Michele Lanza, and Yasutaka Kamei. The human side of fuzzing: Challenges faced by developers during fuzzing activities. ACM Transactions on Software Engineering and Methodology, 33(1):14:1–14:??, January 2024. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3611668.

Nguyen:2014:DDI

[NKWF14] Thanhvu Nguyen, Deepak Kapur, Westley Weimer, and Stephanie Forrest. DIG: a dynamic invariant generator for polynomial and array invariants. ACM Transactions on Software Engineering and Methodology, 23(4):30:1– 30:??, August 2014. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Nie:2011:MFC

[NL11] Changhai Nie and Hareton Leung. The minimal failurecausing schema of combinatorial testing. ACM Transactions on Software Engineering and Methodology, 20(4):15:1– 15:??, September 2011. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Naish:2011:MSB

[NLR11] Lee Naish, Hua Jie Lee, and Kotagiri Ramamohanarao. A model for spectra-based software diagnosis. ACM Transactions on Software Engineering and Methodology, 20(3):11:1– 11:??, August 2011. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Nardone:2023:VGB

[NMA^{+}23] Vittoria Nardone, Biruk Muse, Mouna Abidi, Foutse Khomh, and Massimiliano Di Penta. Video game bad smells: What they are and how developers perceive them. ACM Transactions on Software Engineering and Methodology, 32(4): 88:1–88:??, July 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3563214.

Notkin:2007:Ea

[Not07a] David Notkin. Editorial. ACM Transactions on Software Engineering and Methodology, 16(1): 1:1–1:??, February 2007. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Notkin:2007:Eb

[Not07b] David Notkin. Editorial. ACM Transactions on Software Engineering and Methodology, 16 (2):6:1–6:??, April 2007. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Notkin:2007:Ec

[Not07c] David Notkin. Editorial. ACM Transactions on Software Engineering and Methodology, 17(1): 1:1–1:2, December 2007. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Notkin:2008:Ea

[Not08a] David Notkin. Editorial. ACM Transactions on Software Engineering and Methodology, 17 (2):5:1–5:??, April 2008. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Notkin:2008:Eb

[Not08b] David Notkin. Editorial. ACM Transactions on Software Engineering and Methodology, 17 (3):12:1–12:??, June 2008. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Notkin:2009:E

[Not09] David Notkin. Editorial. ACM Transactions on Software Engineering and Methodology, 18 (3):8:1–8:??, May 2009. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Notkin:2010:E

[Not10] David Notkin. Editorial. ACM Transactions on Software Engineering and Methodology, 19(3): 7:1–7:??, January 2010. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Notkin:2012:E

[Not12] David Notkin. Editorial. ACM Transactions on Software Engineering and Methodology, 21 (2):7:1–7:??, March 2012. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Notkin:2013:ELB

[Not13] David Notkin. Editorial — looking back. ACM Transactions on Software Engineering and Methodology, 22(1):1:1– 1:??, February 2013. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Notkin:2008:ISS

[NP08] David Notkin and Mauro Pezzè.

Introduction to the special section from the ACM international symposium on software testing and analysis (IS-STA 2006). ACM Transactions on Software Engineering and Methodology, 17(2): 6:1–6:??, April 2008. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Ngo:2022:ACE

[NPB22] Chanh Duc Ngo, Fabrizio Pastore, and Lionel Briand. Automated, cost-effective, and update-driven app testing. ACM Transactions on Software Engineering and Methodology, 31(4):61:1–61:??, October 2022. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3502297.

Nguyen:2018:UCM

[NWB^{+}18] Khanh Nguyen, Kai Wang, Yingyi Bu, Lu Fang, and Guoqing Xu. Understanding and combating memory bloat in managed data-intensive systems. ACM Transactions on Software Engineering and Methodology, 26(4):12:1–12:41, February 2018. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Nie:2023:CDD

[NWZ^{+}23] Pengbo Nie, Chengcheng Wan, Jiayu Zhu, Ziyi Lin, Yuting Chen, and Zhendong Su. Coverage-directed differential testing of X.509 certificate validation in SSL/TLS implementations. ACM Transactions on Software Engineering and Methodology, 32(1): 3:1–3:??, January 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3510416.

Ni:2022:JTD

[NXL^{+}22] Chao Ni, Xin Xia, David Lo, Xiaohu Yang, and Ahmed E. Hassan. Just-in-time defect prediction on JavaScript projects: a replication study. ACM Transactions on Software Engineering and Methodology, 31(4): 76:1–76:??, October 2022. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3508479.

Oh:2024:FNO

[OBH24] Jeho Oh, Don Batory, and Rubén Heradio. Finding near-optimal configurations in colossal spaces with statistical guarantees. ACM Transactions on Software Engineering and Methodology, 33(1): 7:1–7:??, January 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3611663.

Ore:2021:EST

[ODE21] John-Paul Ore, Carrick Detweiler, and Sebastian Elbaum. An empirical study on type annotations: Accuracy, speed, and suggestion effectiveness. ACM Transactions on Software Engineering and Methodology, 30(2): 20:1–20:29, March 2021. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3439775.

Ouyang:2009:BPM

[ODV^{+}09] Chun Ouyang, Marlon Dumas, Wil M. P. Van Der Aalst, Arthur H. M. Ter Hofstede, and Jan Mendling. From business process models to processoriented software systems. ACM Transactions on Software Engineering and Methodology, 19 (1):2:1–2:??, August 2009. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Offutt:1992:IST

[Off92] A. Jefferson Offutt. Investigations of the software testing coupling effect. ACM Transactions on Software Engineering and Methodology, 1(1):5–20, January 1992. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

                                        www.acm.org/pubs/articles/
                                        journals/tosem/1992-1-1/p5-
                                        offutt/p5-offutt.pdf;   http:
                                        //www.acm.org/pubs/citations/
                                        journals/tosem/1992-1-1/p5-
                                        offutt/.

Oakes:2024:BDS

[OFS24] Bentley James Oakes, Michalis Famelis, and Houari Sahraoui. Building domain-specific machine learning workflows: a conceptual framework for the state of the practice. ACM Transactions on Software Engineering and Methodology, 33 (4):91:1–91:??, May 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3638243.

Osterweil:2005:E

[OGKW05] Leon Osterweil, Carlo Ghezzi, Jeff Kramer, and Alexander Wolf. Editorial. ACM Transactions on Software Engineering and Methodology, 14(4):381– 382, October 2005. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Ostertag:1992:CSR

[OHDB92] Eduardo Ostertag, James Hendler, Rubén Prieto Díaz, and Christine Braun. Computing similarity in a reuse library system: an AI-based approach. ACM Transactions on Software Engineering and Methodology, 1(3):205–228, July 1992. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1992-1-3/p205-
ostertag/p205-ostertag.pdf;
http://www.acm.org/pubs/
citations/journals/tosem/
1992-1-3/p205-ostertag/.

Ouni:2016:MCC

[OKS^{+}16] Ali Ouni, Marouane Kessentini, Houari Sahraoui, Katsuro Inoue, and Kalyanmoy Deb. Multi-criteria code refactoring using search-based software engineering: an industrial case study. ACM Transactions on Software Engineering and Methodology, 25(3):23:1– 23:??, August 2016. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Oni:2022:AUR

[OL22] Olawole Oni and Emmanuel Letier. Analyzing uncertainty in release planning: a method and experiment for fixed-date release cycles. ACM Transactions on Software Engineering and Methodology, 31(2): 24:1–24:39, April 2022. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3490487.

Offutt:1996:EDS

[OLR^{+}96] A. Jefferson Offutt, Ammei Lee, Gregg Rothermel, Roland H. Untch, and Christian Zapf. An experimental determination of sufficient mutant operators. ACM Transactions on Software Engineering and Methodology, 5(2):99–118, April 1996. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1996-5-2/p99-
offutt/p99-offutt.pdf;   http:
//www.acm.org/pubs/citations/
journals/tosem/1996-5-2/p99-
offutt/.

Olender:1992:ISA

[OO92] Kurt M. Olender and Leon J. Osterweil. Interprocedural static analysis of sequencing constraints. ACM Transactions on Software Engineering and Methodology, 1(1):21–52, January 1992. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1992-1-1/p21-
olender/p21-olender.pdf;
http://www.acm.org/pubs/citations/
journals/tosem/1992-1-1/p21-
olender/.

Ohrndorf:2021:HBM

[OPK^{+}21] Manuel Ohrndorf, Christopher Pietsch, Udo Kelter, Lars Grunske, and Timo Kehrer. History-based model repair recommendations. ACM Transactions on Software Engineering and Methodology, 30(2): 15:1–15:46, March 2021. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3419017.

Ojdanic:2023:MTE

[OSD^{+}23] Milos Ojdanic, Ezekiel Soremekun, Renzo Degiovanni, Mike Papadakis, and Yves Le Traon. Mutation testing in evolving systems: Studying the relevance of mutants to code evolution. ACM Transactions on Software Engineering and Methodology, 32(1):14:1–14:??, January 2023. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3530786.

Orso:2004:CDD

[OSH04] Alessandro Orso, Saurabh Sinha, and Mary Jean Harrold. Classifying data dependences in the presence of pointers for program comprehension, testing, and debugging. ACM Transactions on Software Engineering and Methodology, 13 (2):199–239, April 2004. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Ostroff:1999:CRD

[Ost99] Jonathan S. Ostroff. Composition and refinement of discrete real-time systems. ACM Transactions on Software Engineering and Methodology, 8(1):1–48, January 1999. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1999-8-1/p1-
ostroff/p1-ostroff.pdf;

                                        http://www.acm.org/pubs/citations/
                                        journals/tosem/1999-8-1/p1-
                                        ostroff/.

Pizzolotto:2024:MDB

[PBC24] Davide Pizzolotto, Stefano Berlato, and Mariano Ceccato. Mitigating debugger-based attacks to Java applications with self-debugging. ACM Transactions on Software Engineering and Methodology, 33(4): 107:1–107:??, May 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3631971.

Ponge:2010:AAT

[PBCT10] Julien Ponge, Boualem Benatallah, Fabio Casati, and Farouk Toumani. Analysis and applications of timed service protocols. ACM Transactions on Software Engineering and Methodology, 19(4):11:1–11:??, April 2010. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Paige:2007:MBM

[PBO07] Richard F. Paige, Phillip J. Brooke, and Jonathan S. Ostroff. Metamodel-based model conformance and multiview consistency checking. ACM Transactions on Software Engineering and Methodology, 16(3): 11:1–11:??, July 2007. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Pavese:2016:LME

[PBU16] Esteban Pavese, Víctor Braberman, and Sebastian Uchitel. Less is more: Estimating probabilistic rewards over partial system explorations. ACM Transactions on Software Engineering and Methodology, 25(2): 16:1–16:??, May 2016. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Patnaik:2023:SSS

[PDHR23] Nikhil Patnaik, Andrew Dwyer, Joseph Hallett, and Awais Rashid. SLR: From Saltzer and Schroeder to 2021 … 47 years of research on the development and validation of security API recommendations. ACM Transactions on Software Engineering and Methodology, 32 (3):60:1–60:??, July 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3561383.

Petrenko:1997:CRT

[Pet97]  Alexandre Petrenko.  Com-

ments on “A reduced test suite for protocol conformance testing”. ACM Transactions on Software Engineering and Methodology, 6(3):329–331, July 1997. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1997-6-3/p329-
petrenko/p329-petrenko.pdf;
 http://www.acm.org/pubs/

                                        citations/journals/tosem/
                                        1997-6-3/p329-petrenko/. See

[Ber94].

Pezze:2019:Ea

[Pez19a] Mauro Pezzè. Editorial. ACM Transactions on Software Engineering and Methodology, 28 (2):6:1–6:??, April 2019. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.
                                        acm.org/ft_gateway.cfm?id=
                                        3317953.

Pezze:2019:Eb

[Pez19b] Mauro Pezzé. Editorial. ACM Transactions on Software Engineering and Methodology, 28(3): 12:1–12:??, August 2019. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.
                                        acm.org/ft_gateway.cfm?id=
                                        3339833.

Pezze:2019:Ec

[Pez19c] Mauro Pezzè. Editorial. ACM Transactions on Software Engineering and Methodology, 28(4): 19:1–19:??, October 2019. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.
                                        acm.org/ft_gateway.cfm?id=
                                        3363297.

Pezze:2019:EIE

[Pez19d] Mauro Pezzè. Editorial from the incoming Editor-in-Chief. ACM Transactions on Software Engineering and Methodology, 28(1): 1:1–1:??, February 2019. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Pezze:2020:Ea

[Pez20a] Mauro Pezzè. Editorial. ACM Transactions on Software Engineering and Methodology, 29 (2):8e:1, April 2020. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3383775.

Pezze:2020:Eb

[Pez20b] Mauro Pezzè. Editorial. ACM Transactions on Software Engineering and Methodology, 29 (3):14e:1, July 2020. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3402931.

Pezze:2021:E

[Pez21] Mauro Pezzè. Editorial. ACM Transactions on Software Engineering and Methodology, 30 (3):26e:1, May 2021. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3450737.

Pezze:2022:ERP

[Pez22] Mauro Pezzè. Editorial: a retrospective and prospective reflection. ACM Transactions on Software Engineering and Methodology, 31(2):18e:1–18e:2, April 2022. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3523278.

Pezze:2023:EAS

[Pez23] Mauro Pezzè. Editorial: Announcing six TOSEM issues per year. ACM Transactions on Software Engineering and Methodology, 32(1): 1:1–1:??, January 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3583569.

Pezze:2024:EII

[Pez24a] Mauro Pezzè. Editorial: ICSE and the incredible contradictions of software engineering. ACM Transactions on Software Engineering and Methodology, 33(4):83:1–83:??, May 2024. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3656301.

Pezze:2024:ETF

[Pez24b] Mauro Pezzè. Editorial: Toward the future with eight issues per year. ACM Transactions on Software Engineering and Methodology, 33(1): 1:1–1:??, January 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3637444.

Poshyvanyk:2012:CLU

[PGM12] Denys Poshyvanyk, Malcom Gethers, and Andrian Marcus. Concept location using formal concept analysis and information retrieval. ACM Transactions on Software Engineering and Methodology, 21(4):23:1– 23:??, November 2012. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Pilgun:2020:FGC

[PGZ^{+}20] Aleksandr Pilgun, Olga Gadyatskaya, Yury Zhauniarovich, Stanislav Dashevskyi, Artsiom Kushniarou, and Sjouke Mauw. Fine-grained code coverage measurement in automated black-box Android testing. ACM Transactions on Software Engineering and Methodology, 29(4):23:1–23:35, October 2020. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3395042.

Pant:2024:EAA

[PHS^{+}24] Aastha Pant, Rashina Hoda, Simone V. Spiegler, Chakkrit Tantithamthavorn, and Burak Turhan. Ethics in the age of AI: an analysis of AI practitioners’ awareness and challenges. ACM Transactions on Software Engineering and Methodology, 33(3): 80:1–80:??, March 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3635715.

Payton:2010:SSA

[PJRR10] Jamie Payton, Christine Julien, Gruia-Catalin Roman, and Vasanth Rajamani. Semantic selfassessment of query results in dynamic environments. ACM Transactions on Software Engineering and Methodology, 19 (4):12:1–12:??, April 2010. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Parry:2022:SFT

[PKHM22] Owain Parry, Gregory M. Kapfhammer, Michael Hilton, and Phil McMinn. A survey of flaky tests. ACM Transactions on Software Engineering and Methodology, 31(1): 17:1–17:74, January 2022. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3476105.

Pan:2024:EEA

[PKMY24] Weifeng Pan, Marouane Kessentini, Hua Ming, and Zijiang Yang. EASE: an effort-aware extension of unsupervised key class identification approaches. ACM Transactions on Software Engineering and Methodology, 33(4):84:1–84:??, May 2024. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3635714.

Proksch:2015:ICC

[PLM15] Sebastian Proksch, Johannes Lerch, and Mira Mezini. Intelligent code completion with Bayesian networks. ACM Transactions on Software Engineering and Methodology, 25(1): 3:1–3:??, December 2015. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Perez:2023:HQM

[PLMC23] Francisca Pérez, Raúl Lapeña, Ana Marcén, and Carlos Cetina. How the quality of maintenance tasks is affected by criteria for selecting engineers for collaboration. ACM Transactions on Software Engineering and Methodology, 32(3): 63:1–63:??, July 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3561384.

Pan:2023:PWT

[PLP^{+}23] Minxue Pan, Yifei Lu, Yu Pei, Tian Zhang, and Xuandong Li. Preference-wise testing of Android apps via test amplification. ACM Transactions on Software Engineering and Methodology, 32(1): 4:1–4:??, January 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3511804.

Pantiuchina:2022:WDD

[PLZ^{+}22] Jevgenija Pantiuchina, Bin Lin, Fiorella Zampetti, Massimiliano Di Penta, Michele Lanza, and Gabriele Bavota. Why do developers reject refactorings in open-source projects? ACM Transactions on Software Engineering and Methodology, 31(2): 23:1–23:23, April 2022. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3487062.

Podgurski:1999:ESR

[PMM^{+}99] Andy Podgurski, Wassim Masri, Yolanda McCleese, Francis G. Wolff, and Charles Yang. Estimation of software reliability by stratified sampling. ACM Transactions on Software Engineering and Methodology, 8(3):263–283, July 1999. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

                                        www.acm.org/pubs/articles/
                                        journals/tosem/1999-8-3/p263-
                                        podgurski/p263-podgurski.
                                        pdf;   http://www.acm.org/
                                        pubs/citations/journals/tosem/
                                        1999-8-3/p263-podgurski/.

Pradella:2013:BSC

[PMS13] Matteo Pradella, Angelo Morzenti, and Pierluigi San Pietro. Bounded satisfiability checking of metric temporal logic specifications. ACM Transactions on Software Engineering and Methodology, 22(3):20:1–20:??, July 2013. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Pons:2002:TAC

[Pon02] Alexander P. Pons. Temporal abstract classes and virtual temporal specifications for realtime systems. ACM Transactions on Software Engineering and Methodology, 11(3):291– 308, 2002. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Pacheco:2023:MTD

[PORH23] Michael Pacheco, Gustavo Oliva, Gopi Krishnan Rajbahadur, and Ahmed Hassan. Is my transaction done yet? An empirical study of transaction processing times in the Ethereum blockchain platform. ACM Transactions on Software Engineering and Methodology, 32(3):59:1–59:??, July 2023. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3549542.

Podgurski:1993:RRS

[PP93] Andy Podgurski and Lynn Pierce. Retrieving reusable software by sampling behavior. ACM Transactions on Software Engineering and Methodology, 2(3):286–303, July 1993. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1993-2-3/p286-
podgurski/p286-podgurski.
pdf;   http://www.acm.org/
pubs/citations/journals/tosem/
1993-2-3/p286-podgurski/.

Parisi-Presicce:1994:ATC

[PPP94] Francesco Parisi-Presicce and Alfonso Pierantonio. An algebraic theory of class specification. ACM Transactions on Software Engineering and Methodology, 3(2):166–199, April 1994. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1994-3-2/p166-
parisi-presicce/p166-parisi-
presicce.pdf;   http://www.
acm.org/pubs/citations/journals/
tosem/1994-3-2/p166-parisi-
presicce/.

Picco:2001:RAC

[PRM01] Gian Pietro Picco, Gruia-Catalin Roman, and Peter J. McCann. Reasoning about code mobility with mobile UNITY. ACM Transactions on Software Engineering and Methodology, 10(3):338–395, July 2001. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Porter:1998:USV

[PSMV98] Adam Porter, Harvey Siy, Audris Mockus, and Lawrence Votta. Understanding the sources of variation in software inspections. ACM Transactions on Software Engineering and Methodology, 7(1):41–79, January 1998. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1998-7-1/p41-
porter/p41-porter.pdf;   http:
//www.acm.org/pubs/citations/
journals/tosem/1998-7-1/p41-
porter/.

Perry:2001:PCL

[PSV01] Dewayne E. Perry, Harvey P.

Siy, and Lawrence G. Votta. Parallel changes in large-scale software development: an observational case study. ACM Transactions on Software Engineering and Methodology, 10 (3):308–337, July 2001. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Polyvyanyy:2020:MPR

[PSW^{+}20] Artem Polyvyanyy, Andreas Solti, Matthias Weidlich, Claudio Di Ciccio, and Jan Mendling. Monotone precision and recall measures for comparing executions and specifications of dynamic systems. ACM Transactions on Software Engineering and Methodology, 29(3):17:1– 17:41, July 2020. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3387909.

Paulweber:2021:SIT

[PSZ21] Philipp Paulweber, Georg Simhandl, and Uwe Zdun. Specifying with interface and trait abstractions in abstract state machines: a controlled experiment. ACM Transactions on Software Engineering and Methodology, 30(4):47:1–47:29, July 2021. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3450968.

Pezze:1995:GMR

[PTY95] Mauro Pezzè, Richard N. Taylor, and Michal Young. Graph models for reachability analysis of concurrent programs. ACM Transactions on Software Engineering and Methodology, 4(2):171–213, April 1995. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

                                        www.acm.org/pubs/articles/
                                        journals/tosem/1995-4-2/p171-
                                        pezze/p171-pezze.pdf;   http:
                                        //www.acm.org/pubs/citations/
                                        journals/tosem/1995-4-2/p171-
                                        pezze/.

Polyvyanyy:2017:IDP

[PVHW17] Artem Polyvyanyy, Wil M. P. Van Der Aalst, Arthur H. M. Ter Hofstede, and Moe T. Wynn. Impact-driven process model repair. ACM Transactions on Software Engineering and Methodology, 25(4):28:1– 28:??, May 2017. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Partachi:2023:AMI

[PWB23] Profir-Petru Pârtachi, David R. White, and Earl T. Barr. Aide-mémoire: Improving a project’s collective memory via pull request-issue links. ACM Transactions on Software Engineering and Methodology, 32 (2):32:1–32:??, April 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3542937.

Pohl:1999:PTP

[PWD^{+}99] Klaus Pohl, Klaus Weidenhaupt, Ralf Dömges, Peter Haumer, Matthias Jarke, and Ralf Klamma. PRIME — toward process-integrated modeling environments: 1. ACM Transactions on Software Engineering and Methodology, 8 (4):343–410, October 1999. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1999-8-4/p343-
pohl/p343-pohl.pdf;   http:
//www.acm.org/pubs/citations/
journals/tosem/1999-8-4/p343-
pohl/.

Pan:2014:GTG

[PWX14] Kai Pan, Xintao Wu, and Tao Xie. Guided test generation for database applications via synthesized database interactions. ACM Transactions on Software Engineering and Methodology, 23(2):12:1–12:??, March 2014. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Palepu:2017:DDS

[PXJ17] Vijay Krishna Palepu, Guoqing Xu, and James A. Jones. Dynamic dependence summaries. ACM Transactions on Software Engineering and Methodology, 25(4):30:1–30:??, May 2017. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Pantiuchina:2020:WDR

[PZS^{+}20] Jevgenija Pantiuchina, Fiorella Zampetti, Simone Scalabrino, Valentina Piantadosi, Rocco Oliveto, Gabriele Bavota, and Massimiliano Di Penta. Why developers refactor source code: a mining-based study. ACM Transactions on Software Engineering and Methodology, 29(4): 29:1–29:30, October 2020. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3408302.

Qin:2023:NTE

[QHM^{+}23] Shisong Qin, Fan Hu, Zheyu Ma, Bodong Zhao, Tingting Yin, and Chao Zhang. NSFuzz: Towards efficient and stateaware network service fuzzing. ACM Transactions on Software Engineering and Methodology, 32(6):160:1–160:??, November 2023. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3580598.

Qi:2013:PEB

[QNR13] Dawei Qi, Hoang D. T. Nguyen, and Abhik Roychoudhury. Path exploration based on symbolic output. ACM Transactions on Software Engineering and Methodology, 22(4):32:1– 32:??, October 2013. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Qi:2012:DAD

[QRLV12] Dawei Qi, Abhik Roychoudhury, Zhenkai Liang, and Kapil Vaswani. DARWIN: an approach to debugging evolving programs. ACM Transactions on Software Engineering and Methodology, 21(3):19:1–19:??, June 2012. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Qi:2024:RCN

[QSZG24] Binhang Qi, Hailong Sun, Hongyu Zhang, and Xiang Gao. Reusing convolutional neural network models through modularization and composition. ACM Transactions on Software Engineering and Methodology, 33(3):70:1–70:??, March 2024. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3632744.

Queralt:2012:VVU

[QT12] Anna Queralt and Ernest Teniente. Verification and validation of UML conceptual schemas with OCL constraints. ACM Transactions on Software Engineering and Methodology, 21(2):13:1–13:??, March 2012. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Qi:2023:ABL

[QWG^{+}23] Hua Qi, Zhijie Wang, Qing Guo, Jianlang Chen, Felix Juefei-Xu, Fuyuan Zhang, Lei Ma, and Jianjun Zhao. ArchRepair: Block-level architectureoriented repairing for deep neural networks. ACM Transactions on Software Engineering and Methodology, 32(5):129:1– 129:??, September 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3585005.

Rothermel:2001:MTS

[RBL^{+}01] Gregg Rothermel, Margaret Burnett, Lixin Li, Christopher Dupuis, and Andrei Sheretov. A methodology for testing spreadsheets. ACM Transactions on Software Engineering and Methodology, 10 (1):110–147, January 2001. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/2001-10-1/
p110-rothermel/p110-rothermel.
pdf;   http://www.acm.org/
pubs/citations/journals/tosem/
2001-10-1/p110-rothermel/.

Robol:2023:CVM

[RBPG23] Marco Robol, Travis D. Breaux, Elda Paja, and Paolo Giorgini. Consent verification monitoring. ACM Transactions on Software Engineering and Methodology, 32(1):2:1–2:??, January 2023. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3490754.

Rombaut:2023:TNS

[RCAH23] Benjamin Rombaut, Filipe R.

Cogo, Bram Adams, and Ahmed E. Hassan. There’s no such thing as a free lunch: Lessons learned from exploring the overhead introduced by the greenkeeper dependency bot in Npm. ACM Transactions on Software Engineering and Methodology, 32(1): 11:1–11:??, January 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3522587.

Radoi:2015:ETS

[RD15] Cosmin Radoi and Danny Dig.

Effective techniques for static race detection in Java parallel loops. ACM Transactions on Software Engineering and Methodology, 24(4):24:1– 24:??, August 2015. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Reiss:1999:DE

[Rei99] Steven P. Reiss. The Desert environment. ACM Transactions on Software Engineering and Methodology, 8(4):297–342, October 1999. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1999-8-4/p297-
reiss/p297-reiss.pdf;   http:
//www.acm.org/pubs/citations/
journals/tosem/1999-8-4/p297-
reiss/.

Rothermel:2004:TSC

[REM^{+}04] Gregg Rothermel, Sebastian Elbaum, Alexey G. Malishevsky, Praveen Kallakuri, and Xuemei Qiu. On test suite composition and cost-effective regression testing. ACM Transactions on Software Engineering and Methodology, 13(3): 277–331, July 2004. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Rolland:2024:ASN

[RFDS24] Knut H. Rolland, Brian Fitzgerald, Torgeir Dingsøyr, and Klaas-Jan Stol. Acrobats and safety nets: Problematizing large-scale agile software development. ACM Transactions on Software Engineering and Methodology, 33(2):33:1– 33:??, February 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3617169.

Ramirez:2023:TIA

[RFR23] Aurora Ramírez, Robert Feldt, and José Raúl Romero. A taxonomy of information attributes for test case prioritisation: Applicability, machine learning. ACM Transactions on Software Engineering and Methodology, 32(1):21:1–21:??, January 2023. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3511805.

Rigby:2014:PRO

[RGCS14] Peter C. Rigby, Daniel M.

German, Laura Cowen, and Margaret-Anne Storey. Peer review on open-source software projects: Parameters, statistical models, and theory. ACM Transactions on Software Engineering and Methodology, 23(4): 35:1–35:??, August 2014. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Roychoudhury:2012:SMS

[RGS12] Abhik Roychoudhury, Ankit Goel, and Bikram Sengupta. Symbolic message sequence charts. ACM Transactions on Software Engineering and Methodology, 21(2):12:1–12:??, March 2012. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Rothermel:1997:SER

[RH97] Gregg Rothermel and Mary Jean Harrold. A safe, efficient regression test selection technique. ACM Transactions on Software Engineering and Methodology, 6(2):173–210, April 1997. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1997-6-2/p173-
rothermel/p173-rothermel.
pdf;   http://www.acm.org/
pubs/citations/journals/tosem/
1997-6-2/p173-rothermel/.

Russo:2024:UDWa

[RHV24] Daniel Russo, Paul H. P. Hanel, and Niels Van Berkel. Understanding developers well-being and productivity: a 2-year longitudinal analysis during the COVID-19 pandemic. ACM Transactions on Software Engineering and Methodology, 33(3): 57:1–57:??, March 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3638244.

Russo:2024:UDWb

[RHvB24] Daniel Russo, Paul H. P. Hanel, and Niels van Berkel. Understanding developers wellbeing and productivity: a 2year longitudinal analysis during the COVID-19 Pandemic-RCR report. ACM Transactions on Software Engineering and Methodology, 33(3): 82:1–82:??, March 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3640338.

Reuling:2019:AWP

[RKBL19] Dennis Reuling, Udo Kelter, Johannes Bürdek, and Malte Lochau. Automated N-way program merging for facilitating family-based analyses of variant-rich software. ACM Transactions on Software Engineering and Methodology, 28(3): 13:1–13:??, August 2019. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?id=
3313789.

Robillard:2003:SAS

[RM03] Martin P. Robillard and Gail C.

Murphy. Static analysis to support the evolution of exception structure in objectoriented systems. ACM Transactions on Software Engineering and Methodology, 12(2):191– 221, April 2003. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Robillard:2007:RCS

[RM07] Martin P. Robillard and Gail C.

Murphy. Representing concerns in source code. ACM Transactions on Software Engineering and Methodology, 16(1):3:1– 3:??, February 2007. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Romdhana:2022:DRL

[RMCT22] Andrea Romdhana, Alessio Merlo, Mariano Ceccato, and Paolo Tonella. Deep reinforcement learning for black-box testing of Android apps. ACM Transactions on Software Engineering and Methodology, 31(4): 65:1–65:??, October 2022. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3502868.

Roman:1997:MUR

[RMP97] Gruia-Catalin Roman, Peter J.

McCann, and Jerome Y. Plun.

Mobile UNITY: reasoning and specification in mobile computing. ACM Transactions on Software Engineering and Methodology, 6(3):250–282, July 1997. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1997-6-3/p250-
roman/p250-roman.pdf;   http:
//www.acm.org/pubs/citations/
journals/tosem/1997-6-3/p250-
roman/.

Riesco:2018:PII

[RO18] Adrián Riesco and Kazuhiro Ogata. Prove it! Inferring formal proof scripts from CafeOBJ proof scores. ACM Transactions on Software Engineering and Methodology, 27 (2):6:1–6:??, July 2018. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Robillard:2008:TAS

[Rob08] Martin P. Robillard. Topology analysis of software dependencies. ACM Transactions on Software Engineering and Methodology, 17(4):18:1– 18:??, August 2008. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Rosenblum:2013:ELF

[Ros13a] David S. Rosenblum. Editorial — looking forward. ACM Transactions on Software Engineering and Methodology, 22(1): 2:1–2:??, February 2013. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Rosenblum:2013:MDN

[Ros13b] David S. Rosenblum. In memoriam: David Notkin (1955– 2013). ACM Transactions on Software Engineering and Methodology, 22(3):16:1–16:??, July 2013. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Rosenblum:2014:Ea

[Ros14a] David S. Rosenblum. Editorial.

ACM Transactions on Software Engineering and Methodology, 23(1):1:1–1:??, February 2014. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Rosenblum:2014:Eb

[Ros14b] David S. Rosenblum. Editorial.

ACM Transactions on Software Engineering and Methodology, 23(2):11:1–11:??, March 2014. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Rosenblum:2014:E

[Ros14c] David S. Rosenblum. Editorial.

ACM Transactions on Software Engineering and Methodology, 23(4):27:1–27:??, August 2014. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Rosenblum:2016:E

[Ros16] David S. Rosenblum. Editorial. ACM Transactions on Software Engineering and Methodology, 25(2):11:1–11:??, May 2016. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Rosenblum:2017:E

[Ros17] David S. Rosenblum. Editorial. ACM Transactions on Software Engineering and Methodology, 26(2):6:1–6:??, October 2017. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Rosenblum:2018:Ea

[Ros18a] David S. Rosenblum. Editorial. ACM Transactions on Software Engineering and Methodology, 27(1):1:1–1:??, June 2018. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Rosenblum:2018:Eb

[Ros18b] David S. Rosenblum. Editorial. ACM Transactions on Software Engineering and Methodology, 27(3):9:1–9:??, October 2018. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Rosenblum:2019:FEO

[Ros19] David S. Rosenblum. Farewell editorial from the outgoing Editor-in-Chief. ACM Transactions on Software Engineering and Methodology, 28(1):1:1– 1:??, February 2019. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Rauf:2022:CAS

[RPT^{+}22] Irum Rauf, Marian Petre, Thein Tun, Tamara Lopez, Paul Lunn, Dirk Van Der Linden, John Towse, Helen Sharp, Mark Levine, Awais Rashid, and Bashar Nuseibeh. The case for adaptive security interventions. ACM Transactions on Software Engineering and Methodology, 31(1):9:1–9:52, January 2022. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3471930.

Rahman:2023:SRA

[RR23] Mohammad Masudur Rahman and Chanchal K. Roy. A systematic review of automated query reformulations in source code search. ACM Transactions on Software Engineering and Methodology, 32(6):159:1– 159:??, November 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3607179.

Rahman:2021:SSA

[RRPW21] Akond Rahman, Md Rayhanur Rahman, Chris Parnin, and Laurie Williams. Security smells in Ansible and Chef scripts: a replication study. ACM Transactions on Software Engineering and Methodology, 30(1): 3:1–3:31, January 2021. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3408897.

Rajan:2009:UAO

[RS09] Hridesh Rajan and Kevin J. Sullivan. Unifying aspect- and object-oriented design. ACM Transactions on Software Engineering and Methodology, 19 (1):3:1–3:??, August 2009. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Ryder:2005:ISE

[RSB05] Barbara G. Ryder, Mary Lou Soffa, and Margaret Burnett. The impact of software engineering research on modern programming languages. ACM Transactions on Software Engineering and Methodology, 14(4): 431–477, October 2005. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Rosa:2023:WQA

[RSBO23] Giovanni Rosa, Simone Scalabrino, Gabriele Bavota, and Rocco Oliveto. What quality aspects influence the adoption of Docker images? ACM Transactions on Software Engineering and Methodology, 32(6):142:1– 142:??, November 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3603111.

Rahman:2023:SMO

[RSBP23] Akond Rahman, Shazibul Islam Shamim, Dibyendu Brinto Bose, and Rahul Pandita. Security misconfigurations in open source Kubernetes manifests: an empirical study. ACM Transactions on Software Engineering and Methodology, 32 (4):99:1–99:??, July 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3579639.

Ricca:2014:AES

[RST^{+}14] Filippo Ricca, Giuseppe Scanniello, Marco Torchiano, Gianna Reggio, and Egidio Astesiano. Assessing the effect of screen mockups on the comprehension of functional requirements. ACM Transactions on Software Engineering and Methodology, 24(1):1:1–1:??, September 2014. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Russo:2021:ASM

[Rus21] Daniel Russo. The agile success model: a mixed-methods study of a large-scale agile transformation. ACM Transactions on Software Engineering and Methodology, 30(4): 52:1–52:46, July 2021. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3464938.

Roshandel:2004:MSM

[RVMRM04] Roshanak Roshandel, André Van Der Hoek, Marija Mikic-Rakic, and Nenad Medvidovic. Mae—a system model and environment for managing architectural evolution. ACM Transactions on Software Engineering and Methodology, 13(2):240– 276, April 2004. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Razzaq:2019:SEE

[RWEB19] Abdul Razzaq, Asanka Wasala, Chris Exton, and Jim Buckley. The state of empirical evaluation in static feature location. ACM Transactions on Software Engineering and Methodology, 28(1):2:1–2:??, February 2019. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Ren:2019:NNB

[RXX^{+}19] Xiaoxue Ren, Zhenchang Xing, Xin Xia, David Lo, Xinyu Wang, and John Grundy. Neural network-based detection of self-admitted technical debt: From performance to explainability. ACM Transactions on Software Engineering and Methodology, 28(3): 15:1–15:??, August 2019. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.
                                        acm.org/ft_gateway.cfm?id=
                                        3324916.

Stol:2014:KFA

[SAB^{+}14] Klaas-Jan Stol, Paris Avgeriou, Muhammad Ali Babar, Yan Lucas, and Brian Fitzgerald. Key factors for adopting inner source. ACM Transactions on Software Engineering and Methodology, 23(2):18:1– 18:??, March 2014. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Smaragdakis:2002:MLO

[SB02] Yannis Smaragdakis and Don Batory. Mixin layers: an object-oriented implementation technique for refinements and collaboration-based designs. ACM Transactions on Software Engineering and Methodology, 11(2):215–255, 2002. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Snook:2006:UBF

[SB06] Colin Snook and Michael Butler. UML-B: Formal modeling and design aided by UML. ACM Transactions on Software Engineering and Methodology, 15 (1):92–122, January 2006. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Sayar:2023:DSJ

[SBBL23] Imen Sayar, Alexandre Bartel, Eric Bodden, and Yves Le Traon. An in-depth study of Java deserialization remotecode execution exploits and vulnerabilities. ACM Transactions on Software Engineering and Methodology, 32(1):25:1– 25:??, January 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3554732.

Sobhy:2021:ESA

[SBMK21] Dalia Sobhy, Rami Bahsoon, Leandro Minku, and Rick Kazman. Evaluation of software architectures under uncertainty: a systematic literature review. ACM Transactions on Software Engineering and Methodology, 30(4): 51:1–51:50, July 2021. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3464305.

Sattler:2023:SIP

[SBS^{+}23] Florian Sattler, Sebastian Böhm, Philipp Dominik Schubert, Norbert Siegmund, and Sven Apel. SEAL: Integrating program analysis and repository mining. ACM Transactions on Software Engineering and Methodology, 32(5):121:1–121:??, September 2023. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3585008.

Sinnig:2013:UCT

[SCK13] Daniel Sinnig, Patrice Chalin, and Ferhat Khendek. Use case and task models: an integrated development methodology and its formal foundation. ACM Transactions on Software Engineering and Methodology, 22 (3):27:1–27:??, July 2013. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Sun:2023:DHS

[SCL^{+}23] Xiaoyu Sun, Xiao Chen, Li Li, Haipeng Cai, John Grundy, Jordan Samhi, Tegawendé Bissyandé, and Jacques Klein. Demystifying hidden sensitive operations in Android apps. ACM Transactions on Software Engineering and Methodology, 32 (2):50:1–50:??, April 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3574158.

Sun:2024:IGM

[SDG^{+}24] Chang-Ai Sun, Hepeng Dai, Ning Geng, Huai Liu, Tsong Yueh Chen, Peng Wu, Yan Cai, and Jinqiu Wang. An interleaving guided metamorphic testing approach for concurrent programs. ACM Transactions on Software Engineering and Methodology, 33(1):8:1–8:??, January 2024. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3607182.

Sun:2023:FDM

[SDLC23] Chang-Ai Sun, Hepeng Dai, Huai Liu, and Tsong Yueh Chen. Feedback-directed metamorphic testing. ACM Transactions on Software Engineering and Methodology, 32(1): 20:1–20:??, January 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3533314.

Stolee:2014:SSS

[SED14] Kathryn T. Stolee, Sebastian Elbaum, and Daniel Dobos. Solving the search for source code. ACM Transactions on Software Engineering and Methodology, 23(3):26:1–26:??, May 2014. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Sadeghi:2017:ECA

[SEM17] Alireza Sadeghi, Naeem Esfahani, and Sam Malek. Ensuring the consistency of adaptation through inter- and intracomponent dependency analysis. ACM Transactions on Software Engineering and Methodology, 26(1):2:1–2:??, July 2017. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Song:2024:IPC

[SER24] Qunying Song, Emelie Engström, and Per Runeson. Industry practices for challenging autonomous driving systems with critical scenarios. ACM Transactions on Software Engineering and Methodology, 33 (4):99:1–99:??, May 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3640334.

Stol:2018:ASE

[SF18] Klaas-Jan Stol and Brian Fitzgerald. The ABC of software engineering research. ACM Transactions on Software Engineering and Methodology, 27(3): 11:1–11:??, October 2018. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Sun:2024:EAF

[SFC^{+}24] Weisong Sun, Chunrong Fang, Yuchen Chen, Quanjun Zhang, Guanhong Tao, Yudu You, Tingxu Han, Yifei Ge, Yuling Hu, Bin Luo, and Zhenyu Chen. An extractive-and-abstractive framework for source code summarization. ACM Transactions on Software Engineering and Methodology, 33(3): 75:1–75:??, March 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3632742.

Sherman:2015:DTB

[SGD15] Elena Sherman, Brady J.

Garvin, and Matthew B. Dwyer. Deciding type-based partialorder constraints for pathsensitive analysis. ACM Transactions on Software Engineering and Methodology, 24(3): 15:1–15:??, May 2015. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Sistla:2000:SSB

[SGE00] A. Prasad Sistla, Viktor Gyuris, and E. Allen Emerson. SMC: a symmetry-based model checker for verification of safety and liveness properties. ACM Transactions on Software Engineering and Methodology, 9(2):133–166, April 2000. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/2000-9-2/p133-
sistla/p133-sistla.pdf;
http://www.acm.org/pubs/citations/
journals/tosem/2000-9-2/p133-
sistla/.

Scanniello:2014:IUA

[SGG^{+}14] Giuseppe Scanniello, Carmine Gravino, Marcela Genero, Jose’ A. Cruz-Lemus, and Genoveffa Tortora. On the impact of UML analysis models on source-code comprehensibility and modifiability. ACM Transactions on Software Engineering and Methodology, 23(2): 13:1–13:??, March 2014. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Shen:2023:CSM

[SGHM23] Bowen Shen, Muhammad Ali Gulzar, Fei He, and Na Meng. A characterization study of merge conflicts in Java projects. ACM Transactions on Software Engineering and Methodology, 32 (2):40:1–40:??, April 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3546944.

Sundelin:2022:TAS

[SGhWG22] Anders Sundelin, Javier Gonzalezhuerta, Krzysztof Wnuk, and Tony Gorschek. Towards an anatomy of software craftsmanship. ACM Transactions on Software Engineering and Methodology, 31(1): 6:1–6:49, January 2022. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3468504.

Shonle:2012:FCR

[SGL12] Macneil Shonle, William G.

Griswold, and Sorin Lerner. A framework for the checking and refactoring of crosscutting concepts. ACM Transactions on Software Engineering and Methodology, 21(3):15:1–15:??, June 2012. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Sullivan:2010:MAO

[SGR^{+}10] Kevin Sullivan, William G.

Griswold, Hridesh Rajan, Yuanyuan Song, Yuanfang Cai, Macneil Shonle, and Nishit Tewari. Modular aspect-oriented design with XPIs. ACM Transactions on Software Engineering and Methodology, 20(2):5:1– 5:??, August 2010. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Scanniello:2015:DDP

[SGR^{+}15] Giuseppe Scanniello, Carmine Gravino, Michele Risi, Genoveffa Tortora, and Gabriella Dodero. Documenting designpattern instances: a family of experiments on source-code comprehensibility. ACM Transactions on Software Engineering and Methodology, 24(3): 14:1–14:??, May 2015. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Shreeve:2023:MSU

[SGR^{+}23] Benjamin Shreeve, Catarina Gralha, Awais Rashid, João Araújo, and Miguel Goulão. Making sense of the unknown: How managers make cyber security decisions. ACM Transactions on Software Engineering and Methodology, 32(4): 83:1–83:??, July 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3548682.

Sharafi:2021:TOM

[SHLW21] Zohreh Sharafi, Yu Huang, Kevin Leach, and Westley Weimer. Toward an objective measure of developers’ cognitive activities. ACM Transactions on Software Engineering and Methodology, 30(3): 30:1–30:40, May 2021. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3434643.

Sutton:1995:ALS

[SHO95] Stanley M. Sutton, Dennis Heimbigner, and Leon J. Osterweil. APPL/A: a language for software process programming. ACM Transactions on Software Engineering and Methodology, 4(3):221–286, July 1995. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1995-4-3/p221-
sutton/p221-sutton.pdf;
http://www.acm.org/pubs/citations/
journals/tosem/1995-4-3/p221-
sutton/.

Sinha:2001:ICD

[SHR01] Saurabh Sinha, Mary Jean Harrold, and Gregg Rothermel. Interprocedural control dependence. ACM Transactions on Software Engineering and Methodology, 10(2):209–254, April 2001. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/2001-10-2/
p209-sinha/p209-sinha.pdf;
 http://www.acm.org/pubs/
citations/journals/tosem/
2001-10-2/p209-sinha/.

Sworna:2023:AFA

[SIB23] Zarrin Tasnim Sworna, Chadni Islam, and Muhammad Ali Babar. APIRO: a framework for automated security tools API recommendation. ACM Transactions on Software Engineering and Methodology, 32(1):

24:1–24:??, January 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3512768.

Singh:2010:SWE

[Sin10] Param Vir Singh. The smallworld effect: The influence of macro-level properties of developer collaboration networks on open-source project success. ACM Transactions on Software Engineering and Methodology, 20(2):6:1–6:??, August 2010. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Sahin:2014:CSD

[SKBD14] Dilan Sahin, Marouane Kessentini, Slim Bechikh, and Kalyanmoy Deb. Code-smell detection as a bilevel problem. ACM Transactions on Software Engineering and Methodology, 24(1): 6:1–6:??, September 2014. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Sohn:2023:ASB

[SKY23] Jeongju Sohn, Sungmin Kang, and Shin Yoo. Arachne: Searchbased repair of deep neural networks. ACM Transactions on Software Engineering and Methodology, 32(4):85:1–85:??, July 2023. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3563210.

Sun:2021:TRE

[SLB^{+}21] Xiaoyu Sun, Li Li, Tegawendé F.

Bissyandé, Jacques Klein, Damien Octeau, and John Grundy. Taming reflection: an essential step toward wholeprogram analysis of Android apps. ACM Transactions on Software Engineering and Methodology, 30(3):32:1–32:36, May 2021. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3440033.

Sun:2013:MVH

[SLD^{+}13] Jun Sun, Yang Liu, Jin Song Dong, Yan Liu, Ling Shi, and Étienne André. Modeling and verifying hierarchical real-time systems using stateful timed CSP. ACM Transactions on Software Engineering and Methodology, 22(1):3:1– 3:??, February 2013. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Shao:2024:ESR

[SLWZ24] Changjie Shao, Gaolei Li, Jun Wu, and Xi Zheng. Exploring semantic redundancy using backdoor triggers: a complementary insight into the challenges facing DNN-based software vulnerability detection. ACM Transactions on Software Engineering and Methodology, 33(4):92:1–92:??, May 2024. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3640333.

Strecker:2012:ADC

[SM12] Jaymie Strecker and Atif M. Memon. Accounting for defect characteristics in evaluations of testing techniques. ACM Transactions on Software Engineering and Methodology, 21(3): 17:1–17:??, June 2012. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Siegel:2008:CSE

[SMAC08] Stephen F. Siegel, Anastasia Mironova, George S. Avrunin, and Lori A. Clarke. Combining symbolic execution with model checking to verify parallel numerical programs. ACM Transactions on Software Engineering and Methodology, 17 (2):10:1–10:??, April 2008. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Sobhy:2022:CPS

[SMBK22] Dalia Sobhy, Leandro Minku, Rami Bahsoon, and Rick Kazman. Continuous and proactive software architecture evaluation: an IoT case. ACM Transactions on Software Engineering and Methodology, 31 (3):46:1–46:54, July 2022. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3492762.

Scalabrino:2021:ASB

[SMBO21] Simone Scalabrino, Antonio Mastropaolo, Gabriele Bavota, and Rocco Oliveto. An adaptive search budget allocation approach for search-based test case generation. ACM Transactions on Software Engineering and Methodology, 30(3): 36:1–36:26, May 2021. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3446199.

Schneider:1992:ESF

[SMT92] G. Michael Schneider, Johnny Martin, and W. T. Tsai. An experimental study of fault detection in user requirements documents. ACM Transactions on Software Engineering and Methodology, 1(2):188–204, April 1992. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1992-1-2/p188-
schneider/p188-schneider.
pdf;   http://www.acm.org/
pubs/citations/journals/tosem/
1992-1-2/p188-schneider/.

Song:2019:SEI

[SMY19] Liyan Song, Leandro L. Minku, and Xin Yao. Software effort interval prediction via Bayesian inference and synthetic bootstrap resampling. ACM Transactions on Software Engineering and Methodology, 28(1):5:1– 5:??, February 2019. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Shi:2024:ASE

[SMY^{+}24] Weishi Shi, Heather Moses, Qi Yu, Samuel Malachowsky, and Daniel E. Krutz. ALL: Supporting experiential accessibility education and inclusive software development. ACM Transactions on Software Engineering and Methodology, 33(2): 39:1–39:??, February 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3625292.

Sullivan:1992:REI

[SN92] Kevin J. Sullivan and David Notkin. Reconciling environment integration and software evolution. ACM Transactions on Software Engineering and Methodology, 1(3):229–268, July 1992. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

                                        www.acm.org/pubs/articles/
                                        journals/tosem/1992-1-3/p229-
                                        sullivan/p229-sullivan.pdf;
                                         http://www.acm.org/pubs/
                                        citations/journals/tosem/
                                        1992-1-3/p229-sullivan/.

Snelting:1996:RCB

[Sne96] Gregor Snelting. Reengineering of configurations based on mathematical concept analysis. ACM Transactions on Software Engineering and Methodology, 5(2):146–189, April 1996. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1996-5-2/p146-
snelting/p146-snelting.pdf;
 http://www.acm.org/pubs/
citations/journals/tosem/
1996-5-2/p146-snelting/.

Sintaha:2023:KDS

[SNM23] Mifta Sintaha, Noor Nashid, and Ali Mesbah. Katana: Dual slicing based context for learning bug fixes. ACM Transactions on Software Engineering and Methodology, 32(4): 100:1–100:??, July 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3579640.

Sarro:2018:LPB

[SP18] Federica Sarro and Alessio Petrozziello. Linear programming as a baseline for software effort estimation. ACM Transactions on Software Engineering and Methodology, 27(3): 12:1–12:??, October 2018. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Steimann:2010:TMI

[SPAK10] Friedrich Steimann, Thomas Pawlitzki, Sven Apel, and Christian Kästner. Types and modularity for implicit invocation with implicit announcement. ACM Transactions on Software Engineering and Methodology, 20(1):

1:1–1:??, June 2010. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Siegmund:2021:MVH

[SPAS21] Janet Siegmund, Norman Peitek, Sven Apel, and Norbert Siegmund. Mastering variation in human studies: The role of aggregation. ACM Transactions on Software Engineering and Methodology, 30(1): 2:1–2:40, January 2021. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3406544.

Santhiar:2014:MUT

[SPK14] Anirudh Santhiar, Omesh Pandita, and Aditya Kanade. Mining unit tests for discovery and migration of math APIs. ACM Transactions on Software Engineering and Methodology, 24(1): 4:1–4:??, September 2014. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Sommerville:2005:ESI

[SR05] Ian Sommerville and Jane Ransom. An empirical study of industrial requirements engineering process assessment and improvement. ACM Transactions on Software Engineering and Methodology, 14(1):85–117, January 2005. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Snelting:2006:EPC

[SRK06] Gregor Snelting, Torsten Robschink, and Jens Krinke. Efficient path conditions in dependence graphs for software safety analysis. ACM Transactions on Software Engineering and Methodology, 15(4):410– 457, October 2006. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Scanniello:2017:FFC

[SRTR17] Giuseppe Scanniello, Michele Risi, Porfirio Tramontana, and Simone Romano. Fixing faults in C and Java source code: Abbreviated vs. full-word identifier names. ACM Transactions on Software Engineering and Methodology, 26(2):6:1– 6:??, October 2017. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Schrefl:2002:BCS

[SS02] Michael Schrefl and Markus Stumptner. Behavior-consistent specialization of object life cycles. ACM Transactions on Software Engineering and Methodology, 11(1):92–148, January 2002. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Sinha:2006:HMB

[SS06] Avik Sinha and Carol Smidts.

HOTTest: a model-based test design technique for enhanced testing of domain-specific applications. ACM Transactions on Software Engineering and Methodology, 15(3): 242–278, July 2006. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Soltana:2020:PCS

[SSB20] Ghanem Soltana, Mehrdad Sabetzadeh, and Lionel C. Briand. Practical constraint solving for generating system test data. ACM Transactions on Software Engineering and Methodology, 29(2):11:1– 11:48, April 2020. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3381032.

Sarker:2023:AIT

[STDB23] Jaydeb Sarker, Asif Kamal Turzo, Ming Dong, and Amiangshu Bosu. Automated identification of toxic code reviews using ToxiCR. ACM Transactions on Software Engineering and Methodology, 32(5):118:1– 118:??, September 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3583562.

Shariffdeen:2021:APT

[STGR21] Ridwan Salihin Shariffdeen, Shin Hwei Tan, Mingyuan Gao, and Abhik Roychoudhury. Automated patch transplantation. ACM Transactions on Software Engineering and Methodology, 30(1):6:1–6:36, January 2021. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3412376.

Sharma:2018:RWF

[STS^{+}18] Abhishek Sharma, Yuan Tian, Agus Sulistya, Dinusha Wijedasa, and David Lo. Recommending who to follow in the software engineering Twitter space. ACM Transactions on Software Engineering and Methodology, 27(4):16:1– 16:??, November 2018. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Sim:2011:HWD

[SURL11] Susan Elliott Sim, Medha Umarji, Sukanya Ratanotayanon, and Cristina V. Lopes. How well do search engines support code retrieval on the Web? ACM Transactions on Software Engineering and Methodology, 21(1): 4:1–4:??, December 2011. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Soto-Valero:2023:CBD

[SVDHB23] César Soto-Valero, Thomas Durieux, Nicolas Harrand, and Benoit Baudry. Coverage-based debloating for Java bytecode. ACM Transactions on Software Engineering and Methodology, 32(2):38:1–38:??, April 2023. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3546948.

Shin:2024:GBM

[SWW^{+}24] Jiho Shin, Moshi Wei, Junjie Wang, Lin Shi, and Song Wang. The good, the bad, and the missing: Neural code generation for machine learning tasks. ACM Transactions on Software Engineering and Methodology, 33(2):51:1– 51:??, February 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3630009.

Sun:2024:ALI

[SXX^{+}24] Jiamou Sun, Zhenchang Xing, Xin Xia, Qinghua Lu, Xiwei Xu, and Liming Zhu. Aspectlevel information discrepancies across heterogeneous vulnerability reports: Severity, types and detection methods. ACM Transactions on Software Engineering and Methodology, 33(2): 49:1–49:??, February 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3624734.

Safdar:2021:RFC

[SYA21] Safdar Aqeel Safdar, Tao Yue, and Shaukat Ali. Recommending faulty configurations for interacting systems under test using multi-objective search. ACM Transactions on Software Engineering and Methodology, 30(4):53:1–53:36, July 2021. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3464939.

Sun:2023:RIC

[SYL^{+}23] Weifeng Sun, Meng Yan, Zhongxin Liu, Xin Xia, Yan Lei, and David Lo. Revisiting the identification of the co-evolution of production and test code. ACM Transactions on Software Engineering and Methodology, 32(6):152:1– 152:??, November 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3607183.

Sheng:2019:TPA

[SZH^{+}19] Feng Sheng, Huibiao Zhu, Jifeng He, Zongyuan Yang, and Jonathan P. Bowen. Theoretical and practical aspects of linking operational and algebraic semantics for MDESL. ACM Transactions on Software Engineering and Methodology, 28(3): 14:1–14:??, August 2019. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?id=
3295699.

Suneja:2023:ISA

[SZZ^{+}23] Sahil Suneja, Yufan Zhuang, Yunhui Zheng, Jim Laredo, Alessandro Morari, and Udayan Khurana. Incorporating signal awareness in source code modeling: an application to vulnerability detection. ACM Transactions on Software Engineering and Methodology, 32(6):145:1– 145:??, November 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3597202.

Tramontana:2019:DEO

[TAA^{+}19] Porfirio Tramontana, Domenico Amalfitano, Nicola Amatucci, Atif Memon, and Anna Rita Fasolino. Developing and evaluating objective termination criteria for random testing. ACM Transactions on Software Engineering and Methodology, 28(3): 17:1–17:??, August 2019. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.
                                        acm.org/ft_gateway.cfm?id=
                                        3339836.

Tichy:1995:AR

[TAW95] Walter Tichy, Rolf Adams, and Annette Weinert. Authors’ response. ACM Transactions on Software Engineering and Methodology, 4(2):217–219, April 1995. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

                                        www.acm.org/pubs/articles/
                                        journals/tosem/1995-4-2/p217-
                                        tichy/p217-tichy.pdf;   http:
                                        //www.acm.org/pubs/citations/
                                        journals/tosem/1995-4-2/p217-

tichy/. See [ATW94, Bre95].

Trammell:1992:APC

[TBS92] Carmen J. Trammell, Leon H. Binder, and Cathrine E. Snyder. The automated production control documentation system: a case study in cleanroom software engineering. ACM Transactions on Software Engineering and Methodology, 1(1):81–94, January 1992. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1992-1-1/p81-
trammell/p81-trammell.pdf;
 http://www.acm.org/pubs/
citations/journals/tosem/
1992-1-1/p81-trammell/.

Taymouri:2020:CAW

[TC20] Farbod Taymouri and Josep Carmona. Computing alignments of well-formed process models using local search. ACM Transactions on Software Engineering and Methodology, 29(3): 15:1–15:41, July 2020. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3394056.

Tian:2019:DTC

[TCDZ19] Cong Tian, Chu Chen, Zhenhua Duan, and Liang Zhao. Differential testing of certificate validation in SSL/TLS implementations: an RFC-guided approach. ACM Transactions on Software Engineering and Methodology, 28(4):24:1–24:??, October 2019. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?id=
3355048.

Tip:2001:SBA

[TD01] F. Tip and T. B. Dinesh. A slicing-based approach for locating type errors. ACM Transactions on Software Engineering and Methodology, 10(1):5–55, January 2001. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

                                        www.acm.org/pubs/articles/
                                        journals/tosem/2001-10-1/
                                        p5-tip/p5-tip.pdf; http://
                                        www.acm.org/pubs/citations/
                                        journals/tosem/2001-10-1/
                                        p5-tip/.

Tang:2024:OSV

[TDL^{+}24] Wensheng Tang, Dejun Dong, Shijie Li, Chengpeng Wang, Peisen Yao, Jinguo Zhou, and Charles Zhang. Octopus: Scaling value-flow analysis via parallel collection of realizable path conditions. ACM Transactions on Software Engineering and Methodology, 33(3): 66:1–66:??, March 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3632743.

Traini:2022:HSR

[TDT^{+}22] Luca Traini, Daniele Di Pompeo, Michele Tucci, Bin Lin, Simone Scalabrino, Gabriele Bavota, Michele Lanza, Rocco Oliveto, and Vittorio Cortellessa. How software refactoring impacts execution time. ACM Transactions on Software Engineering and Methodology, 31 (2):25:1–25:23, April 2022. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3485136.

Tilevich:2011:EEP

[TG11] Eli Tilevich and Sriram Gopal.

Expressive and extensible parameter passing for distributed object systems. ACM Transactions on Software Engineering and Methodology, 21(1):3:1– 3:??, December 2011. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Tamasauskaite:2023:DKG

[TG23] Gyte Tamasauskaite and Paul Groth. Defining a knowledge graph development process through a systematic review. ACM Transactions on Software Engineering and Methodology, 32(1):27:1–27:??, January 2023. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3522586.

Taipalus:2024:FSE

[TG24] Toni Taipalus and Hilkka Grahn. Framework for SQL error message design: a datadriven approach. ACM Transactions on Software Engineering and Methodology, 33(1): 9:1–9:??, January 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3607180.

Thiran:2006:WBE

[THHB06] Philippe Thiran, Jean-Luc Hainaut, Geert-Jan Houben, and Djamal Benslimane. Wrapperbased evolution of legacy information systems. ACM Transactions on Software Engineering and Methodology, 15(4):329– 359, October 2006. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Tiwana:2008:ICD

[Tiw08]  Amrit Tiwana.  Impact of

classes of development coordination tools on software development performance: a multinational empirical study. ACM Transactions on Software Engineering and Methodology, 17 (2):11:1–11:??, April 2008. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Tsuchiya:2002:FCE

[TK02] Tatsuhiro Tsuchiya and Tohru Kikuno. On fault classes and error detection capability of specification-based testing. ACM Transactions on Software Engineering and Methodology, 11(1):58–62, January 2002. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Tan:2024:URT

[TLJZ24] Xin Tan, Xinyue Lv, Jing Jiang, and Li Zhang. Understanding real-time collaborative programming: a study of Visual Studio Live Share. ACM Transactions on Software Engineering and Methodology, 33(4): 110:1–110:??, May 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3643672.

Tian:2023:BBW

[TLL^{+}23] Haoye Tian, Kui Liu, Yinghua Li, Abdoul Kader Kaboré, Anil Koyuncu, Andrew Habib, Li Li, Junhao Wen, Jacques Klein, and Tegawendé F. Bissyandé. The best of both worlds: Combining learned embeddings with engineered features for accurate prediction of correct patches. ACM Transactions on Software Engineering and Methodology, 32(4):92:1–92:??, July 2023. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3576039.

Tian:2022:PPC

[TLP^{+}22] Haoye Tian, Yinghua Li, Weiguo Pian, Abdoul Kader Kaboré, Kui Liu, Andrew Habib, Jacques Klein, and Tegawendé F. Bissyandé. Predicting patch correctness based on the similarity of failing test cases. ACM Transactions on Software Engineering and Methodology, 31(4):77:1– 77:??, October 2022. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3511096.

Tappenden:2014:ACC

[TM14] Andrew F. Tappenden and James Miller. Automated cookie collection testing. ACM Transactions on Software Engineering and Methodology, 23(1): 3:1–3:??, February 2014. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Tateishi:2013:PIS

[TPT13] Takaaki Tateishi, Marco Pistoia, and Omer Tripp. Pathand index-sensitive string analysis based on monadic secondorder logic. ACM Transactions on Software Engineering and Methodology, 22(4):33:1– 33:??, October 2013. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Tilevich:2009:JOE

[TS09] Eli Tilevich and Yannis Smaragdakis. J-Orchestra: Enhancing Java programs with distribution capabilities. ACM Transactions on Software Engineering and Methodology, 19(1):1:1– 1:??, August 2009. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Troya:2018:SBF

[TSPRC18] Javier Troya, Sergio Segura, Jose Antonio Parejo, and Antonio Ruiz-Cortés. Spectrumbased fault localization in model transformations. ACM Transactions on Software Engineering and Methodology, 27(3):13:1– 13:??, October 2018. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Tao:2021:SDM

[TTL^{+}21] Yida Tao, Shan Tang, Yepang Liu, Zhiwu Xu, and Shengchao Qin. Speeding up data manipulation tasks with alternative implementations: an exploratory study. ACM Transactions on Software Engineering and Methodology, 30(4): 49:1–49:28, July 2021. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3456873.

Tufano:2019:ESL

[TWB^{+}19] Michele Tufano, Cody Watson, Gabriele Bavota, Massimiliano Di Penta, Martin White, and Denys Poshyvanyk. An empirical study on learning bug-fixing patches in the wild via neural machine translation. ACM Transactions on Software Engineering and Methodology, 28(4): 19:1–19:??, October 2019. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?id=
3340544.

Trinkenreich:2022:WPO

[TWS^{+}22] Bianca Trinkenreich, Igor Wiese, Anita Sarma, Marco Gerosa, and Igor Steinmacher. Women’s participation in open source software: a survey of the literature. ACM Transactions on Software Engineering and Methodology, 31(4):81:1– 81:??, October 2022. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3510460.

Tyszberowicz:1992:OPL

[TY92] Shmuel Tyszberowicz and Amiram Yehudai. OBSERV — a prototyping language and environment. ACM Transactions on Software Engineering and Methodology, 1(3):269–309, July 1992. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1992-1-3/p269-
tyszberowicz/p269-tyszberowicz.
pdf;   http://www.acm.org/
pubs/citations/journals/tosem/
1992-1-3/p269-tyszberowicz/

.

Tian:2024:CSS

[TZD^{+}24] Yongqiang Tian, Xueyan Zhang, Yiwen Dong, Zhenyang Xu, Mengxiao Zhang, Yu Jiang, Shing-Chi Cheung, and Chengnian Sun. On the caching schemes to speed up program reduction. ACM Transactions on Software Engineering and Methodology, 33(1):17:1– 17:??, January 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3617172.

Tian:2023:FDB

[TZW^{+}23] Yongqiang Tian, Wuqi Zhang, Ming Wen, Shing-Chi Cheung, Chengnian Sun, Shiqing Ma, and Yu Jiang. Finding deviated behaviors of the compressed DNN models for image classifications. ACM Transactions on Software Engineering and Methodology, 32(5):128:1– 128:??, September 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3583564.

Tan:2009:CDM

[TZZ09] Hee Beng Kuan Tan, Yuan Zhao, and Hongyu Zhang. Conceptual data model-based software size estimation for information systems. ACM Transactions on Software Engineering and Methodology, 19(2):4:1– 4:??, October 2009. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Tang:2023:SAD

[TZZ^{+}23] Shuncheng Tang, Zhenya Zhang, Yi Zhang, Jixiang Zhou, Yan Guo, Shuang Liu, Shengjian Guo, Yan-Fu Li, Lei Ma, Yinxing Xue, and Yang Liu. A survey on automated driving system testing: Landscapes and trends. ACM Transactions on Software Engineering and Methodology, 32(5):124:1– 124:??, September 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3579642.

Unterkalmsteiner:2014:TRE

[UFG14] M. Unterkalmsteiner, R. Feldt, and T. Gorschek. A taxonomy for requirements engineering and software test alignment. ACM Transactions on Software Engineering and Methodology, 23(2):16:1–16:??, March 2014. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Uddin:2022:ESE

[UGKR22] Gias Uddin, Yann-Gaël Guéhénuc, Foutse Khomh, and Chanchal K. Roy. An empirical study of the effectiveness of an ensemble of stand-alone sentiment detection tools for software engineering datasets. ACM Transactions on Software Engineering and Methodology, 31 (3):48:1–48:38, July 2022. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3491211.

Uchitel:2004:IES

[UKM04] Sebastian Uchitel, Jeff Kramer, and Jeff Magee. Incremental elaboration of scenariobased specifications and behavior models using implied scenarios. ACM Transactions on Software Engineering and Methodology, 13(1):37–85, January 2004. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Uddin:2021:AAU

[UKR21] Gias Uddin, Foutse Khomh, and Chanchal K. Roy. Automatic API usage scenario documentation from technical Q&A sites. ACM Transactions on Software Engineering and Methodology, 30(3): 31:1–31:45, May 2021. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3439769.

Vidal:2018:ARB

[VBZ^{+}18] Santiago Vidal, Iñaki Berra, Santiago Zulliani, Claudia Marcos, and J. Andrés Díaz Pace. Assessing the refactoring of brain methods. ACM Transactions on Software Engineering and Methodology, 27(1):2:1– 2:??, June 2018. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Vandehei:2021:LDL

[VCF21] Bailey Vandehei, Daniel Alencar Da Costa, and Davide Falessi. Leveraging the defects life cycle to label affected versions and defective classes. ACM Transactions on Software Engineering and Methodology, 30(2): 24:1–24:35, March 2021. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3433928.

Venturini:2023:DYY

[VCP^{+}23] Daniel Venturini, Filipe Roseiro Cogo, Ivanilton Polato, Marco A. Gerosa, and Igor Scaliante Wiese. I depended on you and you broke me: an empirical study of manifesting breaking changes in client packages. ACM Transactions on Software Engineering and Methodology, 32(4): 94:1–94:??, July 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3576037.

Vouillon:2013:SCC

[VD13] Jérôme Vouillon and Roberto Di Cosmo. On software component co-installability. ACM Transactions on Software Engineering and Methodology, 22(4): 34:1–34:??, October 2013. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

vandenBrand:1996:GFC

[vdBV96] Mark van den Brand and Eelco Visser. Generation of formatters for context-free languages. ACM Transactions on Software Engineering and Methodology, 5(1):1–41, January 1996. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1996-5-1/p1-
van\_den\_brand/p1-van\_den\
_brand.pdf;   http://www.
acm.org/pubs/citations/journals/
tosem/1996-5-1/p1-van\_den\
_brand/; http://www.acm.org/
pubs/citations/journals/tosem/

1996-5-1/p1-van_den_brand/

.

Vogel-Heuser:2022:MIA

[VHNF22] Birgit Vogel-Heuser, Eva-Maria Neumann, and Juliane Fischer. MICOSE4aPS: Industrially applicable maturity metric to improve systematic reuse of control software. ACM Transactions on Software Engineering and Methodology, 31(1): 5:1–5:24, January 2022. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3467896.

VanDenBrand:2003:TRT

[VKV03] Mark G. J. Van Den Brand, Paul Klint, and Jurgen J. Vinju. Term rewriting with traversal functions. ACM Transactions on Software Engineering and Methodology, 12(2):152– 190, April 2003. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

VonRhein:2018:VAS

[VLJ^{+}18] Alexander Von Rhein, Jörg Liebig, Andreas Janker, Christian Kästner, and Sven Apel. Variability-aware static analysis at scale: an empirical study. ACM Transactions on Software Engineering and Methodology, 27(4):18:1–18:??, November 2018. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Verwijs:2023:TST

[VR23] Christiaan Verwijs and Daniel Russo. A theory of Scrum team effectiveness. ACM Transactions on Software Engineering and Methodology, 32(3): 74:1–74:??, July 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3571849.

Venkatasubramanian:2004:FMR

[VTA04] Nalini Venkatasubramanian, Carolyn Talcott, and Gul A. Agha. A formal model for reasoning about adaptive QoSenabled middleware. ACM Transactions on Software Engineering and Methodology, 13 (1):86–147, January 2004. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Wallach:2000:SSM

[WAF00] Dan S. Wallach, Andrew W. Appel, and Edward W. Felten. SAFKASI: a security mechanism for language-based systems. ACM Transactions on Software Engineering and Methodology, 9(4):341–378, October 2000. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/2000-9-4/p341-
wallach/p341-wallach.pdf;
 http://www.acm.org/pubs/
citations/journals/tosem/
2000-9-4/p341-wallach/.

Waga:2023:PTP

[WAH23] Masaki Waga, Étienne André, and Ichiro Hasuo. Parametric timed pattern matching. ACM Transactions on Software Engineering and Methodology, 32(1): 10:1–10:??, January 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3517194.

Walkinshaw:2013:ACS

[WB13] Neil Walkinshaw and Kirill Bogdanov. Automated comparison of state-based software models in terms of their language and structure. ACM Transactions on Software Engineering and Methodology, 22(2):13:1– 13:??, March 2013. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Wolter:2023:OSL

[WBRH23] Thomas Wolter, Ann Barcomb, Dirk Riehle, and Nikolay Harutyunyan. Open source license inconsistencies on GitHub. ACM Transactions on Software Engineering and Methodology, 32(5): 110:1–110:??, September 2023. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3571852.

Wu:2023:AIU

[WC23] Jianwei Wu and James Clause.

Automated identification of uniqueness in JUnit tests. ACM Transactions on Software Engineering and Methodology, 32(1): 18:1–18:??, January 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3533313.

Wang:2020:KDI

[WCB^{+}20] Guanhua Wang, Sudipta Chattopadhyay, Arnab Kumar Biswas, Tulika Mitra, and Abhik Roychoudhury. KLEESpectre: Detecting information leakage through speculative cache attacks via symbolic execution. ACM Transactions on Software Engineering and Methodology, 29(3):14:1– 14:31, July 2020. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3385897.

Wu:2021:WAA

[WCG^{+}21] Bozhi Wu, Sen Chen, Cuiyun Gao, Lingling Fan, Yang Liu, Weiping Wen, and Michael R. Lyu. Why an Android app is classified as malware: Toward malware classification interpretation. ACM Transactions on Software Engineering and Methodology, 30(2): 21:1–21:29, March 2021. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3423096.

Wen:2023:RPS

[WCJL23] Jinfeng Wen, Zhenpeng Chen, Xin Jin, and Xuanzhe Liu.

Rise of the planet of serverless computing: a systematic review. ACM Transactions on Software Engineering and Methodology, 32(5):131:1– 131:??, September 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3579643.

Watson:2022:SLR

[WCP^{+}22] Cody Watson, Nathan Cooper, David Nader Palacio, Kevin Moran, and Denys Poshyvanyk. A systematic literature review on the use of deep learning in software engineering research. ACM Transactions on Software Engineering and Methodology, 31(2):32:1–32:58, April 2022. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3485275.

Weyuker:1996:UFC

[Wey96] Elaine J. Weyuker. Using failure cost information for testing and reliability assessment. ACM Transactions on Software Engineering and Methodology, 5(2):87–98, April 1996. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1996-5-2/p87-
weyuker/p87-weyuker.pdf;
http://www.acm.org/pubs/citations/
journals/tosem/1996-5-2/p87-
weyuker/.

Wagner:2019:SQR

[WFF^{+}19] Stefan Wagner, Daniel Méndez Fernández, Michael Felderer, Antonio Vetrò, Marcos Kalinowski, Roel Wieringa, Dietmar Pfahl, Tayana Conte, Marie-Therese Christiansson, Desmond Greer, Casper Lassenius, Tomi Männistö, Maleknaz Nayebi, Markku Oivo, Birgit Penzenstadler, Rafael Prikladnicki, Guenther Ruhe, André Schekelmann, Sagar Sen, Rodrigo Spínola, Ahmed Tuzcu, Jose Luis De La Vara, and Dietmar Winkler. Status quo in requirements engineering: a theory and a global family of surveys. ACM Transactions on Software Engineering and Methodology, 28(2): 9:1–9:??, April 2019. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.
                                        acm.org/ft_gateway.cfm?id=
                                        3306607.

Wang:2020:UED

[WFW^{+}20] Zhendong Wang, Yang Feng, Yi Wang, James A. Jones, and David Redmiles. Unveiling elite developers’ activities in open source projects. ACM Transactions on Software Engineering and Methodology, 29(3):16:1– 16:35, July 2020. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3387111.

Wursch:2013:EQF

[WGG13] Michael Würsch, Emanuel Giger, and Harald C. Gall. Evaluating a query framework for software evolution data. ACM Transactions on Software Engineering and Methodology, 22(4): 38:1–38:??, October 2013. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Wassermann:2007:SCD

[WGSD07] Gary Wassermann, Carl Gould, Zhendong Su, and Premkumar Devanbu. Static checking of dynamically generated queries in database applications. ACM Transactions on Software Engineering and Methodology, 16 (4):14:1–14:??, September 2007. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Wang:2023:SCJ

[WHP^{+}23] Chao Wang, Hao He, Uma Pal, Darko Marinov, and Minghui Zhou. Suboptimal comments in Java projects: From independent comment changes to commenting practices. ACM Transactions on Software Engineering and Methodology, 32 (2):45:1–45:??, April 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3546949.

Weyns:2023:AFF

[WI23] Danny Weyns and Usman M.

Iftikhar. ActivFORMS: a formally founded model-based approach to engineer self-adaptive systems. ACM Transactions on Software Engineering and Methodology, 32(1):12:1– 12:??, January 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3522585.

Whittle:2010:SHS

[WJ10] Jon Whittle and Praveen K. Jayaraman. Synthesizing hierarchical state machines from expressive scenario descriptions. ACM Transactions on Software Engineering and Methodology, 19(3):8:1–8:??, January 2010. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Wu:2023:RAK

[WJZ^{+}23] Di Wu, Xiao-Yuan Jing, Hongyu Zhang, Yang Feng, Haowen Chen, Yuming Zhou, and Baowen Xu. Retrieving API knowledge from tutorials and Stack Overflow based on natural language queries. ACM Transactions on Software Engineering and Methodology, 32(5): 109:1–109:??, September 2023. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3565799.

Wang:2020:MTN

[WLS^{+}20] Wenhan Wang, Ge Li, Sijie Shen, Xin Xia, and Zhi Jin. Modular tree network for source code representation learning. ACM Transactions on Software Engineering and Methodology, 29(4):31:1–31:23, October 2020. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3409331.

Wang:1993:DRT

[WME93] Farn Wang, Aloysius K. Mok, and E. Allen Emerson. Distributed real-time system specification and verification in APTL. ACM Transactions on Software Engineering and Methodology, 2(4):346–378, October 1993. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1993-2-4/p346-
wang/p346-wang.pdf;   http:
//www.acm.org/pubs/citations/
journals/tosem/1993-2-4/p346-
wang/.

Wang:2023:SSA

[WMW^{+}23] Huaijin Wang, Pingchuan Ma, Shuai Wang, Qiyi Tang, Sen Nie, and Shi Wu. sem2vec: Semantics-aware assembly tracelet embedding. ACM Transactions on Software Engineering and Methodology, 32(4):90:1–90:??, July 2023. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3569933.

Whigham:2015:BMS

[WOM15] Peter A. Whigham, Caitlin A.

Owen, and Stephen G. Macdonell. A baseline model for software effort estimation. ACM Transactions on Software Engineering and Methodology, 24 (3):20:1–20:??, May 2015. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Whittaker:1993:MAS

[WP93] James A. Whittaker and J. H. Poore. Markov analysis of software specifications. ACM Transactions on Software Engineering and Methodology, 2(1):93–106, January 1993. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

                                        www.acm.org/pubs/articles/
                                        journals/tosem/1993-2-1/p93-
                                        whittaker/p93-whittaker.pdf;
                                         http://www.acm.org/pubs/
                                        citations/journals/tosem/
                                        1993-2-1/p93-whittaker/.

Wang:2019:OTS

[WPB19] Chunhui Wang, Fabrizio Pastore, and Lionel Briand. Oracles for testing software timeliness with uncertainty. ACM Transactions on Software Engineering and Methodology, 28(1): 1:1–1:??, February 2019. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Wang:2023:XEC

[WPX^{+}23] Chong Wang, Xin Peng, Zhenchang Xing, Yue Zhang, Mingwei Liu, Rong Luo, and Xiujie Meng. XCoS: Explainable code search based on query scoping and knowledge graph. ACM Transactions on Software Engineering and Methodology, 32(6): 140:1–140:??, November 2023. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3593800.

Williams:2022:GFI

[WRD^{+}22] Ryan Williams, Tongwei Ren, Lorenzo De Carli, Long Lu, and Gillian Smith. Guided feature identification and removal for resource-constrained firmware. ACM Transactions on Software Engineering and Methodology, 31(2):28:1–28:25, April 2022. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3487568.

Wei:2023:HTB

[WSG^{+}23] Hongwei Wei, Xiaohong Su, Cuiyun Gao, Weining Zheng, and Wenxin Tao. A hypothesis testing-based framework for software cross-modal retrieval in heterogeneous semantic spaces. ACM Transactions on Software Engineering and Methodology, 32(5):123:1– 123:??, September 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3591868.

Weiss:2024:ATSa

[WT24a] Michael Weiss and Paolo Tonella. Adopting two supervisors for efficient use of large-scale remote deep neural networks. ACM Transactions on Software Engineering and Methodology, 33(1):28:1– 28:??, January 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3617593.

Weiss:2024:ATSb

[WT24b] Michael Weiss and Paolo Tonella. Adopting two supervisors for efficient use of largescale remote deep neural networks — RCR report. ACM Transactions on Software Engineering and Methodology, 33(1): 29:1–29:??, January 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3617594.

Wei:2023:DMD

[WWAC23] Zhengyuan Wei, Haipeng Wang, Imran Ashraf, and Wing-Kwong Chan. DeepPatch: Maintaining deep learning model programs to retain standard accuracy with substantial robustness improvement. ACM Transactions on Software Engineering and Methodology, 32(6):150:1– 150:??, November 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3604609.

Wu:2024:PEM

[WWJ^{+}24] Jiarong Wu, Lili Wei, Yanyan Jiang, Shing-Chi Cheung, Luyao Ren, and Chang Xu. Programming by example made easy. ACM Transactions on Software Engineering and Methodology, 33(1):4:1–4:??, January 2024. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

URL https://dl.acm.org/
doi/10.1145/3607185.

Wang:2023:PIM

[WWL^{+}23] Shangwen Wang, Ming Wen, Bo Lin, Yepang Liu, Tegawendé F. Bissyandé, and Xiaoguang Mao. Pre-implementation method name prediction for objectoriented programming. ACM Transactions on Software Engineering and Methodology, 32(6): 157:1–157:??, November 2023. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3597203.

Wang:2023:AFP

[WWY^{+}23] Chengpeng Wang, Wenyang Wang, Peisen Yao, Qingkai Shi, Jinguo Zhou, Xiao Xiao, and Charles Zhang. Anchor: Fast and precise value-flow analysis for containers via memory orientation. ACM Transactions on Software Engineering and Methodology, 32(3): 66:1–66:??, July 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3565800.

Wu:2022:TTI

[WWZ^{+}22] Siwei Wu, Lei Wu, Yajin Zhou, Runhuai Li, Zhi Wang, Xiapu Luo, Cong Wang, and Kui Ren. Time-travel investigation: Toward building a scalable attack detection framework on Ethereum. ACM Transactions on Software Engineering and Methodology, 31(3): 54:1–54:33, July 2022. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3505263.

Wang:2021:CAR

[WXL^{+}21] Haoye Wang, Xin Xia, David Lo, Qiang He, Xinyu Wang, and John Grundy. Contextaware retrieval-based deep commit message generation. ACM Transactions on Software Engineering and Methodology, 30 (4):56:1–56:30, July 2021. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3464689.

Wang:2024:BAE

[WYC^{+}24] Han Wang, Sijia Yu, Chunyang Chen, Burak Turhan, and Xiaodong Zhu. Beyond accuracy: an empirical study on unit testing in open-source deep learning projects. ACM Transactions on Software Engineering and Methodology, 33(4): 104:1–104:??, May 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3638245.

Wang:2020:IIC

[WYMW20] Junjie Wang, Ye Yang, Tim Menzies, and Qing Wang. iSENSE2.0: Improving completionaware crowdtesting management with duplicate tagger and sanity checker. ACM Transactions on Software Engineering and Methodology, 29(4): 24:1–24:27, October 2020. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3394602.

Wang:2022:CFA

[WYW^{+}22] Junjie Wang, Ye Yang, Song Wang, Jun Hu, and Qing Wang. Context- and fairnessaware in-process crowdworker recommendation. ACM Transactions on Software Engineering and Methodology, 31(3): 35:1–35:31, July 2022. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3487571.

Xu:2023:DTB

[XAY23] Qinghua Xu, Shaukat Ali, and Tao Yue. Digital twinbased anomaly detection with curriculum learning in cyberphysical systems. ACM Transactions on Software Engineering and Methodology, 32(5):113:1– 113:??, September 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3582571.

Xu:2010:PCC

[XCCY10] Chang Xu, S. C. Cheung, W. K. Chan, and Chunyang Ye. Partial constraint checking for context consistency in pervasive computing. ACM Transactions on Software Engineering and Methodology, 19(3):9:1– 9:??, January 2010. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Xie:2013:TAR

[XCKX13] Xiaoyuan Xie, Tsong Yueh Chen, Fei-Ching Kuo, and Baowen Xu. A theoretical analysis of the risk evaluation formulas for spectrum-based fault localization. ACM Transactions on Software Engineering and Methodology, 22(4):31:1– 31:??, October 2013. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Xu:2022:MSC

[XCS^{+}22] Hui Xu, Zhuangbin Chen, Mingshen Sun, Yangfan Zhou, and Michael R. Lyu. Memorysafety challenge considered solved? An in-depth study with all Rust CVEs. ACM Transactions on Software Engineering and Methodology, 31(1): 3:1–3:25, January 2022. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3466642.

Xu:2023:LLI

[XGF^{+}23] Sihan Xu, Ya Gao, Lingling Fan, Zheli Liu, Yang Liu, and Hua Ji. LiDetector: License incompatibility detection for open source software. ACM Transactions on Software Engineering and Methodology, 32(1): 22:1–22:??, January 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3518994.

Xiang:2024:ATS

[XHL^{+}24] Yi Xiang, Han Huang, Sizhe Li, Miqing Li, Chuan Luo, and Xiaowei Yang. Automated test suite generation for software product lines based on qualitydiversity optimization. ACM Transactions on Software Engineering and Methodology, 33(2): 46:1–46:??, February 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3628158.

Xue:2020:MOI

[XL20] Yinxing Xue and Yan-Fu Li. Multi-objective integer programming approaches for solving the multi-criteria test-suite minimization problem: Towards sound and complete solutions of a particular search-based softwareengineering problem. ACM Transactions on Software Engineering and Methodology, 29(3): 20:1–20:50, July 2020. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3392031.

Xie:2024:SCS

[XLD^{+}24] Yutao Xie, Jiayi Lin, Hande Dong, Lei Zhang, and Zhonghai Wu. Survey of code search based on deep learning. ACM Transactions on Software Engineering and Methodology, 33(2): 54:1–54:??, February 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3628161.

Xie:2022:NNP

[XLW^{+}22] Xiaofei Xie, Tianlin Li, Jian Wang, Lei Ma, Qing Guo, Felix Juefei-Xu, and Yang Liu. NPC: Neuron path coverage via characterizing decision logic of deep neural networks. ACM Transactions on Software Engineering and Methodology, 31 (3):47:1–47:27, July 2022. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3490489.

Xie:2007:DCA

[XM07] Qing Xie and Atif M. Memon. Designing and comparing automated test oracles for GUIbased software applications. ACM Transactions on Software Engineering and Methodology, 16(1):4:1–4:??, February 2007. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Xie:2008:UPS

[XM08] Qing Xie and Atif M. Memon.

Using a pilot study to derive a GUI model for automated testing. ACM Transactions on Software Engineering and Methodology, 18(2):7:1–7:??, November 2008. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Xu:2014:SRB

[XMA^{+}14] Guoqing Xu, Nick Mitchell, Matthew Arnold, Atanas Rountev, Edith Schonberg, and Gary Sevitsky. Scalable runtime bloat detection using abstract dynamic slicing. ACM Transactions on Software Engineering and Methodology, 23(3):23:1– 23:??, May 2014. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Xu:2013:PML

[XR13] Guoqing Xu and Atanas Rountev. Precise memory leak detection for Java software using container profiling. ACM Transactions on Software Engineering and Methodology, 22 (3):17:1–17:??, July 2013. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Xiao:2024:MEC

[XSA^{+}24] Ya Xiao, Wenjia Song, Salman Ahmed, Xinyang Ge, Bimal Viswanath, Na Meng, and Danfeng (Daphne) Yao. Measurement of embedding choices on cryptographic API completion tasks. ACM Transactions on Software Engineering and Methodology, 33(3): 56:1–56:??, March 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3625291.

Xu:2023:SFC

[XTW^{+}23] Xiong Xu, Jean-Pierre Talpin, Shuling Wang, Bohua Zhan, and Naijun Zhan. Semantics foundation for cyber-physical systems using higher-order UTP. ACM Transactions on Software Engineering and Methodology, 32(1): 9:1–9:??, January 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3517192.

Xu:2022:ICG

[XVN22] Frank F. Xu, Bogdan Vasilescu, and Graham Neubig. In-IDE code generation from natural language: Promise and challenges. ACM Transactions on Software Engineering and Methodology, 31(2): 29:1–29:47, April 2022. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3487569.

Xiong:2022:LFS

[XW22] Yingfei Xiong and Bo Wang. L2S: a framework for synthesizing the most probable program under a specification. ACM Transactions on Software Engineering and Methodology, 31 (3):34:1–34:45, July 2022. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3487570.

Xiang:2018:CSP

[XZZL18] Yi Xiang, Yuren Zhou, Zibin Zheng, and Miqing Li. Configuring software product lines by combining many-objective optimization and SAT solvers. ACM Transactions on Software Engineering and Methodology, 26(4): 14:1–14:??, February 2018. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Yu:2020:SCR

[YABLR20] Xiao Liang Yu, Omar Al-Bataineh, David Lo, and Abhik Roychoudhury. Smart contract repair. ACM Transactions on Software Engineering and Methodology, 29(4):27:1– 27:32, October 2020. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3402450.

Yuan:2020:TBE

[YB20] Yuan Yuan and Wolfgang Banzhaf. Toward better evolutionary program repair: an integrated approach. ACM Transactions on Software Engineering and Methodology, 29(1):5:1– 5:53, February 2020. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3360004.

Yue:2013:FTU

[YBL13] Tao Yue, Lionel C. Briand, and Yvan Labiche. Facilitating the transition from use case models to analysis models: Approach and experiments. ACM Transactions on Software Engineering and Methodology, 22(1): 5:1–5:??, February 2013. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Yue:2015:AAF

[YBL15] Tao Yue, Lionel C. Briand, and Yvan Labiche. aToucan: an automated framework to derive UML analysis models from use case models. ACM Transactions on Software Engineering and Methodology, 24(3):13:1–13:??, May 2015. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Yanez:2021:AIT

[YBZK21] Wendy Yánez, Rami Bahsoon, Yuqun Zhang, and Rick Kazman. Architecting Internet of Things systems with blockchain: a catalog of tactics. ACM Transactions on Software Engineering and Methodology, 30 (3):35:1–35:46, May 2021. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3442412.

Yang:2024:APE

[YDX^{+}24] Shouguo Yang, Chaopeng Dong, Yang Xiao, Yiran Cheng, Zhiqiang Shi, Zhi Li, and Limin Sun. Asteria-Pro: Enhancing deep learning-based binary code similarity detection by incorporating domain knowledge. ACM Transactions on Software Engineering and Methodology, 33(1): 1:1–1:??, January 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3604611.

Yang:2023:TIG

[YFD^{+}23] Jia Yang, Cai Fu, Fengyang Deng, Ming Wen, Xiaowei Guo, and Chuanhao Wan. Toward interpretable graph tensor convolution neural network for code semantics embedding. ACM Transactions on Software Engineering and Methodology, 32(5): 115:1–115:??, September 2023. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3582574.

Yoo:2013:FLP

[YHC13] Shin Yoo, Mark Harman, and David Clark. Fault localization prioritization: Comparing information-theoretic and coverage-based approaches. ACM Transactions on Software Engineering and Methodology, 22(3):19:1–19:??, July 2013. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Yang:2024:TTI

[YHG^{+}24] Yuanhang Yang, Wei He, Cuiyun Gao, Zenglin Xu, Xin Xia, and Chuanyi Liu. TopicAns: Topic-informed architecture for answer recommendation on technical Q&A site. ACM Transactions on Software Engineering and Methodology, 33(1): 20:1–20:??, January 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3607189.

Yu:2022:AIE

[YHL^{+}22] Hao Yu, Xing Hu, Ge Li, Ying Li, Qianxiang Wang, and Tao Xie. Assessing and improving an evaluation dataset for detecting semantic code clones via deep learning. ACM Transactions on Software Engineering and Methodology, 31(4): 62:1–62:??, October 2022. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3502852.

Yang:1992:PIA

[YHR92] Wuu Yang, Susan Horwitz, and Thomas Reps. A program integration algorithm that accommodates semantics-preserving transformations. ACM Transactions on Software Engineering and Methodology, 1(3):310–354, July 1992. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

                                        www.acm.org/pubs/articles/
                                        journals/tosem/1992-1-3/p310-

yang/p310-yang.pdf;   http:
//www.acm.org/pubs/citations/
journals/tosem/1992-1-3/p310-
yang/.

Yang:2024:LWC

[YHXY24] Yanming Yang, Xing Hu, Xin Xia, and Xiaohu Yang. The lost world: Characterizing and detecting undiscovered test smells. ACM Transactions on Software Engineering and Methodology, 33(3):59:1–59:??, March 2024. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3631973.

Yan:2020:AGS

[YJW^{+}20] Gaogao Yan, Li Jiao, Shuling Wang, Lingtai Wang, and Naijun Zhan. Automatically generating SystemC code from HCSP formal models. ACM Transactions on Software Engineering and Methodology, 29(1):4:1– 4:39, February 2020. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3360002.

Yang:2023:SCP

[YKY^{+}23] Zhen Yang, Jacky Wai Keung, Xiao Yu, Yan Xiao, Zhi Jin, and Jingyu Zhang. On the significance of category prediction for code-comment synchronization. ACM Transactions on Software Engineering and Methodology, 32(2):30:1–30:??, April 2023. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3534117.

Yang:2023:SWE

[YLM^{+}23] Deheng Yang, Yan Lei, Xiaoguang Mao, Yuhua Qi, and Xin Yi. Seeing the whole elephant: Systematically understanding and uncovering evaluation biases in automated program repair. ACM Transactions on Software Engineering and Methodology, 32(3): 65:1–65:??, July 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3561382.

Yang:2014:DIS

[YPRK14] Guowei Yang, Suzette Person, Neha Rungta, and Sarfraz Khurshid. Directed incremental symbolic execution. ACM Transactions on Software Engineering and Methodology, 24(1): 3:1–3:??, September 2014. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Yi:2015:SCC

[YQTR15] Jooyong Yi, Dawei Qi, Shin Hwei Tan, and Abhik Roychoudhury. Software change contracts. ACM Transactions on Software Engineering and Methodology, 24(3):18:1–18:??, May 2015. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Young:1995:CAT

[YTL^{+}95] Michal Young, Richard N. Taylor, David L. Levine, Kari A. Nies, and Debra Brodbeck. A concurrency analysis tool suite for Ada programs: rationale, design, and preliminary experience. ACM Transactions on Software Engineering and Methodology, 4(1):65–106, January 1995. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1995-4-1/p65-
young/p65-young.pdf;   http:
//www.acm.org/pubs/citations/
journals/tosem/1995-4-1/p65-
young/.

Yu:2023:STQ

[Yu23] Nengkun Yu. Structured theorem for quantum programs and its applications. ACM Transactions on Software Engineering and Methodology, 32(4): 103:1–103:??, July 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3587154.

Yang:2016:HPP

[YWC16] Chunbai Yang, Shangru Wu, and W. K. Chan. Hierarchical program paths. ACM Transactions on Software Engineering and Methodology, 25(3): 27:1–27:??, August 2016. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Yoo:2017:HCG

[YXK^{+}17] Shin Yoo, Xiaoyuan Xie, Fei-Ching Kuo, Tsong Yueh Chen, and Mark Harman. Human competitiveness of genetic programming in spectrum-based fault localisation: Theoretical and empirical analysis. ACM Transactions on Software Engineering and Methodology, 26 (1):4:1–4:??, July 2017. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Yang:2022:PMS

[YXL^{+}22] Yanming Yang, Xin Xia, David Lo, Tingting Bi, John Grundy, and Xiaohu Yang. Predictive models in software engineering: Challenges and opportunities. ACM Transactions on Software Engineering and Methodology, 31(3): 56:1–56:72, July 2022. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3503509.

Yin:2023:SEV

[YXM^{+}23] Zijing Yin, Yiwen Xu, Fuchen Ma, Haohao Gao, Lei Qiao, and Yu Jiang. Scanner++: Enhanced vulnerability detection of Web applications with attack intent synchronization. ACM Transactions on Software Engineering and Methodology, 32(1): 7:1–7:??, January 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3517036.

Yang:2023:TPB

[YXX^{+}23] Shouguo Yang, Zhengzi Xu, Yang Xiao, Zhe Lang, Wei Tang, Yang Liu, Zhiqiang Shi, Hong Li, and Limin Sun. Towards practical binary code similarity detection: Vulnerability verification via patch semantic analysis. ACM Transactions on Software Engineering and Methodology, 32(6):158:1– 158:??, November 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3604608.

Yitagesu:2023:EPB

[YXZ^{+}23] Sofonias Yitagesu, Zhenchang Xing, Xiaowang Zhang, Zhiyong Feng, Xiaohong Li, and Linyi Han. Extraction of phrasebased concepts in vulnerability descriptions through unsupervised labeling. ACM Transactions on Software Engineering and Methodology, 32(5):112:1– 112:??, September 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3579638.

Yang:2022:DDR

[YZP^{+}22] Wenhua Yang, Chong Zhang, Minxue Pan, Chang Xu, Yu Zhou, and Zhiqiu Huang. Do developers really know how to use Git commands? A large-scale study using Stack Overflow. ACM Transactions on Software Engineering and Methodology, 31(3):44:1–44:29, July 2022. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3494518.

Yang:2024:HIG

[YZY^{+}24] Guang Yang, Yu Zhou, Wenhua Yang, Tao Yue, Xiang Chen, and Taolue Chen. How important are good method names in neural code generation? A model robustness perspective. ACM Transactions on Software Engineering and Methodology, 33(3):60:1–60:??, March 2024. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3630010.

Zhang:2022:AHS

[ZA22] Man Zhang and Andrea Arcuri. Adaptive hypermutation for search-based system test generation: a study on REST APIs with EvoMaster. ACM Transactions on Software Engineering and Methodology, 31(1): 2:1–2:52, January 2022. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3464940.

Zhang:2023:OPF

[ZA23] Man Zhang and Andrea Arcuri. Open problems in fuzzing RESTful APIs: a comparison of tools. ACM Transactions on Software Engineering and Methodology, 32(6):144:1– 144:??, November 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3597205.

Zhang:2023:WBF

[ZAL^{+}23] Man Zhang, Andrea Arcuri, Yonggang Li, Yang Liu, and Kaiming Xue. White-box fuzzing RPC-based APIs with EvoMaster: an industrial case study. ACM Transactions on Software Engineering and Methodology, 32(5):122:1– 122:??, September 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3585009.

Zave:2004:ATT

[Zav04] Pamela Zave. Address translation in telecommunication features. ACM Transactions on Software Engineering and Methodology, 13(1):1–36, January 2004. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Zeil:1992:DLE

[ZAW92] Steven J. Zeil, Faten H. Afifi, and Lee J. White. Detection of linear errors via domain testing. ACM Transactions on Software Engineering and Methodology, 1(4):422–451, October 1992. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/

                                        journals/tosem/1992-1-4/p422-
                                        zeil/p422-zeil.pdf;   http:
                                        //www.acm.org/pubs/citations/
                                        journals/tosem/1992-1-4/p422-
                                        zeil/.

Zhu:2013:ADP

[ZB13] Hong Zhu and Ian Bayley. An algebra of design patterns. ACM Transactions on Software Engineering and Methodology, 22 (3):23:1–23:??, July 2013. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Zhang:2023:JSH

[ZBA23] Man Zhang, Asma Belhadi, and Andrea Arcuri. JavaScript SBST heuristics to enable effective fuzzing of NodeJS web APIs. ACM Transactions on Software Engineering and Methodology, 32(6):139:1– 139:??, November 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3593801.

Zhou:2022:PTP

[ZCH22] Jianyi Zhou, Junjie Chen, and Dan Hao. Parallel test prioritization. ACM Transactions on Software Engineering and Methodology, 31(1): 8:1–8:50, January 2022. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3471906.

Zhao:2024:ADU

[ZCL^{+}24] Zhe Zhao, Guangke Chen, Tong Liu, Taishan Li, Fu Song, Jingyi Wang, and Jun Sun. Attack as detection: Using adversarial attack methods to detect abnormal examples. ACM Transactions on Software Engineering and Methodology, 33(3): 68:1–68:??, March 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3631977.

Zheng:2018:MAI

[ZCT18] Yongjie Zheng, Cuong Cu, and Richard N. Taylor. Maintaining architecture-implementation conformance to support architecture centrality: From single system to product line development. ACM Transactions on Software Engineering and Methodology, 27(2): 8:1–8:??, July 2018. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Zhang:2024:CDW

[ZCW^{+}24] Yixuan Zhang, Shangtong Cao, Haoyu Wang, Zhenpeng Chen, Xiapu Luo, Dongliang Mu, Yun Ma, Gang Huang, and Xuanzhe Liu. Characterizing and detecting WebAssembly runtime bugs. ACM Transactions on Software Engineering and Methodology, 33(2):37:1– 37:??, February 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3624743.

Zheng:2023:SES

[ZCZ^{+}23] Zibin Zheng, Weili Chen, Zhijie Zhong, Zhiguang Chen, and Yutong Lu. Securing the Ethereum from smart Ponzi schemes: Identification using static features. ACM Transactions on Software Engineering and Methodology, 32(5):130:1– 130:??, September 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3571847.

Zhang:2014:ATV

[ZE14] Pingyu Zhang and Sebastian Elbaum. Amplifying tests to validate exception handling code: an extended study in the mobile application domain. ACM Transactions on Software Engineering and Methodology, 23(4): 32:1–32:??, August 2014. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Zhang:2022:TRD

[ZFL^{+}22] Huangzhao Zhang, Zhiyi Fu, Ge Li, Lei Ma, Zhehao Zhao, Hua’an Yang, Yizhe Sun, Yang Liu, and Zhi Jin. Towards robustness of deep program processing models-detection, estimation, and enhancement. ACM Transactions on Software Engineering and Methodology, 31(3):50:1–50:40, July 2022. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3511887.

Zhang:2024:SLB

[ZFM^{+}24] Quanjun Zhang, Chunrong Fang, Yuxiang Ma, Weisong Sun, and Zhenyu Chen. A survey of learning-based automated program repair. ACM Transactions on Software Engineering and Methodology, 33 (2):55:1–55:??, February 2024. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3631974.

Zirak:2024:IAP

[ZH24] Armin Zirak and Hadi Hemmati. Improving automated program repair with domain adaptation. ACM Transactions on Software Engineering and Methodology, 33(3): 65:1–65:??, March 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3631972.

Zhu:2024:CAT

[ZHC24] Mingxuan Zhu, Dan Hao, and Junjie Chen. Compiler autotuning through multiplephase learning. ACM Transactions on Software Engineering and Methodology, 33(4): 100:1–100:??, May 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3640330.

Zhang:2018:ESM

[ZHO^{+}18] Yuanyuan Zhang, Mark Harman, Gabriela Ochoa, Guenther Ruhe, and Sjaak Brinkkemper. An empirical study of meta- and hyper-heuristic search for multi-objective release planning. ACM Transactions on Software Engineering and Methodology, 27(1): 3:1–3:??, June 2018. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Zhang:2023:DDA

[ZHS^{+}23] Xueling Zhang, John Heaps, Rocky Slavin, Jianwei Niu, Travis Breaux, and Xiaoyin Wang. DAISY: Dynamicanalysis-induced source discovery for sensitive data. ACM Transactions on Software Engineering and Methodology, 32 (4):89:1–89:??, July 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3569936.

Zhang:2023:DBR

[ZHV^{+}23] Ting Zhang, Donggyun Han, Venkatesh Vinayakarao, Ivana Clairine Irsan, Bowen Xu, Ferdian Thung, David Lo, and Lingxiao Jiang. Duplicate bug report detection: How far are we? ACM Transactions on Software Engineering and Methodology, 32 (4):97:1–97:??, July 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3576042.

Zhou:2024:AFC

[ZHXZ24] Anshunkang Zhou, Yikun Hu, Xiangzhe Xu, and Charles Zhang. ARCTURUS: Full coverage binary similarity analysis with reachability-guided emulation. ACM Transactions on Software Engineering and Methodology, 33(4):96:1–96:??, May 2024. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3640337.

Zhao:2021:PPA

[ZHZ^{+}21] Guoliang Zhao, Safwat Hassan, Ying Zou, Derek Truong, and Toby Corbin. Predicting performance anomalies in software systems at run-time. ACM Transactions on Software Engineering and Methodology, 30 (3):33:1–33:33, May 2021. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3440757.

Zave:1993:CC

[ZJ93] Pamela Zave and Michael Jackson. Conjunction as composition. ACM Transactions on Software Engineering and Methodology, 2(4):379–411, October 1993. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1993-2-4/p379-
zave/p379-zave.pdf;   http:
//www.acm.org/pubs/citations/

journals/tosem/1993-2-4/p379-
zave/.

Zave:1997:FDC

[ZJ97] Pamela Zave and Michael Jackson. Four dark corners of requirements engineering. ACM Transactions on Software Engineering and Methodology, 6(1):1–30, January 1997. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1997-6-1/p1-
zave/p1-zave.pdf;   http:/
/www.acm.org/pubs/citations/
journals/tosem/1997-6-1/p1-
zave/.

Zambonelli:2003:DMS

[ZJW03] Franco Zambonelli, Nicholas R. Jennings, and Michael Wooldridge. Developing multiagent systems: The Gaia methodology. ACM Transactions on Software Engineering and Methodology, 12 (3):317–370, July 2003. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Zhang:2023:FCPa

[ZKW^{+}23a] Zenong Zhang, George Klees, Eric Wang, Michael Hicks, and Shiyi Wei. Fuzzing configurations of program options. ACM Transactions on Software Engineering and Methodology, 32 (2):53:1–53:??, April 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3580597.

Zhang:2023:FCPb

[ZKW^{+}23b] Zenong Zhang, George Klees, Eric Wang, Michael Hicks, and Shiyi Wei. Fuzzing configurations of program options — RCR report. ACM Transactions on Software Engineering and Methodology, 32(2): 55:1–55:??, April 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3580601.

Zheng:2013:PRP

[ZL13] Zibin Zheng and Michael R.

Lyu. Personalized reliability prediction of Web services. ACM Transactions on Software Engineering and Methodology, 22(2):12:1–12:??, March 2013. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Zhang:2023:AIR

[ZLL^{+}23] Jingxuan Zhang, Junpeng Luo, Jiahui Liang, Lina Gong, and Zhiqiu Huang. An accurate identifier renaming prediction and suggestion approach. ACM Transactions on Software Engineering and Methodology, 32(6): 148:1–148:??, November 2023. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3603109.

Zhu:2024:DBE

[ZLP^{+}24] Tingwei Zhu, Zhong Li, Minxue Pan, Chaoxuan Shi, Tian Zhang, Yu Pei, and Xuandong Li. Deep is better? An empirical comparison of information retrieval and deep learning approaches to code summarization. ACM Transactions on Software Engineering and Methodology, 33(3): 67:1–67:??, March 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3631975.

Zhang:2023:IGL

[ZLS^{+}23] Zhuo Zhang, Yan Lei, Ting Su, Meng Yan, Xiaoguang Mao, and Yue Yu. Influential global and local contexts guided trace representation for fault localization. ACM Transactions on Software Engineering and Methodology, 32(3):78:1–78:??, July 2023. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3576043.

Zhang:2022:TCO

[ZLT^{+}22] Yuxia Zhang, Hui Liu, Xin Tan, Minghui Zhou, Zhi Jin, and Jiaxin Zhu. Turnover of companies in OpenStack: Prevalence and rationale. ACM Transactions on Software Engineering and Methodology, 31(4): 75:1–75:??, October 2022. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3510849.

Zhao:2021:ISD

[ZLW^{+}21] Yanjie Zhao, Li Li, Haoyu Wang, Haipeng Cai, Tegawendé F. Bissyandé, Jacques Klein, and John Grundy. On the impact of sample duplication in machinelearning-based Android malware detection. ACM Transactions on Software Engineering and Methodology, 30(3): 40:1–40:38, May 2021. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3446905.

Zhou:2024:CVD

[ZMC^{+}24] Yuanhang Zhou, Fuchen Ma, Yuanliang Chen, Meng Ren, and Yu Jiang. CLFuzz: Vulnerability detection of cryptographic algorithm implementation via semantic-aware fuzzing. ACM Transactions on Software Engineering and Methodology, 33(2):45:1–45:??, February 2024. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3628160.

Zhou:2016:IRO

[ZMM^{+}16] Minghui Zhou, Audris Mockus, Xiujuan Ma, Lu Zhang, and Hong Mei. Inflow and retention in OSS communities with commercial involvement: a case study of three hybrid projects. ACM Transactions on Software Engineering and Methodology, 25(2):13:1–13:??, May 2016. CODEN ATSMER.

ISSN 1049-331X (print), 1557-7392 (electronic).

Zuo:2023:TME

[ZNZ^{+}23] Zhiqiang Zuo, Xintao Niu, Siyi Zhang, Lu Fang, Siau Cheng Khoo, Shan Lu, Chengnian Sun, and Guoqing Harry Xu. Toward more efficient statistical debugging with abstraction refinement. ACM Transactions on Software Engineering and Methodology, 32(2): 36:1–36:??, April 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3544790.

Zarir:2021:DCE

[ZOJH21] Abdullah A. Zarir, Gustavo A. Oliva, Zhen M. (Jack) Jiang, and Ahmed E. Hassan. Developing cost-effective blockchainpowered applications: a case study of the gas usage of smart contract transactions in the Ethereum blockchain platform. ACM Transactions on Software Engineering and Methodology, 30(3):28:1–28:38, May 2021. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3431726.

Zdun:2023:MSM

[ZQS^{+}23] Uwe Zdun, Pierre-Jean Queval, Georg Simhandl, Riccardo Scandariato, Somik Chakravarty, Marjan Jelic, and Aleksandar Jovanovic. Microservice security metrics for secure communication, identity management, and observability. ACM Transactions on Software Engineering and Methodology, 32(1): 16:1–16:??, January 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3532183.

Zohdinasab:2023:EEF

[ZRGT23] Tahereh Zohdinasab, Vincenzo Riccio, Alessio Gambi, and Paolo Tonella. Efficient and effective feature space exploration for testing deep learning systems. ACM Transactions on Software Engineering and Methodology, 32(2): 49:1–49:??, April 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3544792.

Zeller:1997:UVT

[ZS97] Andreas Zeller and Gregor Snelting. Unified versioning through feature logic. ACM Transactions on Software Engineering and Methodology, 6 (4):398–441, October 1997. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

www.acm.org/pubs/articles/
journals/tosem/1997-6-4/p398-
zeller/p398-zeller.pdf;
http://www.acm.org/pubs/citations/
journals/tosem/1997-6-4/p398-
zeller/.

Zhang:2020:MTC

[ZSHD20] Xihui Zhang, Thomas F. Stafford, Tao Hu, and Hua Dai. Measuring task conflict and person conflict in software testing. ACM Transactions on Software Engineering and Methodology, 29(4):25:1–25:19, October 2020. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3395029.

Zhang:2013:CDC

[ZSL^{+}13] Wei Zhang, Chong Sun, Junghee Lim, Shan Lu, and Thomas Reps. ConMem: Detecting crash-triggering concurrency bugs through an effectoriented approach. ACM Transactions on Software Engineering and Methodology, 22(2):10:1– 10:??, March 2013. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Zhao:2022:RAE

[ZSL^{+}22] Yu Zhao, Ting Su, Yang Liu, Wei Zheng, Xiaoxue Wu, Ramakanth Kavuluru, William G. J. Halfond, and Tingting Yu. ReCDroid+: Automated endto-end crash reproduction from bug reports for Android apps. ACM Transactions on Software Engineering and Methodology, 31(3):36:1–36:33, July 2022. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3488244.

Zhou:2022:SAI

[ZSW^{+}22] Yaqin Zhou, Jing Kai Siow, Chenyu Wang, Shangqing Liu, and Yang Liu. SPI: Automated identification of security patches via commits. ACM Transactions on Software Engineering and Methodology, 31(1): 13:1–13:27, January 2022. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3468854.

Zhang:2023:TIT

[ZSWS23] Mengdi Zhang, Jun Sun, Jingyi Wang, and Bing Sun. TestSGD: Interpretable testing of neural networks against subtle group discrimination. ACM Transactions on Software Engineering and Methodology, 32(6):137:1– 137:??, November 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3591869.

Zampetti:2023:CID

[ZTP^{+}23] Fiorella Zampetti, Damian Tamburri, Sebastiano Panichella, Annibale Panichella, Gerardo Canfora, and Massimiliano Di Penta. Continuous integration and delivery practices for cyberphysical systems: an interviewbased study. ACM Transactions on Software Engineering and Methodology, 32(3): 73:1–73:??, July 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3571854.

Zaremski:1995:SMT

[ZW95] Amy Moormann Zaremski and Jeannette M. Wing. Signature matching: a tool for using software libraries. ACM Transactions on Software Engineering and Methodology, 4(2):146–170, April 1995. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

                                        www.acm.org/pubs/articles/
                                        journals/tosem/1995-4-2/p146-
                                        zaremski/p146-zaremski.pdf;
                                         http://www.acm.org/pubs/
                                        citations/journals/tosem/
                                        1995-4-2/p146-zaremski/.

Zaremski:1997:SMS

[ZW97] Amy Moormann Zaremski and Jeannette M. Wing. Specification matching of software components. ACM Transactions on Software Engineering and Methodology, 6(4):333–369, October 1997. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL http://

                                        www.acm.org/pubs/articles/
                                        journals/tosem/1997-6-4/p333-
                                        zaremski/p333-zaremski.pdf;
                                         http://www.acm.org/pubs/
                                        citations/journals/tosem/
                                        1997-6-4/p333-zaremski/.

Zhang:2021:CSO

[ZWCH21] Haoxiang Zhang, Shaowei Wang, Tse-Hsun (Peter) Chen, and Ahmed E. Hassan. Are comments on Stack Overflow well organized for easy retrieval by developers? ACM Transactions on Software Engineering and Methodology, 30(2): 22:1–22:31, March 2021. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3434279.

Zou:2022:BSP

[ZWGX22] Changwei Zou, Xudong Wang, Yaoqing Gao, and Jingling Xue. Buddy stacks: Protecting return addresses with efficient thread-local storage and runtime re-randomization. ACM Transactions on Software Engineering and Methodology, 31(2): 35e:1–35e:37, April 2022. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3494516.

Zhang:2022:MRE

[ZWL^{+}22] Peng Zhang, Yang Wang, Xutong Liu, Yanhui Li, Yibiao Yang, Ziyuan Wang, Xiaoyu Zhou, Lin Chen, and Yuming Zhou. Mutant reduction evaluation: What is there and what is missing? ACM Transactions on Software Engineering and Methodology, 31(4): 69:1–69:??, October 2022. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3522578.

Zhang:2024:AET

[ZWL^{+}24] Peng Zhang, Yang Wang, Xutong Liu, Zeyu Lu, Yibiao Yang, Yanhui Li, Lin Chen, Ziyuan Wang, Chang-Ai Sun, Xiao Yu, and Yuming Zhou. Assessing effectiveness of test suites: What do we know and what should we do? ACM Transactions on Software Engineering and Methodology, 33 (4):86:1–86:??, May 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3635713.

Zhu:2024:SAG

[ZWT^{+}24] Hengcheng Zhu, Lili Wei, Valerio Terragni, Yepang Liu, Shing-Chi Cheung, Jiarong Wu, Qin Sheng, Bing Zhang, and Lihong Song. StubCoder: Automated generation and repair of stub code for Mock objects. ACM Transactions on Software Engineering and Methodology, 33(1): 16:1–16:??, January 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3617171.

Zou:2021:IAM

[ZWY^{+}21] Deqing Zou, Yueming Wu, Siru Yang, Anki Chauhan, Wei Yang, Jiangying Zhong, Shihan Dou, and Hai Jin. Int-Droid: Android malware detection based on API intimacy analysis. ACM Transactions on Software Engineering and Methodology, 30(3):39:1–39:32, May 2021. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3442588.

Zheng:2024:CLS

[ZWZ^{+}24] Xiaoye Zheng, Zhiyuan Wan, Yun Zhang, Rui Chang, and David Lo. A closer look at the security risks in the Rust ecosystem. ACM Transactions on Software Engineering and Methodology, 33(2):34:1– 34:??, February 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3624738.

Zhou:2014:DSP

[ZXLC14] Yuming Zhou, Baowen Xu, Hareton Leung, and Lin Chen. An in-depth study of the potentially confounding effect of class size in fault prediction. ACM Transactions on Software Engineering and Methodology, 23(1): 10:1–10:??, February 2014. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Zhi:2024:SST

[ZXS^{+}24] Yuhan Zhi, Xiaofei Xie, Chao Shen, Jun Sun, Xiaoyu Zhang, and Xiaohong Guan. Seed selection for testing deep neural networks. ACM Transactions on Software Engineering and Methodology, 33(1):23:1– 23:??, January 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3607190.

Zhou:2018:HFW

[ZYL^{+}18] Yuming Zhou, Yibiao Yang, Hongmin Lu, Lin Chen, Yanhui Li, Yangyang Zhao, Junyan Qian, and Baowen Xu. How far we have progressed in the journey? An examination of crossproject defect prediction. ACM Transactions on Software Engineering and Methodology, 27 (1):1:1–1:??, June 2018. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Zeng:2023:DEV

[ZYL^{+}23] Chen Zeng, Yue Yu, Shanshan Li, Xin Xia, Zhiming Wang, Mingyang Geng, Linxiao Bai, Wei Dong, and Xiangke Liao. deGraphCS: Embedding variable-based flow graph for neural code search. ACM Transactions on Software Engineering and Methodology, 32 (2):34:1–34:??, April 2023. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3546066.

Zhang:2023:PQA

[ZZC^{+}23] Yedi Zhang, Zhe Zhao, Guangke Chen, Fu Song, and Taolue Chen. Precise quantitative analysis of binarized neural networks: a BDD-based approach. ACM Transactions on Software Engineering and Methodology, 32(3):62:1–62:??, July 2023. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3563212.

Zhao:2006:STS

[ZZL^{+}06] Wei Zhao, Lu Zhang, Yin Liu, Jiasu Sun, and Fuqing Yang. SNIAFL: Towards a static noninteractive approach to feature location. ACM Transactions on Software Engineering and Methodology, 15(2):195– 226, April 2006. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Zhou:2024:DDR

[ZZL^{+}24] Yu Zhou, Weilin Zhan, Zi Li, Tingting Han, Taolue Chen, and Harald Gall. DRIVE: Dockerfile rule mining and violation detection. ACM Transactions on Software Engineering and Methodology, 33(2):30:1– 30:??, February 2024. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3617173.

Zhou:2022:ARD

[ZZS^{+}22] Yu Zhou, Xiaoqing Zhang, Juanjuan Shen, Tingting Han, Taolue Chen, and Harald Gall. Adversarial robustness of deep code comment generation. ACM Transactions on Software Engineering and Methodology, 31(4): 60:1–60:??, October 2022. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3501256.

Zhu:2023:CSU

[ZZW^{+}23] Chenguang Zhu, Mengshi Zhang, Xiuheng Wu, Xiufeng Xu, and Yi Li. Client-specific upgrade compatibility checking via knowledge-guided discovery. ACM Transactions on Software Engineering and Methodology, 32(4):98:1–98:??, July 2023. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3582569.

Zou:2021:IDL

[ZZX^{+}21] Deqing Zou, Yawei Zhu, Shouhuai Xu, Zhen Li, Hai Jin, and Hengkai Ye. Interpreting deep learning-based vulnerability detector predictions based on heuristic searching. ACM Transactions on Software Engineering and Methodology, 30(2): 23:1–23:31, March 2021. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3429444.

Zhang:2021:UWR

[ZZY^{+}21] Huihui Zhang, Man Zhang, Tao Yue, Shaukat Ali, and Yan Li. Uncertainty-wise requirements prioritization with search. ACM Transactions on Software Engineering and Methodology, 30(1): 4:1–4:54, January 2021. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3408301.